A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1980–1984

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.04

Title word cross-reference

#11812 [1033]. #17627 [300].

(+ + −−) [591, 1129, 1288]. (−∞, ∞) [2203]. (1 − X^{2}) [1832]. (2 + 1) [344]. (4 + 1) [1674]. (ad^{⊗k})_{s} [2109]. (Δ − m^{2})u = 0 [834]. (JJ^{∗})^{−1} [870]. (∂/∂x)[a(u + b)^{−2??}(∂u/∂x)] − (∂u/∂t) = 0 [170]. (Ψ(z) − z)^{−1} [998]. (Sp(4): 4) [1039]. (x, t) [2015]. ∗ [1325, 658]. 0 [601]. 1 [492, 678]. 1 + 1 [1602, 1895, 968]. 1 + 3 [968]. 1 − exp[−a(r − r_{e})] [1342]. 1/2 [1692, 637, 101, 1752, 238, 404, 467, 529, 1793, 601, 1478]. 1/N [285, 1845, 410]. 1/r [419, 1942]. 2 [160, 1690, 957, 678]. 2 + 1 [2093]. 2 + 2 [1371]. 2ω [650]. 2 × 2 [2161]. 2 × 2 × N [1557]. 2 × N [1734]. 3 [2104, 13, 466, 917, 1391, 109, 245]. 3 + 1 [2085, 1623]. 3 × N [1223]. 4 [1221, 1875, 561, 591, 1129, 1288, 1050, 944]. 5 [1876]. 6 [2104, 1811, 1350, 109, 245]. 6j [2071]. 9 [801]. [SU(2)]^{3} [2125, 2124]. ^{+} [643]. − [1452]. ^{−1} [721]. _{2} [558]. A [1009, 385]. A(0, n) [208]. A(m, n) [764]. A_{1}, [1627]. A_{2}, [1627]. A_{3}, [1627]. αx^{2} + βx^{4} [1840]. B^{∗} [1683]. B_{2}, [1627]. β [1981]. c [323]. C^{∗} [799]. E_{6} ⊗ U(1) [1097]. H [544, 565, 487]. HH [544, 1099, 912]. CC^{∗} [902]. : exp ξ:_{d} [1795]. CP^{N} [494]. CP_{2}^{n−1} [1464]. D [1220, 709, 1974, 913, 1118, 1643, 2213, 1973, 1196, 1934]. d = 11 [2158]. D = 8 [1915]. r̈ = l̂_{r}{[Ü(t)/U(t)]r − [μ_{0}/U(t)]r^{−2}} [1542]. E_{1}(x) [1236]. E_{6} [726, 2110, 2071]. E_{7} [726]. E_{8} [726]. exp[i(ω/c)(r − ct)]/r [1869]. F(4) [1188]. F_{4} [726, 1811]. G [837]. G(3) [1188]. g − 2 [610]. G_{2} [1082, 855, 1197, 1627, 301, 796]. GL(4, R) [2208]. GL(n, R) [859]. Gln?? [55]. Gsl(2) [1884]. G ⊇ [SU(2)]^{n} [1881]. H [985, 82, 255, 1209, 509, 1262, 1158]. i^{∘} [2085]. III ⊗ III [167]. II ⊗ II [165]. II ⊗ III [166]. ∫_{0}^{1} r|v(r)| dr + ∫_{1}^{∞} |v(r)| dr < ∞ [214]. I ⊗ I [162]. I ⊗ II [163]. I ⊗ III [164]. ISO(n) [19, 686]. IU(n) [583]. J [906, 1742, 2104, 1875, 1811, 801, 1391, 1350, 1876, 2184, 109, 245]. J^{αβ} [592]. jm [2104]. K [424]. K_{3} [500]. l [485, 1681, 271, 1803, 1472, 1473]. L^{2} [267]. L_{w}^{l/2}(R^{l}) [599]. λ [913]. λP (φ) [341]. LCRG [1238]. m [409]. M_{4} [380, 171, 536]. N [1175, 1839, 1962, 1312, 2021, 842, 1267, 963, 1857, 396, 318, 1597, 1535, 45, 542, 1942, 610, 1491, 1497, 2113, 896, 1085]. n + 1 [2006]. N = 2 [1101]. n_{t} = (n^{−1}n_{x})_{x} [1043]. N ≥ 1 [804]. O(3) [1667, 802, 803, 1433]. O(3, 3) [380]. O(4) [1432, 196]. O(6) ⊇ SO(2)SU(3) ⊇ SO(3) [1907]. O(n) [86, 805, 1435]. O(p) ↓ O(p − 2) × O(2) [1534]. O(p, q) [351, 56]. O^{+}(2, 2) [2000]. O_{l}^{0} [795, 293, 294, 1484]. osp(1, 2) [949]. Osp(2, 1) [471]. OSP(3, 2) [2177]. P [122, 168]. P(1, 4) [2176]. P (h, T) [2212]. P (φ) [705]. P (φ)_{2} [1184]. P_{l}^{0} [1082]. p_{α} [592]. ∂^{2}u/∂x∂t = f(u)and∂u/∂t = ∂^{2}u/∂x^{2} + f(u) [1730]. {∂_{xx} + 2i∂_{t} − 2g_{2}(t)x^{2} − 2g_{1}(t)x − 2g_{0}(t)}Ψ(x, t) = 0 [901]. πN [1079]. P n3′n [305]. R [444, 1437, 1027]. R(2λ + 1) [397]. R(3) [91, 397, 796]. R(5) [292]. R(5)O_{l}^{0} [398]. R(5)SU(2) ⊗ SU(2) [795]. R(7) [302, 1483, 1484]. R + R^{2} [1686]. R^{4} [1106]. R^{n} [1964, 1656]. ρ [1828]. S [1297, 1376, 1471, 1689, 1470, 385, 2142, 836, 1365, 223]. S(f_{1} + f_{2}) ⊇ S(f_{1}) × S(f_{2}) [438]. S^{∗} [497]. S_{L} [1348]. σ [1753, 1501]. Sl(2) ∧ T_{2} [861]. SL(2, C) [1077, 1624]. SL(2, R) [925, 117]. SL(3, R) [210, 12, 436]. SL_{4}(R) [205]. SO(10) [2110]. SO(2) [1890, 1989]. SO(2, 1) [111]. SO(2N) [1388, 1762]. SO(2n) → SO(2n − 2) ⊗ U(1) [2002]. SO(3) [1814]. SO(3) ⊗ SO(3) [1811]. SO(3, 1) [585, 1568]. SO(3, 2) [687]. SO(4) [1432, 1923]. so(4, 2) [308]. SO(4, 2) [687]. SO(5) [1431]. SO(7) [1610, 1877, 2125, 2124]. SO(N) [1267, 1878, 1395, 859, 261]. SO(n + 1) [686]. SO(n, 1) [116, 686]. SO^{∗}(2n) [1762]. Sp(2n) ↓ Sp(2n − 2) [115]. sp(3, R) [154]. Sp(3, R) × O(N − 1) [2210]. sp(4N; r)[gl(4N; r)] [804]. Sp(6) [1880, 854]. Sp(6) ⊇ Sp(2) × O(3) [2105]. Sp(6) ⊇ U(3) [1880]. Sp(n) [1319, 1762]. Sp(n, R) [1319]. Sp(p) × Sp(q) [1810]. Sp(p + q) [1810]. Sp(p, q) [1810]. Spin(2, 1) [926]. Spin(2, 2) [926]. spl(2, 1) [206, 207]. SU(1, 1) [1195]. SU(2) [197, 1132, 1379, 749, 89, 1076, 473, 2028, 1812]. SU(2) ⊗ U(N) [804]. SU(2) × SU(2) [1431]. SU(2n) [1762]. SU(3) [684, 1194, 802, 803, 1108, 1763, 1164, 1732, 1025, 306, 1199, 1433]. SU(3) ⊇ R(3) [1396]. SU(3) ⊇ SO(3) [353]. SU(3) × U(1) [854]. SU(5) [1080].

SU(5) ⊇ SU(3) × SU(2) × U(1) [1569]. SU(8) [2158]. SU(M) [1231]. SU(m/n) [1607, 1307]. SU(mn) ⊇ SU(m) × SU(n) [438]. SU(N) [1273, 180, 755, 1263, 2013, 372, 2104, 859, 1753, 800, 352]. SU(N) × SU(N) [198]. SU(N + M) [1087]. SU(N/M) [1087]. SU(n|1) [1935]. SU(p, q) [1999]. SU^{∗}(2n) [1762]. SU_{2} [84]. SU_{2} × SU_{2} [1946]. SUn [1651, 1348]. T [727, 1663, 1065, 1256, 920, 1472, 1473, 1455, 2133, 2132, 2220, 1079, 502, 503]. T^{∗}Q ⊕ TQ [1659]. U(1) [675]. U(2) [754]. U(2)σC^{2×2} [303]. u(3) [1637]. U(4) [1490]. U(6) ⊇ SU(3) ⊇ SO(3) ⊇ SO(2) [676]. U(6) ⊇ U(3) ⊇ U(2) ⊇ U(1) [676]. U(6) ∧ SU(3) [1756]. U(N) [1987, 406, 1263, 410, 2104, 86, 474, 440, 247, 1319, 1762]. U(N/M) [1987]. U(p) × U(q) [1109]. U(p + q) [1109]. U(p + q) ↓ U(p) ⊗ U(q) [1086]. U(p, q) [1109]. V [2098]. V (r) = (sin r)/r^{β}, 1/3 < β ≤ 1/2 [1018]. V^{∗} [2077]. ε [97, 1670]. φ^{3} [719]. φ^{4} [1499]. φ_{4}^{4} [1270]. x^{2} + λx^{2}/(1 + gx^{2}) [422]. x^{2N} [1322]. x^{α} [592]. xy [347]. y_{LM}[(r_{1} ⋀ r_{2}) ⋀ r_{3}] [1440]. y_{t} + y_{xxx} − 6y^{2}y_{x} + 6λy_{x} = 0 [1899]. Z [683, 63]. Z_{2} [386]. Z_{2} × Z_{2} [386]. Z_{2} × Z_{2} × Z_{2} [386]. Z_{n} [1650]. Z_{N} × Z_{N} [1200]. |x|^{−(n+1)/2} [1964]. |x|^{ν} [1161, 903].

* [1058]. *-quantization [1058].

-algebra [799, 1683]. -algebras [658, 2077]. -anharmonic [1322]. -basis [500]. -bell [1981]. -body [1175, 1962, 1312, 963]. -boson [1637]. -bubble [610]. -bug [542]. -classes [1501]. -component [917]. -de [1674]. -dependent [267]. -depending [2015]. -dimensional [344, 2113, 1839, 1267, 1535, 45]. -equation [255, 1209, 509, 1158]. -equations [82]. -field [1027]. -fields [385]. -forms [957]. -gravitating [567]. -invariant [1077, 1624]. -manifolds [591, 1129, 1288]. -matrix [1471, 1689, 1470, 1663, 385, 1065, 1256, 2142, 836, 906, 1742]. -metacyclic [683]. -model [1753]. -momenta [1221]. -number [323]. -parameter [160]. -photon [409]. -potentials [599]. -proper [1009]. -separable [444]. -separation [1437]. -space [561, 487]. -spaces [565]. -supermanifolds [1828]. -surfaces [424]. -symbol [1811, 2071]. -theorem [985, 1262]. -theory [1270]. -topology [678]. -type [122, 168]. -vector [944]. -wave [13, 466, 1365].

15015b [581]. 18 [1033]. 19 [204, 158, 300].

20 [203, 370, 436, 243, 466, 624, 51, 581, 244, 17, 105, 348, 52, 53]. 20009b [1194]. 21 [613, 242, 350, 614, 947, 798, 934, 467, 437, 894, 536, 537, 1004, 349, 649]. 22 [1160, 1720, 1566, 1194, 893, 1034, 1129, 1005, 1084, 1130, 1162, 895, 1006]. 23 [1281, 1886, 1161, 1428, 1429, 1608, 1085]. 24 [1606, 1722, 2214, 2102, 2222, 2218, 1607, 1430, 2147, 2224]. 25 [2216, 2180, 2213, 2219, 2223, 2221, 2179, 2215, 2217, 2148, 2181, 2220].

55 [1033]. 56 [300].

80b [105]. 80f [624]. 80g [244]. 80i [436]. 80k [243, 349]. 80m [242]. 81066b [537, 1004]. 81a [894]. 81d [466]. 81f [649]. 81g [613]. 81j [467, 581]. 81k [934]. 81m [798, 537, 1004]. 82045a [105]. 82b [947]. 82d [1720]. 82f [1162]. 82j [1129]. 82m [1161]. 83a [536]. 83b [1194]. 83j [1566]. 83m [1886, 1608]. 84b [1722]. 84g [2218]. 84j [2214]. 84m [2224]. 85d [2215]. 85e [2222, 2223]. 85i [2219]. 85m [2220]. 86e [2221].

Abel [484]. abelian [1419, 881, 1856, 856, 857, 1571, 1715, 504]. abelizing [2045]. abruptly [48]. absence [1703, 1815, 361]. absolutely [1815]. Absorption [258, 409]. ac [1468]. accelerated [909, 910]. Accelerating [1789, 964, 929]. acceleration [1370, 547]. accidental [1409]. according [353]. accurate [420]. acoustic [1963, 1500, 1968]. acoustical [2162]. acoustics [1, 1033, 886]. action [1406, 1218, 1429, 1120, 531, 1784, 1892]. action-principle [1218, 1429]. actions [1379, 413]. activated [48]. active [1271]. adapted [1282, 1283]. Addendum [895, 934]. Addition [1200, 249, 418]. additional [864]. additive [1172]. adequate [90]. adiabatic [8, 370, 2053]. Adjoint [1294, 1293, 862, 1785, 269, 798, 755, 1633, 139, 1613]. adjointness [520, 194, 2154]. Adjoints [693, 1084]. admitting [1856, 1070, 2024]. admixtures [920]. advantages [265]. Aesthetic [187, 527]. Affine [501, 1766, 1058]. affinely [1816]. affinely-connected [1816]. against [1481]. Aharonov [551, 702, 452]. algebra [930, 804, 246, 799, 1354, 2045, 1228, 1570, 763, 702, 2108, 1434, 1884, 1960, 754, 1039, 308, 1792, 1726, 468, 210, 470, 762, 1349, 87, 680, 488, 1683, 55]. Algebraic [773, 326, 950, 1676, 1090, 185, 235, 710, 989]. Algebraically [670, 751, 893, 544, 43, 1099, 1978]. Algebras [758, 759, 1760, 248, 1629, 574, 575, 1487, 1688, 590, 2077, 1206, 1140, 1882, 1883, 2033, 1532, 1766, 1988, 2109, 790, 650, 2183, 1616, 1111, 658, 2066, 1765, 106, 107, 897, 1675, 354, 977, 1571, 1488, 469, 896, 1085, 1809, 1670, 2036, 2154, 739, 1814, 2034]. algorithm [2002, 398, 1245, 1289]. aligned [1974]. allowed [1361]. almost [633, 1815, 530, 1737]. almost-Mathieu [1815]. along [1875, 1876, 2184]. alpha [2148, 1871]. alpha-function [2148, 1871]. alternative [340, 664, 24, 1176]. ambiguities [898]. amplitude [445, 496, 125, 1964, 284, 68, 331, 71]. Amplitudes [607, 1692, 664, 46, 484, 905, 714, 324, 262, 195, 1824, 1270, 1990]. analog [816, 879]. analogue [417]. Analysis [579, 272, 1315, 1394, 804, 670, 892, 1570, 863, 2090, 1681, 1127, 17, 7, 105, 1423, 1889, 1480, 1687, 2154, 1904, 1827]. Analytic [97, 1880, 983, 1166, 997, 810, 1108, 1840, 372, 2044, 1927, 1407, 1584, 2165, 1637, 47, 560, 1585]. Analytical [507, 540, 847, 794, 727, 1899, 4, 204, 578, 1774]. Analyticity [341, 2093, 1366, 2164, 2064, 745, 1079, 1990]. angle [1352, 1353]. angles [1327, 820, 2152]. Angular [1366, 1267, 287, 326, 552, 911, 249, 581, 271, 1573, 1621, 2224, 1485]. Anharmonic [2047, 2123, 1843, 1449, 1322, 191, 978, 1506, 741, 1020, 510, 2127, 712, 1095, 889, 1408, 90, 1451]. anharmonicity [510]. anholonomic [275]. Anisotropic [1672, 2168, 1952, 453, 129, 899, 2060, 567, 463, 612]. annihilation [2126]. annihilators [2108]. Anomalies [976, 1801, 410]. anomalous [1991]. Anomaly [1273, 2149, 1759, 800]. Anomaly-free [1273]. ansatz [1381]. antenna [392, 1032, 435]. antennas [1155, 1359]. anti [1540]. anti-Hermitian [1540]. anticommuting [758, 759, 428, 429, 430, 1760]. antiexact [785]. antilinear [653, 1390]. antiunitary [808, 1724]. any [804, 785, 2, 158]. aperiodic [2140]. aperture [435]. apparent [1596]. applicability [2064]. applicable [1031]. Application [1094, 615, 978, 2106, 1561, 213, 63, 886, 1580, 752, 1886, 1739, 1740, 1396, 1060, 209, 581, 1345, 1009, 194, 179, 218, 2010, 1236, 710, 1089, 1608, 2008, 1017, 1647, 1997, 998]. Applications [700, 1504, 2181, 1903, 286, 784, 151, 54, 421, 1618, 1612, 1359, 1723, 175, 106, 107, 1874, 239, 1078, 649, 1493, 803, 2125, 1604]. applied [1242, 342, 343, 1176, 1867]. approach [930, 1092, 1717, 540, 863, 1736, 1778, 1582, 1037, 600, 127, 1296, 1299, 211, 1201, 609, 921, 290, 86, 1343, 884, 1842, 1575, 1958, 550, 2094, 1663, 462, 1249, 2100, 1925, 875, 1795, 1170, 1042, 1065, 1256, 318, 524, 32, 30, 488, 989, 630, 929, 1290, 1611]. approaches [192, 1208, 1182]. approaching [2169]. approximants [1291, 509, 1224]. Approximate [1972, 75, 1209, 850, 319, 1962]. approximated [578]. approximating [887, 1158]. approximation [1104, 350, 88, 445, 1312, 486, 1513, 6, 624, 404, 467, 529, 794, 83, 437, 35, 908, 1510, 1965, 63, 450, 1994, 2081, 1623]. approximations [1887, 988, 1617, 265, 1128, 20, 213, 1236, 557, 1263, 625, 2080, 1771]. arbitrarily [846, 1050]. arbitrary [246, 1693, 1254, 1119, 1156, 1875, 1041, 812, 1681, 1732, 1864, 450, 1074, 1876, 2184, 1982, 547, 33, 1685, 1342, 1581]. argument [891]. arguments [2178]. arising [931, 579, 1524, 2206, 1567, 646, 822, 2193]. Arnol′d [1001]. arrangement [704, 1367]. array [290, 713]. arrays [1981, 982]. arrow [1808]. arterial [1426]. aspect [2221, 2225]. Aspects [1309, 2136]. assembly [291, 614]. associate [1579]. associated [816, 1340, 1895, 785, 1797, 1058, 1616, 1212, 1708, 765, 1832]. asymmetric [698, 1333, 1857, 744]. asymmetry [1341]. Asymptotic [637, 1043, 1449, 24, 20, 381, 328, 862, 1785, 1310, 1340, 1205, 1322, 541, 277, 1291, 938, 1046, 1583, 63, 1508, 1445, 2142, 1078, 907, 90]. Asymptotically [565, 40, 41, 2167, 545, 632]. Asymptotics [643]. atlas [937]. atmosphere [366, 572, 694]. atmospheric [395]. atom [1015, 1062, 4, 204, 921, 2079, 1927, 1060, 2098, 644, 221]. atomic [643, 4, 204, 1561, 577]. atoms [623]. attached [1469]. attractive [2070]. attractors [1758]. autocatalysis [391]. automorphism [1709, 1227]. automorphisms [1767, 1567]. auxiliary [1498]. averages [1051]. axial [103]. axial-symmetric [103]. Axially [2207, 1731, 1372, 456]. axiomatic [1183, 1778, 532, 481, 1182]. axis [1875, 820, 1359, 1876, 2184, 2203]. axisymmetric [365, 840].

B [1886]. back [55]. background [1297, 530]. Bäcklund [1943, 825, 840, 1800, 1753, 1595, 970, 621, 1113, 1133, 59, 691, 815, 587, 252, 2013, 1838, 1730, 1494]. backscatter [1972]. Backscattering [1901]. backward [331]. Baker [197, 1650]. balance [1791]. ballooning [1757]. bands [2115, 1361]. Bargmann [1010, 1690, 641, 1560]. barotropic [339]. barriers [1593]. Bars [377]. basal [1760]. based [485, 792, 1755, 957, 1600]. bases [1811]. basic [206]. basis [1354, 326, 1880, 1907, 961, 1696, 802, 803, 1610, 2125, 1108, 414, 854, 859, 1346, 1109, 1319, 1762, 1810, 1878, 1805, 1621, 2224, 2210, 676, 500, 721, 1086, 1433, 2124, 353]. Baxter [1376]. BBGKY [233]. BCS [432]. be [349, 21]. beach [1157]. beams [677]. behavior [1138, 1043, 1449, 1631, 631, 1676, 1315, 149, 277, 1289, 1506, 1291, 381, 262, 195, 1565, 770, 1270, 2142, 547]. Belinskiı [993]. bell [1981]. Bender [1322]. Benjamin [23, 1837]. Benson [1561]. Bergmann [780]. Berryman [689]. Bessel [1886, 1287, 1886, 763, 2005, 933, 2178, 2120, 2219, 303, 20, 661, 1941, 1089, 1608, 443, 1233]. Bethe [1027, 1381, 2062, 747, 2173, 2174, 1338, 554]. Bethe-ansatz [1381]. between [122, 168, 130, 1432, 2043, 101, 1916, 1919, 643, 1579, 365, 993, 399, 970, 1416, 822, 354, 1363, 2131, 1809, 1314, 1348, 1545, 2126]. Beyond [811]. bi [952]. bi-Hamiltonian [952]. Bianchi [144, 228, 535, 2101, 1369]. bichromatic [2098]. Biconical [1032]. Bifurcation [1798, 1937, 898]. Bilateral [159]. bilinear [130, 1853]. binary [1638, 984]. Binding [1994, 577]. biorthogonal [951]. Bipolar [560]. bipolaron [346]. birth [1738]. bisolitons [1579]. bistable [1333]. black [672, 1003, 723, 1789, 1857, 1479]. Bloch [1361, 434]. blood [1426]. blooming [578]. Blow [2046]. Blow-up [2046]. bodies [984, 985]. body [1260, 1175, 178, 325, 1756, 1962, 945, 1312, 698, 290, 1985, 177, 176, 963, 141, 14, 244, 93, 1770, 71, 1338, 1639]. Bogoliubov [340]. Bogomolny [1731]. Bohm [551, 702, 452]. Boltzmann [190, 945, 985, 879]. Bondi [1854]. Bonnet [228]. boomerons [388]. Borchers [1792]. Borel [817, 2192, 60]. Born [1500]. Bose [467, 1334, 1831, 1522, 281, 133, 404, 1434, 1179, 1039, 878, 360, 1540, 883, 1061, 263, 447, 44, 877]. Boson [949, 790, 1149, 75, 1763, 639, 1378, 1259, 1761, 1303, 1637, 2126, 1073, 1196]. Boson-fermion [949, 790]. Bosonic [502]. bosonization [968]. Bound [666, 2021, 828, 601, 2203, 496, 1327, 1908, 1456, 4, 204, 1035, 1452, 1213, 904, 1361, 266, 990, 668, 533, 1847, 223]. bound-bound [4, 204]. bound-state [1456, 904]. boundaries [966, 322]. Boundary [605, 1375, 2092, 1648, 2222, 102, 395, 1531, 1114, 1284, 2100, 629, 1605, 1383, 1002, 885, 1982, 1167]. boundary-free [1002]. bounded [1115, 703, 1062, 692, 1893]. Bounds [2156, 1083, 814, 869, 1368, 1355, 178, 1955, 125, 1161, 903, 362, 1141, 2068, 1636, 176, 327, 1729, 138, 577]. bracket [2109]. brackets [569]. bradyon [1478]. Bradyons [1869]. Branching [1087, 2002, 855, 2009]. branchings [1231]. Brans [604, 1486, 339, 1177, 995, 456, 2101]. Bravais [1375]. breakdown [432]. breaker [714]. breakers [284]. breaking [1703, 198, 64, 1028, 1560]. breathers [155]. Bremmer [1505]. Brillouin [1128]. broadening [321]. Broken [1695, 1419, 1559]. Broken-symmetry [1695]. Brown [1745]. Brownian [917, 32, 1149]. BRS [1026, 431]. bubble [610]. Bubnov [869]. bug [542]. Building [992]. built [1814]. bundle [665, 858, 127, 1124, 2141, 380, 524, 480]. bundles [215, 947, 1299, 413, 1265, 384, 1336, 14, 244, 1120, 1755, 1399, 523, 1630]. Burniston [2044]. by-product [1866].

C. [1588]. Calapso [659]. calculating [2056]. calculation [1041, 703, 933, 1255, 1824, 885, 2148, 1871, 223]. Calculations [46, 4, 204, 578, 1123, 463, 1771]. calculus [1242, 46, 731, 175, 1669, 649, 1892]. Campbell [197, 1650]. Can [1444, 269, 798]. cannot [349, 21]. Canonical [1464, 333, 682, 764, 117, 670, 1911, 1677, 1094, 2114, 220, 320, 934, 1410, 2218, 2079, 308, 2160, 822, 1540, 2073, 1227, 1896]. Carleman [1235]. carrying [911]. Cartan [934, 1541, 320, 230, 228, 955, 1198]. Cartesian [1389, 1051, 1049, 1050]. Cascaded [1247]. case [10, 203, 505, 670, 2105, 1926, 1030, 2117, 421, 719, 1699, 662, 601, 1588]. cases [1635, 1442, 1968]. Casimir [976, 725, 726, 1629, 1652, 1532, 1536, 2176, 897, 1728, 1670, 1671, 885, 1982]. catastrophe [490]. Cauchy [994, 279, 1885, 1971, 275, 1686]. Causal [1868, 1972, 2206, 1709]. causality [633]. Cayley [2034]. cell [2200]. censorship [671]. center [2078, 359, 1765, 475, 1588]. center-of-mass [2078]. centered [210]. central [861, 282, 223]. centrally [703]. centripetal [690]. Certain [49, 931, 761, 251, 2196, 1285, 810, 1237, 1825, 469, 951, 595, 1227, 1135]. Chain [691, 2128, 526, 1932, 2105, 1733, 2088]. chains [889]. Chandler [870]. Chandrasekhar [82, 255, 1209, 509, 1485, 1158]. change [242, 22, 656, 847, 1444]. changes [812]. channel [704, 1456, 1367, 1513, 1783, 553, 420, 502, 503]. Chapman [614, 291]. Character [1533, 1987, 1626, 1627, 617, 1007, 1390, 352]. Characteristic [15, 51, 1296, 1532, 1536, 809, 581, 72]. characteristics [1648, 2222, 591, 1129]. Characterization [1540, 1135]. characters [855]. Charge [150, 1764, 96, 783, 1041, 911, 754, 2107, 941, 1029, 379]. Charged [1516, 1385, 777, 1639, 566, 1373, 1296, 1640, 1863, 465, 914, 846, 193, 828, 1806, 1912]. Charged-particle [1639]. charges [1424, 958, 240, 2204]. Charging [332]. charmonium [1017]. chemical [1758, 2070]. Chern [1106]. Chew [2209]. chiral [2067, 1381, 1946, 2029]. chiral-invariant [1381]. Chitre [184, 225]. choice [1911, 874]. choosing [172]. chromodynamics [850, 1185, 1993]. chronoprojective [1541]. Chung [1896]. circle [654]. circuit [284, 714, 1359]. circular [392, 2051, 885]. class [1972, 2135, 190, 724, 2045, 1629, 212, 2138, 1328, 1364, 1103, 2196, 1469, 1910, 1895, 1877, 1614, 1913, 2077, 54, 1504, 1072, 2003, 1335, 2120, 2219, 359, 2205, 119, 1333, 435, 1680, 510, 2127, 922, 1046, 1583, 2139, 416, 1573, 659, 1219, 2075, 1233, 1600, 2193, 2203, 1746, 317, 1135]. classes [864, 159, 27, 1110, 1864, 1501, 1697, 72]. Classical [1635, 236, 984, 985, 767, 776, 288, 806, 1961, 1833, 461, 1720, 1916, 219, 576, 1735, 675, 1137, 1401, 824, 960, 674, 1696, 570, 2190, 1487, 2078, 1171, 539, 1861, 529, 639, 1848, 549, 807, 2020, 1921, 917, 1759, 986, 89, 1923, 1249, 695, 1745, 2017, 2014, 592, 1995, 468, 2088, 531, 118, 977, 1488, 346, 1489, 1363, 828, 979, 2073, 1658, 1502, 2126, 548, 1498, 1560]. Classification [18, 243, 698, 118, 823, 2151, 353, 91, 293, 294, 302, 398, 1082, 881, 185, 839, 710, 469, 896, 1085, 739, 292, 301, 397, 795, 796, 1483, 1484, 1187, 1346]. classified [1907]. Clebsch [372, 1354, 683, 1036, 439, 162, 163, 164, 165, 166, 167, 1567, 1569, 2110]. Clifford [1085, 930, 2033, 1988, 650, 896, 1809, 1630, 2034]. closed [2105, 1758, 1199, 95, 1454, 1472, 1473]. closer [669]. closest [2100]. Closing [2045]. cloud [1634]. Cluster [1644, 2106, 918, 1121]. clustering [1260]. cnoidal [2187, 2188]. coefficient [1593, 230, 2115]. coefficients [372, 1354, 683, 1036, 1673, 439, 162, 163, 164, 165, 166, 167, 1567, 805, 1396, 1569, 2110, 1236, 463, 954, 1298, 2008, 1415, 109, 245]. Cohen [1405]. coherence [1386, 408, 318, 612, 241]. Coherent [220, 2082, 781, 1693, 517, 552, 2035, 1823, 1334, 1195, 1010, 89, 156, 157, 681, 1934, 447]. Cohomology [1765, 2066]. coinciding [1298]. coinciding-pole [1298]. cold [201, 464]. collapse [465]. collective [1274, 1756, 534, 1030, 1196, 1934, 682]. Collectivity [2105, 1926, 1925]. collineations [967]. collision [997, 1638]. collisionless [848, 1034, 923, 924, 1603, 1604]. collisions [984, 2069]. collocation [1311, 1722]. Color [2154, 1532, 754, 1444]. combinatoric [963]. Combinatorics [93, 740, 597, 598]. Combined [1992, 711]. Comment [1123, 2221, 2225]. Comments [1317, 624, 649, 2004, 1886, 1164, 1168, 547]. common [1639]. commutant [1234]. Commutants [2157]. Commutation [160, 1227]. commutative [799]. commutator [169]. commutators [1655]. commuting [91, 1093, 1841, 1881]. Compact [1801, 1626, 545, 809, 1940, 591, 1129, 1288, 504, 1392, 1982]. Compactification [1306]. compactness [1122]. Comparison [1605, 82, 1653]. Compatibility [613, 1458, 1502, 62]. compatible [1788]. Complementary [1395, 971]. Complete [1274, 2212, 1080, 2016, 1606, 1482, 685, 269, 798, 91, 1719, 66, 1769, 1859]. Completely [1577, 1169, 874, 928]. Completeness [871, 586, 1566, 725, 214, 69, 25, 1019]. Complex [1379, 58, 1450, 1468, 1008, 763, 1919, 1327, 57, 1487, 46, 1273, 1679, 1812, 249, 820, 444, 1929, 744, 977, 1571, 1165, 1971, 514, 836, 1651, 1352, 1353, 2208]. Complex-potential [1450]. complex-scalar [514]. complexification [1245]. complexified [58]. component [1186, 1187, 917]. components [1433]. Composite [1920, 325, 358, 1114, 636, 1734, 1997]. compressible [1835]. Computation [1676, 1553, 428, 429, 71]. computer [1031]. computing [369]. concept [404, 467, 2197, 845]. Concerning [1051]. concise [420]. concomitant [150]. concrete [1088]. condensation [1073]. condensed [281]. condition [291, 614, 1458, 1861, 619, 1724, 1791, 2100, 1605, 907, 1526, 877]. Conditional [1146, 1205, 1322]. conditioning [2122]. Conditions [1328, 2092, 605, 2155, 1186, 356, 76, 657, 1986, 555, 395, 757, 413, 1465, 710, 629, 1491]. conducting [1314, 1014]. conduction [48]. conductivity [1606, 1482, 1719, 1083]. conductors [1167]. cone [1553, 1642, 196, 1641, 1461]. configuration [2098, 2095]. configurational [2174]. configurations [401]. confined [2078, 201, 464]. confinement [850, 1185, 1993]. confining [1000, 1366, 1428, 1224]. Conformal [1691, 189, 1465, 1497, 1186, 1187, 146, 1119, 2206, 4, 204, 921, 2168, 697, 1060, 150, 1414, 1329, 1535]. conformal-invariance [1186, 1187]. Conformally [604, 2027, 1594, 566, 2170, 221, 1977]. Confrontation [534]. Congruence [306, 1110]. Congruences [1971, 1856, 1786, 1519]. conjecture [1772, 602, 603, 1782]. conjugate [12, 436]. conjugation [2107, 1651]. connected [1449, 702, 1816, 1940, 1284, 562, 1813]. connecting [825]. Connection [1579, 970, 122, 168, 2043, 1503, 1559, 399, 1956, 2131]. connections [788, 333, 2141, 588]. consequence [1092]. Consequences [148, 730, 2001]. Conservation [1157, 23, 695, 1356, 1172, 1950, 627, 150, 2053, 1397, 1586]. conservative [722]. Conserved [10, 203, 505, 952, 1625, 415, 416, 919]. considerations [312]. Consistency [994]. consistent [398, 1651, 1348]. constant [445, 1552, 1957, 341, 1220, 709, 1974, 1643, 842, 2213, 1973, 1928, 98, 2008, 1597]. constant-amplitude [445]. constants [2026, 1582, 1882, 986, 1542, 647, 2221, 1841, 1543, 2198]. constituent [1848]. constrained [1833, 2048, 1896, 435, 1733]. constraint [551, 668]. Constraints [2206, 724, 2045, 1012, 198, 1842, 235, 1079, 331, 991, 1243]. constructed [1529]. Constructing [1174, 1173, 2023, 792, 322]. Construction [1906, 125, 1052, 267, 234, 26, 1938, 623, 1793, 870, 1535, 724, 1735, 424, 1188, 1357, 380, 990, 1795, 2173, 1990, 560, 1318, 2034]. constructions [1919]. constructive [1299]. Contacts [521]. containing [1496, 161, 661, 2017]. continua [2163]. continual [1620]. continuation [1166, 2156, 997, 1584, 47]. Continued [1769, 1096, 771, 1005, 1451, 1584]. Continuity [1511, 1503, 1122]. continuous [2099, 1563, 775, 1815, 1794, 1733, 1704, 117, 1135]. continuously [253]. Continuum [175, 649, 1242, 199, 917, 907]. contraction [1229]. contractions [1118, 1776]. contribution [610]. contributions [749]. control [1261, 1047, 596]. controllability [1662]. convention [1651]. conventions [1348]. Convergence [1821, 1330, 1663, 689, 1065, 1256, 831, 1893, 748, 1981, 1224]. convergent [2093]. Converging [577]. conversion [848, 1034, 1931]. Convex [722]. convolution [1941]. cooperative [2137]. Coordinate [186, 812, 1443, 196, 937, 965, 381, 645, 1646, 1412]. coordinate-free [1443, 937, 1412]. coordinates [483, 874, 545, 1530, 444, 792, 1013, 843, 1006]. copies [881, 2175]. core [1775]. corepresentations [162, 163, 164, 165, 166, 167, 615, 1347, 1724, 1723, 1998, 1007]. cores [177]. correlation [1330, 1775, 2088, 721, 429]. correlations [637, 2212, 2137, 1093, 1261, 748, 2019]. correspondence [700, 1360, 792, 1363]. corresponding [130]. coset [389, 1650, 2195]. cosine [2068]. cosmic [671]. Cosmological [1976, 1672, 1649, 2214, 1220, 709, 1974, 1643, 842, 2213, 1973, 1975, 535, 425, 1597, 2169, 2171]. cosmologies [424, 1317, 1193, 298, 299]. cosmology [1218, 1429, 1068, 2017, 2101]. Cotangent [127]. Coulomb [483, 1839, 1094, 1887, 961, 1681, 2068, 772, 484, 905, 329, 1248, 1624, 1870, 343, 1510, 1965, 682, 1190, 193, 828, 936, 71, 1153, 920, 1382, 1743, 2074, 1803, 818, 1454, 1639, 1472, 1473, 1455, 2133, 2132, 2220]. Coulomb-distorted [772]. Coulomb-like [193]. Coulomb-modified [1153, 1382, 1743, 1803]. Coulomb-nuclear [920]. Coulomb-plus-separable [1803]. Coulomb-type [484, 905]. Coulombic [810, 1141, 119]. counterexamples [671]. counterintuitive [1503]. counting [1493]. Coupled [1979, 942, 2161, 829, 830, 558, 1016, 1184, 1335, 150, 1922, 235, 2211, 2093, 1885, 489, 1162, 1538, 2004]. couplers [2180, 1967]. coupling [1475, 1476, 12, 436, 341, 1330, 1456, 708, 189, 1506, 581, 889, 239, 55]. Covariance [1325, 357, 1387, 715, 1601, 1623, 1630]. Covariant [512, 384, 1589, 999, 1443, 867, 1145, 722, 1183]. covariants [1131, 952, 2182]. coverage [1981]. covering [85, 364]. coverings [1229]. creation [457, 1181, 2126]. criteria [2099]. criterion [725]. Critical [2031, 1289, 1565, 1304, 579, 97, 2068, 887]. criticality [291, 614, 694]. critiques [1778]. Cross [1692, 2037, 533, 991, 1243]. cross-field [2037]. crossing [1821, 1744, 140, 622]. crossing-symmetric [1821, 622]. Crystal [2032, 1894, 49, 48, 846, 2211]. crystallographic [1347]. crystals [1695, 946]. CS [2069]. cube [1281, 1038]. cubic [978, 510, 383, 1875, 2184]. cumulants [1415, 1874]. Current [923, 924, 1603, 1604, 113, 702, 48]. currents [1602, 1625, 449, 919]. curvature [1890, 564, 188, 189, 619, 1616, 967, 98, 2054, 488, 226]. curvature-zero [619]. Curvatures [144, 744]. curve [2093]. curved [605, 272, 1752, 1946, 100, 1992, 1796, 669]. curves [1370, 1531, 1113, 953, 14, 244]. cuspoid [2114]. cut [335]. cutoffs [2080, 2081]. cuts [497, 1642, 1641]. Cycle [1523]. cyclic [1688, 2108]. cylinder [79, 1709]. cylinders [1459, 102, 42]. cylindrical [1869, 1836, 2051, 1155, 645, 1909, 1747]. Cylindrically [2084, 1517, 201].

damped [323, 1450, 1959, 2140]. damping [1957, 123]. Darboux [1705]. data [1908, 1166, 797, 1963, 15, 51, 1857, 1193, 2162, 1902]. Debye [1953]. Decay [815, 637, 1261, 962, 200, 721, 1279]. Decaying [173, 222, 748]. declining [414]. Decomposition [1228, 1696, 754, 663, 491, 1371, 1136]. decompositions [1825]. Decoupling [2008, 528, 782, 1918, 519]. decreasing [1964]. defined [364, 856, 857, 693, 1084]. definite [1664, 1231]. definition [730, 480]. Definitions [1362]. Deformation [1286, 1543, 1280, 2143, 825, 997, 1067, 1612, 867, 834, 196, 1616]. deformations [1488]. degeneracies [508, 636, 1939, 2113]. degeneracy [2, 158, 1409, 1734, 1377]. Degenerate [451, 544, 222, 1109, 1810, 1099, 1393, 110]. degree [1615, 1954, 1433]. degrees [873, 1052]. delay [1512, 1850]. delta [747, 1302, 1654]. delta-function [747]. Demianski [777, 1177]. Demiański-type [1177]. Democracy [882]. Dense [2, 158, 1688]. Densities [999, 10, 203, 505, 1527, 201, 464, 415, 416]. density [797, 1963, 564, 1561, 722]. dependence [1071, 224, 935, 729, 1790, 1142]. dependent [1677, 1449, 1115, 1735, 1957, 1147, 664, 2163, 421, 240, 2011, 1842, 267, 986, 1542, 732, 1622, 324, 66, 510, 1773, 1249, 922, 846, 355, 822, 1053, 2014, 1968, 629, 1966, 736, 1841, 448, 849, 1130, 2166, 739, 901, 1098, 124]. Dependent- [124]. depending [2015, 1008, 1491]. Derivation [1739, 1740, 1645, 554, 2129, 664, 307, 1897, 1412, 1460, 1528]. derivative [2064, 1555, 680, 1494]. derivatives [1491]. derived [129, 2025]. described [130]. describing [1931]. Description [1954, 2085, 1450, 1590, 2055, 1884, 746, 1239, 263, 1528, 2140, 1572]. descriptions [2043, 1797]. destruction [997]. detailed [1791]. determinants [134, 765, 1198]. Determination [1875, 1876, 2184, 1086, 2044, 1277, 398, 1082, 1462, 319, 1079, 1484, 1997]. deterministic [875]. development [878]. developments [1045]. deviation [1246, 1412]. Diagonalizability [749]. diagonalization [1991]. diagonalizations [16, 581]. diagram [610]. diagrams [1609, 2201, 2216, 1532, 1575, 2194, 1415]. dichotomous [1751]. Dicke [456, 604, 1486, 339, 1177, 995, 644, 2101]. Dickson [2034]. dielectric [79]. diffeomorphisms [1401]. difference [371, 1698, 597, 598, 2008]. different [825, 399]. Differentiable [1936, 199, 1299]. Differential [2114, 122, 168, 582, 871, 1986, 952, 1237, 309, 1576, 618, 1207, 1357, 1632, 1816, 1862, 867, 655, 1400, 646, 1931, 1236, 1805, 124, 2151, 1538, 1358, 1494, 2004, 814]. diffraction [1492, 794, 578, 2083, 601, 668]. Diffusion [1861, 129, 280, 586, 1566, 212, 1043, 170, 391, 767, 126, 1001, 1333, 2037, 942, 1525, 1167]. diffusion-reaction [391]. diffusions [1830, 1048]. dilated [1679]. dilation [1547, 2102, 1850]. Dimension [617, 666, 1707, 904, 706, 92, 1592, 2130, 2072, 2217, 1885, 434]. dimensional [1260, 1727, 1894, 1667, 1281, 1038, 2155, 1354, 1629, 1071, 1839, 1228, 728, 950, 2105, 1926, 1907, 1156, 2129, 1908, 1963, 1986, 2117, 1901, 1035, 491, 285, 2108, 2011, 1825, 177, 2057, 1051, 1267, 1039, 308, 917, 234, 696, 1339, 444, 621, 1940, 787, 1302, 66, 2127, 1780, 2176, 1249, 1872, 179, 1258, 2136, 713, 1239, 1980, 1193, 636, 1361, 1981, 852, 2014, 1749, 2215, 344, 34, 894, 1968, 527, 815, 1112, 962, 1929, 1621, 2224, 1042, 1074, 1447, 2013, 1790, 1834, 2131, 1397, 2113, 1049, 1050, 823, 1528, 1535, 1820, 1278, 45, 182, 135, 1525, 2203]. dimensional [2177, 1835, 1829, 317, 2061, 1498, 1460, 72, 259, 1030, 1114]. dimensionality [1150]. Dimensions [977, 1933, 1125, 1115, 1602, 1119, 13, 466, 1895, 1196, 1934, 984, 985, 968, 359, 1945, 650, 254, 1133, 70, 268, 537, 771, 1004, 1005, 990, 1408, 1497, 896, 1085, 1549, 1623, 1463]. dimer [1375]. dipole [486, 1754]. Dirac [1518, 507, 1656, 46, 2063, 1960, 1077, 1624, 1870, 1958, 1856, 1625, 342, 343, 194, 388, 1824, 1726, 969, 753, 828, 1550, 1991, 1796, 1647, 1425, 589, 919, 2030, 964, 1630, 2074, 139, 181]. Dirac-spinor [46]. Direct [1216, 987, 2097, 1290, 690, 772, 1249, 1775, 2075, 862]. direction [867, 756]. discontinuities [868]. discontinuity [756]. Discontinuous [1204, 959, 1233]. Discrete [417, 516, 1932, 816, 1225, 1296, 394, 695, 351, 361, 974]. Discriminant [2115]. discussion [669]. disk [103, 392]. disordered [891]. Dispersion [733, 63, 625, 2093, 1769, 1404]. dispersion-free [1404]. dispersive [1661]. displaced [475]. dissipative [1137, 54, 971, 1528, 2140, 1543]. dissociating [200]. dissociation [2100]. distance [2100]. distant [2019]. distinct [2201, 2216]. Distorted [1003, 772]. distortion [1636]. distribution [340, 1041, 1373, 1732, 2018, 1729, 819, 328, 258, 1644, 1253, 1502, 226, 2007]. distribution-valued [226]. Distributional [714]. Distributionlike [658]. distributions [1496, 1656, 132, 2041, 1093, 24, 1102, 2060, 941, 1912, 139]. divergent [2038, 2202]. divisible [493]. do [1550]. Documentation [265]. domain [2131, 480]. domains [1334, 322]. dominant [77]. Donder [1867]. don’t [1741]. Doppler [1741]. Double [1761, 2080, 973, 2187, 2188, 1280, 2143, 147, 615, 2106, 1176, 889, 2081, 104, 1705]. Double-Gel’fand [1761]. double-quadratic [104]. double-well [147, 1176, 1705]. doubling [1289]. doubly [26]. drift [403, 402]. driven [283, 1751]. Dual [838, 1066, 1919, 1328, 1379, 1800, 1076, 1664, 1466, 1467, 1384, 951, 2007, 1069]. Dual-mass [838]. duality [1915, 2022, 1799, 1651, 1348, 1526]. duct [202, 756]. due [11, 348]. dumbbells [1557, 713, 982, 1948]. Durgut [377]. Dust [42, 1373]. Dyadic [1053, 1443]. Dynamic [889, 104]. Dynamical [1582, 1030, 1196, 1934, 64, 1862, 732, 2014, 737, 1159, 379, 135, 114, 1104, 1735, 2118, 1861, 1195, 1842, 986, 1542, 180, 1126, 1423, 499, 1751, 1507, 559, 1023]. Dynamically [640]. Dynamics [841, 427, 799, 1717, 97, 1401, 1563, 1271, 1544, 1150, 1928, 89, 712, 569, 1048, 1897, 562, 1633, 2073, 774, 1659, 1660, 1543, 663]. Dynkin [1609]. dyon [512]. Dyson [1763, 90, 1023].

earth [797]. Eckart [246]. Edward [624]. Edwards [1580]. Effect [2100, 551, 1741, 2032, 702, 174, 1315, 1143, 1745]. effective [1851, 1781, 6, 624, 1150, 1083, 2128]. effects [1459, 2051, 1375, 1190]. eigenenergies [239]. eigenfunctions [269, 798, 561, 1679, 423, 366, 1708, 271, 377, 1408, 322]. eigenproblems [688]. eigenstate [398, 1082, 1484]. eigenstates [294, 2052, 264, 360]. Eigenvalue [538, 286, 1355, 398, 1082, 1055, 2052, 1714, 423, 1584, 1585, 1484, 1808]. Eigenvalues [1843, 1652, 293, 741, 1020, 1708, 2113, 1671, 1485, 869, 976, 1355, 178, 1225, 703, 1035, 1700, 2223, 176, 327, 2057, 1506, 271, 328, 2195, 1408, 369, 2203, 795, 1742]. eigenvectors [1456]. eight [2176]. eight-dimensional [2176]. eikonal [1031]. Einstein [51, 1129, 1162, 1100, 1066, 1328, 229, 1220, 336, 709, 1914, 230, 711, 1101, 2023, 1179, 38, 39, 333, 334, 15, 277, 1102, 278, 332, 1217, 1518, 1709, 2084, 1664, 591, 1288, 1853, 1665, 568, 1372, 1976, 489, 1858, 456, 1685, 2171, 72]. elastic [626, 988, 394, 2163, 2051, 1114, 1997]. elasticity [2199]. Electric [783, 1280, 2143, 1928, 2086, 941]. electrically [1296]. electrodiffusion [646]. Electrodynamics [2163, 868, 520, 1562, 1015, 512, 357, 486, 1890, 549, 234, 282, 1095, 2039, 67, 1678]. electrohydrostatic [1280]. electrohydrostatics [2143]. electrolyte [2100]. Electromagnetic [79, 143, 2051, 2022, 941, 1014, 65, 1276, 1693, 1475, 1476, 1979, 1952, 1292, 835, 1385, 1802, 1081, 1091, 793, 1928, 278, 813, 1601, 201, 464, 1872, 852, 1048, 1968, 1189, 1154, 1750, 644, 778, 1275, 78, 31, 844, 1167]. electromagnetics [1, 1033]. Electromagnetism [734, 1071, 240, 1191, 1755, 1790, 1384]. Electron [1276, 1406, 2, 158, 342, 343, 1995]. electron-gas [2, 158]. electronic [1561]. Electrons [2211, 1190, 434]. Electrostatic [890, 1316]. Electrovac [2213, 1973, 1101, 709, 1974, 911, 913, 1643, 225, 338]. element [1083]. Elementary [253, 1201, 1848, 807, 349, 21]. elements [111, 1880, 961, 1844, 1487, 4, 204, 440, 805, 1812, 1248, 758, 1109, 1627, 271, 1488, 676, 154, 1637, 759, 1670, 2000, 1342, 1760, 2074]. ellipsoids [102]. elliptic [1439, 1355, 1966]. elliptical [102]. Ellis [1317]. embedded [2057]. embedding [1235, 1265, 2136]. Embeddings [1231, 1147]. Emden [689]. emergence [96]. end [1932]. end-to-end [1932]. energies [577, 1451]. Energy [564, 514, 286, 1515, 631, 496, 1161, 903, 1366, 485, 1703, 703, 134, 884, 1332, 687, 1506, 1191, 2086, 971, 262, 355, 822, 710, 941, 138, 68, 1854, 1002, 1982, 1078, 2203, 1742, 498, 360]. energy-dependent [355, 822]. energy-entropy [1703]. energy-momentum [1191, 498]. Energy-particle-number [514]. entire [1239, 1996]. entities [1906]. Entropy [674, 1916, 1703, 1732, 482, 2042, 722]. envelope [1680]. envelopes [1564]. Enveloping [2108]. epoch [1804]. equation [2161, 850, 666, 291, 614, 2167, 190, 1043, 422, 579, 170, 1524, 956, 2187, 2188, 2189, 1186, 1187, 99, 1817, 1031, 2026, 340, 23, 1156, 251, 1324, 224, 2196, 1285, 988, 998, 690, 269, 798, 1067, 1640, 985, 767, 102, 1815, 1739, 1088, 1899, 129, 1937, 1863, 719, 1699, 2185, 2063, 545, 927, 769, 1960, 981, 1595, 327, 2057, 1077, 1624, 123, 497, 1213, 2115, 255, 1209, 987, 1622, 2127, 1645, 254, 1133, 1931, 904, 1126, 1423, 1668, 1053, 1054, 509, 98, 750, 737, 1845, 2015, 70, 34, 894, 1591, 1708, 629, 1523, 691, 815, 1112, 962, 941]. equation [328, 2144, 875, 1983, 587, 377, 124, 202, 1480, 679, 253, 1969, 2145, 476, 1819, 1865, 1021, 1365, 1796, 2054, 1837, 1976, 1838, 1528, 2140, 1078, 1017, 1647, 1820, 1600, 589, 1202, 1412, 944, 901, 964, 1158, 442, 980, 1323, 811, 128, 826, 1402, 1902, 2200, 554, 1222, 1493, 1498, 2074, 1290]. equations [10, 203, 505, 1356, 122, 168, 79, 1972, 371, 2180, 1967, 586, 1566, 1355, 1175, 311, 212, 766, 1100, 1578, 1943, 1183, 1514, 1311, 1721, 1722, 151, 816, 217, 1328, 441, 574, 575, 950, 229, 1119, 2067, 1962, 945, 1836, 251, 825, 1147, 57, 58, 2114, 13, 466, 1895, 1617, 672, 1220, 336, 709, 1974, 403, 812, 1800, 1200, 785, 1914, 191, 871, 1986, 561, 776, 126, 1091, 169, 152, 211, 1201, 2058, 952, 1732, 1072, 1237, 711, 1698, 407, 309, 1576, 1818, 1101, 2023, 660, 333, 1441, 1357, 1632, 1816, 1862, 717, 334, 1380]. equations [15, 51, 1102, 899, 1783, 399, 278, 813, 970, 2076, 1437, 1856, 986, 1542, 82, 384, 1625, 581, 332, 1217, 1326, 1518, 2191, 1616, 787, 1264, 1302, 1398, 1923, 1539, 2084, 646, 1145, 335, 1690, 1664, 689, 1922, 1236, 355, 235, 1864, 415, 416, 2027, 1799, 1665, 59, 1731, 568, 1968, 553, 1372, 990, 1203, 1841, 1798, 124, 597, 598, 2008, 94, 1885, 659, 250, 979, 1633, 411, 641, 1397, 2159, 2075, 489, 1162, 1634, 2151, 1658, 2080, 2039, 2193, 951, 260, 919, 2030, 1730, 701, 1538, 1253, 1685, 1900, 2091, 135, 739, 901, 996, 1835]. equations [1358, 1494, 2016, 2171, 2004, 1023, 1580, 1705, 814]. equilibria [17, 7, 105]. Equilibrium [74, 73, 75, 570, 1891, 401]. Equivalence [373, 1846, 457, 492, 477, 1208, 2139, 1202, 844, 2049, 906]. equivalency [485]. Equivalent [824, 662, 1891, 1950, 1403]. ergodic [1779]. Ermakov [448, 250, 823]. Ermakov-type [823]. Ernst [892, 840, 993, 1067, 1091, 1595, 813, 587]. Ernst-equation [1595]. Erratum [203, 613, 1160, 242, 1720, 350, 614, 1281, 2216, 2180, 1566, 370, 1033, 1194, 1606, 1721, 1722, 436, 2214, 243, 947, 2102, 466, 1161, 2222, 1428, 798, 893, 467, 2213, 204, 2218, 2219, 158, 1429, 2223, 51, 1607, 437, 581, 1034, 244, 2225, 17, 105, 1129, 2179, 300, 2215, 894, 348, 536, 1608, 537, 1004, 1005, 1430, 2147, 2224, 349, 2217, 1084, 1130, 1085, 1162, 2148, 2181, 52, 1006, 53, 2220]. Error [1400, 1214]. errors [1166]. essential [194]. Estabrook [1595]. estimates [757, 1400, 1214, 68]. Estimation [2007]. Euclidean [204, 1066, 1379, 4, 921, 149, 1122, 787, 1076, 2065, 2093, 40, 768, 181]. Euler [1129, 591, 1352, 1353, 2152]. Eulerian [2043]. evaluated [349, 21]. evaluating [1655]. Evaluation [1832, 931, 2002, 1947, 2212, 810, 119, 175, 661, 1046, 1583, 1866, 649, 19, 583]. Even [203, 2212, 239, 1078, 1647, 10]. even-power [239]. event [543]. evidence [1782]. Evolution [1356, 212, 816, 1680, 10, 203, 505, 122, 168, 1116, 766, 1240, 1180, 441, 191, 211, 1698, 1172, 769, 660, 1107, 396, 1398, 1539, 355, 415, 416, 2144, 1203, 1841, 1445, 2145, 2075, 1446, 1730, 1900]. evolutions [54]. Ewald [119, 811]. Ewald-type [811]. Exact [1933, 2180, 1932, 1563, 4, 204, 1300, 1975, 1333, 339, 2127, 922, 1872, 636, 982, 271, 1344, 644, 1408, 1279, 1991, 1900, 630, 901, 5, 53, 1493, 1771, 873, 2067, 2212, 1913, 1072, 711, 2011, 1102, 1518, 510, 1249, 2017, 1734, 98, 750, 1372, 875, 679, 1751, 1967]. Exactly [2062, 423, 2061, 286, 606, 211, 177, 477, 1549]. example [949, 432, 1737, 206, 1528, 2140, 487]. examples [1503, 269, 798, 1351, 504, 1618]. exceptional [248, 1188, 913, 1572]. exchange [297, 473]. excitations [1756, 347]. excitons [241]. excluded [525]. exclusion [296]. Exhaustive [1974]. exist [1550]. Existence [214, 1456, 1302, 1291, 1305, 1021, 345, 595, 586, 1566, 1328, 1579, 439, 363, 344, 718, 1847, 843, 1006]. expanding [457, 1181, 492, 912]. Expansion [475, 518, 1496, 1869, 1987, 542, 97, 289, 1297, 1310, 1322, 978, 772, 285, 622, 833, 918, 819, 632, 297, 1121, 645, 2093, 1638, 1078, 1858, 1644, 560, 1859, 90, 256]. Expansions [1698, 1341, 507, 1205, 541, 37, 954, 1508, 1445, 1769, 239, 2146]. experiments [33]. Explicit [931, 973, 1942, 1947, 1119, 1401, 1188, 1772, 1362, 1866, 19, 583, 224, 729, 1982, 108]. explicitly [359, 1841]. explode [962]. explode-decay [962]. exponent [1815]. Exponential [738, 1045, 2068, 2033, 1335, 366, 572, 2010, 694, 1142, 1411]. exponentially [414]. exponentials [1701]. exponents [1235]. expressed [1050]. Expression [1440, 2112, 496, 1974, 1443]. expressions [372, 973, 1942, 1880, 2212, 4, 204, 578]. Extended [217, 236, 237, 238, 273, 527, 1374, 1451, 1739, 1740, 1151, 1989, 2086, 1414, 1605, 774, 2091, 1623, 460]. Extension [2099, 1681, 769, 2190, 802, 803, 1245, 861, 597, 598, 469, 1433]. extensions [2121, 1853, 210, 864]. extensive [1570]. Exterior [635, 1114, 680]. External [1383, 1145, 235, 1074, 1303, 1991, 1167, 1586]. extrema [1940]. extremal [1938]. extremum [134, 198].

F. [1474]. factor [435, 1155]. Factorization [2161, 376, 620, 928, 2186, 1579, 1200, 618, 380, 171, 536]. Factorized [1376, 1895, 1739]. factors [438, 560]. Faddeev [1175, 871, 1124, 716, 2175, 431]. fall [2206, 1011]. families [1357]. family [646, 1372, 1786, 2157, 1613, 1040]. far [1891]. faster [1964]. feedback [1480]. Fermi [1496, 507, 1016, 1312, 1139, 1831, 1945, 819, 261, 876, 1540, 883, 106, 107, 1061, 478, 1644, 153]. Fermion [1260, 1830, 1602, 639, 949, 1228, 54, 1378, 1259, 790, 1377, 2126]. Fermion-boson [639, 1378, 1259, 2126]. Fermionic [681, 100, 503]. fermions [1302, 1939, 2087, 1334]. ferroelectric [182]. ferromagnetic [867]. ferromagnets [296]. Feshbach [906]. few [290, 176, 141]. few-body [290, 176, 141]. Feynman [2216, 2201, 960, 719, 1699, 1985, 284, 714, 1403, 1208, 1893, 1339, 175, 262, 195, 1784, 1117, 1270, 649, 90]. Fiber [523, 858, 384, 1336, 14, 244, 645, 524]. fibers [1399]. Field [390, 1558, 717, 94, 1594, 670, 829, 830, 752, 873, 461, 1720, 1693, 579, 1125, 272, 1183, 99, 229, 1301, 1481, 2067, 2032, 1516, 1147, 1285, 1027, 1844, 1220, 336, 835, 916, 709, 1974, 881, 1688, 785, 1802, 1914, 561, 74, 1527, 1890, 608, 492, 1689, 1470, 609, 1075, 36, 1921, 189, 1101, 2023, 1122, 1599, 477, 15, 277, 51, 368, 793, 1267, 1421, 899, 1928, 1792, 148, 278, 2170, 332, 1143, 141, 918, 1601, 1789, 1146, 1264, 1872, 1145, 882, 1664, 2037, 528, 782, 1918, 282, 1944, 273, 2139, 706, 2017, 2098, 1070, 573]. field [1799, 638, 1263, 187, 527, 11, 348, 1793, 1190, 1795, 1074, 519, 40, 41, 644, 193, 828, 2073, 514, 523, 524, 1991, 1418, 78, 759, 1977, 1528, 2140, 774, 595, 1549, 1525, 434, 128, 1402, 548, 1182, 181, 1867]. fields [605, 419, 142, 1276, 426, 374, 1241, 219, 143, 1186, 1187, 915, 2138, 1919, 1137, 2158, 1582, 1379, 18, 243, 1340, 1103, 858, 1266, 657, 236, 237, 238, 454, 1091, 968, 1334, 1831, 1848, 1920, 149, 705, 413, 1024, 2022, 1529, 1124, 2141, 627, 628, 851, 813, 809, 378, 1657, 745, 787, 1465, 1076, 946, 381, 748, 1239, 1661, 235, 344, 316, 70, 1189, 941, 1929, 535, 531, 385, 1466, 1467, 1154, 716, 2024, 1750, 338, 778, 1275, 489, 1162, 1106, 1858, 456, 1747, 1730, 31, 735, 560, 1866, 1069, 1374, 1859, 844, 2028, 663]. fields [1167, 455, 1560, 386]. Fierz [1545]. fifth [952]. fifth- [952]. filter [1653]. filters [773, 1286, 1653]. finding [178, 1249]. Finite [372, 684, 1194, 99, 1128, 1768, 1965, 2086, 1190, 56, 2177, 371, 862, 1785, 1629, 1228, 761, 1186, 1187, 457, 161, 1379, 1156, 1108, 1164, 1075, 491, 2108, 1825, 1110, 48, 2010, 2139, 1627, 1591, 1083, 531, 1795, 1999, 349, 21, 519, 849, 1130, 1435, 1002, 1809, 31]. Finite- [2177]. finite-action [531]. finite-component [1186, 1187]. finite-dimensional [1629, 1228, 491, 2108, 1825]. Finite-size [1190]. Finitely [1471]. finiteness [1225]. Finsler [627, 628]. Finslerian [867, 594]. First [1950, 2160, 1773, 1658, 724, 2045, 1246, 1986, 2058, 1237, 1204, 729, 1500, 923, 924, 1603, 1604, 1236, 1538, 2004, 259, 1715]. first- [1715]. First-order [1950, 1658, 1986, 2058, 1204, 1236, 1538, 2004, 259]. Five [437, 35, 1763]. Five-term [437, 35]. Fixed [2003, 1684, 1951, 473, 1079, 128, 826, 1402]. FKG [1598]. flat [604, 566, 632, 2221, 965, 565, 145, 221, 1977, 2198, 929]. flip [533, 991, 1243]. flow [1732, 1426, 1976, 420, 1718, 1314, 1525]. flows [395, 200, 1497, 1002]. Floyd [624]. Fluctuation [1522, 350, 88, 445, 1138]. Fluctuations [705, 390, 1386, 1387, 283, 746, 1745]. fluid [566, 1672, 1301, 1481, 1102, 1596, 2060, 1427, 778, 1275, 721, 1002, 1313, 2101, 1278, 2039, 1912, 1835]. fluids [1910, 1852, 2205, 2168, 73, 401, 1855, 971, 567, 707, 844, 1746]. fluorescence [2098]. Fock [217, 1294, 358, 208, 345, 681]. Fokker [698, 1863, 1072, 1323]. foliated [2118]. force [1891, 141, 282, 338, 885]. forced [1957, 1737, 1738]. forces [1586]. forever [1221]. form [670, 1018, 1587, 871, 1853, 721, 722, 912, 560, 907, 1454, 1472, 1473]. Formal [1462, 400, 1214, 1813, 2126]. formalism [1752, 840, 232, 513, 230, 1403, 839, 573]. formats [655]. former [265]. forms [1487, 1398, 957, 1540]. formula [197, 1439, 1492, 1322, 1200, 1655, 1288, 1199, 154, 1637, 1982, 2175]. Formulas [2195, 617, 945, 1205, 952, 2149, 119, 1573, 1650, 1342]. Formulation [667, 1659, 131, 1057, 173, 865, 326, 108, 1914, 561, 1590, 240, 708, 2079, 1792, 789, 1855, 2065, 1555, 1983, 1784, 940, 909, 910, 559, 1448]. formulations [1071, 1790, 1715]. forward [331, 533]. foundations [1632, 1816]. Four [4, 204, 921, 1281, 1038, 1071, 1821, 696, 444, 462, 27, 527, 1790, 1243, 72, 2132]. four-dimensional [1281, 1038, 1071, 696, 444, 527, 1790, 72]. four-nucleon [462]. four-point [1821]. Fourier [407, 2111]. fourth [1615, 1931, 860, 897, 612]. fourth-degree [1615]. fourth-order [1931, 897, 612]. Fowler [689]. Fractals [1889]. fraction [1769, 1096]. Fractional [1887, 1236]. fractions [1584, 1451]. frame [454]. frames [127, 1206, 380, 964]. framework [532]. Fredholm [765]. Free [186, 2167, 1839, 2138, 2206, 1910, 1852, 785, 1273, 1802, 1443, 205, 482, 937, 1601, 779, 1404, 1303, 1002, 2140, 1412, 1746, 1136, 1678]. free-particle [1839]. Freed [1580]. freedom [873, 1052, 550, 1465]. frequency [1340, 1901, 1268, 1269, 2131, 781]. frequency-domain [2131]. friction [323]. Frictionally [2140]. Friedmann [1192, 1425]. Froissart [1327]. Fröman [265]. Fuda [1681]. fully [1186, 881, 1744]. fully-reducible [1186]. function [79, 1932, 483, 1735, 1310, 340, 1296, 2021, 747, 1794, 1863, 1744, 2120, 2219, 1453, 140, 622, 1632, 530, 1248, 1927, 1386, 1387, 1624, 1870, 1302, 661, 1284, 213, 63, 1775, 2056, 625, 875, 597, 598, 475, 2209, 478, 721, 2036, 1446, 1295, 2148, 1871, 612, 1654, 90]. Functional [1258, 2179, 1905, 1313, 548, 931, 1843, 406, 863, 232, 978, 134, 749, 2144, 1042, 2145, 1491, 318, 153]. Functions [2165, 2044, 817, 526, 1886, 507, 1131, 1839, 1226, 956, 1887, 1516, 2047, 1844, 1330, 1849, 933, 414, 1873, 2106, 1504, 855, 2003, 2178, 545, 303, 1407, 2018, 26, 435, 2182, 20, 1346, 2037, 1239, 1053, 2088, 1941, 1089, 1608, 1996, 1203, 1805, 473, 349, 21, 1769, 383, 1491, 2209, 318, 936, 1418, 1050, 428, 429, 1233, 1644, 47, 1485, 1685, 322, 19, 583, 686, 1771, 818, 1040, 1438, 814, 1090]. fundamental [1492, 690, 1489, 504, 1231]. Further [547, 546].

G [2221]. Galerkin [869, 1514, 1721, 1436, 1836]. Galilean [1541, 1266, 1689, 1470, 1126, 1668, 1944]. Galilean-invariant [1126]. Galilei [685, 407]. Galileian [108]. Galois [618]. gamma [1438]. Gamow [865]. gap [493]. Gardner [506]. gas [1522, 75, 1406, 281, 1732, 2, 158, 530, 1303, 1634, 44, 877]. gases [340, 200]. Gauge [219, 432, 356, 477, 524, 2092, 752, 461, 1720, 406, 2045, 1475, 1419, 784, 915, 1379, 18, 243, 76, 1648, 2222, 881, 1986, 1559, 708, 285, 1921, 413, 1024, 1267, 1124, 851, 550, 150, 1337, 1336, 715, 1601, 2191, 1465, 1923, 1076, 227, 1120, 1755, 1416, 573, 880, 1716, 1799, 1263, 2093, 716, 1801, 1106, 2175, 2208, 498, 386, 1715, 1660]. gauges [1183, 1464]. Gauss [228, 2091]. Gaussian [931, 956, 890, 1657, 175, 748, 213, 468, 1042, 1341, 375, 781, 61, 649]. Gel’fand [798, 894, 973, 269, 2106, 247, 34, 1591, 1708, 990, 1761, 679, 1969]. General [1946, 1459, 609, 474, 581, 1724, 841, 7, 2015, 2107, 1112, 1824, 1675, 1728, 571, 1635, 670, 868, 873, 1911, 426, 2135, 2172, 1652, 1370, 1103, 1910, 453, 993, 1974, 564, 1360, 563, 992, 1699, 1207, 2168, 1437, 760, 1345, 1347, 1723, 937, 1596, 1217, 1857, 632, 1616, 17, 105, 231, 648, 2103, 839, 967, 743, 710, 1048, 276, 2072, 2217, 838, 338, 909, 910, 1909, 1671, 239, 1078, 1647, 2123, 275, 522, 1912, 61, 1519, 844, 2203, 1630, 780, 1746, 1925]. Generalization [851, 2044, 1405, 1101, 1599, 1421, 306, 845, 944, 964, 1318]. generalizations [1285, 823]. Generalized [1667, 752, 1677, 1477, 763, 1132, 1915, 2192, 831, 303, 1827, 1828, 1826, 1163, 1500, 1521, 1620, 782, 1774, 106, 936, 1659, 1660, 55, 33, 412, 1323, 256, 1019, 1984, 1287, 1376, 2138, 1310, 945, 1036, 858, 2047, 1205, 1474, 2068, 700, 1360, 1863, 981, 1300, 15, 277, 51, 150, 2183, 1413, 89, 1645, 1305, 822, 1288, 2179, 1905, 2066, 1329, 815, 1941, 1232, 990, 56, 1750, 1408, 1219, 1050, 30, 1412, 2152, 996]. generalizing [1380]. generated [2045, 2191, 1133, 680]. Generating [1226, 855, 1805, 473, 406, 1131, 1137, 365, 1567, 2003, 1248, 2182, 2209]. generation [1943, 1023]. generator [352]. generators [1533, 1880, 1626, 440, 805, 2029, 1627, 1726, 1061]. Geodesic [1246, 729, 1582, 1220, 709, 1856, 382, 779, 1412]. geodesics [95]. Geometric [1357, 926, 974, 975, 1092, 1299, 1960, 773, 1207, 1090, 1632, 1612, 1816, 794, 1956, 1374, 1859]. Geometrical [1984, 76, 2122, 593, 1218, 1429, 1413, 716, 431, 1325, 1717, 315, 1332, 411, 2001]. geometrically [2078]. geometries [672]. Geometry [1842, 146, 1541, 2105, 1926, 251, 1915, 1642, 1749, 2215, 1925, 955, 1497, 501]. geophysical [2162]. Gerlach [33]. Geroch [602, 1787, 1460]. ghost [1754, 431, 386]. Ghosts [269, 798]. Gibbs [2146]. Gibson [870]. give [269, 798]. given [2041, 1075]. GKS [1598]. Global [1240, 2118, 1758, 344, 730, 1002, 1335, 856, 857, 216, 584, 412, 929]. globally [242, 22]. glueball [2093]. gluon [1185, 1993]. gluons [1920]. Gödel [1301, 425]. Gödel-like [425]. Gödel-type [1301]. goes [1272]. good [335]. good-cut [335]. Gordan [372, 1354, 683, 1036, 439, 162, 163, 164, 165, 166, 167, 1567, 1569, 2110]. Gordon [666, 1943, 958, 251, 1088, 529, 639, 711, 174, 1783, 970, 2052, 254, 1423, 98, 344, 70, 124, 1885, 2054, 1499, 1078, 701, 944, 2016, 1290]. Graded [1337, 1669, 1884, 1039, 790, 210, 1650, 1308]. Gradient [1169, 1937, 1942, 1573, 1654]. gradings [386]. Gram [1198]. grand [2073]. graph [719, 1699]. graphical [1133]. graphs [853]. Grassmann [1548, 106, 107]. gravitating [2205, 2060, 1855, 567, 1427, 489, 1162]. Gravitation [227, 1413, 2086, 1120, 1755, 994, 279, 780]. Gravitational [278, 670, 873, 426, 1979, 1788, 1552, 1103, 1147, 365, 916, 708, 189, 2207, 1553, 2170, 745, 381, 337, 1745, 1716, 876, 228, 40, 41, 1750, 338, 523, 524, 1550, 1747, 735]. graviting [567]. gravity [915, 1481, 1300, 746, 1414, 880, 1555, 1686, 1461, 1460]. gravothermal [490]. Green [79, 526, 483, 1839, 1873, 530, 1248, 1624, 1870, 1053, 383, 322, 90, 814]. Gribov [898, 2175]. Griffiths [891]. Gross [1381]. Ground [1161, 903, 643, 1644]. Group [413, 1754, 1423, 989, 2074, 1879, 246, 1092, 136, 1281, 1038, 1432, 2153, 1304, 111, 585, 1598, 948, 183, 85, 685, 1037, 1875, 1547, 2102, 365, 1395, 302, 1030, 1196, 1934, 2035, 801, 1936, 786, 2106, 1118, 4, 921, 1195, 113, 805, 2022, 1286, 717, 1797, 1058, 1059, 820, 1856, 209, 1940, 1725, 66, 2176, 227, 100, 748, 1674, 1761, 500, 1568, 1628, 1876, 2184, 504, 499, 1028, 2082, 1134, 1646, 1806, 1801, 1086, 1394, 895, 616, 1352, 1353, 2152, 2000, 1813, 1866, 292, 301, 796, 110, 1460, 1611]. group-invariant [1940, 1866]. group-theoretic [365]. Group-theoretical [1423]. groups [1533, 976, 1987, 1477, 1476, 1131, 761, 683, 85, 2158, 161, 1673, 1767, 1696, 1935, 304, 615, 160, 539, 1626, 204, 86, 859, 1632, 1797, 1695, 2150, 1759, 809, 1282, 760, 1345, 1346, 1347, 1724, 1940, 1723, 1998, 1319, 1762, 1810, 2183, 1923, 856, 857, 2179, 1905, 472, 1627, 2107, 470, 762, 1007, 1349, 1390, 1430, 1320, 56, 1489, 411, 584, 524, 469, 1809, 1392, 1231, 926, 974, 975, 1308, 522, 1764, 412, 397, 1227, 1198]. growth [200]. Guichard [659]. gyrotropic [1154].

H [204, 4, 921, 1452]. H-atom [921]. Haar [2119]. Hadamard [1712]. hadrons [499]. half [2165]. half-plane [2165]. Hall [176]. Hamilton [1735, 1324, 2048, 30, 1867]. Hamiltonian [2223, 1169, 1911, 516, 219, 1578, 1577, 1495, 866, 983, 570, 1772, 1676, 1055, 952, 708, 290, 2011, 928, 927, 1842, 1700, 1531, 1001, 1530, 1250, 1855, 1773, 922, 1013, 355, 822, 1983, 2173, 2174, 1991, 1659, 1660, 735, 1543, 259, 1586]. Hamiltonians [257, 520, 1257, 1846, 1364, 1456, 1679, 2031, 1615, 2197, 359, 1052, 1468, 27, 1249, 450, 264, 630, 1498, 2128]. Hammerstein [1524]. Hanbury [1745]. handles [911]. Hankel [1941]. happens [1272]. hard [177, 1775]. hard-core [1775]. Hardy [1504]. Hardy-class [1504]. Harmonic [1844, 1435, 2112, 829, 830, 1839, 326, 552, 323, 1957, 664, 1450, 703, 1959, 696, 1059, 396, 324, 66, 1733, 738, 1584, 1585, 1621, 2224, 1170, 2166, 1096, 1904]. harmonics [1574, 418, 1132, 1875, 943, 2195, 1876, 2184, 2113]. Harris [1742]. Hartree [217, 1294, 345, 1994]. Hausdorff [197, 1650]. Hauser [393, 840, 993]. having [1874]. Hawking [364]. heat [767, 1976, 1314, 1202]. heavy [1136]. Heiles [983, 28, 546]. Heisenberg [917, 378, 1733, 98, 750, 347, 907, 296]. Heisenberg- [347]. helicity [1692, 827, 1936]. helium [643]. Helmholtz [418, 356, 1739, 1213, 1437, 476, 1819, 1820, 701, 442, 980]. Hénon [983, 28, 546]. hereditary [1055]. Hermitian [1577, 491, 1540]. Hertz [1154, 1748]. Hestenes [1960]. heterogeneous [1083]. hidden [1756, 1404]. Hierarchical [191]. hierarchy [1578, 1495, 233]. Higgs [1421, 2141, 1417, 1716, 755, 1560]. High [631, 1340, 262, 1508, 1078, 1356, 286, 1144, 496, 1901, 2197, 1268, 1269, 2136]. high-dimensional [2136]. High-energy [631, 262, 1078, 496]. high-frequency [1901]. high-order [2197]. Higher [1027, 1617, 6, 624, 1749, 2215, 1074, 1633, 1835, 1574, 1943, 265, 1986, 2149, 593, 1945, 1123, 1336, 2191, 254, 70, 1555, 979, 241]. Higher-dimensional [1749, 2215, 1986]. Higher-order [1617, 1633, 1835, 1943, 265, 2149, 1123, 2191, 979, 241]. Hilbert [1160, 1084, 821, 173, 865, 1962, 1056, 900, 1590, 1504, 36, 1544, 1251, 184, 225, 1362, 462, 213, 638, 1232, 330, 693, 1683]. Hill [2115]. Hinshelwood [391]. Hirzebruch [1288]. Hitchin [1288]. Hodge [680]. Hoenselaers [455]. hole [672, 1789, 1857]. holes [1003, 723, 1479]. hollow [1459]. holomorphic [734]. Holonomy [2158, 1028]. Homogeneous [2079, 1498, 229, 1171, 1443, 454, 1218, 1429, 184, 225, 298, 299, 535, 1399, 597, 338, 888, 1279, 733, 1014]. Homothetic [393, 568, 38, 39]. Hooft [1421, 410]. Hooke [1477]. horizon [543]. horizons [1596]. horseshoe [1930]. Hubbard [580]. Hull [618]. Hwa [1896]. hydrodynamic [1649, 2214]. hydrodynamics [1897]. hydrogen [1015, 1062, 2079, 1927, 1060, 221, 1771]. hyperangular [810, 1141].

Hyperbolic [146, 1577, 1618, 1619, 427, 2208, 117]. hyperfine [1570]. hypergeometric [2047, 423, 1472]. hypergraphs [1874]. hyperoctahedral [1432, 2153]. hyperspherical [1574, 1844]. hypersurfaces [24, 144]. Hypervirial [703, 555, 933, 1062]. Hypervirial-perturbational [1062]. hypothesis [1304, 747, 671, 1404, 554].

ice [182]. Icosahedral [1876]. ideal [2039, 44, 877]. identification [853]. Identities [1726, 1915, 1532, 1536, 1212, 228, 1545, 1369]. identity [1492, 2129, 955]. II [581, 537, 505, 168, 136, 830, 1785, 1185, 1057, 111, 1476, 367, 1183, 1187, 108, 864, 575, 121, 1926, 1910, 998, 237, 1321, 293, 803, 2125, 1196, 403, 304, 1802, 985, 1740, 620, 1360, 639, 968, 309, 1619, 149, 39, 1816, 793, 1624, 1827, 1230, 83, 1020, 343, 1345, 1601, 1269, 1663, 194, 1022, 924, 648, 381, 779, 1583, 262, 857, 1585, 762, 268, 1004, 124, 41, 2174, 598, 2145, 1063, 2162, 1275, 1876, 910, 318, 1050, 759, 429, 447, 974, 1660, 503, 1353, 1382, 157, 901, 1623, 245, 980, 1494, 826, 460, 1611]. III [613, 62, 1993, 585, 212, 2105, 945, 1917, 238, 294, 1934, 1334, 1051, 1828, 550, 1346, 1603, 1349, 771, 1005, 400, 2184, 430, 2140, 975, 1743, 1433, 1448, 1402, 256]. Ilamed [1318]. illumination [1105]. Image [322, 2097, 2096]. imaginary [1415]. Impact [1254]. impenetrable [872, 1268]. imperfect [1522, 281]. Implementation [1227]. importance [549, 1159]. important [469, 223]. Impossibility [1636]. imprimitivity [806]. improper [1346]. Improved [1993]. Improvement [1327, 176, 60]. impurities [1894]. impurity [1592]. incident [1951, 128, 826, 1402]. included [1081]. inclusion [1997]. incomplete [1769]. Incompressibility [775, 1458]. Incompressible [721, 1002, 1313]. inconsistency [1181]. Indecomposable [1434, 1187, 1812, 717, 1535]. Indefinite [1985, 36, 368, 793, 1792, 638, 1879]. Independent [446, 141, 291, 614, 1016, 988, 1617, 124]. independent-variable [124]. index [161, 1216, 1288, 756, 2166]. indicated [1253]. indices [2149, 860, 897, 1675, 1728]. indirect [297]. indistinguishable [1223, 1557, 508, 1948]. induced [1552, 403, 113, 402, 1493, 373]. induction [1725]. inelastic [1215]. Inequalities [1537, 888, 1598, 1703, 2178, 647, 1561, 2088, 2133, 2132, 2220]. inequality [1288, 1779, 514]. inequivalent [1864]. inertial [2206]. Infeld [618]. Infinite [1039, 79, 558, 950, 657, 727, 1848, 1035, 970, 1998, 712, 468, 2088, 1042, 500, 2177, 538, 1438, 1073]. Infinite-dimensional [1039, 950, 1035, 1042, 2177]. Infinitely [1841, 493, 2029]. Infinitesimal [1727, 1625, 1762, 1810, 86, 2028]. Infinities [415]. infinity [142, 1257, 2085, 1642, 1641, 745, 416, 40, 41, 939, 940, 770, 2059]. infrared [67]. inhomogeneities [1277]. Inhomogeneous [95, 79, 1952, 1107, 17, 7, 105, 1193, 1489, 598, 733]. Initial [1193, 78, 630, 1167, 540, 15, 51, 1857, 994, 279, 1555]. Initial-boundary-value [1167]. Initial-value [630, 540]. inner [62, 312, 613, 1407, 1792, 1070, 1430, 1320, 2157, 701]. inner-product [1430, 1320]. Inönü [1229]. instability [1481, 155, 1493]. instanton [530].

Instantons [1782, 45, 518]. integer [804]. Integrability [1615, 1834, 1772, 1632, 1816, 1961, 2030, 2028, 2016]. Integrable [1442, 1169, 1577, 983, 657, 606, 1614, 928, 1766, 1001, 1052, 477, 1737, 1250, 953, 1418, 1697]. Integral [647, 79, 406, 311, 579, 1839, 1240, 1180, 1514, 1311, 1721, 1722, 1947, 125, 960, 232, 313, 1321, 320, 934, 785, 1740, 6, 624, 1195, 1255, 1959, 719, 1699, 1343, 1248, 1958, 1208, 1893, 1123, 749, 89, 1302, 48, 1054, 753, 349, 21, 318, 2080, 2081, 2193, 559, 223, 1832, 1822, 2119, 111]. Integrals [2012, 478, 410, 1496, 931, 1104, 1439, 551, 656, 1246, 2114, 265, 1174, 1205, 1322, 810, 1141, 786, 933, 1128, 1237, 1410, 2218, 2178, 2120, 2219, 2197, 1179, 1702, 1339, 1359, 729, 1173, 1589, 556, 2160, 1773, 175, 661, 846, 1046, 1583, 2179, 1905, 3, 300, 765, 1784, 1117, 1313, 2111, 428, 429, 430, 1233, 153, 649]. Integrated [668]. Integration [350, 88, 1986, 1287, 1974, 1406, 1359, 661, 1258, 736, 172, 1866]. intense [1385]. Intensity [1387]. interacting [272, 75, 1563, 396, 2098, 385, 1170, 1991, 78, 735]. Interaction [367, 1326, 1478, 1016, 1570, 810, 1141, 747, 621, 141, 297, 924, 1603, 644, 156, 157, 2095]. Interaction-set [1326]. interactions [526, 1512, 1511, 2117, 484, 1335, 1619, 1510, 1965, 1442, 1152]. interface [657]. Interior [1114, 562, 688, 914]. intermediate [987, 1837, 1838]. Internal [1393, 640, 405, 653, 1521, 882, 227]. Interpolation [821, 1160, 1075]. interpretation [1682, 133, 1413, 778, 1275, 411, 2054, 832]. interrelations [27]. interval [2010, 1591]. intrachain [1932]. Intrinsic [522, 251]. Invariance [151, 1325, 1186, 1187, 1570, 675, 356, 413, 1151, 1465, 716, 1384]. invariances [1984]. Invariant [788, 1058, 701, 1594, 1933, 374, 1292, 812, 1381, 320, 934, 786, 1197, 2011, 1921, 1818, 2022, 717, 1077, 1624, 384, 1940, 510, 116, 1416, 1126, 2027, 1435, 1866, 135, 1097, 498]. invariants [248, 1520, 1131, 1449, 1137, 1532, 1619, 1536, 732, 27, 2176, 1249, 922, 1872, 218, 2014, 897, 1728, 1489, 979, 1098, 1498]. Inverse [1277, 2180, 1967, 699, 1707, 611, 1453, 959, 588, 1668, 463, 70, 268, 92, 537, 771, 1004, 1005, 990, 1592, 1996, 679, 1298, 2162, 2096, 128, 826, 1402, 1509, 1242, 1492, 1, 1033, 816, 1401, 2129, 797, 993, 1216, 1914, 2058, 769, 884, 731, 1238, 935, 987, 459, 1711, 388, 1212, 408, 1968, 2130, 1344, 400, 1214, 1954, 2083, 2131, 2142, 571, 1105, 182, 2007, 1832, 1902, 2200, 554, 1785]. inverse-energy-weighted [884]. inverse-source [408]. inversion [1492, 1500, 2162]. inversions [187]. invertibility [1220, 709, 1178]. investigation [1100]. involution [1841]. involutional [16, 342, 343, 2103]. involutions [209]. involving [1254, 1516, 933, 2178, 1864, 1941, 1233, 951, 1685, 1832]. ion [1452, 646]. ionic [1258]. IR [1226]. irreducibility [615, 1724]. Irreducible [653, 861, 1230, 1607, 1307, 820, 1989, 372, 1432, 925, 1228, 161, 760, 207, 1864, 1674, 1199, 1628, 1049, 1086, 1231, 110]. Irregular [1365]. Irreversible [2137, 1544, 997, 998, 2062, 2061]. irrotational [339]. isentropic [1835].

Ising [637, 1598, 2032, 2212, 147, 1416, 1074, 891, 2146]. Isobars [367]. isolated [2207]. isolation [502, 503]. Isometries [1788, 1980, 160, 2022, 1143]. isometry [1856, 522, 929]. Isomonodromy [619]. isoscalar [438]. isospectral [825, 1055, 1613]. isospin [1946, 2106, 1079]. isotropic [326, 552, 2060, 2127, 770, 888]. isotropy [1695, 2150, 2169]. isovector [367]. iterated [1492, 1180]. Iteration [553, 125, 82]. iterative [689]. IV [2181, 312, 1962, 1831, 1576, 1347, 1604, 175, 990, 1214, 301, 1903, 649, 117]. IX [535, 796].

J [203, 613, 1160, 242, 1720, 350, 614, 1281, 1886, 2216, 2180, 1566, 370, 1033, 1194, 1606, 1722, 436, 2214, 243, 947, 2102, 466, 1161, 2222, 1428, 798, 934, 624, 893, 467, 2213, 204, 2218, 2219, 158, 1429, 2223, 51, 1607, 437, 581, 1034, 244, 2221, 17, 105, 1129, 2179, 300, 2215, 894, 348, 536, 1608, 537, 1004, 1005, 1430, 2147, 2224, 349, 2217, 1084, 1130, 1085, 1162, 895, 2148, 2181, 52, 1006, 649, 53, 2220]. Jacobi [1582, 1324, 2048, 1035, 30, 1136, 1867]. Janis [1245]. Jet [215, 947]. Joint [1093, 1253, 1955, 132, 1502, 1831]. Jordan [915, 1487, 1140, 1488, 1097]. Jost [1849, 61, 270, 1454]. Jump [657]. JWKB [265].

Kac [1609]. kaleidoscopes [274]. Kaluza [915, 2025, 1554, 1860, 1710, 1329, 2199, 1399, 2087]. Kantowski [2169]. KdV [2161, 1898, 1836, 1668, 2015, 691, 815, 1983, 87]. Kepler [114, 697, 1060, 986]. kernel [1947, 1205, 34, 894, 1742]. kernels [1027, 1508, 1445]. Kerr [1979, 143, 1851, 995, 337, 1069]. Kerr-like [995, 337]. Killing [142, 1066, 543, 1684, 634, 912, 1978]. kinematical [739]. Kinematics [2069, 1544]. Kinetic [2089, 340, 2090, 879, 1284]. kink [104]. Kinks [388, 1499]. Kinnersley [455, 184, 225]. Klein [711, 98, 344, 1885, 666, 915, 174, 2025, 1783, 1554, 1860, 1710, 1329, 2199, 1399, 2054, 1078, 701, 2087]. KMS [1861, 877]. Kolmogorov [614, 291]. Konhauser [951]. Korteweg [1422, 956, 2187, 2188, 2189, 958, 970, 1113, 1291]. Kostant [1165, 2036]. Kovalevski [1971]. Kovalevski-like [1971]. Kowalewski [1816]. Kreın [1045]. Kronecker [539]. Krook [190]. Kruskal [2134]. Kruskal-like [2134].

L [649]. label [115, 1534, 1881]. labels [1393]. Ladder [764, 719]. Lagrange [667, 1634, 2007]. Lagrangian [2043, 1699, 1531, 1207, 1690, 1995]. Lagrangians [934, 2046, 320, 1921, 1210, 662, 1950]. Laguerre [1941]. Lamb [1682]. Landau [1417, 2191]. Langer [1213]. Langevin [1817, 403, 1645, 1222, 1493]. Langmuir [391, 156, 157, 81]. Laplace [1205, 1322, 102, 541, 175, 649]. Laplacian [2195]. Large [1506, 1270, 239, 721]. Lateral [395]. Lattice [526, 1150, 383, 1939, 406, 424, 417, 341, 786, 404, 467, 609, 1075, 285, 1375, 119, 247, 1223, 1557, 1737, 1738, 2094, 636, 1734, 1263, 2088, 2093, 1074, 1948, 2013, 1377, 2073, 2175, 182, 2146, 525]. lattices [1906, 607, 1331, 2062, 2137, 2173, 2174, 383, 932, 2061, 458]. law [391, 2164, 399, 833, 721]. laws [1356, 1157, 23, 1172, 1950, 627, 2053, 695, 1586]. Lax [2116, 1494]. Lax-pairs [2116]. layer [1280, 2143]. layered [797]. LCAO [1805]. LCAO-type [1805]. Leading [1924]. least [1784, 1892]. Lee [1896, 2147, 998, 1666, 649]. left [497, 145]. left-flat [145]. Legendre [1027, 1184, 1832, 818, 1040, 1438]. Lenz [2121]. level [1563, 1229, 2098]. levels [1366, 138, 1073]. Levitan [798, 894, 269, 34, 1591, 1708, 990, 679, 1969]. Lewis [64]. Lie [804, 248, 246, 617, 976, 1629, 1652, 949, 763, 574, 575, 950, 1767, 1487, 1676, 1882, 1883, 133, 1626, 86, 805, 2108, 1884, 1766, 981, 1286, 1632, 1039, 2109, 790, 970, 1940, 2183, 1616, 1398, 1110, 1111, 472, 2066, 1765, 1627, 2107, 897, 1675, 208, 764, 1061, 118, 354, 977, 1571, 411, 584, 1244, 1392, 1231, 1670, 1671, 2036, 926, 974, 975, 1308, 1730, 1572, 1764, 1630, 1814, 2119, 1198]. Lie-algebraic [950]. Lifshitz [2191]. lifting [1229]. Light [1642, 196, 1641]. light-cone [196]. Lightlike [1776]. like [2134, 1846, 1147, 995, 914, 337, 193, 425, 1971]. limit [558, 147, 692, 329, 149, 917, 1268, 1269, 1855, 180, 500, 1994, 1384, 1837, 1502]. Limitations [1385]. limited [1370]. limiting [1681, 1204, 3, 300]. limits [1330, 714, 479, 1341]. line [1044, 212]. Linear [945, 1917, 812, 1067, 1250, 848, 1034, 572, 673, 1022, 1098, 2181, 1903, 1727, 526, 1652, 766, 1952, 1579, 1895, 220, 785, 1986, 1168, 653, 2063, 773, 1286, 1826, 1500, 366, 1236, 137, 1653, 1390, 875, 253, 979, 1392, 1671, 223, 733, 597, 598]. linear-antilinear [653]. Linearization [242, 22, 152, 489, 1162, 1612, 1351, 2075, 1813]. Linearized [17, 7, 105, 1979, 916]. Linearly [531]. lines [146, 1238, 1011, 1341]. Linkages [1257, 563]. Liouville [688, 1133, 569, 2013, 2159, 989]. Lipkin [517]. Lippmann [1640, 899]. liquids [942]. little [1797]. Livsic [222]. Lobachevsky [1822]. Local [1186, 1187, 1470, 1921, 1414, 573, 743, 1943, 1556, 113, 702, 359, 1592, 276, 1537, 584, 412, 1526, 1135, 2208, 181]. localization [1901, 550]. localizations [1868]. Localized [391, 193, 1041, 595]. Locally [2070, 2179, 1905]. located [703]. logarithmic [2192]. logics [2122, 1251]. logs [1924]. Long [972, 219, 1512, 1849, 987, 1966, 1837, 1838]. Long-range [972, 219, 1512, 1849, 1966]. longitudinal [1276, 749, 721]. look [774]. Lorentz [804, 111, 585, 1354, 1424, 486, 46, 1443, 1271, 549, 717, 1151, 196, 209, 880, 1953, 941, 56, 768, 1568, 1628, 1806, 2152, 1623]. Lorentz-transformations [46]. Lorentzian [1469, 1341]. Low [1515, 2146, 149, 1940, 1834, 68, 1073]. low-dimensional [1940, 1834]. Low-energy [1515]. low-lying [1073]. Löwdin [2148, 1871]. Lower [138, 869, 1355, 178, 1955, 362, 1141, 1636, 176, 327, 668, 577]. lowering [948, 86, 474]. lowest [1355]. lump [621]. lumps [287]. Lund [1219]. Lyapunov [1235, 1815, 2144, 2145]. lying [1073].

Maartens [1317]. Machian [1486, 1459]. macroscopic [534, 899]. Magnetic [101, 286, 1131, 579, 788, 1481, 1852, 454, 240, 723, 580, 1723, 1998, 1789, 941, 1007, 1798, 519, 1029, 1550, 1991, 379, 434, 1167]. Magnetization [1517]. magnetized [1931]. magnetogasdynamics [80]. magnetohydrodynamic [1314]. magnetohydrodynamics [722, 1694]. magnetohydrostatic [1909]. magnon [904]. Maluzhinetz [179]. Mandelstam [850, 1185, 2209, 1990]. Manifestly [1292, 1443, 1145]. manifold [2118, 767, 1299, 545, 221, 1011]. manifolds [1541, 1919, 1171, 199, 1896, 2040, 1816, 937, 1589, 591, 1129, 1288, 72]. many [1260, 325, 1756, 1471, 359, 2023, 2029, 1841, 93, 112]. many-body [1260, 325, 1756, 93]. many-center [359]. many-parameter [2023]. map [1045, 559]. mapping [2033, 319, 1182]. mappings [728, 1009]. maps [130, 1117, 1829]. Marchenko [2180, 1967, 690, 990]. marginals [2041]. Markov [1335, 1751]. Markovian [998, 1862, 1323]. Martin [1327]. Mass [1142, 1520, 605, 493, 1477, 1516, 1851, 2078, 368, 793, 1272, 1714, 1518, 195, 1864, 2093, 838, 1021, 642, 455, 1023]. mass-eigenvalue [1714]. mass-spectrum [1477]. masses [2012]. Massive [492, 101, 2050, 2021, 1548, 749, 1690, 1384, 1813]. Massless [2087, 829, 830, 101, 1797, 1339, 2204, 2027, 969]. master [1523]. master-equation [1523]. Matching [229, 2053, 498, 458]. Material [1860, 513, 959]. materials [1083]. Math [203, 613, 1160, 242, 1720, 350, 614, 1281, 1886, 2216, 2180, 1566, 370, 1033, 1194, 1606, 1722, 436, 2214, 243, 947, 2102, 466, 1161, 2222, 1428, 798, 934, 624, 893, 467, 2213, 204, 2218, 2219, 158, 1429, 2223, 51, 1607, 437, 581, 1034, 244, 2221, 17, 105, 1129, 2179, 300, 2215, 894, 348, 536, 1608, 537, 1004, 1005, 1430, 2147, 2224, 349, 2217, 1084, 1130, 1085, 1162, 895, 2148, 2181, 52, 1006, 649, 53, 2220]. Mathematical [1310, 997, 998, 1321, 704, 1090, 1612, 1284, 909, 910]. Mathieu [1815, 123]. Matrices [582, 1490, 1376, 1283, 389, 1008, 1447, 433, 1808, 2132]. Matrix [654, 1812, 676, 2161, 1933, 111, 1297, 417, 816, 441, 1880, 961, 1844, 727, 414, 1471, 1689, 1470, 4, 204, 1035, 440, 805, 1441, 1248, 222, 16, 581, 378, 1109, 1663, 1212, 271, 1824, 385, 1065, 1256, 154, 1637, 2142, 836, 1600, 2000, 920, 1342, 1997, 2074, 538, 906, 1742, 1472, 1473, 1455, 2133, 2220, 16]. matrix/inclusion [1997]. matter [339, 95, 1747, 735]. Maupertuis [1892]. Maximal [1571, 1700, 2223, 739]. maximality [549]. Maximum [2042, 991, 1243, 194]. Maxwell [711, 51, 426, 1476, 1183, 336, 1974, 1914, 36, 2023, 1529, 15, 277, 278, 332, 1518, 344, 1858, 456, 260, 1069]. Mayer [1121, 1415]. Mean [1890, 1263, 74, 1259, 1400, 2073]. meaning [1029, 968]. means [1492, 46, 1799, 623]. Measure [1589, 175, 1866, 649]. measured [1252]. Measurement [1140, 515, 1172, 832]. Measurements [1251, 1636]. measures [1174, 1122, 1173, 2179, 1905, 1042, 2173, 2174]. measuring [1704]. mechanical [1016, 516, 1018, 1012, 1547, 2102, 1207, 1662, 1930, 2165, 872, 832, 2019, 90]. mechanics [1520, 446, 173, 865, 219, 1094, 1955, 490, 1736, 1778, 824, 997, 998, 606, 1781, 984, 985, 767, 1171, 199, 704, 1367, 1590, 359, 1332, 2089, 550, 2170, 1409, 789, 378, 175, 1620, 695, 1791, 1668, 1774, 957, 316, 138, 1047, 1363, 221, 32, 1874, 428, 429, 430, 515, 1278, 1658, 1253, 1097, 2049, 649, 596, 518, 1782, 451, 1816]. mechanism [1181, 1716]. Mechanisms [1378]. media [1917, 394, 1081, 775, 513, 129, 421, 1114, 1500, 155, 959, 673, 1022, 1661, 253, 1154, 1105, 733]. medium [390, 1952, 611, 1963, 512, 1901, 769, 899, 1107, 2144, 2145, 612, 321]. Mehler [1205]. Melnick [1144]. Melnikov [1001]. Memory [875, 1932, 1947, 2163, 1203]. memory-dependent [2163]. meson [367, 237, 1977]. mesons [367]. metacyclic [683]. metals [297]. metaplectic [471]. method [2044, 1843, 2038, 65, 311, 178, 183, 766, 1514, 1311, 1721, 1722, 217, 1436, 1836, 1875, 997, 77, 1505, 1703, 2023, 769, 2069, 1001, 1213, 2076, 16, 581, 792, 175, 1645, 689, 1046, 1176, 213, 63, 1212, 2186, 2056, 1083, 1605, 1876, 2184, 182, 442, 980, 649, 256, 255]. methods [1676, 539, 421, 530, 1209, 138, 1805, 2131, 322, 811, 548, 906]. metric [2134, 1979, 1071, 143, 892, 1851, 1266, 393, 36, 1985, 777, 1177, 995, 591, 1129, 1288, 967, 638, 1790, 1646, 9, 52, 1879]. metrics [1551, 1300, 1517, 632, 298, 736, 1069, 2208]. Mickelsson [948]. microscopic [1274, 534, 1030, 1196, 1934]. microstrip [103, 392]. Micu [248]. Micu-type [248]. Mills [1421, 1922, 99, 675, 76, 1340, 1266, 1800, 785, 751, 893, 333, 1380, 1529, 1148, 1420, 749, 787, 1264, 1923, 1665, 70, 531, 1798, 1466, 1467, 2159, 1866, 2028]. minima [1417]. Minimal [1475, 1476, 210]. minimizing [1706]. minimum [1491, 447, 1847]. Minkowski [146, 58, 787, 1264, 1918, 1467, 1776, 1134, 481]. Minkowskian [41]. miracle [2129]. misleading [1924]. Missing [115, 1534, 1881]. Miura [376]. Mixed [1935, 1483, 663]. mixing [883]. mixtures [1258]. Mobility [29]. modal [202]. mode [1931, 815, 962]. model [1894, 829, 830, 1562, 1667, 1984, 637, 1672, 190, 286, 558, 579, 517, 1419, 391, 1012, 1713, 2032, 2212, 147, 1184, 998, 2190, 1563, 1381, 494, 1679, 2021, 1548, 358, 1464, 280, 1753, 144, 1259, 793, 1421, 917, 580, 1891, 972, 2010, 1416, 876, 753, 2210, 1565, 730, 644, 1219, 295, 832, 1478, 595, 886, 1949]. Modeling [1631]. Models [198, 1351, 1933, 1598, 516, 1376, 1274, 534, 1649, 2214, 453, 75, 1844, 2090, 2122, 74, 1558, 290, 1975, 477, 2029, 339, 2110, 231, 648, 535, 2173, 2174, 338, 1909, 430, 2101, 1600, 182, 2169, 1705, 2128]. modes [103, 392, 2070, 530, 645, 718]. Modified [766, 822, 897, 958, 1640, 6, 624, 1595, 20, 846, 63, 377, 1838, 2101, 1153, 1382, 1743, 1803]. modular [2189, 1309]. modulation [781]. module [1683]. modules [2108, 1938]. modulus [797]. molecular [2069, 792, 1897]. molecules [623]. Molien [1131]. Möller [214, 1178]. Moment [509, 1991]. momenta [1221]. moments [101, 819, 2042]. momentum [287, 326, 552, 1955, 1366, 1936, 911, 149, 1267, 249, 1191, 581, 1872, 360, 1966, 271, 1573, 1621, 2224, 1270, 500, 498, 1454, 2133, 2132, 2220]. momentum-dependent [1966]. Monodromy [834]. monoenergetic [1863]. monopole [1380, 1267, 1080]. monopoles [788, 2086, 770, 1550, 45]. monotone [831]. morphisms [1178]. Morse [2166, 1507, 1342]. Morse-oscillator [1342]. most [1810, 110]. motion [1276, 1246, 1582, 2206, 2012, 2078, 1259, 917, 986, 1542, 2221, 1145, 1733, 852, 941, 1666, 2147, 1841, 250, 641, 32, 1806, 2140, 1658, 1149, 1543, 2198, 1136]. motions [445, 393, 1458, 38, 39, 2168, 150, 14, 244, 682, 1136]. Moving [1206, 1241, 1656, 1589, 11, 348, 2211, 2072, 2217]. MR [613, 242, 1720, 1886, 1566, 1194, 1722, 436, 2214, 243, 947, 466, 1161, 2222, 798, 934, 624, 467, 2218, 2219, 2223, 581, 244, 2221, 105, 1129, 2215, 894, 536, 1608, 537, 1004, 2224, 349, 1162, 649, 2220, 1033, 300]. MRT [2173]. multi [646]. multi-ion [646]. multichannel [289]. multicomponent [561]. Multidimensional [662, 2058, 1711, 328, 597, 598, 1295, 630, 1290]. Multifrequency [1902, 2200, 318]. multigroup [291, 614, 82, 459]. multimode [409]. multiparticle [740]. multiphoton [921]. Multiple [394, 1336, 1081, 1343, 1442, 962]. multiplets [755]. multiplication [1234]. multiplicative [5, 53]. multiplicities [1226, 1864]. multiplicity [2002, 763, 205]. multiplicity-free [205]. multipliers [667, 2007]. multiplying [291, 614]. Multipole [632, 1041, 293, 294, 302, 398, 1082, 513, 213, 63, 1798, 1457, 1858, 31, 1859, 292, 301, 397, 795, 796, 1483, 1484, 1014]. multipole-phonon [293, 294, 302, 398, 292, 301, 397, 795, 796, 1483, 1484]. multipole-photon [1082]. Multipoles [241]. multiquadratic [889]. multiquark [642]. multisoliton [1291]. multispecies [2039]. Multispinor [1878]. multispinors [641]. multitime [1330]. multivariable [1687]. multivariate [2041]. Mutual [1386].

N. [265]. Narrow [286]. natural [1777, 754]. naturality [502, 503]. Nature [314, 2206, 506]. Navier [1694, 1313]. near [350, 88, 119, 233]. near-Coulombic [119]. near-resonant [350, 88]. nearest [526, 508, 636, 1734]. nearest-neighbor [1734]. Necessary [2155]. neighbor [526, 508, 636, 1734]. Nel [1317]. Nelson [1135]. nested [821, 1160]. net [911]. networks [449, 678]. Neugebauer [1101]. Neumann [607, 1331]. neutrino [1856, 1143, 337, 299, 718]. neutrino-gravitational [337]. neutrinos [1865]. neutron [1225, 82, 1480, 1383]. Neveu [1381]. Newberger [1886]. Newman [1979, 143, 1851, 1245, 839]. Newton [1552, 307, 255]. Newtonian [2121, 1702, 1855, 1461, 2059]. next [526]. Nikodým [831, 1501]. nilpotent [2185]. NLS [2026]. no [613, 242, 1886, 1566, 1033, 1194, 436, 2214, 243, 947, 466, 1161, 2222, 798, 934, 624, 467, 2218, 2219, 2223, 581, 244, 2221, 17, 105, 1129, 300, 2215, 894, 536, 1608, 537, 1004, 2224, 349, 1162, 52, 649, 53, 2220]. Noether [732]. noise [403, 402, 1751, 1222, 1493]. noise-induced [403, 402, 1493]. Non [862, 1785, 1548, 1486, 269, 798, 881, 1375, 1862, 504, 1323, 1613]. non-abelian [881, 504]. non-Bravais [1375]. Non-Grassmann [1548]. non-Machian [1486]. non-Markovian [1862, 1323]. Non-self-adjoint [862, 1785, 269, 798, 1613]. nonabelian [372, 417, 1495, 1648, 2222, 1024, 1267, 1539]. nonabsorbing [959]. nonaxial [392]. nonbijective [1094]. noncanonical [1013, 1170]. Nonclassical [1241, 1635]. noncommuting [2056]. noncompact [574, 575, 2179, 1905, 2082]. Noncomplex [808]. nonconservative [1736, 2049]. nonconserved [1424]. noncovariant [1183, 1464]. noncrystallographic [1347]. nondensely [693, 1084]. nondifferentiable [1774]. Nondiffractive [274]. Nonequilibrium [283, 1261]. Nonexistence [2167, 1817, 76, 1217]. nonexpanding [382, 779]. nonghost [2017]. Nonglobal [603]. nonglobally [427]. nonhomogeneous [1917, 673, 1022]. noninertial [127]. nonintegrable [983, 1895]. nonintegral [1150]. Noniterative [2023]. Nonlinear [1449, 405, 1107, 1539, 1427, 979, 1528, 2140, 10, 203, 505, 122, 168, 868, 2167, 1355, 212, 1043, 579, 170, 1240, 1180, 1606, 1482, 626, 540, 863, 864, 441, 1564, 2067, 1436, 251, 224, 2196, 57, 13, 466, 1676, 1946, 494, 152, 2058, 952, 1899, 1818, 769, 660, 981, 1357, 2076, 732, 655, 1625, 987, 2191, 1398, 1773, 922, 282, 355, 1126, 1944, 415, 416, 98, 750, 29, 1112, 962, 875, 1203, 730, 1351, 193, 828, 448, 250, 514, 2075, 2151, 677, 1600, 2193, 919, 2030, 1538, 1813, 1900, 156, 157, 887, 554, 2004, 1222]. nonlinear [1493, 1580, 863]. nonlinearities [1579]. nonlocal [600, 727, 2163, 1807, 1689]. nonmanifold [742]. nonnegative [1295]. nonnegativity [1021]. Nonorthogonal [444]. Nonperiodic [81]. Nonperturbative [850, 1185, 1993]. nonquantum [1631]. Nonrecurrence [1015]. Nonrelativistic [483, 520, 1476, 325, 362, 1385, 807, 2079, 1409, 223, 67, 451]. nonrotating [1245, 1747]. Nonscalar [802, 803, 1433, 397, 796]. nonsimply [702]. Nonsingular [1640, 1974]. nonsingularity [1556]. nonspherical [984, 985, 1268, 1269, 1623]. nonstandard [1889, 1042]. nonstatic [1793]. nonstationary [192]. nonsymmetric [1554, 1860, 1710, 994, 279]. nontrivial [96, 1765, 2024]. nontwisting [842, 43, 1597]. nonuniform [391, 769, 1105]. Nonuniqueness [1, 1033, 1954]. nonvanishing [2205]. Nordtvedt [780]. Normal [1587, 1487, 2070]. Normalization [1381, 1338, 266]. Note [594, 9, 52, 799, 422, 827, 28, 546, 335, 2103, 1745, 145, 109, 245, 25, 525]. novel [2120, 2219, 1714]. nuclear [1274, 534, 290, 2210, 1153, 920, 1382, 1743]. nucleon [462]. Null [142, 1285, 1553, 634, 1369, 1748, 2059, 1257, 892, 1220, 709, 1118, 1643, 1642, 1856, 1420, 1641, 940, 1970, 2024, 1971, 1786, 1384, 1069, 1461]. null-path [1420]. null-plane [1384]. Number [325, 2201, 2216, 323, 1225, 1908, 1035, 1452, 1759, 306, 800, 902, 514, 1106, 1073]. numbers [721]. Numerical [853, 661, 71, 1742, 1782, 217]. Numerov [217]. NUT [1221].

object [1151, 852, 1623]. Objects [895, 616, 1299, 384, 273, 2091]. obliquely [1951]. observability [33]. observable [1259, 1990]. Observables [1252, 1955, 999, 1093, 592, 1704]. observations [1620]. Observer [454]. observers [1469, 909, 910, 929, 1581]. obstacle [1115]. obtaining [1055, 795]. occupation [1557, 713, 1948]. occupational [1223]. ocean [886]. octonion [1749, 2215]. Octonionic [1572, 1915]. Odd [505]. Off [1473, 600, 727, 1849, 1681, 329, 201, 464, 71, 1639, 1455]. Off-shell [1473, 600, 727, 1849, 1681, 71, 1639, 1455]. Ohm [1694]. old [2076]. old-new [2076]. On-shell [1849]. One [945, 1260, 1727, 666, 1156, 1901, 359, 2205, 177, 2057, 917, 1684, 1302, 66, 1249, 179, 1258, 904, 713, 706, 1193, 636, 1361, 1981, 1968, 1592, 2130, 1074, 1447, 2072, 2217, 1885, 644, 1397, 1528, 2203, 442, 980, 1835, 1829, 434, 2061, 1498, 2095, 259, 1030, 92]. one- [1968, 2095]. One-body [945]. one-dimensional [1260, 1727, 1156, 1901, 177, 2057, 917, 1302, 66, 1249, 179, 1258, 713, 1193, 636, 1361, 1981, 1074, 1447, 1397, 1528, 2203, 1835, 1829, 2061, 1498, 259, 1030]. one-fixed-point [1684]. one-particle [359]. one-space [666]. one-time [666]. Ono [23, 1837]. Operator [1505, 421, 1403, 153, 2161, 862, 1785, 1570, 1888, 1579, 293, 294, 302, 398, 802, 803, 1082, 2125, 2077, 2117, 414, 452, 1267, 307, 396, 1234, 1117, 519, 361, 902, 680, 1149, 292, 301, 397, 795, 796, 2071, 1433, 1483, 1484, 2124, 1580, 256]. operator-valued [1117]. operatorial [2045]. Operators [1294, 1293, 1621, 2224, 976, 1045, 725, 726, 1629, 115, 948, 1652, 1534, 325, 1511, 582, 1895, 997, 269, 798, 91, 302, 363, 214, 1763, 590, 692, 376, 620, 1794, 86, 474, 928, 1700, 2223, 618, 1826, 1431, 1039, 1536, 249, 1856, 1762, 1810, 599, 1510, 1965, 194, 822, 1212, 1361, 743, 2107, 2056, 360, 351, 1966, 1390, 1805, 479, 1061, 693, 1084, 676, 1671, 263, 2157, 447, 2116, 33, 642, 2126, 1764, 292, 301, 397, 796, 1814, 1433, 1483, 1881, 1526, 2074, 2095, 139, 1613]. opposite [1368]. optical [2180, 1967, 1513, 647, 2010]. optics [65, 1564, 794, 1386, 1387, 49, 1970, 677]. Optimal [1261, 1047, 2132]. Orbit [1417, 1940, 1643, 986, 1542, 1392]. Orbital [287]. orbitals [1362]. orbits [728, 698, 1118, 977, 1829]. order [10, 203, 505, 1356, 1943, 1606, 1482, 1875, 1617, 265, 810, 1141, 485, 1986, 6, 2058, 624, 2021, 952, 2149, 1237, 593, 2197, 1204, 1950, 1039, 580, 1123, 1336, 1998, 2191, 923, 924, 1603, 1604, 1719, 972, 1931, 1236, 1627, 897, 1508, 1876, 979, 1633, 318, 936, 1658, 433, 1538, 612, 1835, 241, 2004, 259, 814, 1715]. ordered [1701, 1415]. ordering [153]. orders [1574]. ordinary [122, 168, 371, 655, 2151]. oriented [1050]. orthocomplementations [1778]. Orthogonal [1220, 709, 1825, 2010, 1371, 1168, 654, 541, 26, 380, 1008]. Orthogonality [1778]. orthogonalization [555]. orthogonally [1856]. orthonormal [623, 353]. orthosymplectic [1652, 1061, 1671, 2177]. oscillating [1481, 363]. oscillations [1436, 81]. oscillator [829, 830, 111, 585, 8, 370, 1449, 1839, 326, 552, 323, 1907, 961, 1957, 1450, 703, 1840, 978, 1959, 64, 696, 1059, 741, 1020, 66, 738, 1584, 1585, 2186, 682, 360, 1621, 2224, 1095, 1408, 447, 2123, 1096, 1342, 1098, 90]. Oscillators [375, 1843, 1016, 2047, 1322, 664, 1506, 396, 324, 510, 2127, 78, 1451]. Oscillatory [1314]. other [1985, 1442, 2007]. other-than-Feynman [1985]. outer [1394]. overall [1997]. overdense [201, 464]. overlapping [497]. overview [2187]. packets [482]. Padé [1291, 213, 63, 625]. Page [1485]. Painlevé [1772, 169, 1201, 2197, 1816, 1398, 1834, 2159, 1358, 1494, 1697]. Pair [1294, 1293, 713, 1856, 443, 1600, 1097, 1881]. pairs [2116, 1494, 458]. Palatini [780]. paper [1886, 2221, 2225, 649]. Para [261, 107, 133, 1434, 1039, 790, 360, 1061, 263, 447, 1061]. para-Bose [133, 1434, 1039, 360, 263, 447, 1061]. para-boson [790]. Para-Fermi [261, 107, 1061]. Para-Grassmann [107]. Parabolic [988, 1354, 483, 1617, 1739, 1740]. paradigm [626]. parallel [1116, 1105]. parameter [1254, 160, 1118, 2023, 1856, 729, 1684, 668, 1600]. parameters [2068, 2183, 1008, 1803]. parametrization [761, 1352, 1353]. Parametrizations [389, 1650]. parametrized [1702, 1146]. parastatistics [1521]. Paraxial [677]. parities [1138]. parity [1292, 1452, 1272]. part [2166, 2073, 2028]. Partial [62, 312, 613, 586, 1566, 297, 371, 484, 37, 1357, 1407, 124, 597, 598, 2008, 2157, 1358, 1494, 2016, 1454, 1472, 1455]. Partial-range [586, 1566]. partial-wave [1472]. Partially [2035]. Particle [419, 1419, 113, 1833, 1693, 1717, 2050, 1839, 1240, 1180, 1524, 1515, 1422, 457, 1752, 340, 1271, 1863, 359, 1797, 1958, 986, 1872, 846, 1145, 179, 271, 969, 753, 1633, 514, 1991, 1806, 112, 1647, 1859, 157, 2095, 353, 1678, 1639, 282]. particles [1692, 101, 325, 1296, 1385, 2078, 1640, 421, 288, 807, 358, 368, 793, 508, 378, 1268, 1269, 2204, 1690, 1211, 1944, 2027, 1981, 527, 1170, 2072, 2217, 1425, 431, 1374, 1152]. particular [2077, 271, 1588]. Partition [290, 740, 77, 2021, 428, 2036]. Partitioning [869, 1368]. partitions [93, 1874]. partly [1306]. passage [1545]. passive [1525]. Passivity [570]. Path [1104, 551, 1406, 240, 1195, 1343, 1702, 89, 736, 753, 1947, 215, 947, 960, 1174, 1255, 1410, 1959, 2218, 1958, 729, 1420, 1173, 1589, 556, 175, 1046, 1583, 3, 300, 1784, 1117, 1701, 318, 2111, 559, 1866, 649, 1740]. Path-dependent [240]. path-ordered [1701]. path-parameter [729]. paths [1287, 960, 1122]. pattern [435, 1155]. patterns [391]. Pauli [1364, 750]. peak [601, 668]. Peierls [891]. pendulum [626, 908]. penetrable [1770]. Penrose [839]. Percolation [458, 1416]. Percus [1156]. Perfect [2101, 566, 1301, 1910, 1852, 1102, 1855, 1746]. Perimeter [542]. Periodic [346, 551, 728, 956, 2117, 1088, 280, 1603, 1361, 1592, 1427, 872, 1064, 1829, 1290]. permeability [611]. permittivity [611]. permutation [1037, 1673, 1761, 1231, 1611]. Permutational [1036, 112]. permuting [290]. Perturbation [289, 1215, 1297, 2188, 530, 139, 2202, 666, 766, 1821, 593, 1218, 1429, 884, 1575, 1013, 1585, 2100, 1845, 954, 1605, 1244, 1546]. perturbational [1062]. perturbations [1979, 1001, 738, 748, 1383, 451]. Perturbative [1783, 1176, 2156, 1924, 774]. Perturbed [866, 1680, 1341]. Pfaffian [1612]. Phase [1911, 238, 550, 3, 300, 1348, 223, 1368, 136, 1104, 1138, 2155, 390, 631, 125, 1156, 265, 236, 237, 600, 281, 6, 624, 1410, 2218, 2018, 1695, 1589, 89, 1416, 876, 264, 651, 271, 969, 233, 936, 1079, 30, 1411, 2080, 2081, 1651, 1904]. Phase-integral [223, 2080, 2081]. phase-integrals [265]. phase-plane [2018]. phase-shifts [271]. Phase-space [550, 1410, 2218]. phenomena [96, 1304, 49, 887]. phonon [91, 293, 294, 302, 398, 292, 301, 397, 795, 796, 1483, 1484]. phonon-states [91]. photoelectron [1493]. photon [1181, 1082, 4, 204, 2006, 409]. Phys [203, 613, 1160, 242, 1720, 350, 614, 1281, 1886, 2216, 2180, 1566, 370, 1033, 1194, 1606, 1722, 436, 2214, 243, 947, 2102, 466, 1161, 2222, 1428, 798, 934, 624, 893, 467, 2213, 204, 2218, 2219, 158, 1429, 2223, 51, 1607, 437, 581, 1034, 244, 2221, 17, 105, 1129, 2179, 300, 2215, 894, 348, 536, 1608, 537, 1004, 1005, 1430, 2147, 2224, 349, 2217, 1084, 1130, 1085, 1162, 895, 2148, 2181, 52, 1006, 649, 53, 2220]. physical [1906, 65, 871, 1632, 1816, 1530, 638, 2001, 832]. physically [469, 223]. physics [931, 130, 1886, 1242, 1888, 674, 1090, 1612, 557, 1089, 1608, 93]. Picard [618]. picture [323, 1558, 1527, 2141]. piecewise [960]. piezoelectricity [1807]. pinch [757]. Planar [1823, 180, 1443, 83]. Planck [349, 1863, 1072, 21, 1323]. Plane [2143, 1373, 791, 229, 611, 1385, 2058, 1128, 772, 1975, 2018, 342, 201, 464, 567, 2017, 1665, 1929, 2165, 1384, 836, 420, 1105, 1718, 322]. plane-parallel [1105]. Plane-surface [2143]. plane-symmetric [2017]. plasma [1886, 1606, 1482, 1127, 848, 1034, 923, 924, 1603, 1604, 1719, 17, 7, 105, 1931, 63, 625, 1089, 1608, 1769]. plasmas [1757, 201, 464, 1603, 1949]. plate [1951]. plates [1314]. Plebański [839]. plethysms [860, 473]. plus [1803, 1639, 1473]. Poincaré [934, 136, 320, 1936, 407, 2182, 2176, 227, 880, 1716, 1799, 1674, 500, 1813]. point [1131, 1963, 1875, 1821, 615, 2117, 1207, 760, 1345, 1346, 1347, 1723, 1998, 1684, 1668, 2009, 1784, 2072, 2217, 1876, 2184, 1418, 525]. pointlike [2050]. points [1520, 1166, 932, 172, 887]. Poiseuille [420]. Poisson [960, 2006, 119, 2109, 569, 1247, 2009, 2039, 2154, 811]. polarizability [889, 104]. polarization [1777]. polarized [1692]. polaron [346, 295]. pole [688, 1298]. poles [865, 289, 836, 1444, 907]. polycnoidal [2189]. polymer [1932, 1580]. Polynomial [2026, 2221, 2198, 1610, 1688, 1237, 1615, 2182, 935, 2160, 1236, 415, 416, 1941, 377, 2123]. polynomials [493, 582, 120, 121, 2106, 654, 541, 247, 2010, 1761, 1435, 951, 1832]. polytensors [1389]. Polyzou [562]. Popov [1124, 716, 2175, 431]. porous [2144, 2145]. position [1955, 482, 307, 1272, 360]. Positive [132, 1664, 692, 687, 557, 361]. Positive-definite [1664]. Positivity [2018, 1854, 1815, 148]. possessing [1118]. possible [1795]. Post [2121, 176]. Post-Newtonian [2121]. Postnikov [784]. Potential [1527, 1368, 862, 1785, 422, 1156, 57, 58, 690, 1041, 1781, 1593, 1406, 2068, 1450, 2063, 1927, 1870, 935, 1940, 1302, 1176, 137, 271, 1252, 1592, 1170, 1303, 1256, 193, 828, 1142, 239, 1078, 1647, 2123, 1847, 1507, 223, 1705]. potentials [1475, 8, 370, 656, 1018, 631, 824, 1964, 1161, 903, 600, 727, 1000, 1366, 1428, 2164, 363, 214, 1849, 1681, 6, 1679, 624, 329, 1615, 37, 2089, 1210, 1333, 267, 1417, 399, 278, 986, 599, 355, 1774, 1775, 1361, 2186, 1953, 1048, 1966, 1447, 1154, 448, 1365, 1017, 1411, 1224, 1748, 61, 270, 739, 1803, 1639, 1473]. Power [622, 2038, 2164, 399, 833, 239, 1078, 1647]. power-law [399, 833]. powers [2138, 819, 1342, 1832]. pp [1913]. pp-wave [1913]. Practical [1324, 2064]. Pre [937, 900]. Pre-atlas [937]. pre-Hilbert [900]. precession [1806]. predictive [1530]. presence [1172]. present [1804]. preserve [1646]. preserving [1636, 834]. presymplectic [1896]. primordial [1804]. principal [413, 1753, 380, 975]. principle [1833, 1952, 457, 199, 492, 1218, 1429, 194, 879, 310, 780, 1892]. principles [2043, 1441, 667, 378, 971, 1120, 1774, 1537, 1063, 2151]. printed [1359]. probabilistic [131, 1057, 1687, 1448]. Probability [2041, 2040, 2173, 2174, 1631, 1056, 713]. probes [1552]. Problem [1293, 114, 1785, 1242, 286, 1477, 178, 1, 1033, 1240, 1180, 1094, 763, 417, 816, 1401, 1015, 797, 1312, 1474, 698, 134, 747, 639, 1375, 1619, 1985, 184, 225, 177, 176, 731, 1530, 697, 1060, 1238, 935, 963, 2052, 1714, 1409, 423, 1857, 459, 28, 546, 509, 1584, 1585, 235, 463, 994, 408, 2042, 279, 1968, 187, 1232, 1344, 694, 1885, 2083, 932, 78, 2142, 331, 1838, 275, 1202, 1507, 2097, 2096, 1686, 1808, 2049, 128, 826, 1402, 1902, 2200, 1167, 538, 67]. problems [931, 862, 1839, 540, 1952, 997, 998, 1067, 102, 1055, 1831, 1114, 987, 141, 959, 1711, 179, 1284, 2037, 355, 70, 682, 1996, 2174, 400, 1214, 1954, 2162, 1971, 571, 1105, 2116, 2203]. procedure [1706, 1339, 3, 300]. process [1015, 2070, 1204, 1135, 409]. processes [257, 516, 960, 2062, 2137, 1861, 921, 1768, 1620, 1247, 2009, 1704, 1751, 375, 781, 296, 1323, 5, 53, 2061]. processing [2097, 2096]. product [1906, 62, 312, 613, 2112, 801, 1407, 1792, 1282, 661, 207, 1070, 1430, 1320, 1199, 1675, 2157, 172, 1394, 1866, 397, 796, 1483]. Production [1479]. products [1886, 1915, 2005, 539, 687, 116, 1089, 1608, 1231, 701, 818, 1040, 1438]. profile [212]. profiles [797, 611, 1213]. projected [1472]. Projection [1888, 1311, 1722, 1875, 2108, 1876, 2184]. projections [491, 1134]. Projective [85, 1833, 685, 220, 615, 1345, 1724, 1723, 1998, 479]. projectors [1027]. Prolongation [224, 1206, 2185, 14, 244, 87]. prolongations [837, 1539, 2015]. Proof [2002, 2164, 602, 1918, 1439, 772, 1599, 603, 361, 1588]. propagating [1898, 321]. Propagation [1958, 756, 612, 1193, 202, 886, 733, 1014, 259]. propagator [1185, 1993, 1957, 450]. Propagators [1839, 1985, 630, 1546]. Propensities [2020]. proper [1009, 768, 1568, 1628, 433]. Properties [1175, 1713, 403, 364, 1051, 2182, 1121, 1049, 1050, 402, 292, 1483, 1906, 1984, 817, 1424, 97, 151, 1522, 315, 950, 1310, 1036, 727, 1321, 590, 160, 871, 1679, 1075, 2031, 1090, 1362, 1357, 1862, 1259, 2018, 959, 231, 648, 360, 1966, 1850, 2093, 2072, 2217, 1165, 716, 476, 1628, 69, 734, 447, 680, 2028, 1498, 2095, 1715, 302, 2173, 301, 397, 796]. property [1914, 1937, 1335, 2020, 1632, 1816, 148, 1501, 1834, 1358, 1494, 1697, 1613]. protons [1338]. proven [1306]. proving [1703]. Prüfer [2203]. pseudo [998, 768]. pseudo-Euclidean [768]. pseudo-Markovian [998]. pseudobidimensional [982]. pseudodifferential [1966]. pseudopotential [1945]. Pseudopotentials [981, 2015]. Pseudoscalar [1016]. pseudospin [2031]. Pure [2024, 851].

QCA [1997]. QCD [786, 1924]. QED [1691]. Quadratic [27, 264, 1579, 986, 1872, 889, 630, 104, 1483, 1222, 1498, 1104]. quadratics [1250]. quadrupole [91, 293, 294, 795]. quadrupole-phonon [294]. Quality [435, 1155]. quantal [806, 807, 1098]. quantities [1990, 223, 1639]. Quantization [1092, 1777, 652, 968, 1334, 1831, 1060, 1780, 1325, 1875, 313, 1503, 1321, 835, 916, 1139, 1548, 1959, 2006, 1842, 1058, 1039, 1403, 1309, 261, 1956, 753, 1170, 1029, 1876, 2184, 926, 974, 975, 550, 1601]. quantized [829, 830, 2026, 36]. Quantum [1116, 516, 1018, 323, 1056, 1171, 1703, 608, 1172, 368, 793, 1944, 1704, 2166, 221, 1415, 872, 1192, 495, 182, 296, 1904, 596, 409, 1104, 520, 1562, 96, 850, 1185, 1993, 131, 1057, 1016, 1916, 272, 446, 173, 865, 1183, 1631, 1955, 457, 1778, 1962, 132, 1964, 1547, 2102, 997, 998, 1781, 1844, 674, 232, 1563, 2122, 1688, 704, 1367, 1527, 1093, 492, 1861, 1590, 240, 404, 467, 529, 639, 1848, 549, 1636, 1544, 2020, 847, 359, 1251, 1768, 1332, 2089, 1961, 1662, 234, 1792, 148, 550, 908, 1409, 789, 396, 175, 882, 1620, 746]. quantum [1791, 782, 180, 273, 2139, 638, 1793, 138, 1795, 1047, 2165, 1779, 1363, 2073, 449, 499, 32, 1418, 832, 774, 1549, 1253, 2146, 1097, 2019, 1760, 1448, 649, 554, 67, 1182, 90, 518, 1782, 451]. Quantum-mechanical [872, 1016, 1662, 2019]. quantum-thermodynamic [1916]. quark [1310, 1366, 754]. quartic [1095]. Quasi [881, 2059]. Quasi-abelian [881]. quasi-Newtonian [2059]. Quasiclassical [1693]. Quasicontinual [609, 1075]. Quasigroup [724]. quasilinear [1618, 1619, 124]. quasiperiodic [1259]. quasistatic [1949]. Quasisteady [1426, 395]. quaternion [1749, 2215, 1568]. quaternionic [1683]. quaternions [758, 1163]. quotient [221].

R [624]. Rac [608]. Racah [1570, 470, 762, 1349, 55]. radial [214, 1255, 399, 1365]. Radiation [1928, 715, 1601, 1562, 1788, 1890, 2207, 1553, 1745, 2024]. Radiative [873, 369]. Radon [831, 1501]. raising [86, 474]. Raman [847]. ramified [834]. Random [449, 932, 390, 394, 1406, 2137, 1081, 1901, 129, 421, 705, 1150, 1891, 1657, 748, 511, 2211, 258, 612, 525, 321, 317, 548]. randomly [1341]. range [586, 1566, 219, 1512, 1156, 1849, 37, 1510, 1965, 972, 1966]. rank [727, 1965]. Ranks [1050]. rate [1738]. rates [454]. Rational [1331, 1516, 47, 1639]. Ray [218]. Rayleigh [1481]. rays [43, 382, 779]. reaction [391, 1928, 870, 562, 678]. reactions [290]. Real [1359, 1264, 1379, 439, 2035, 1487, 1039, 118, 977, 1571, 433]. reality [2104]. Realization [1657, 469, 111, 585, 1946]. Realizations [1811, 108, 308, 2109, 764, 2071]. receiving [1032]. Reciprocal [80, 1397]. Reciprocal-type [80]. reciprocity [639, 1378]. reconstruction [638]. recovery [2155]. rectangular [1948]. Recurrence [954, 560, 1645, 1411]. recursion [2161, 952, 1734, 982, 154, 2116, 1771]. Recursive [1342, 1943, 623]. Redish [562]. reduced [858, 1870, 581, 128, 826, 1402, 1902, 2200, 1460]. Reducible [1848, 1186]. Reduction [441, 494, 2149, 1799, 1430, 1320, 969, 1231, 434, 670, 115, 1534, 2118, 2006, 1818, 697, 696, 1059, 1923, 1980, 795]. reductions [925]. redundant [907]. reference [964]. refinement [821, 1160, 1312]. reflection [50, 1298]. reflectionless [1224]. reflectivity [1453]. refraction [1216]. refractive [756]. Regge [1428, 289, 1000, 2164, 1219, 239, 1078]. regimes [983]. region [2156, 31]. regions [669]. Regular [593, 545]. regularity [605, 877]. Regularization [257, 2046, 1794, 1339]. regularizations [24]. regularized [714]. regulator [714]. Reid [218]. reinterpretation [431]. rejection [1687]. related [8, 370, 313, 1321, 1406, 2058, 963, 389, 1442, 1048]. relating [682, 2034]. relation [1916, 1919, 643, 808, 543, 1416, 982, 354, 896, 1085, 1411, 669, 2095, 818]. relations [62, 613, 1432, 1567, 1681, 555, 2064, 1645, 954, 560, 2126, 1227, 733, 1019, 1771, 1455]. Relationship [993, 542, 1027, 822, 1809]. Relationships [365]. relatives [529, 639, 1848]. Relativistic [1125, 357, 235, 499, 32, 1806, 868, 1717, 192, 2050, 1577, 1606, 1482, 574, 575, 1458, 698, 775, 152, 807, 943, 174, 1985, 1179, 550, 789, 73, 401, 49, 231, 648, 1423, 1864, 852, 1190, 1769, 200, 722, 1418, 1152, 1867, 1476]. relativistically [1818, 707]. Relativity [186, 873, 1911, 1370, 58, 1910, 453, 993, 564, 563, 992, 2168, 594, 550, 937, 1596, 1217, 841, 1857, 632, 839, 967, 743, 710, 276, 838, 338, 909, 910, 1909, 275, 522, 1912, 1519, 844, 1746]. relaxation [1127]. relevance [682]. relevant [1187]. remaining [675]. remarkable [170, 317]. Remarks [1778, 1410, 2218, 1024, 1551, 2130, 2083, 951, 72, 461, 1720, 1997, 342, 343]. Renormalization [1304, 748]. Renormalized [1244, 262]. repeated [1864]. Representation [1490, 1271, 809, 1819, 447, 498, 752, 2134, 1254, 1422, 85, 853, 1037, 961, 1587, 220, 1763, 414, 169, 1075, 358, 754, 303, 1077, 1624, 471, 89, 261, 755, 1674, 1929, 1154, 1568, 1134, 1149, 989, 2000, 19, 583, 686, 1454, 1039]. Representations [720, 683, 702, 859, 471, 1319, 360, 500, 1308, 372, 136, 1281, 1038, 1432, 2153, 925, 949, 1839, 761, 685, 161, 1696, 269, 798, 1610, 1877, 1196, 1934, 1273, 1936, 539, 1740, 205, 653, 113, 1434, 1812, 1884, 687, 247, 1286, 1362, 717, 861, 1431, 1230, 1607, 1307, 1039, 790, 820, 808, 1282, 209, 760, 1345, 1940, 1725, 1109, 1762, 1810, 1133, 1110, 48, 206, 207, 1864, 658, 1674, 351, 1430, 1320, 1878, 1199, 1675, 1728, 208, 764, 800, 1999, 1351, 1628, 2073, 2111, 1801, 1086, 1231, 1393, 480, 1535, 926, 1683, 989, 1813, 2177, 1227, 1654, 117, 110]. representations [1611, 2132, 373, 1879]. representatives [1388]. reproducing [1742]. reptating [1932]. repulsive [1302]. reservoir [2211]. Resolution [2200]. resolvent [256]. Resonance [865, 103, 392, 2098]. Resonances [174, 330, 101, 286, 1215, 269, 798, 1504, 2055, 1407, 310, 1063, 2080, 2081]. resonant [350, 88, 621, 155, 69, 1019]. resonators [103]. respect [1511, 1395, 414, 549]. Response [1552, 945, 2, 158, 2181, 1903, 1997]. responses [864, 923, 924, 1603, 1604]. rest [605]. Restricted [1442, 435]. Restrictions [707]. result [963]. results [1031, 76, 2156, 1360, 1139, 1333, 266, 1751]. resummations [1638]. retarded [545]. reversal [470, 762, 1349]. reversibility [1586]. revisited [62, 613, 1911, 945, 932, 2181, 1903]. revolution [50]. Reynolds [721]. RHS [2055]. Riccati [927, 1441, 1600]. Ricci [839]. Riemann [1619, 967, 1070, 228, 955]. Riemannian [767, 1171, 333, 2025, 444, 729, 1589, 1749, 2215, 1329, 1011]. Riesenfeld [64]. Riesz [1474]. rigged [173, 865, 1590, 1504, 330]. right [497, 1099]. Rigid [14, 244, 1458]. Rigorous [1562, 1180, 2129]. ripplons [1112]. RKKY [297]. Robinson [2185]. role [96, 574, 575, 1956]. Root [1138]. Roots [202]. rosette [36]. rosettes [36]. Rotating [1459, 757, 1757, 73, 401, 1584, 1585, 849, 1130, 707, 1096, 95, 1747, 635]. rotation [454, 2022, 820, 1352, 1353]. rotational [357, 1937, 668, 942]. rotationally [135]. rotations [820, 1799, 768]. Ruelle [2099]. rule [1886, 2201, 2216, 1089, 1608]. rules [1087, 576, 1515, 120, 121, 305, 304, 1802, 855, 884, 1538, 55, 2004]. Runge [2121].

S [1886]. Sachs [2169]. Salingaros [1318]. Salpeter [1027, 1338]. sample [1122, 1135]. sample-wise [1135]. sandwich [603]. SAS [1914]. satisfied [1726]. satisfy [2178]. satisfying [214]. Sato [1551]. saturation [391, 1981]. Scalar [1977, 1152, 1594, 829, 830, 426, 2135, 2172, 367, 1119, 1301, 1103, 1699, 188, 1920, 1599, 899, 1940, 1518, 116, 511, 1929, 1154, 514, 489, 1162, 456, 1525]. scalar-tensor [1103]. scalars [91]. Scale [149, 1535]. scaling [1468]. scattered [1239]. Scattering [1894, 1547, 2102, 1367, 1081, 1268, 1269, 511, 1005, 1770, 1064, 1368, 79, 862, 65, 1277, 2180, 1967, 8, 370, 1846, 1492, 1254, 1419, 699, 576, 1515, 367, 1297, 1115, 1018, 631, 1514, 1311, 1721, 1722, 1511, 496, 1327, 1887, 125, 1962, 1707, 2129, 1908, 611, 1964, 993, 362, 600, 394, 1216, 1914, 1558, 1527, 2058, 1590, 905, 329, 1453, 37, 1343, 769, 452, 773, 1238, 935, 987, 141, 740, 1326, 201, 464, 1663, 1009, 588, 1951, 1212, 1668, 1054, 463, 852, 70, 1966, 1850, 268, 92, 537, 771, 1004, 990, 1592, 2130, 1996]. scattering [1605, 330, 679, 1344, 1298, 519, 400, 1214, 68, 1065, 1256, 2162, 2131, 872, 2142, 331, 1838, 274, 1153, 1382, 1743, 554, 1019, 1509, 1803, 1639]. Scheme [1294, 823]. schemes [1475, 1476, 82]. Schild [1069]. Schrödinger [894, 665, 752, 2167, 422, 1817, 417, 816, 1364, 1366, 767, 2117, 981, 1362, 327, 2057, 1783, 935, 399, 599, 1622, 2191, 2127, 355, 822, 1126, 1212, 1944, 1361, 737, 1845, 34, 553, 629, 1112, 962, 328, 1508, 519, 1885, 361, 1365, 1142, 1528, 2140, 135, 739, 901, 554, 1705]. Schrödinger-type [2191, 1528, 2140]. Schwarzian [1494]. Schwarzschild [52, 424, 1147, 1642, 1596, 914, 1479, 495, 9]. Schwarzschild-like [1147, 914]. Schwinger [1713, 1640, 2021, 899, 1023]. screened [1681, 2068, 329, 271]. screening [2068]. screw [757]. screw-pinch [757]. sea [1605]. Second [860, 1606, 1482, 1184, 1237, 923, 924, 1603, 1604, 1719, 936, 1433, 814, 1715]. second-order [1606, 1482, 1237, 1719, 936, 814, 1715]. section [533, 991, 1243]. sections [1692]. sector [754, 749]. Sectors [1802]. see [1741]. Segal [1139]. Selection [305, 304]. Self [520, 398, 2205, 2060, 567, 1466, 1467, 1069, 862, 1785, 1100, 1066, 1919, 1328, 12, 436, 1379, 1649, 2214, 269, 798, 1800, 194, 1076, 1664, 2144, 2145, 1427, 1633, 1279, 489, 1162, 677, 2154, 139, 1613]. self-adjoint [1633, 139]. Self-adjointness [520, 194, 2154]. self-conjugate [12, 436]. Self-consistent [398]. Self-dual [1466, 1467, 1069, 1066, 1919, 1328, 1379, 1800, 1076, 1664]. Self-gravitating [2205, 2060, 1427, 489, 1162]. Self-graviting [567]. self-similar [1100, 1649, 2214]. self-similarity [2144, 2145, 1279]. self-trapped [677]. Semiclassical [1255, 557, 266, 1278, 1682, 556, 746, 1771]. semidirect [1282]. semigroups [1104, 1861]. semimartingales [1135]. Semiregular [1725]. semisimple [763, 1883, 1626, 86, 805, 2108, 354, 926, 974, 975]. Semiunitary [685]. Sensitivity [863, 864, 2090, 1687]. separability [1220, 709, 124]. Separable [1796, 727, 444, 2037, 1803, 1639, 1473]. Separating [126]. Separation [1856, 1577, 1437, 472]. sequence [970]. series [2038, 2202, 956, 2188, 763, 1712, 1821, 1505, 1899, 719, 1699, 622, 2029, 1891, 1109, 349, 21, 2111, 974, 975, 951, 117, 90, 44, 818, 1438]. sesquidual [2158]. set [552, 1274, 91, 1688, 414, 134, 1080, 1326, 479, 932, 2152]. sets [623]. setting [2170]. seven [1118]. seven-parameter [1118]. seventh [952]. seventh-order [952]. several [252, 2132]. Shabat [862, 1785, 822, 1509]. shallow [1157]. Shape [2037, 547]. sharp [1936]. Sharpening [1779]. Shear [1852, 797, 1910, 779, 1159, 1746]. Shear-free [1852, 1910, 779, 1746]. shearfree [1856, 382]. shearing [913]. shell [600, 727, 1849, 1681, 2210, 71, 1639, 1473, 1455]. shells [2105, 1747]. Shift [293, 294, 302, 398, 1082, 351, 292, 301, 397, 795, 796, 1483, 1484, 1682, 802, 803, 2125, 1431, 936, 1341, 1433, 2124, 1433]. Shift-operator [293, 292]. shifts [1368, 631, 271, 1079, 1411]. shock [568, 722]. short [37, 1510, 1965]. short-range [37, 1965]. Shubnikov [304, 615, 1347]. side [1099]. sided [625, 68, 1135]. Siegel [1588]. Siewert [2044]. sigma [1667, 1984, 494, 730, 1600]. Signals [868]. signature [591, 1129, 1288]. Signatures [1999]. similar [1100, 1649, 2214, 497, 124]. Similarity [813, 919, 1091, 2144, 2145, 1279]. Simple [2102, 1618, 1619, 307, 962, 2148, 1871, 248, 1439, 367, 269, 798, 1913, 545, 1940, 1616, 1110, 1627, 2107, 319, 897, 1675, 118, 2036, 1764, 669, 1198, 1588, 1473, 1547]. simplification [1911, 2080]. Simplifications [403]. Simplified [263, 1460]. Simultaneity [2104, 276]. simultaneous [843, 1006]. simultaneously [1252]. Sine [958, 1943, 251, 1088, 529, 639, 970, 2052, 254, 1423, 70, 124, 1499, 944, 2016, 1290]. sine-Gordon [1943, 251, 1088, 529, 639, 970, 2052, 254, 1423, 70, 124, 1499, 944, 2016, 1290]. Single [1054, 398, 1974, 553, 756]. single- [553]. single-step [756]. singletons [608]. Singular [966, 953, 1731, 242, 22, 656, 766, 631, 2046, 320, 934, 1330, 24, 366, 846, 2100, 759]. singularities [1241, 657, 998, 1849, 519, 843, 1006]. sinusoidal [1605]. site [713]. Sitter [2134, 1714, 1709, 1674]. six [1907, 308]. six-dimensional [1907, 308]. size [1190]. skeletonizations [1589]. Slater [134]. slightly [1268, 1269, 511]. slope [533]. sloping [1157]. Slow [212, 350, 88, 445, 1138]. slow-fluctuation [350, 88, 445, 1138]. slowly [73, 401, 748]. small [748]. smooth [1982, 1718]. smoothly [329]. solid [2112, 890, 1316]. solitary [1817, 2084]. Soliton [1076, 1898, 1578, 1919, 365, 1738, 1680]. soliton-theoretic [365]. solitons [1422, 529, 639, 1848, 14, 244, 2053, 1944, 1112, 411]. soluble [1260, 286, 358, 359]. Solution [102, 935, 497, 255, 459, 1711, 1647, 811, 371, 2155, 2180, 1967, 1932, 1240, 217, 1486, 340, 1974, 785, 1459, 1863, 1118, 711, 847, 1101, 334, 1710, 1209, 914, 2017, 408, 2100, 279, 815, 1203, 679, 2174, 644, 420, 1820, 1900, 2171, 1493]. Solutions [13, 466, 672, 1118, 254, 1931, 1665, 1017, 1685, 1933, 1972, 2135, 2167, 190, 586, 1566, 1175, 1180, 1100, 1144, 956, 675, 1817, 1328, 229, 1119, 2067, 1340, 1147, 2196, 57, 1285, 365, 1220, 709, 1800, 1200, 1849, 1913, 871, 776, 1091, 1201, 751, 893, 1899, 1643, 2213, 1973, 1072, 1698, 1204, 2023, 333, 1753, 1357, 794, 1380, 1300, 1975, 1102, 1783, 1114, 834, 339, 813, 621, 1080, 1217, 1025, 1518, 787, 1264, 1400, 1680, 1302, 2127, 1133, 1076, 2084, 1291, 1664, 337, 2037, 1922, 706, 98, 750, 1731, 568, 962, 1372, 2144, 1798, 1344, 346, 253]. solutions [400, 1214, 1969, 2145, 476, 1819, 193, 425, 659, 1219, 250, 1021, 295, 1365, 345, 1597, 571, 1976, 2193, 919, 701, 595, 61, 270, 901, 887, 1158, 5, 53, 1949, 1580, 1290]. solvable [1894, 816, 441, 1364, 2062, 211, 177, 423, 1549, 2061, 1705]. Solving [1845, 1514, 1311, 1721, 1722, 2076, 442, 980]. Some [461, 1720, 817, 1045, 784, 2067, 76, 453, 590, 900, 1164, 1168, 1679, 2178, 1396, 1090, 1380, 1529, 2029, 1518, 231, 648, 1390, 1165, 338, 504, 1909, 2001, 1751, 680, 1997, 1498, 1438, 1727, 136, 656, 1100, 1031, 1778, 1456, 952, 1335, 1464, 1357, 1259, 1265, 790, 1773, 179, 689, 1787, 730, 2072, 2217, 476, 295, 1971, 1786, 1874, 1418, 1747, 1900, 887, 5, 53, 818, 1867, 1321, 1360, 1049]. Sommerfeld [1359, 179]. sound [1277, 11, 348]. Source [485, 1, 1033, 699, 1963, 1155, 408, 1968, 11, 348, 1954]. sources [367, 751, 893, 2207, 1068, 1860, 355, 1798]. Space [238, 2150, 382, 779, 1203, 488, 387, 226, 799, 136, 605, 96, 1911, 666, 291, 614, 604, 2172, 374, 1045, 272, 173, 865, 146, 424, 1752, 224, 2206, 58, 1056, 236, 237, 874, 304, 1946, 900, 1914, 561, 129, 1590, 1504, 966, 1410, 2218, 36, 1544, 188, 702, 1251, 364, 225, 1362, 1223, 1557, 1407, 2029, 1155, 1417, 550, 1856, 729, 1420, 1337, 1714, 1345, 1723, 1143, 462, 1589, 89, 787, 1264, 14, 244, 42, 2221, 1780, 1076, 43, 938, 100, 521, 1992, 1918, 742, 1734, 98, 3, 300, 344, 264, 651, 527, 1793, 145, 969, 208, 1650]. space [1042, 41, 693, 1084, 1948, 1467, 1885, 768, 1971, 1776, 32, 1796, 1134, 30, 2054, 1977, 481, 2157, 1457, 1371, 1600, 487, 495, 2198, 1978, 929, 1822, 427, 1760, 1904, 669, 2095, 2133, 2132, 2220]. space-independent [291, 614]. spacelike [98, 1519, 843, 1006]. spaces [130, 62, 312, 613, 821, 1160, 1228, 1066, 1577, 1735, 1962, 544, 36, 549, 807, 38, 39, 333, 1825, 627, 628, 444, 1940, 565, 389, 2136, 636, 638, 1749, 2215, 1329, 1674, 330, 40, 2195, 1099, 1134, 912, 701, 1879]. Spacetime [837, 1266, 566, 593, 1024, 1163, 965, 734, 845]. Spacetimes [1070, 840, 1709, 718, 843, 1006, 635]. sparse [1902]. Spatial [2137, 1148, 2085, 13, 466, 564, 745, 1133, 706, 187, 40, 41, 940, 253]. Spatially [299, 956, 454]. Special [1817, 1628, 1635, 2189, 58, 751, 893, 1396, 43, 1941, 1633, 1671, 1535, 1978, 1090]. Specializations [1766]. specification [485]. spectra [1562, 1794, 2098]. Spectral [8, 370, 1088, 757, 1966, 321, 862, 417, 816, 441, 2156, 997, 269, 798, 590, 871, 1831, 935, 388, 355, 1508, 1165, 1341, 2116]. spectroscopy [1017, 1647]. spectrum [520, 1477, 1563, 1296, 1815, 1757, 1361, 2186, 2072, 2217, 361]. speed [1277]. sphere [652, 1216, 1245, 917, 1596]. spheres [453, 914]. Spherical [48, 1912, 1869, 1492, 1132, 772, 465, 661, 846, 645, 2113, 1050, 1654]. Spherically [188, 1710, 2135, 1887, 1317, 1800, 1210, 1102, 464, 31]. Spin [529, 1793, 1833, 1692, 637, 101, 108, 1752, 2212, 232, 238, 492, 2106, 230, 404, 467, 867, 1272, 1173, 556, 972, 1690, 856, 857, 1864, 210, 1565, 1074, 411, 601, 533, 991, 1243, 1478, 33]. Spin- [1793, 1692, 637, 101, 1752, 238, 492, 1478]. spin-extensions [210]. spin-flip [533, 991, 1243]. spin-isospin [2106]. spin-variable [1173]. spinning [2050, 1516, 1995, 1634, 1152, 455]. Spinor [374, 1068, 315, 1877, 46, 968, 1334, 1831, 1386, 1387, 1465, 2017, 750, 535, 193]. spinorial [1824]. spinorization [2091]. Spinors [1463, 1714, 209, 1824, 380, 896, 1085, 387, 1545, 758]. spins [1254]. Splines [1514, 1311, 1721, 1722]. Splitting [1559, 2095]. spontaneous [580, 1028]. spontaneously [1419]. spreading [482]. spurious [403, 871, 402]. squared [1698, 2052, 375]. Stability [445, 50, 1293, 1480, 2099, 242, 22, 542, 728, 1758, 757, 2115, 2053, 489, 1162, 9, 52, 2200]. stable [212, 47, 1829]. standard [1631, 319]. star [1575, 2194, 680]. Stark [1846, 174, 1468]. Stark-like [1846]. stars [635]. State [807, 799, 666, 579, 1735, 643, 1161, 903, 1934, 1456, 36, 2020, 1289, 1462, 89, 904, 266, 479, 2082, 1644, 1525, 2203]. state-property [2020]. statements [1503]. states [973, 1693, 517, 1419, 173, 552, 1274, 1908, 643, 75, 220, 91, 293, 294, 302, 398, 1082, 1030, 1196, 1934, 2035, 1108, 1381, 1936, 74, 1823, 1590, 2055, 854, 1195, 1452, 1618, 247, 1362, 1010, 1213, 1793, 1523, 468, 990, 2165, 1404, 828, 69, 112, 447, 1847, 223, 2146, 681, 292, 301, 397, 795, 796, 1483, 1484, 353, 1454]. Static [1594, 426, 566, 2135, 2172, 1103, 1922, 604, 1952, 229, 1317, 486, 1102, 2170, 1217, 42, 231, 648, 427]. Stationary [1564, 1104, 365, 840, 336, 711, 1517, 278, 332, 745, 632, 965, 1372, 1457, 1858, 456, 923]. Statistical [490, 606, 2170, 2054, 931, 1310, 997, 998, 884, 1259, 1261, 1791, 138, 1874, 428, 429, 430, 1278]. Statistics [902, 133, 113, 847, 1223, 1557, 1523, 1948, 409]. steady [1523, 1525]. steady-state [1525]. steep [201, 464]. stellar [231, 648]. Stenschke [176]. step [398, 756]. Stern [33]. Stiefel [1265, 1787]. Stieltjes [1362]. stiff [766, 339]. Stochastic [835, 916, 282, 1981, 316, 1047, 1784, 1701, 1751, 1678, 799, 1015, 2090, 1862, 1400, 1620, 557, 319, 1095, 1262, 515, 886, 781, 296, 5, 53, 596, 1892]. Stokes [1044, 851, 1313, 1694]. Strange [706, 1536]. strata [977, 1392]. stratified [611, 899]. streets [50]. strength [1421, 1847]. String [957]. strings [1970, 1971, 1786, 2159, 1369, 1748]. Strong [1386, 1387, 1510, 1300, 2098]. strongly [2202, 558, 1335, 2093]. Structural [1695, 1811, 890, 1316, 2071]. Structure [1281, 1038, 2153, 1882, 1139, 1109, 938, 1211, 1666, 2147, 40, 41, 233, 461, 1720, 1598, 1495, 315, 983, 224, 1266, 2206, 704, 952, 905, 1544, 2020, 1464, 773, 1207, 1927, 1124, 1762, 1810, 651, 716, 584, 1457, 2039, 735, 996, 181, 1639]. structures [1541, 215, 837, 947, 1469, 593, 14, 244, 1392]. Study [1189, 579, 690, 24, 497, 194, 235, 1329, 1996, 69, 850]. Sturm [688]. subalgebras [1883, 1571]. subduction [1283]. subgroup [925, 161, 1108, 855, 1435]. Subgroups [472, 372, 684, 1194, 1164, 2150, 2176, 56, 1352, 1353]. submanifolds [1656]. submerged [1951]. submersion [2025]. submitted [375]. submodules [1938]. subrepresentations [2109]. subspace [1266]. substances [867]. substructure [592]. Subsystems [130, 141]. Sufficient [1491]. Sum [1886, 576, 120, 121, 1515, 1128, 884, 1089, 1608]. summability [484, 2192, 60]. summable [817]. Summation [2202, 37, 2038, 2005, 119]. Summational [1520]. summing [719, 1699]. sums [119, 1040]. Super [584, 1767, 969, 2154]. superalgebra [1110, 207, 208, 764, 2177]. superalgebraic [133]. superalgebras [1652, 949, 1188, 1532, 1230, 1536, 206, 764, 1061, 118, 1671, 1572, 2126]. superconductivity [579]. superconductor [432]. Superconformal [100]. Superfiber [1124]. superfield [2065]. superfields [1989]. superfluid [11, 348]. supergravity [2158, 1913, 776, 1414, 386]. supergroup [1087]. supergroups [197, 617, 720, 976, 1987, 2119]. superluminal [1189, 1425]. Supermanifolds [1026, 1767, 1827, 1828, 1826, 216, 504]. Superoperator [1546]. superposable [531]. superposed [1017]. Superposition [1441, 1753, 1538, 1112, 448, 250, 979, 2151, 2004]. Superselection [96, 1802]. superspace [314, 2094, 1414, 523, 524, 501, 681]. superspaces [1827, 1826]. supersymmetric [288, 1337, 2183, 70, 753, 1152]. supersymmetrical [882]. supersymmetries [532]. Supersymmetry [1988, 1111, 2094, 2065]. supertableaux [1609, 1935]. superunitary [1935]. supplementary [975]. supported [1481]. surface [1241, 1280, 2143, 1963, 1218, 1429, 660, 917, 1596, 543, 1951, 511, 1605, 1982, 1105]. surface-event [543]. surfaces [424, 1315, 50, 15, 51, 1113, 953, 872, 1064]. survival [1804]. suspensions [1401]. symbol [1811, 1350, 2071]. symbols [2104, 801]. Symmetric [1712, 1914, 185, 2135, 229, 1887, 103, 392, 1317, 1821, 1373, 1610, 874, 1800, 1200, 2106, 1744, 188, 333, 622, 1210, 2207, 1975, 2029, 2141, 1102, 1517, 1213, 1710, 2084, 180, 2017, 1731, 1372, 2195, 722, 456, 1600, 1394, 895, 616, 31]. symmetrical [676, 1136]. Symmetries [665, 309, 1576, 332, 298, 1372, 641, 1597, 260, 589, 1829, 1586, 640, 114, 1582, 1055, 952, 1471, 1689, 1470, 653, 981, 1024, 1163, 970, 986, 1542, 1625, 1146, 953, 27, 882, 695, 1992, 1414, 856, 857, 737, 1329, 59, 1391, 1350, 1841, 2024, 2001, 412, 135, 109, 245, 1835, 1760, 1475]. symmetrized [1128, 1675]. Symmetry [1606, 1482, 1567, 1818, 1719, 1377, 739, 901, 2095, 1727, 1092, 1777, 1281, 1038, 1419, 1144, 432, 405, 1756, 12, 436, 1946, 1703, 211, 1559, 1937, 1732, 2003, 198, 64, 50, 140, 2029, 1059, 1695, 1856, 808, 792, 841, 66, 567, 1521, 1665, 883, 470, 762, 1349, 473, 1028, 1909, 1134, 112, 1231, 379, 1507, 2208, 1560]. Symplectic [1736, 789, 1533, 948, 2035, 1540, 1008, 1571, 2082, 110]. synchronized [1469]. synchronous [843, 1006]. synthesis [435]. system [1260, 542, 766, 2118, 57, 606, 1593, 1563, 191, 529, 639, 1531, 1090, 1612, 1542, 1737, 1930, 556, 381, 1126, 1423, 1442, 468, 479, 2008, 481, 1644, 1694, 460]. system-theoretic [1090]. Systematic [176, 2080]. systems [2099, 1727, 1104, 1169, 1677, 350, 88, 1016, 311, 784, 1577, 1735, 540, 863, 864, 1137, 866, 1756, 12, 436, 997, 998, 232, 570, 2048, 1614, 1772, 2012, 1676, 1513, 54, 1140, 1758, 1055, 2058, 2106, 1937, 1139, 404, 467, 1848, 806, 2020, 1896, 2011, 1204, 1618, 1619, 928, 927, 1766, 1315, 1842, 1702, 773, 1207, 1632, 1612, 1816, 1001, 1662, 1530, 497, 986, 222, 732, 655, 196, 1891, 1250, 462, 1589, 2191, 1400, 283, 1773, 965, 922, 878, 180, 712, 856, 857, 282, 822, 1442, 29, 2014, 1523, 1540, 2088, 875, 106, 107]. systems [346, 347, 1834, 448, 250, 2166, 2073, 891, 1796, 823, 1528, 2140, 774, 2151, 1295, 153, 642, 1543, 2146, 1697, 1526, 1509, 259, 1586].

’t [410, 721, 1421]. Tabensky [1144]. table [1007]. tableau [539]. tachyon [1192, 1478]. tachyons [1869, 1479, 1804, 495]. tadpoles [1123]. Tani [358]. Taylor [1481, 1127]. Tchebycheff [1362]. technique [1138, 441, 2125, 1595, 1403, 1176, 597, 598, 2080, 2124]. Techniques [502, 503, 1888, 365, 993, 293, 294, 302, 398, 802, 803, 1082, 2108, 1351, 292, 301, 397, 795, 796, 1433, 1483, 1484]. temperature [1729, 2139, 1480, 2146, 2007]. temperatures [557]. Tensor [687, 946, 2071, 1906, 670, 426, 2135, 2172, 1656, 1606, 1482, 1103, 858, 1852, 1168, 943, 2205, 1051, 1536, 1191, 84, 1719, 207, 967, 1070, 1853, 743, 2195, 2113, 1231, 1669, 560, 1814, 1433, 1654, 2074, 498]. Tensorial [1570]. tensors [1292, 1051, 2182, 839, 941, 744, 1049, 1050, 1545, 226]. term [2021, 437, 35, 1771]. terms [1692, 1763, 1218, 1429, 820, 1239, 1996, 349, 21, 1154, 2054]. ternary [2184]. Test [178, 143, 1687, 1374, 1859]. tetrad [892]. Tetrads [1581, 1630]. th [1942, 2021, 318]. th-order [318]. their [1943, 761, 837, 1027, 952, 529, 639, 1848, 2178, 549, 1396, 2150, 808, 760, 1345, 1723, 27, 1133, 264, 682, 1491, 1341, 1384, 896, 1085, 2091, 739, 2128]. theorem [246, 605, 1169, 418, 985, 1168, 1556, 491, 1896, 1405, 851, 1893, 792, 1009, 1680, 528, 782, 1918, 1021, 1262, 1392, 60, 2095, 1588]. theorems [1636, 831, 2194, 1787, 1002]. theoretic [1919, 365, 1090, 477, 2074]. Theoretical [961, 721, 1423]. theories [2092, 1635, 752, 461, 1720, 2135, 1125, 784, 915, 1103, 1986, 1527, 1739, 1740, 1559, 1689, 1470, 1921, 1702, 1599, 1124, 148, 1337, 1336, 1569, 749, 1146, 2136, 2139, 706, 638, 1329, 1263, 228, 2093, 1399, 716, 1418, 2175, 595, 1686, 1715, 1867]. Theory [1094, 1807, 342, 343, 1641, 1222, 1635, 821, 1160, 1562, 666, 131, 1057, 1304, 426, 604, 2172, 406, 8, 370, 1846, 1044, 272, 1419, 183, 1240, 1180, 1524, 219, 699, 1183, 1631, 1215, 1018, 99, 1514, 1311, 1721, 1722, 763, 325, 1292, 863, 864, 1225, 1486, 1962, 945, 340, 76, 341, 1547, 2102, 1648, 2222, 1587, 1056, 362, 394, 2048, 1385, 1505, 1802, 1367, 1127, 1558, 608, 1088, 230, 404, 467, 609, 1172, 708, 285, 593, 1218, 1429, 174, 660, 1122, 884, 1332, 1754, 618, 1612, 2025, 303, 277, 368, 793, 1575, 867, 627, 628, 1792]. theory [339, 1554, 1860, 1710, 1437, 809, 1420, 1500, 209, 581, 760, 1345, 1346, 1347, 1723, 1998, 918, 740, 1663, 1009, 1177, 995, 1413, 2086, 459, 17, 7, 105, 1013, 1284, 882, 227, 1755, 898, 528, 782, 1918, 742, 1585, 273, 994, 408, 573, 880, 1716, 1845, 279, 1653, 187, 527, 1966, 1793, 1850, 1995, 1390, 1573, 2199, 1795, 330, 1779, 519, 1270, 870, 562, 1383, 1065, 216, 514, 888, 1279, 1244, 2082, 30, 1499, 774, 1096, 2193, 1549, 156, 157, 780, 2181, 1903, 2019, 1546, 1760, 1448, 548, 2087, 1182, 596, 1580, 1019, 1509, 458, 1867]. Theory [802, 1618, 2124]. Thermal [578, 746, 347, 1107, 2174]. thermodynamic [1916, 460]. Thermodynamics [460, 192, 54, 1302]. Theta [956]. thin [603]. thin-sandwich [603]. things [992]. third [1615, 433]. third- [1615]. third-order [433]. Thirring [2190, 1548]. Thomas [1312, 876, 1806]. Thompson [2221]. Thorpe [1288]. those [1087]. Three [852, 537, 1005, 990, 1771, 1894, 1354, 178, 1125, 1115, 2105, 2129, 13, 466, 1563, 2117, 2178, 359, 754, 177, 1051, 758, 621, 462, 787, 2127, 1780, 1133, 1872, 179, 1211, 904, 2098, 1442, 34, 894, 1968, 2072, 2217, 1408, 2131, 1049, 1050, 331, 991, 759, 1549, 1338, 135, 353, 818, 924, 1603, 268, 771, 1004]. three- [462]. three-body [178, 177, 1338]. Three-dimensional [852, 1894, 1354, 2105, 2129, 2117, 1051, 621, 787, 2127, 1780, 1872, 34, 894, 1968, 2131, 1049, 1050, 135]. three-level [1563, 2098]. three-particle [179, 353]. three-quark [754]. Three-term [1771]. three-twistor [1211]. three-wave [621, 1442, 924, 1603]. Tight [2068, 327]. Time [1512, 1115, 1147, 986, 1542, 1857, 396, 878, 1850, 1445, 849, 1130, 1446, 605, 96, 1911, 1677, 666, 604, 2172, 374, 1449, 272, 1735, 424, 1752, 1957, 2206, 58, 988, 1617, 664, 874, 698, 129, 421, 545, 2011, 1842, 732, 1891, 1143, 1622, 324, 66, 510, 1773, 1249, 922, 846, 1053, 742, 98, 2014, 344, 1968, 629, 527, 1966, 736, 470, 762, 1349, 1203, 1841, 1467, 2166, 1776, 2131, 258, 32, 1415, 2054, 481, 1371, 488, 495, 739, 901, 929, 1098, 1760, 669, 1586]. time- [2131]. Time-asymmetric [1857, 698]. Time-dependent [1115, 1147, 986, 1542, 849, 1130, 1677, 1449, 1735, 1957, 664, 421, 2011, 1842, 732, 1622, 324, 66, 510, 1773, 1249, 922, 846, 1053, 2014, 1968, 629, 1966, 736, 1841, 2166, 739, 901, 1098]. time-independent [988, 1617]. time-ordered [1415]. time-series [1891]. Timelike [1370, 939, 940, 95]. times [966, 188, 225, 1856, 729, 42, 43, 382, 779, 521, 1992, 1793, 145, 41, 1971, 1796, 1457, 226, 1978, 427]. Tjon [190, 996]. Toda [417, 1495, 2012, 1737, 1738, 2013]. Tomimatsu [1551]. tomography [1492]. tool [497, 1329]. tools [1985]. top [1267, 1136]. topics [900]. Topological [371, 1648, 2222, 749, 1029, 312]. topologically [2201, 2216]. topology [96, 465, 1148, 584, 791, 678]. tops [1136]. toroidal [1127]. torsion [2158, 858, 1686, 1630, 2087]. torsional [797]. torus [2092]. total [2201, 2216, 1851, 1021, 533]. Trace [1212, 46]. traceless [743]. Trajectories [836, 419, 1000, 1428, 1332, 239, 1078]. trajectory [1693]. trajectory-coherent [1693]. Transfer [1447, 369, 1314]. transform [1287, 816, 441, 1184, 313, 1321, 2058, 1010, 1684, 175, 388, 1212, 1941, 443, 649]. Transformation [1673, 846, 1439, 1424, 1943, 2189, 1271, 621, 16, 506, 179, 1953, 691, 815, 1824, 2013, 1397, 1838, 2203, 1822, 1494, 554, 1705]. transformations [1727, 1677, 374, 1475, 1094, 1026, 1868, 825, 840, 220, 1800, 320, 934, 46, 1443, 1168, 1201, 376, 1410, 2218, 1896, 1753, 1595, 184, 225, 150, 16, 342, 343, 581, 332, 1398, 2160, 2103, 822, 59, 682, 587, 124, 252, 768, 80, 1646, 1535, 1660, 431, 501]. transforms [111, 1027, 541, 407, 117]. Transient [1661]. transients [1167]. Transition [324, 1156, 664, 4, 204, 1416, 876, 319, 233]. transitions [847, 890, 1316, 1695]. transitive [1856]. transitivity [1220, 709, 144]. translation [414]. translational [1570, 852, 942]. Transmission [1593, 2081, 1238, 2115, 1593]. Transmutation [1706]. transonic [1718]. Transparent [1269]. transport [1116, 183, 1240, 1180, 1524, 1225, 1917, 129, 1863, 82, 366, 572, 673, 459, 1022, 1711, 1284, 463, 1480, 253, 1383, 571, 2193]. Transverse [155, 743]. transverse-traceless [743]. trap [258]. trapped [677]. Trautman [2185]. Treatment [656, 1062, 1840, 2069, 556, 1258, 832, 686, 2074]. tree [280, 1623]. triality [306]. triangle [2178, 1123, 800]. Triangular [2212, 1801]. triaxial [1136]. tricritical [1315]. tridiagonal [538]. trilinear [1853]. trimers [1223]. triple [2112]. triviality [1599]. tropospheric [202, 756]. turbulence [721, 888, 1279, 1313]. turbulent [1606, 1482]. Twenty [2132]. Twenty-four [2132]. Twiss [1745]. twist [779]. twist-free [779]. Twisted [2092, 1292]. twistfree [382]. twisting [913]. Twistor [1120, 1755, 938, 1211, 100]. twistors [930, 94]. Two [192, 1908, 943, 2011, 2194, 1239, 625, 68, 1525, 1667, 2155, 586, 1566, 390, 183, 1515, 728, 1926, 1962, 340, 1707, 1963, 75, 1640, 984, 985, 1513, 698, 126, 4, 204, 285, 359, 1985, 1052, 1289, 1506, 234, 1783, 1208, 1856, 207, 2014, 2056, 553, 815, 1112, 962, 1929, 1621, 2224, 1170, 2013, 1418, 78, 823, 1820, 1278, 1600, 71, 182, 317, 1460, 1136, 1135, 1463, 2095, 1639, 1114, 141, 759]. two-body [698, 1985, 71, 1639, 141]. two-channel [1513, 1783, 553]. Two-dimensional [1908, 2011, 1239, 1525, 1667, 2155, 728, 1926, 1963, 285, 234, 2014, 815, 1112, 962, 1929, 1621, 2224, 2013, 823, 1820, 1278, 182, 317, 1460, 1114]. two-group [183]. two-parameter [1856, 1600]. two-particle [1515, 340, 2095]. two-phase [390]. two-photon [4]. two-point [1418]. Two-sided [625, 68, 1135]. two-state [1289]. two-way [586, 1566, 126].

Type [1282, 1283, 122, 168, 248, 1629, 1524, 915, 1301, 1220, 709, 1974, 304, 913, 1118, 1643, 2213, 1973, 484, 905, 1230, 1177, 2191, 381, 34, 894, 228, 1805, 2199, 535, 118, 80, 1597, 823, 1528, 2140, 2101, 1483, 811, 1509]. type- [1597]. Type-adapted [1282, 1283]. type-I [2101].

UIR [116]. UIRs [974, 975]. unbounded [590, 2077]. uncertainty [1537, 447]. uncorrelated [1502]. unequal [1166, 2012]. unidimensional [1732]. Unidirectional [259]. Unification [1595, 1750, 2110]. Unified [1717, 237, 238, 1755, 915, 1266, 1201, 1884, 760, 1345, 1346, 1347, 1723, 1998, 940, 562, 686]. uniform [1596, 852, 1445]. unique [2155, 1864]. unique-mass-spin [1864]. uniqueness [1456, 1191, 401, 1305, 880]. unit [654, 84]. unitarily [1933]. Unitarity [386, 125, 497, 533, 991, 1243, 818, 1455]. Unitary [1846, 1674, 1533, 925, 205, 389, 1430, 1320, 1086]. Universal [655, 1229]. universality [2206]. universe [1181, 1301, 1159, 1192, 1425]. universes [457, 454, 492, 144, 723, 849, 1130, 95]. unmagnetized [1606, 1482, 848, 1034]. unnatural [1452]. Unrenormalized [1023]. Unsolved [430]. Upper [1355, 362, 1141, 533, 496, 1035, 1361]. use [1570, 1324, 1055, 1845, 1653, 428, 429, 430]. useful [1655]. using [1676, 1595, 435, 839, 1966].

V [302, 1723]. vacua [96, 1148]. Vacuum [911, 671, 461, 1720, 229, 1220, 1974, 1101, 1517, 632, 1372, 718, 260]. vaguely [1966]. Validity [747, 2064]. valley [1344]. Valleys [1011]. Value [1636, 540, 102, 1114, 1284, 994, 279, 1555, 78, 630, 1167]. valued [1117, 226]. values [862, 1785, 71]. vanishing [188]. Variable [600, 640, 1216, 943, 1173, 124, 428, 429, 430]. Variable-phase [600]. variables [242, 22, 656, 446, 1577, 126, 1856, 1668, 472, 252, 1404]. variation [692, 623]. Variational [1952, 1041, 1729, 378, 1063, 1833, 1843, 2043, 183, 667, 1774, 1805, 310, 331, 668, 2154, 780]. variations [1242, 731, 1892]. various [1112]. varying [1898, 253]. Vector [1137, 130, 2112, 2121, 418, 1915, 1825, 1542, 1724, 1070, 623, 1042, 1865, 1388, 1730, 944, 663]. vectorlike [783]. vectors [865, 1066, 1688, 1936, 185, 1938, 634, 912, 433, 1978]. Velocity [1963, 1898, 452, 48, 448]. velocity-dependent [448]. Verification [1335]. Verma [1938]. Versal [1488]. versus [85, 265, 1961]. Vertex [140, 1821, 1744, 622, 1123]. vertical [1951]. Vessiot [618]. VI [1998, 397]. via [291, 614, 544, 667, 16, 209, 581, 749, 2160, 1170, 2203]. vibration [797]. vibrations [792]. view [1207, 1668, 1784, 2059]. viewpoint [1088, 898, 2199]. VII [398, 332]. VIII [795]. virial [1168, 1415, 44]. virtual [199, 1590, 2055]. viscosity [1873]. Viscous [778, 1275, 1672, 1002, 844]. vision [1031]. visits [932]. Vlasov [1719]. volume [558, 1795]. Vortex [1715, 1531]. vortices [50, 1305, 770]. vorticity [1159]. Vries [1422, 956, 2187, 2188, 2189, 958, 970, 1113, 1291]. vs [639].

Waals [972, 1565, 1365]. Waerden [1197]. Wagoner [780]. Wahlquist [1595]. walks [1150, 449, 932, 525, 317]. warm [848, 1034]. water [1157]. Watson [443, 60]. Wave [156, 157, 1869, 1972, 2187, 2188, 1186, 1187, 1511, 574, 575, 950, 1887, 1564, 2196, 1285, 13, 466, 988, 1617, 672, 363, 214, 1385, 1913, 1739, 1740, 152, 2106, 772, 484, 407, 545, 482, 37, 189, 1619, 1927, 1386, 1387, 1928, 621, 987, 1025, 848, 1034, 1510, 1965, 297, 2084, 1951, 1053, 1054, 235, 1864, 1193, 2027, 1442, 511, 1929, 202, 1865, 1365, 1837, 1838, 1017, 2080, 1253, 733, 128, 826, 1402, 1902, 2200, 1771, 259, 1472, 924, 1603]. wavefunction [1338]. wavefunctions [833, 137, 266, 1142, 907]. waveguide [2051, 886]. wavenumber [836]. Waves [545, 1869, 1157, 2189, 1817, 1436, 1081, 1128, 1901, 772, 421, 1618, 1619, 342, 343, 201, 464, 852, 568, 200, 1427, 791, 722, 547, 156, 157, 81, 321, 1454, 1455]. way [586, 1566, 126]. Weak [1404, 1949, 1330, 2197, 1122, 1745, 2100, 1303, 200, 1491]. weak-Painlevé [2197]. weakening [148]. Weakly [156, 157, 1184, 1213, 2211]. Weight [1388, 2010]. weighted [884]. weights [1742]. Weingarten [2091]. Weinhold [1646]. well [147, 1176, 1705]. Weyl [670, 1852, 1503, 700, 1360, 2205, 1797, 1217, 1922, 750, 744, 2054, 1867]. which [1629, 1646]. whole [2203]. wholespace [1014]. whose [1070]. Wick [493, 2138]. width [1341]. Wien [2100]. Wiener [656, 1205, 1322, 765]. Wightman [1868, 148]. Wigner [246, 801, 1229, 805, 890, 1316, 1396, 307, 1690, 1170, 641, 1446, 1295, 1560]. wild [1930]. wind [280]. wind-tree [280]. wise [1135]. Wit [410]. within [669]. without [751, 893, 841, 3, 300, 1801, 1982, 1686]. Witten [1021]. WKB [311, 1044, 1505, 1176, 450]. WKBJ [437, 35, 908]. Wobbling [1499]. work [199, 1816]. world [146]. Wu [190, 1322, 1380, 996].

X [1082]. Xanthopoulos [455]. XI [1483]. XII [1484].

Yakubovskiı [1175]. Yakubovskiı-Faddeev [1175]. Yang [333, 1665, 2147, 99, 151, 675, 76, 1340, 1266, 1800, 785, 751, 893, 747, 1380, 1529, 1421, 1148, 1420, 749, 787, 1264, 1923, 1922, 59, 70, 1666, 531, 1798, 1466, 1467, 2159, 1866, 2028]. Yevick [1156]. Young [539, 1532]. Young-tableau [539]. Yukawa [558, 1316, 1411, 270].

Zak [1010]. Zakharov [862, 1785, 993, 822, 1509]. Zero [195, 1933, 605, 1852, 1873, 619, 530, 1272, 809, 1518, 1616, 1070, 718]. zero-curvature [1616]. Zeromass [195]. zero-modes [718]. zero-rest-mass [605]. zeros [2044, 120, 121, 1811, 1239, 1996, 1666, 2147, 2071]. zeroth [810, 1141]. zeta [1794]. zone

References

Bleistein:1977:NIS

[1] Norman Bleistein and Jack K. Cohen. Nonuniqueness in the inverse source problem in acoustics and electromagnetics. Journal of Mathematical Physics, 18(2):194–201, February 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v18/i2/p194_s1. See erratum [1033].

Gouedard:1978:DEG

[2] C. Gouedard and C. Deutsch. Dense electron-gas response at any degeneracy. Journal of Mathematical Physics, 19(1):32–38, January 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/i1/p32_s1.

See erratum [158].

Mizrahi:1978:PSP

[3] M. M. Mizrahi. Phase space path integrals, without limiting procedure. Journal of Mathematical Physics, 19(1):298–307, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [300].

Gazeau:1978:FEC

[4] J. P. Gazeau. Four Euclidean conformal group in atomic calculations: Exact analytical expressions for the bound-bound two-photon transition matrix elements in the H atom. Journal of Mathematical Physics, 19(5): 1041–1048, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v19/i5/p1041_s1. See erratum [204].

Wodkiewicz:1979:ESS

[5] K. Wódkiewicz. Exact solutions of some multiplicative stochastic processes. Journal of Mathematical Physics, 20(1):45–48, January 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [53].

Floyd:1979:HOM

[6] Edward R. Floyd. Higher order modified potentials for the effective phase integral approximation. Journal of Mathematical Physics, 20(1):83–85, January 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [624].

Lewis:1979:LAI

[7] H. Ralph Lewis and Keith R. Symon. Linearized analysis of inhomogeneous plasma equilibria: General theory. Journal of Mathematical Physics, 20(3):413–436, March 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v20/i3/p413_s1. See errata [17, 105].

Ben-Artzi:1979:SST

[8] Matania Ben-Artzi and Allen Devinatz. Spectral and scattering theory for the adiabatic oscillator and related potentials. Journal of Mathematical Physics, 20(4):594–607, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [370].

Wald:1979:NSS

[9] Robert M. Wald. Note on the stability of the Schwarzschild metric. Journal of Mathematical Physics, 20(6):1056–1058, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [52].

Abellanas:1979:CDN

[10] L. Abellanas and A. Galindo. Conserved densities for nonlinear evolution equations. I. Even order case. Journal of Mathematical Physics, 20(6): 1239–1243, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v20/i6/p1239_s1. See erratum [203].

Murray:1979:SFD

[11] J. C. Murray. On the sound field due to a moving source in a superfluid. Journal of Mathematical Physics, 20(7):1409–1412, July 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/i7/p1409_

s1. See erratum [348].

CastilhoAlcaras:1979:CSC

[12] J. A. Castilho Alcarás and L. C. Biedenharn. On the coupling of selfconjugate systems with \overline{SL(3, R)} symmetry. Journal of Mathematical Physics, 20(8):1615–1621, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [436].

Cornille:1979:SNW

[13] H. Cornille. Solutions of the nonlinear 3-wave equations in three spatial dimensions. Journal of Mathematical Physics, 20(8):1653–1666, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [466].

Lakshmanan:1979:RBM

[14] M. Lakshmanan. Rigid body motions, space curves, prolongation structures, fiber bundles, and solitons. Journal of Mathematical Physics, 20 (8):1667–1672, August 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [244].

Horndeski:1979:CSC

[15] Gregory Walter Horndeski. Characteristic surfaces and characteristic initial data for the generalized Einstein–Maxwell field equations. Journal of Mathematical Physics, 20(8):1745–1751, August 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [51].

Kim:1979:NMMa

[16] Shoon K. Kim. A new method of matrix transformation. I. Matrix diagonalizations via involutional transformations. Journal of Mathematical Physics, 20(11):2153–2158, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [581].

Lewis:1979:ELA

[17] H. Ralph Lewis and Keith R. Symon. Erratum: “Linearized analysis of inhomogeneous plasma equilibria: general theory”, [J. Math. Phys. 20 (1979), no. 3, 413–436]. Journal of Mathematical Physics, 20(11):2372, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [7].

Cho:1979:CGF

[18] Y. M. Cho. Classification of gauge fields. Journal of Mathematical Physics, 20(12):2605–2612, December 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v20/i12/p2605_s1. See erratum [243].

Wong:1980:EER

[19] M. K. F. Wong and Hsin Yang Yeh. Explicit evaluation of the representation functions of ISO(n). Journal of Mathematical Physics, 21(1): 1–5, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i1/p1_s1.

Kasperkovitz:1980:AAM

[20] P. Kasperkovitz. Asymptotic approximations for modified Bessel functions. Journal of Mathematical Physics, 21(1):6–13, January 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p6_s1.

Pavelle:1980:PIC

[21] Richard Pavelle. The Planck integral cannot be evaluated in terms of a finite series of elementary functions. Journal of Mathematical Physics, 21 (1):14, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i1/p14_s1. See erratum [349].

Arms:1980:LSG

[22] Judith M. Arms. Linearization stability and a globally singular change of variables. Journal of Mathematical Physics, 21(1):15–18, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v21/i1/p15_s1. See erratum [242].

Chen:1980:CLB

[23] H. H. Chen and D. J. Kaup. Conservation laws of the Benjamin–Ono equation. Journal of Mathematical Physics, 21(1):19–20, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/
p19_s1.

Gravel:1980:ARA

[24] P. Gravel. Asymptotic regularizations as an alternative to distributions for the study of singular hypersurfaces. Journal of Mathematical Physics, 21(1):21–23, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p21_s1.

Weber:1980:NC

[25] T. A. Weber and C. L. Hammer. A note on completeness. Journal of Mathematical Physics, 21(1):24–27, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p24_s1.

Jarem:1980:CDO

[26] John M. Jarem. Construction of doubly orthogonal functions. Journal of Mathematical Physics, 21(1):28–31, January 1980. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p28_s1.

Leach:1980:QHF

[27] P. G. L. Leach. Quadratic Hamiltonians: The four classes of quadratic invariants, their interrelations and symmetries. Journal of Mathematical Physics, 21(1):32–37, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p32_s1.

Leach:1980:NHH

[28] P. G. L. Leach. A note on the Hénon–Heiles problem. Journal of Mathematical Physics, 21(1):38–43, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i1/p38_s1.

Mielnik:1980:MNS

[29] Bogdan Mielnik. Mobility of nonlinear systems. Journal of Mathematical Physics, 21(1):44–54, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p44_s1.

Samuel:1980:PSA

[30] J. Samuel. A phase space approach to generalized Hamilton–Jacobi theory. Journal of Mathematical Physics, 21(1):55–59, January 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p55_s1.

Steiger:1980:EMF

[31] Arno D. Steiger. On electromagnetic multipole fields in a finite, spherically symmetric region. Journal of Mathematical Physics, 21(1):60– 70, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i1/p60_s1.

Roy:1980:RBM

[32] Sisir Roy. Relativistic Brownian motion and the space–time approach to quantum mechanics. Journal of Mathematical Physics, 21(1):71–76, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i1/p71_s1.

Swift:1980:GSG

[33] Arthur R. Swift and Ron Wright. Generalized Stern–Gerlach experiments and the observability of arbitrary spin operators. Journal of Mathematical Physics, 21(1):77–82, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i1/p77_s1.

Moses:1980:KGL

[34] H. E. Moses. A kernel of Gel’fand–Levitan type for the three-dimensional Schrödinger equation. Journal of Mathematical Physics, 21(1):83–89, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i1/p83_s1. See erratum [894].

Kesarwani:1980:FTW

[35] R. N. Kesarwani and Y. P. Varshni. Five-term WKBJ approximation. Journal of Mathematical Physics, 21(1):90–92, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p90_s1. See

erratum [437].

Gessner:1980:RRH

[36] W. Gessner and V. Ernst. The rosette of rosettes of Hilbert spaces in the indefinite metric state space of the quantized Maxwell field. Journal of Mathematical Physics, 21(1):93–111, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p93_s1.

Grinstein:1980:SPW

[37] Fernando F. Grinstein. Summation of partial wave expansions in the scattering by short-range potentials. Journal of Mathematical Physics, 21(1):112–119, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p112_s1.

Halford:1980:ESHa

[38] W. D. Halford and R. P. Kerr. Einstein spaces and homothetic motions. I. Journal of Mathematical Physics, 21(1):120–128, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p120_s1.

Halford:1980:ESHb

[39] W. D. Halford. Einstein spaces and homothetic motions. II. Journal of Mathematical Physics, 21(1):129–134, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p129_s1.

Persides:1980:SGFa

[40] S. Persides. Structure of the gravitational field at spatial infinity. I. asymptotically Euclidean spaces. Journal of Mathematical Physics, 21(1):135–141, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p135_s1.

Persides:1980:SGFb

[41] S. Persides. Structure of the gravitational field at spatial infinity. II. asymptotically Minkowskian space–times. Journal of Mathematical Physics, 21(1):142–151, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i1/p142_s1.

Lathrop:1980:DCS

[42] J. D. Lathrop and M. S. Orsene. Dust cylinders in static space–times. Journal of Mathematical Physics, 21(1):152–153, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p152_s1.

Ludwig:1980:ASS

[43] Garry Ludwig. On algebraically special space–times with nontwisting rays. Journal of Mathematical Physics, 21(1):154–160, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/
p154_s1.

Ziff:1980:VSI

[44] Robert M. Ziff and John M. Kincaid. The virial series of the ideal Bose gas. Journal of Mathematical Physics, 21(1):161–165, January 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i1/p161_
s1.

Tchrakian:1980:DIM

[45] D. H. Tchrakian. N-dimensional instantons and monopoles. Journal of Mathematical Physics, 21(1):166–169, January 1980. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p166_s1.

Eeg:1980:CDS

[46] J. O. Eeg. Calculations of Dirac-spinor amplitudes by means of complex Lorentz-transformations and trace calculus. Journal of Mathematical Physics, 21(1):170–174, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i1/p170_s1.

Stefanescu:1980:SAC

[47] I. Sabba Stefanescu. On the stable analytic continuation with rational functions. Journal of Mathematical Physics, 21(1):175–188, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i1/p175_s1.

Lim:1980:SIR

[48] Chee-Seng Lim. Spherical integral representations for finite velocity conduction of an abruptly activated current through a crystal. Journal of Mathematical Physics, 21(1):189–194, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i1/p189_s1.

Lim:1980:CRP

[49] Chee-Seng Lim. Certain relativistic phenomena in crystal optics. Journal of Mathematical Physics, 21(1):195–210, January 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i1/p195_s1.

Hally:1980:SSV

[50] David Hally. Stability of streets of vortices on surfaces of revolution with a reflection symmetry. Journal of Mathematical Physics, 21(1):211– 217, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i1/p211_s1.

Horndeski:1980:ECS

[51] Gregory Walter Horndeski. Erratum: “Characteristic surfaces and characteristic initial data for the generalized Einstein–Maxwell field equations”, [J. Math. Phys. 20, 1745 (1979)]. Journal of Mathematical Physics, 21(1):218, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i1/p218_s3. See [15].

Wald:1980:ENS

[52] Robert M. Wald. Erratum: “Note on the stability of the Schwarzschild metric”, [J. Math. Phys. 20 (1979), no. 6, 1056–1058]. Journal of Mathematical Physics, 21(1):218, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i1/p218_s2. See [9].

Wodkiewicz:1980:EES

[53] K. Wódkiewicz. Erratum: “Exact solutions of some multiplicative stochastic processes”, [J. Math. Phys. 20 (1979), no. 1, 45–48]. Journal of Mathematical Physics, 21(1):218, January 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i1/p218_s1. See [5].

Fannes:1980:CDE

[54] M. Fannes and F. Rocca. A class of dissipative evolutions with applications in thermodynamics of fermion systems. Journal of Mathematical Physics, 21(2):221–226, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i2/p221_s1.

Sullivan:1980:GBC

[55] John J. Sullivan. Generalized back coupling rules for the Racah algebra of Gln??. Journal of Mathematical Physics, 21(2):227–233, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p227_s1.

Patera:1980:FSG

[56] J. Patera, Y. Saint-Aubin, and H. Zassenhaus. Finite subgroups of the generalized Lorentz groups O(p, q). Journal of Mathematical Physics, 21(2):234–239, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i2/p234_s1.

Collins:1980:ASN

[57] C. B. Collins. All solutions to a nonlinear system of complex potential equations. Journal of Mathematical Physics, 21(2):240–248, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p240_s1.

Collins:1980:CPE

[58] C. B. Collins. Complex potential equations, special relativity, and complexified Minkowski space–time. Journal of Mathematical Physics, 21(2):249–255, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i2/p249_s1.

Morris:1980:BTS

[59] Hedley C. Morris. Bäcklund transformations and the symmetries of the Yang equations. Journal of Mathematical Physics, 21(2):256–260, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p256_s1.

Sokal:1980:IWT

[60] Alan D. Sokal. An improvement of Watson’s theorem on Borel summability. Journal of Mathematical Physics, 21(2):261–263, February 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i2/p261_
s1.

Trlifaj:1980:JSG

[61] L. Trlifaj. The Jost solutions for general Gaussian potentials. Journal of Mathematical Physics, 21(2):264–268, February 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i2/p264_s1.

Antoine:1980:PIPa

[62] J.-P. Antoine. Partial inner product spaces. III. Compatibility relations revisited. Journal of Mathematical Physics, 21(2):268–279, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v21/i2/p268_s1. See erratum [613].

Martin:1980:MAP

[63] Pablo Martín, Guillermo Donoso, and Jorge Zamudio-Cristi. A modified asymptotic Padé method. application to multipole approximation for the plasma dispersion function Z. Journal of Mathematical Physics, 21(2):280–285, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i2/p280_s1.

Gunther:1980:DSB

[64] Neil J. Günther. Dynamical symmetry breaking in the Lewis–Riesenfeld oscillator. Journal of Mathematical Physics, 21(2):286–289, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p286_s1.

Asvestas:1980:POM

[65] John S. Asvestas. The physical optics method in electromagnetic scattering. Journal of Mathematical Physics, 21(2):290–299, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i2/
p290_s1.

Leach:1980:CSG

[66] P. G. L. Leach. The complete symmetry group of the one-dimensional time-dependent harmonic oscillator. Journal of Mathematical Physics, 21(2):300–304, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i2/p300_s1.

Wreszinski:1980:IPN

[67] W. F. Wreszinski. On the infrared problem in nonrelativistic quantum electrodynamics. Journal of Mathematical Physics, 21(2):305–307, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i2/p305_s1.

Ramm:1980:TSE

[68] A. G. Ramm. Two-sided estimates of the scattering amplitude for low energy. Journal of Mathematical Physics, 21(2):308–310, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i2/
p308_s1.

Romo:1980:SCP

[69] W. J. Romo. A study of the completeness properties of resonant states. Journal of Mathematical Physics, 21(2):311–326, February 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i2/p311_
s1.

Morris:1980:ISP

[70] H. C. Morris. Inverse scattering problems in higher dimensions: Yang– Mills fields and the supersymmetric sine-Gordon equation. Journal of Mathematical Physics, 21(2):327–333, February 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i2/p327_s1.

Sramek:1980:NCS

[71] Stanley J. Sramek. Numerical computation of off-shell two-body Coulomb amplitude values. Journal of Mathematical Physics, 21(2):334– 342, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i2/p334_s1.

Xin:1980:RCC

[72] Y. L. Xin. Remarks on characteristic classes of four-dimensional Einstein manifolds. Journal of Mathematical Physics, 21(2):343–346, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p343_s1.

Kunzle:1980:ESR

[73] H. P. Künzle and J. R. Savage. Equilibrium of slowly rotating relativistic fluids. Journal of Mathematical Physics, 21(2):347–354, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p347_s1.

Fannes:1980:ESM

[74] M. Fannes, H. Spohn, and A. Verbeure. Equilibrium states for mean field models. Journal of Mathematical Physics, 21(2):355–358, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p355_s1.

Critchley:1980:AES

[75] Robert Critchley. Approximate equilibrium states for two models of an interacting boson gas. Journal of Mathematical Physics, 21(2):359– 363, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i2/p359_s1.

Chodos:1980:GGC

[76] Alan Chodos and Vincent Moncrief. Geometrical gauge conditions in Yang–Mills theory: Some nonexistence results. Journal of Mathematical Physics, 21(2):364–371, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i2/p364_s1.

Dixon:1980:DPM

[77] Robert M. Dixon and Edward F. Redish. The dominant partition method. Journal of Mathematical Physics, 21(2):372–377, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i2/p372_s1.

Rzazewski:1980:IVP

[78] K. Rzazewski and W. Zakowicz. Initial value problem for two oscillators interacting with electromagnetic field. Journal of Mathematical Physics, 21(2):378–388, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i2/p378_s1.

Alvarez-Estrada:1980:ESI

[79] R. F. Alvarez-Estrada and M. L. Calvo. Electromagnetic scattering by an infinite inhomogeneous dielectric cylinder: New Green’s function and integral equations. Journal of Mathematical Physics, 21(2):389– 394, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i2/p389_s1.

Rogers:1980:RTT

[80] C. Rogers, J. G. Kingston, and W. F. Shadwick. Reciprocal-type transformations in magnetogasdynamics. Journal of Mathematical Physics, 21(2):395–397, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i2/p395_s1.

Wang:1980:NOL

[81] P. K. C. Wang. Nonperiodic oscillations of Langmuir waves. Journal of Mathematical Physics, 21(2):398–407, February 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i2/p398_s1.

Kelley:1980:CIS

[82] C. T. Kelley. A comparison of iteration schemes for Chandrasekhar Hequations in multigroup neutron transport. Journal of Mathematical Physics, 21(2):408–409, February 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i2/p408_s1.

Itzykson:1980:PAI

[83] C. Itzykson and J. B. Zuber. The planar approximation. II. Journal of Mathematical Physics, 21(3):411–421, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i3/p411_s1.

Kibler:1980:UT

[84] Maurice Kibler and Geneviève Grenet. On the SU_{2} unit tensor. Journal of Mathematical Physics, 21(3):422–439, March 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i3/p422_s1.

Carinena:1980:PCG

[85] José F. Cariñena and Mariano Santander. Projective covering group versus representation groups. Journal of Mathematical Physics, 21(3): 440–443, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i3/p440_s1.

Gould:1980:IAS

[86] M. D. Gould. On an infinitesimal approach to semisimple Lie groups and raising and lowering operators of O(n) and U(n). Journal of Mathematical Physics, 21(3):444–453, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i3/p444_s1.

Shadwick:1980:KPA

[87] W. F. Shadwick. The KdV prolongation algebra. Journal of Mathematical Physics, 21(3):454–461, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i3/p454_s1.

Augusteijn:1980:INR

[88] Marijke F. Augusteijn and Ernst Breitenberger. Integration of nearresonant systems in slow-fluctuation approximation. Journal of Mathematical Physics, 21(3):462–471, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i3/p462_s1. See erratum [350].

Kuratsuji:1980:PIR

[89] Hiroshi Kuratsuji and Tōru Suzuki. Path integral in the representation of SU(2) coherent state and classical dynamics in a generalized phase space. Journal of Mathematical Physics, 21(3):472–476, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p472_s1.

Yeung:1980:FDS

[90] Wai Bong Yeung. Is the Feynman–Dyson series adequate for the asymptotic expansion of the Green’s function of the quantum mechanical anharmonic oscillator? Journal of Mathematical Physics, 21(3):477– 485, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i3/p477_s1.

DeMeyer:1980:SCO

[91] H. E. De Meyer and G. Vanden Berghe. A set of commuting operators and R(3) scalars for the complete classification of quadrupole phonon-states. Journal of Mathematical Physics, 21(3):486–492, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p486_s1.

Newton:1980:ISO

[92] Roger G. Newton. Inverse scattering. I. One dimension. Journal of Mathematical Physics, 21(3):493–505, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i3/p493_s1.

Polyzou:1980:CPM

[93] Wayne N. Polyzou. Combinatorics, partitions, and many-body physics. Journal of Mathematical Physics, 21(3):506–513, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p506_s1.

Qadir:1980:FET

[94] Asghar Qadir. Field equations in twistors. Journal of Mathematical Physics, 21(3):514–520, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i3/p514_s1.

Soares:1980:IRU

[95] I. Damião Soares. Inhomogeneous rotating universes with closed timelike geodesics of matter. Journal of Mathematical Physics, 21(3):521– 525, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i3/p521_s1.

Ashtekar:1980:RST

[96] Abhay Ashtekar and Amitabha Sen. On the role of space–time topology in quantum phenomena: Superselection of charge and emergence of nontrivial vacua. Journal of Mathematical Physics, 21(3):526–533, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i3/p526_s1.

Bhattacharjee:1980:APC

[97] Jayanta K. Bhattacharjee and Richard A. Ferrell. Analytic properties of the critical dynamics ε expansion. Journal of Mathematical Physics, 21(3):534–542, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i3/p534_s1.

Mielke:1980:ESN

[98] Eckehard W. Mielke. On exact solutions of the nonlinear Heisenberg– Klein–Gordon equation in a space–time of constant spacelike curvature. Journal of Mathematical Physics, 21(3):543–546, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p543_s1.

Brandt:1980:FFE

[99] Richard A. Brandt, Wing-Chiu Ng, and Wai-Bong Yeung. Finite field equation of Yang–Mills theory. Journal of Mathematical Physics, 21(3): 547–560, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i3/p547_s1.

Lukierski:1980:SGC

[100] J. Lukierski. Superconformal group and curved fermionic twistor space. Journal of Mathematical Physics, 21(3):561–567, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p561_s1.

Barut:1980:MRB

[101] A. O. Barut. Magnetic resonances between massive and massless spin-1/2 particles with magnetic moments. Journal of Mathematical Physics, 21(3):568–570, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i3/p568_s1.

Eyges:1980:SBV

[102] Leonard Eyges. Solution of boundary value problems with Laplace’s equation for ellipsoids and elliptical cylinders. Journal of Mathematical Physics, 21(3):571–581, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i3/p571_s1.

Chew:1980:RAS

[103] W. C. Chew and J. A. Kong. Resonance of the axial-symmetric modes in microstrip disk resonators. Journal of Mathematical Physics, 21(3): 582–591, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i3/p582_s1.

Trullinger:1980:DPD

[104] S. E. Trullinger. Dynamic polarizability of the double-quadratic kink. Journal of Mathematical Physics, 21(3):592–598, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i3/p592_s1.

Lewis:1980:ELA

[105] H. Ralph Lewis and Keith R. Symon. Erratum: “Linearized analysis of inhomogeneous plasma equilibria: general theory”, [J. Math. Phys. 20 (1979), no. 3, 413–436; MR 80b:82045a]. Journal of Mathematical Physics, 21(3):599, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i3/p599_s1. See [7].

Ohnuki:1980:GGA

[106] Y. Ohnuki and S. Kamefuchi. Generalized Grassmann algebras with applications to Fermi systems. Journal of Mathematical Physics, 21(4): 601–608, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p601_s1.

Ohnuki:1980:PGA

[107] Y. Ohnuki and S. Kamefuchi. Para-Grassmann algebras with applications to para-Fermi systems. Journal of Mathematical Physics, 21(4): 609–616, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p609_s1.

Brooke:1980:GFS

[108] J. A. Brooke. A Galileian formulation of spin. II. Explicit realizations. Journal of Mathematical Physics, 21(4):617–621, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p617_s1.

Venkatesh:1980:NSCa

[109] K. Venkatesh. A note on the symmetries of the 3j and 6j coefficients. I. Journal of Mathematical Physics, 21(4):622–629, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p622_s1.

Wong:1980:MDI

[110] M. K. F. Wong and Hsin Yang Yeh. The most degenerate irreducible representations of the symplectic group. Journal of Mathematical Physics, 21(4):630–635, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p630_s1.

Basu:1980:LGO

[111] Debabrata Basu and D. Mitra. The Lorentz group in the oscillator realization. II. Integral transforms and matrix elements of SO(2, 1). Journal of Mathematical Physics, 21(4):636–637, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p636_s1.

Sarma:1980:PSM

[112] C. R. Sarma and G. G. Sahasrabudhe. Permutational symmetry of many particle states. Journal of Mathematical Physics, 21(4):638–649, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p638_s1.

Goldin:1980:PSI

[113] G. A. Goldin, R. Menikoff, and D. H. Sharp. Particle statistics from induced representations of a local current group. Journal of Mathematical Physics, 21(4):650–664, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p650_s1.

Akyildiz:1980:DSK

[114] Yılmaz Akyıldız. On the dynamical symmetries of the Kepler problem. Journal of Mathematical Physics, 21(4):665–670, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p665_s1.

Bincer:1980:MLO

[115] Adam M. Bincer. Missing label operators in the reduction Sp(2n) ↓ Sp(2n − 2). Journal of Mathematical Physics, 21(4):671–674, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/
p671_s1.

Maekawa:1980:ISP

[116] Takayoshi Maekawa. On the invariant scalar products and the UIR of SO(n, 1). Journal of Mathematical Physics, 21(4):675–679, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/
p675_s1.

Wolf:1980:CTI

[117] Kurt Bernardo Wolf. Canonical transforms. IV. Hyperbolic transforms: continuous series of SL(2, R) representations. Journal of Mathematical Physics, 21(4):680–688, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p680_s1.

Parker:1980:CRS

[118] Monique Parker. Classification of real simple Lie superalgebras of classical type. Journal of Mathematical Physics, 21(4):689–697, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/
p689_s1.

Hall:1980:EEP

[119] George L. Hall. Ewald evaluation and Poisson summation formulas for a class of near-Coulombic lattice sums. Journal of Mathematical Physics, 21(4):698–701, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p698_s1.

Case:1980:SRZa

[120] K. M. Case. Sum rules for zeros of polynomials. I. Journal of Mathematical Physics, 21(4):702–708, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i4/p702_s1.

Case:1980:SRZb

[121] K. M. Case. Sum rules for zeros of polynomials. II. Journal of Mathematical Physics, 21(4):709–714, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i4/p709_s1.

Ablowitz:1980:CBNa

[122] M. J. Ablowitz, A. Ramani, and H. Segur. A connection between nonlinear evolution equations and ordinary differential equations of P -type. I. Journal of Mathematical Physics, 21(4):715–721, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p715_s1.

Hsieh:1980:MED

[123] D. Y. Hsieh. On Mathieu equation with damping. Journal of Mathematical Physics, 21(4):722–725, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i4/p722_s1.

Osborne:1980:SSG

[124] Anthony D. Osborne and Allan E. G. Stuart. On the separability of the sine-Gordon equation and similar quasilinear partial differential equations. II. Dependent- and independent-variable transformations. Journal of Mathematical Physics, 21(4):726–734, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p726_s1.

Chan:1980:CPS

[125] S. K. Chan and I. A. Sakmar. Construction of the phase of the scattering amplitude by iteration of bounds from unitarity integral. Journal of Mathematical Physics, 21(4):735–739, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p735_s1.

Fisch:1980:SVT

[126] Nathaniel J. Fisch and Martin D. Kruskal. Separating variables in twoway diffusion equations. Journal of Mathematical Physics, 21(4):740– 750, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p740_s1.

DeFacio:1980:CBA

[127] B. DeFacio and D. G. Retzloff. Cotangent bundle approach to noninertial frames. Journal of Mathematical Physics, 21(4):751–757, April 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p751_
s1.

Weston:1980:IPR

[128] V. H. Weston. Inverse problem for the reduced wave equation with fixed incident field. Journal of Mathematical Physics, 21(4):758–764, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p758_s1.

Furutsu:1980:DED

[129] K. Furutsu. Diffusion equation derived from the space–time transport equation in anisotropic random media. Journal of Mathematical Physics, 21(4):765–777, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p765_s1.

Aerts:1980:SPD

[130] Dirk Aerts. Subsystems in physics described by bilinear maps between the corresponding vector spaces. Journal of Mathematical Physics, 21(4): 778–788, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p778_s1.

Bach:1980:PFQ

[131] Alexander Bach. A probabilistic formulation of quantum theory. Journal of Mathematical Physics, 21(4):789–793, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p789_s1.

Cohen:1980:PQJ

[132] Leon Cohen and Y. I. Zaparovanny. Positive quantum joint distributions. Journal of Mathematical Physics, 21(4):794–796, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p794_s1.

Ganchev:1980:LSI

[133] A. Ch. Ganchev and T. D. Palev. A Lie superalgebraic interpretation of the para-Bose statistics. Journal of Mathematical Physics, 21(4):797– 799, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p797_s1.

Fonte:1980:EPE

[134] Giacomo Fonte. On the extremum problem of the energy functional on the set of Slater determinants. Journal of Mathematical Physics, 21(4): 800–806, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p800_s1.

Truax:1980:DSR

[135] D. R. Truax. Dynamical symmetries of rotationally invariant, threedimensional, Schrödinger equations. Journal of Mathematical Physics, 21(4):807–817, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p807_s1.

Ali:1980:SRP

[136] S. Twareque Ali. On some representations of the Poincaré group on phase space. II. Journal of Mathematical Physics, 21(4):818–829, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p818_s1.

Maurone:1980:LPW

[137] Philip A. Maurone and Alain J. Phares. The linear potential wavefunctions. Journal of Mathematical Physics, 21(4):830–833, April 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p830_
s1.

Nitti:1980:LBQ

[138] L. Nitti, M. Pellicoro, and M. Villani. Lower bounds for quantum energy levels through statistical mechanics methods. Journal of Mathematical Physics, 21(4):834–839, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p834_s1.

Zorbas:1980:PSA

[139] J. Zorbas. Perturbation of self-adjoint operators by Dirac distributions. Journal of Mathematical Physics, 21(4):840–847, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p840_s1.

Heiss:1980:VFC

[140] W. D. Heiss. Vertex function and crossing symmetry. Journal of Mathematical Physics, 21(4):848–853, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i4/p848_s1.

Komarov:1980:ISP

[141] V. V. Komarov, A. M. Popova, and V. L. Shablov. Independent subsystems in the problems of few-body interaction. Two-body scattering in force field. Journal of Mathematical Physics, 21(4):854–861, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p854_s1.

Ashtekar:1980:NIK

[142] Abhay Ashtekar and B. G. Schmidt. Null infinity and Killing fields. Journal of Mathematical Physics, 21(4):862–867, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p862_s1.

Bose:1980:ETF

[143] S. K. Bose. Electromagnetic test fields in the Kerr–Newman metric. Journal of Mathematical Physics, 21(4):868–869, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p868_s1.

Harvey:1980:CHT

[144] Alex Harvey. Curvatures of hypersurfaces of transitivity of Bianchi model universes. Journal of Mathematical Physics, 21(4):870–873, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/
p870_s1.

Newman:1980:NLF

[145] E. T. Newman and K. P. Tod. A note on left-flat space–times. Journal of Mathematical Physics, 21(4):874–877, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p874_s1.

Briginshaw:1980:CGM

[146] A. J. Briginshaw. The conformal geometry of Minkowski space: Hyperbolic world lines. Journal of Mathematical Physics, 21(4):878–880, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p878_s1.

Constantinescu:1980:ILD

[147] Florin Constantinescu and Berthold Ströter. The Ising limit of the double-well model. Journal of Mathematical Physics, 21(4):881–890, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p881_s1.

Johnson:1980:CWP

[148] Glenn E. Johnson and Daniel I. Fivel. Consequences of weakening the positivity property of Wightman quantum field theories. Journal of Mathematical Physics, 21(4):891–895, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p891_s1.

Haba:1980:SLL

[149] Z. Haba. Scale limit and low momentum behavior of Euclidean fields. II. Journal of Mathematical Physics, 21(4):896–901, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p896_s1.

Katzin:1980:CCC

[150] Gerald H. Katzin and Jack Levine. Charge conservation as a concomitant of conformal motions coupled to generalized gauge transformations.

Journal of Mathematical Physics, 21(4):902–908, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p902_s1.

Brihaye:1980:IPY

[151] Y. Brihaye and J. Nuyts. Invariance properties of Yang equations and applications. Journal of Mathematical Physics, 21(4):909–912, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p909_s1.

Flato:1980:LRN

[152] Moshé Flato and Jacques Simon. Linearization of relativistic nonlinear wave equations. Journal of Mathematical Physics, 21(4):913–917, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p913_s1.

Suzuki:1980:OOF

[153] T. Suzuki. Operator ordering and functional integrals for Fermi systems. Journal of Mathematical Physics, 21(4):918–923, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i4/p918_s1.

Rosensteel:1980:RFM

[154] G. Rosensteel. A recursion formula for sp(3, R) matrix elements. Journal of Mathematical Physics, 21(4):924–927, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i4/p924_s1.

Kodama:1980:TIB

[155] Yuji Kodama and Mark J. Ablowitz. Transverse instability of breathers in resonant media. Journal of Mathematical Physics, 21(4):928–931, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i4/p928_s1.

Trocheris:1980:WNTa

[156] M. Trocheris. Weakly nonlinear theory of coherent Langmuir waves. I. wave wave interaction. Journal of Mathematical Physics, 21(4):932– 940, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p932_s1.

Trocheris:1980:WNTb

[157] M. Trocheris. Weakly nonlinear theory of coherent Langmuir waves. II. wave particle interaction. Journal of Mathematical Physics, 21(4): 941–948, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i4/p941_s1.

Gouedard:1980:EDE

[158] C. Gouedard and C. Deutsch. Erratum: “Dense electron-gas response at any degeneracy”, [J. Math. Phys. 19, 32 (1978)]. Journal of Mathematical Physics, 21(4):949, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i4/p949_s1. See erratum [2].

Hasselbarth:1980:BC

[159] W. Hässelbarth, E. Ruch, D. J. Klein, and T. H. Seligman. Bilateral classes. Journal of Mathematical Physics, 21(5):951–953, May 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p951_
s1.

Ernotte:1980:CPP

[160] P. Ernotte. Commutation properties of 2-parameter groups of isometries. Journal of Mathematical Physics, 21(5):954, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p954_s1.

Castrigiano:1980:IRG

[161] D. P. L. Castrigiano and U. Mutze. On the irreducible representations of groups containing a subgroup of finite index. Journal of Mathematical Physics, 21(5):955–960, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i5/p955_s1.

Dirl:1980:CGCa

[162] R. Dirl. Clebsch–Gordan coefficients for corepresentations. I ⊗ I. Journal of Mathematical Physics, 21(5):961–967, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p961_s1.

Dirl:1980:CGCb

[163] R. Dirl. Clebsch–Gordan coefficients for corepresentations. I⊗II. Journal of Mathematical Physics, 21(5):968–974, May 1980. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p968_s1.

Dirl:1980:CGCc

[164] R. Dirl. Clebsch–Gordan coefficients for corepresentations. I⊗III. Journal of Mathematical Physics, 21(5):975–982, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p975_s1.

Dirl:1980:CGCd

[165] R. Dirl. Clebsch–Gordan coefficients for corepresentations. II⊗II. Journal of Mathematical Physics, 21(5):983–988, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p983_s1.

Dirl:1980:CGCe

[166] R. Dirl. Clebsch–Gordan coefficients for corepresentations. II⊗III. Journal of Mathematical Physics, 21(5):989–996, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p989_s1.

Dirl:1980:CGCf

[167] R. Dirl. Clebsch–Gordan coefficients for corepresentations. III⊗III. Journal of Mathematical Physics, 21(5):997–1005, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p997_s1.

Ablowitz:1980:CBNb

[168] M. J. Ablowitz, A. Ramani, and H. Segur. A connection between nonlinear evolution equations and ordinary differential equations of P -type. II. Journal of Mathematical Physics, 21(5):1006–1015, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1006_s1.

Flaschka:1980:CRP

[169] H. Flaschka. A commutator representation of Painlevé equations. Journal of Mathematical Physics, 21(5):1016–1018, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1016_s1.

Bluman:1980:RND

[170] George Bluman and Sukeyuki Kumei. On the remarkable nonlinear diffusion equation (∂/∂x)[a(u + b)^{−2??}(∂u/∂x)] − (∂u/∂t) = 0. Journal of Mathematical Physics, 21(5):1019–1023, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1019_s1.

Nash:1980:TF

[171] Patrick L. Nash. Towards a factorization of M_{4}. Journal of Mathematical Physics, 21(5):1024–1031, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1024_s1. See erratum [536].

Sloan:1980:CPP

[172] Ian H. Sloan. On choosing the points in product integration. Journal of Mathematical Physics, 21(5):1032–1039, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p1032_s1.

Bohm:1980:DSR

[173] A. Bohm. Decaying states in the rigged Hilbert space formulation of quantum mechanics. Journal of Mathematical Physics, 21(5):1040–1043, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i5/p1040_s1.

Graffi:1980:RKG

[174] S. Graffi, V. Grecchi, and M. Maioli. Resonances in the Klein–Gordon theory of the relativistic Stark effect. Journal of Mathematical Physics, 21(5):1044–1054, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i5/p1044_s1.

Lee:1980:CCI

[175] L. L. Lee. Continuum calculus. IV. the Laplace transform method in the evaluation of the Feynman path integrals with a Gaussian measure and applications in quantum mechanics. Journal of Mathematical Physics, 21(5):1055–1069, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1055_s1. See comment [649].

Hill:1980:SIH

[176] Robert Nyden Hill. Systematic improvement of Hall–Post–Stenschke lower bounds to eigenvalues in the few-body problem. Journal of Mathematical Physics, 21(5):1070–1082, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i5/p1070_s1.

Hill:1980:ESO

[177] Robert Nyden Hill. An exactly solvable one-dimensional three-body problem with hard cores. Journal of Mathematical Physics, 21(5):1083– 1085, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1083_s1.

Bernard:1980:TMF

[178] James E. Bernard, Robert Nyden Hill, and Maurice V. Barnhill III. Test of a method for finding lower bounds to eigenvalues of the threebody problem. Journal of Mathematical Physics, 21(5):1086–1091, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i5/p1086_s1.

Lipszyc:1980:ASM

[179] K. Lipszyc. On the application of the Sommerfeld–Maluzhinetz transformation to some one-dimensional three-particle problems. Journal of Mathematical Physics, 21(5):1092–1102, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i5/p1092_s1.

Marchesini:1980:PLS

[180] G. Marchesini and E. Onofri. Planar limit for SU(N) symmetric quantum dynamical systems. Journal of Mathematical Physics, 21(5):1103– 1110, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1103_s1.

deFalco:1980:LSE

[181] Diego de Falco and Francesco Guerra. On the local structure of the Euclidean Dirac field. Journal of Mathematical Physics, 21(5):1111– 1114, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1111_s1.

Thacker:1980:QIM

[182] H. B. Thacker. Quantum inverse method for two-dimensional ice and ferroelectric lattice models. Journal of Mathematical Physics, 21(5):1115– 1117, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1115_s1.

Blenski:1980:VMT

[183] Tomasz Bleński. A variational method in two-group transport theory. Journal of Mathematical Physics, 21(5):1118–1125, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1118_s1.

Hauser:1980:HHPa

[184] Isidore Hauser and Frederick J. Ernst. A homogeneous Hilbert problem for the Kinnersley–Chitre transformations. Journal of Mathematical Physics, 21(5):1126–1140, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1126_s1.

Leibowitz:1980:SVA

[185] Elhanan Leibowitz. Symmetric vectors and algebraic classification. Journal of Mathematical Physics, 21(5):1141–1148, May 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1141_s1.

Carlson:1980:CFR

[186] Paul Carlson. Coordinate free relativity. Journal of Mathematical Physics, 21(5):1149–1154, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1149_s1.

Muraskin:1980:AFT

[187] M. Muraskin. Aesthetic field theory; the problem of spatial inversions. Journal of Mathematical Physics, 21(5):1155–1158, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1155_s1.

Goenner:1980:SSS

[188] Hubert Goenner and Peter Havas. Spherically symmetric space–times with vanishing curvature scalar. Journal of Mathematical Physics, 21(5): 1159–1167, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1159_s1.

Grishchuk:1980:CCG

[189] L. P. Grishchuk and V. M. Yudin. Conformal coupling of gravitational wave field to curvature. Journal of Mathematical Physics, 21(5):1168– 1175, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1168_s1.

Barnsley:1980:CSK

[190] M. Barnsley and H. Cornille. On a class of solutions of the Krook– Tjon–Wu model of the Boltzmann equation. Journal of Mathematical Physics, 21(5):1176–1193, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1176_s1.

Esposito:1980:HEE

[191] R. Esposito and M. Pulvirenti. Hierarchical equations of evolution of an anharmonic system. Journal of Mathematical Physics, 21(5):1194– 1200, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1194_s1.

Bampi:1980:TAN

[192] F. Bampi and A. Morro. Two approaches to nonstationary relativistic thermodynamics. Journal of Mathematical Physics, 21(5):1201–1204, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i5/p1201_s1.

Ranada:1980:LSC

[193] Antonio F. Rañada and Juan M. Usón. Localized solutions of a charged nonlinear spinor field in a Coulomb-like potential. Journal of Mathematical Physics, 21(5):1205–1209, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i5/p1205_s1.

Landgren:1980:AMP

[194] J. J. Landgren, P. A. Rejto, and Martin Klaus. An application of the maximum principle to the study of essential self-adjointness of Dirac operators. II. Journal of Mathematical Physics, 21(5):1210–1217, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i5/p1210_s1.

Manoukian:1980:ZMB

[195] Edward B. Manoukian. Zero-mass behavior of Feynman amplitudes. I. Journal of Mathematical Physics, 21(5):1218–1223, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1218_s1.

Kim:1980:LDL

[196] Y. S. Kim, Marilyn E. Noz, and S. H. Oh. Lorentz deformation in the O(4) and light-cone coordinate systems. Journal of Mathematical Physics, 21(5):1224–1228, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1224_s1.

Aragone:1980:BCH

[197] C. Aragone, J. Chela-Flores, and A. Restuccia. The Baker–Campbell– Hausdorff formula for the SU(2) supergroups. Journal of Mathematical Physics, 21(5):1229–1233, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i5/p1229_s1.

Gherman:1980:MSB

[198] O. Gherman and L. Saliu. Models for SU(N)×SU(N) symmetry breaking with extremum constraints. Journal of Mathematical Physics, 21(5): 1234–1242, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i5/p1234_s1.

Epstein:1980:DMP

[199] Marcelo Epstein and Reuven Segev. Differentiable manifolds and the principle of virtual work in continuum mechanics. Journal of Mathematical Physics, 21(5):1243–1245, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i5/p1243_s1.

Ram:1980:GDW

[200] Rishi Ram, H. N. Singh, and Madhukar Gaur. The growth and decay of weak waves in relativistic flows of dissociating gases. Journal of Mathematical Physics, 21(5):1246–1250, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i5/p1246_s1.

Kriegsmann:1980:SPE

[201] Gregory A. Kriegsmann. On the scattering of plane electromagnetic waves off cylindrically confined cold plasmas with overdense and steep densities. Journal of Mathematical Physics, 21(5):1251–1255, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/
p1251_s1.

Ott:1980:RME

[202] R. H. Ott. Roots of the modal equation for EM wave propagation in a tropospheric duct. Journal of Mathematical Physics, 21(5):1256–1266, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i5/p1256_s1.

Abellanas:1980:ECD

[203] L. Abellanas and A. Galindo. Erratum: “Conserved densities for nonlinear evolution equations. I. Even order case”, [J. Math. Phys. 20, 1239 (1979)]. Journal of Mathematical Physics, 21(5):1267, May 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1267_

s2. See [10].

Gazeau:1980:EFE

[204] J. P. Gazeau. Erratum: “Four Euclidean conformal groups in atomic calculations: Exact analytical expressions for the bound-bound two photon transition matrix elements in the H atom”, [J. Math. Phys. 19, 1041 (1978)]. Journal of Mathematical Physics, 21(5):1267, May 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1267_

s1. See [4].

Friedman:1980:UMF

[205] John L. Friedman and Rafael D. Sorkin. The unitary multiplicity-free representations of \overline{SL_{4}(R)}. Journal of Mathematical Physics, 21(6):1269– 1276, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1269_s1.

Marcu:1980:RER

[206] M. Marcu. The representations of spl(2, 1)—an example of representations of basic superalgebras. Journal of Mathematical Physics, 21(6):

1277–1283, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1277_s1.

Marcu:1980:TPT

[207] M. Marcu. The tensor product of two irreducible representations of the spl(2, 1) superalgebra. Journal of Mathematical Physics, 21(6):1284– 1292, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1284_s1.

Palev:1980:FSR

[208] Tchavdar D. Palev. Fock space representations of the Lie superalgebra A(0, n). Journal of Mathematical Physics, 21(6):1293–1298, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1293_s1.

Kim:1980:TSI

[209] Shoon K. Kim. The theory of spinors via involutions and its application to the representations of the Lorentz group. Journal of Mathematical Physics, 21(6):1299–1311, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i6/p1299_s1.

Neeman:1980:MCG

[210] Y. Ne’eman and Dj. Šijački. Minimal and centered graded spin-extensions of the SL(3, R) algebra. Journal of Mathematical Physics, 21(6):1312– 1317, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1312_s1.

Fokas:1980:SAE

[211] A. S. Fokas. A symmetry approach to exactly solvable evolution equations. Journal of Mathematical Physics, 21(6):1318–1325, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1318_
s1.

Berryman:1980:ESP

[212] James G. Berryman. Evolution of a stable profile for a class of nonlinear diffusion equations. III. slow diffusion on the line. Journal of Mathematical Physics, 21(6):1326–1331, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i6/p1326_s1.

Martin:1980:HPM

[213] Pablo Martín, Jorge Zamudio-Cristi, and Guillermo Donoso. A Hilbert– Padé method for multipole approximations. application to the Gaussian function. Journal of Mathematical Physics, 21(6):1332–1335, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1332_s1.

Dollard:1980:ECM

[214] J. D. Dollard and C. N. Friedman. Existence and completeness of the Möller wave operators for radial potentials satisfying ∫_{0}^{1} r|v(r)| dr + ∫_{1}^{∞} |v(r)| dr < ∞. Journal of Mathematical Physics, 21(6):1336–1339, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1336_s1.

Coleman:1980:JBP

[215] R. A. Coleman and H. Korte. Jet bundles and path structures. Journal of Mathematical Physics, 21(6):1340–1351, June 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1340_s1. See er-

ratum [947].

Rogers:1980:GTS

[216] Alice Rogers. A global theory of supermanifolds. Journal of Mathematical Physics, 21(6):1352–1365, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i6/p1352_s1.

Burke:1980:ENM

[217] Edward A. Burke. Extended Numerov method for the numerical solution of the Hartree–Fock equations. Journal of Mathematical Physics, 21(6): 1366–1369, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1366_s1.

Lutzky:1980:ARR

[218] M. Lutzky. An application of Ray–Reid invariants. Journal of Mathematical Physics, 21(6):1370–1371, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i6/p1370_s1.

Boisseau:1980:GTH

[219] B. Boisseau. Gauge theory in Hamiltonian classical mechanics: The long-range fields. Journal of Mathematical Physics, 21(6):1372–1376, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1372_s1.

Daubechies:1980:CSP

[220] Ingrid Daubechies. Coherent states and projective representation of the linear canonical transformations. Journal of Mathematical Physics, 21(6):1377–1389, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i6/p1377_s1.

Ringwood:1980:HAQ

[221] G. A. Ringwood and J. T. Devreese. The hydrogen atom: Quantum mechanics on the quotient of a conformally flat manifold. Journal of Mathematical Physics, 21(6):1390–1392, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i6/p1390_s1.

Katznelson:1980:DSD

[222] E. Katznelson. Decaying systems with degenerate Livšic matrix. Journal of Mathematical Physics, 21(6):1393–1407, June 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1393_s1.

Thide:1980:PIC

[223] Bo Thidé. Phase-integral calculation of physically important quantities for nonrelativistic bound s states of the linear central potential. Journal of Mathematical Physics, 21(6):1408–1415, June 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1408_s1.

Chowdhury:1980:PSN

[224] A. R. Chowdhury and T. Roy. Prolongation structure for a nonlinear equation with explicit space dependence. Journal of Mathematical Physics, 21(6):1416–1417, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i6/p1416_s1.

Hauser:1980:HHPb

[225] Isidore Hauser and Frederick J. Ernst. A homogeneous Hilbert problem for the Kinnersley–Chitre transformations of electrovac space–times. Journal of Mathematical Physics, 21(6):1418–1422, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1418_s1.

Taub:1980:STD

[226] A. H. Taub. Space–times with distribution-valued curvature tensors. Journal of Mathematical Physics, 21(6):1423–1431, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1423_s1.

Luehr:1980:GIG

[227] C. P. Luehr and M. P. Rosenbaum. Gravitation as an internal gauge theory of the Poincaré group. Journal of Mathematical Physics, 21(6): 1432–1438, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1432_s1.

Nieh:1980:GBB

[228] H. T. Nieh. Gauss–Bonnet and Bianchi identities in Riemann–Cartan type gravitational theories. Journal of Mathematical Physics, 21(6): 1439–1441, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1439_s1.

Carlson:1980:MPS

[229] George T. Carlson, Jr. and John L. Safko. Matching of the plane symmetric static and homogeneous vacuum solutions of the Einstein field equations. Journal of Mathematical Physics, 21(6):1442–1448, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1442_
s1.

Gambini:1980:ECT

[230] R. Gambini and L. Herrera. Einstein–Cartan theory in the spin coefficient formalism. Journal of Mathematical Physics, 21(6):1449–1454, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1449_s1.

Lindblom:1980:SPS

[231] Lee Lindblom. Some properties of static general relativistic stellar models. Journal of Mathematical Physics, 21(6):1455–1459, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1455_
s1.

Dacol:1980:FIF

[232] Dalcio Kisling Dacol. A functional integral formalism for quantum spin systems. Journal of Mathematical Physics, 21(6):1460–1464, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/
p1460_s1.

Ramanathan:1980:SBH

[233] G. V. Ramanathan and C. Jedrzejek. Structure of the BBGKY hierarchy near phase transition. Journal of Mathematical Physics, 21(6):1465– 1472, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1465_s1.

Ito:1980:CTD

[234] Keiichi R. Ito. Construction of two-dimensional quantum electrodynamics. Journal of Mathematical Physics, 21(6):1473–1494, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1473_
s1.

Mathews:1980:RWE

[235] P. M. Mathews, T. R. Govindarajan, M. Seetharaman, and J. Prabhakar. Relativistic wave equations coupled to external fields: An algebraic study of the problem of constraints. Journal of Mathematical Physics, 21(6): 1495–1505, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1495_s1.

Das:1980:EPSa

[236] A. Das. Extended phase space. I. classical fields. Journal of Mathematical Physics, 21(6):1506–1512, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i6/p1506_s1.

Das:1980:EPSb

[237] A. Das. Extended phase space. II. unified meson fields. Journal of Mathematical Physics, 21(6):1513–1520, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i6/p1513_s1.

Das:1980:EPSc

[238] A. Das. Extended phase space. III. unified spin-1/2 fields. Journal of Mathematical Physics, 21(6):1521–1532, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i6/p1521_s1.

Sharma:1980:LCE

[239] L. K. Sharma, Jyoti Choubey, and H. J. W. Müller-Kirsten. Large coupling expansions for eigenenergies and Regge trajectories of the general even-power potential with applications. Journal of Mathematical Physics, 21(6):1533–1538, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i6/p1533_s1.

Gambini:1980:PDQ

[240] R. Gambini and A. Trias. Path-dependent quantum formulation of electromagnetism with magnetic charges. Journal of Mathematical Physics, 21(6):1539–1545, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i6/p1539_s1.

Weiszmann:1980:MEH

[241] Andrei N. Weiszmann. Multipoles and excitons in higher-order coherence. Journal of Mathematical Physics, 21(6):1546–1551, June 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i6/p1546_
s1.

Arms:1980:ELS

[242] Judith M. Arms. Erratum: “Linearization stability and a globally singular change of variables”, [J. Math. Phys. 21 (1980), no. 1, 15–18; MR 80m:35007]. Journal of Mathematical Physics, 21(6):1552, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i6/p1552_s1. See [22].

Cho:1980:ECG

[243] Yongmin Cho. Erratum: “Classification of gauge fields”, [J. Math. Phys. 20 (1979), no. 12, 2605–2612; MR 80k:81086]. Journal of Mathematical Physics, 21(6):1552, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i6/p1552_s3. See [18].

Lakshmanan:1980:ERB

[244] M. Lakshmanan. Erratum: “Rigid body motions, space curves, prolongation structures, fiber bundles, and solitons”, [J. Math. Phys. 20 (1979), no. 8, 1667–1672; MR 80g:58003]. Journal of Mathematical Physics, 21 (6):1552, June 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i6/p1552_s2. See [14].

Venkatesh:1980:NSCb

[245] K. Venkatesh. A note on the symmetries of the 3j and 6j coefficients. II. Journal of Mathematical Physics, 21(7):1555–1561, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1555_s1.

Agrawala:1980:WET

[246] Vishnu K. Agrawala. Wigner–Eckart theorem for an arbitrary group or Lie algebra. Journal of Mathematical Physics, 21(7):1562–1565, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1562_s1.

Henrich:1980:GSR

[247] Christopher J. Henrich. On Gel’fand states of representations of U(n) and the Gel’fand lattice polynomials. Journal of Mathematical Physics, 21(7): 1566–1576, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1566_s1.

Agrawala:1980:MTI

[248] Vishnu K. Agrawala. Micu-type invariants of exceptional simple Lie algebras. Journal of Mathematical Physics, 21(7):1577–1578, July 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1577_
s1.

Joshi:1980:ACA

[249] K. N. Joshi and B. S. Rajput. Addition of complex angular momentum operators. Journal of Mathematical Physics, 21(7):1579–1582, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1579_s1.

Reid:1980:ESN

[250] James L. Reid and John R. Ray. Ermakov systems, nonlinear superposition, and solutions of nonlinear equations of motion. Journal of Mathematical Physics, 21(7):1583–1587, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i7/p1583_s1.

Chinea:1980:IGC

[251] F. J. Chinea. On the intrinsic geometry of certain nonlinear equations: The sine-Gordon equation. Journal of Mathematical Physics, 21(7):1588– 1592, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1588_s1.

Payne:1980:BTS

[252] Dean A. Payne. Bäcklund transformations in several variables. Journal of Mathematical Physics, 21(7):1593–1602, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1593_s1.

Pomraning:1980:ESL

[253] G. C. Pomraning and E. W. Larsen. Elementary solutions of the linear transport equation for continuously varying spatial media. Journal of Mathematical Physics, 21(7):1603–1612, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1603_s1.

Leibbrandt:1980:SSG

[254] George Leibbrandt, Rudolf Morf, and Shein Shion Wang. Solutions of the sine-Gordon equation in higher dimensions. Journal of Mathematical Physics, 21(7):1613–1624, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1613_s1.

Kelley:1980:SCE

[255] C. T. Kelley. Solution of the Chandrasekhar H-equation by Newton’s Method. Journal of Mathematical Physics, 21(7):1625–1628, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/
p1625_s1.

Znojil:1980:GMR

[256] M. Znojil. Generalized method of a resolvent operator expansion. III. Journal of Mathematical Physics, 21(7):1629–1635, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1629_s1.

Albeverio:1980:RHP

[257] S. Albeverio, R. Høegh-Krohn, and L. Streit. Regularization of Hamiltonians and processes. Journal of Mathematical Physics, 21(7):1636– 1642, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1636_s1.

Rosenstock:1980:ATR

[258] Herbert B. Rosenstock. Absorption time by a random trap distribution. Journal of Mathematical Physics, 21(7):1643–1645, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1643_s1.

vanGroesen:1980:UWP

[259] E. W. C. van Groesen. Unidirectional wave propagation in onedimensional first-order Hamiltonian systems. Journal of Mathematical Physics, 21(7):1646–1655, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1646_s1.

Steeb:1980:SVM

[260] W.-H. Steeb. Symmetries and vacuum Maxwell’s equations. Journal of Mathematical Physics, 21(7):1656–1658, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1656_s1.

Maekawa:1980:PFQ

[261] T. Maekawa and N. Noguchi. Para-Fermi quantization in the representation of SO(n). Journal of Mathematical Physics, 21(7):1659–1661, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1659_s1.

Manoukian:1980:HEB

[262] Edward B. Manoukian. High-energy behavior of renormalized Feynman amplitudes. II. Journal of Mathematical Physics, 21(7):1662–1664, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1662_s1.

Schmutz:1980:SBD

[263] M. Schmutz. Simplified Bose description of para-Bose operators. Journal of Mathematical Physics, 21(7):1665–1666, July 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1665_s1.

Moshinsky:1980:QHP

[264] M. Moshinsky and P. Winternitz. Quadratic Hamiltonians in phase space and their eigenstates. Journal of Mathematical Physics, 21(7):1667–1682, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1667_s1.

Dammert:1980:FPI

[265] Örjan Dammert and Per Olof Fröman. N. Fröman’s phase-integrals versus JWKB: Documentation of the advantages of the former in higherorder approximations. Journal of Mathematical Physics, 21(7):1683– 1687, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1683_s1.

Moxhay:1980:SRN

[266] Peter Moxhay and Jonathan L. Rosner. Semiclassical results on normalization of bound state wavefunctions. Journal of Mathematical Physics, 21(7):1688–1694, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1688_s1.

Hooshyar:1980:CDP

[267] M. A. Hooshyar. Construction of L^{2}-dependent potentials. Journal of Mathematical Physics, 21(7):1695–1697, July 1980. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1695_s1.

Newton:1980:ISI

[268] Roger G. Newton. Inverse scattering. II. Three dimensions. Journal of Mathematical Physics, 21(7):1698–1715, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1698_s1. See errata [537,

1005].

DeFacio:1980:GLE

[269] B. DeFacio and H. E. Moses. The Gel’fand–Levitan equation can give simple examples of non-self-adjoint operators with complete eigenfunctions and spectral representations. I. ghosts and resonances. Journal of Mathematical Physics, 21(7):1716–1723, July 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1716_s1. See er-

ratum [798].

Trlifaj:1980:JSY

[270] L. Trlifaj. The Jost solutions for Yukawa potentials. Journal of Mathematical Physics, 21(7):1724–1731, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i7/p1724_s1.

Myhrman:1980:EEE

[271] Ulla Myhrman. Exact eigenvalues, eigenfunctions, matrix elements and phase-shifts for a particle of angular momentum l in a particular screened potential. Journal of Mathematical Physics, 21(7):1732–1739, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/
p1732_s1.

Birrell:1980:AIQ

[272] N. D. Birrell and J. G. Taylor. Analysis of interacting quantum field theory in curved space–time. Journal of Mathematical Physics, 21(7): 1740–1760, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1740_s1.

Matsumoto:1980:EOQ

[273] H. Matsumoto, G. Semenoff, H. Umezawa, and M. Umezawa. Extended objects in quantum field theory. Journal of Mathematical Physics, 21(7):

1761–1769, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1761_s1.

Sutherland:1980:NSS

[274] Bill Sutherland. Nondiffractive scattering: Scattering from kaleidoscopes. Journal of Mathematical Physics, 21(7):1770–1775, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1770_s1.

Stachel:1980:ACP

[275] John Stachel. The anholonomic Cauchy problem in general relativity. Journal of Mathematical Physics, 21(7):1776–1782, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1776_s1.

Olivert:1980:LSG

[276] J. Olivert. On the local simultaneity in general relativity. Journal of Mathematical Physics, 21(7):1783–1785, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1783_s1.

Horndeski:1980:ABG

[277] Gregory Walter Horndeski. The asymptotic behavior of the generalized Einstein–Maxwell field theory. Journal of Mathematical Physics, 21(7): 1786–1789, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1786_s1.

Jones:1980:GEP

[278] Thaddeus C. Jones. Gravitational and electromagnetic potentials of the stationary Einstein–Maxwell field equations. Journal of Mathematical Physics, 21(7):1790–1797, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1790_s1.

Moffat:1980:SCI

[279] J. W. Moffat. A solution of the Cauchy initial value problem in the nonsymmetric theory of gravitation. Journal of Mathematical Physics, 21(7): 1798–1801, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1798_s1.

Hardy:1980:DPW

[280] J. Hardy and J. Weber. Diffusion in a periodic wind-tree model. Journal of Mathematical Physics, 21(7):1802–1808, July 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1802_s1.

Fannes:1980:CPI

[281] M. Fannes and A. Verbeure. The condensed phase of the imperfect Bose gas. Journal of Mathematical Physics, 21(7):1809–1818, July 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1809_
s1.

Marshall:1980:SEN

[282] T. W. Marshall and P. Claverie. Stochastic electrodynamics of nonlinear systems. I. Particle in a central field of force. Journal of Mathematical Physics, 21(7):1819–1825, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1819_s1.

Lavenda:1980:NFD

[283] B. H. Lavenda. Nonequilibrium fluctuations in driven systems. Journal of Mathematical Physics, 21(7):1826–1832, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1826_s1.

Hampton:1980:FAC

[284] Richard R. Hampton. Feynman amplitude circuit breakers. Journal of Mathematical Physics, 21(7):1833–1841, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i7/p1833_s1.

Goldschmidt:1980:ETD

[285] Yadin Y. Goldschmidt. 1/N expansion in two-dimensional lattice gauge theory. Journal of Mathematical Physics, 21(7):1842–1850, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/
p1842_s1.

Barut:1980:NRE

[286] A. O. Barut, M. Berrondo, and G. García-Calderón. Narrow resonances as an eigenvalue problem and applications to high energy magnetic resonances: An exactly soluble model. Journal of Mathematical Physics, 21(7):1851–1855, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i7/p1851_s1.

Ahmed:1980:OAM

[287] S. Ahmed and H. Schiff. Orbital angular momentum of “lumps”. Journal of Mathematical Physics, 21(7):1856–1862, July 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1856_s1.

Galvao:1980:CSP

[288] Carlos A. P. Galva o and Claudio Teitelboim. Classical supersymmetric particles. Journal of Mathematical Physics, 21(7):1863–1880, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1863_s1.

Bosanac:1980:PEM

[289] S. Bosanac. Perturbation expansion of multichannel Regge poles. Journal of Mathematical Physics, 21(7):1881–1887, July 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1881_s1.

Goldflam:1980:PPA

[290] R. Goldflam, K. L. Kowalski, and W. Tobocman. Partition permuting array approach to few-body Hamiltonian models of nuclear reactions. Journal of Mathematical Physics, 21(7):1888–1896, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1888_s1.

Aviram:1980:MCC

[291] I. Aviram and S. Goshen. A multigroup criticality condition for a spaceindependent multiplying assembly via the Chapman–Kolmogorov equation. Journal of Mathematical Physics, 21(7):1897–1901, July 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1897_

s1. See erratum [614].

VandenBerghe:1980:SOTa

[292] G. Vanden Berghe and H. E. De Meyer. Shift-operator techniques for the classification of multipole-phonon states. I. properties of shift operators in theR(5) group. Journal of Mathematical Physics, 21(7):1902– 1905, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1902_s1.

DeMeyer:1980:SOTa

[293] H. E. De Meyer and G. Vanden Berghe. Shift-operator techniques for the classification of multipole-phonon states. II. eigenvalues of the quadrupole shift operator O_{l}^{0}. Journal of Mathematical Physics, 21(7): 1906–1912, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1906_s1.

DeMeyer:1980:SOTb

[294] H. E. De Meyer and G. Vanden Berghe. Shift operator techniques for the classification of multipole-phonon states. III. the quadrupole-phonon eigenstates of O_{l}^{0}. Journal of Mathematical Physics, 21(7):1913–1917, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1913_s1.

Ringwood:1980:SSP

[295] G. A. Ringwood and J. T. Devreese. On some solutions of the polaron model. Journal of Mathematical Physics, 21(7):1918–1920, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1918_s1.

Thomas:1980:QHF

[296] Lawrence E. Thomas. Quantum Heisenberg ferromagnets and stochastic exclusion processes. Journal of Mathematical Physics, 21(7):1921– 1924, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i7/p1921_s1.

Larsen:1980:PWE

[297] Ulf Larsen. Partial wave expansion of the RKKY indirect exchange interaction in metals. Journal of Mathematical Physics, 21(7):1925–1937, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1925_s1.

Melvin:1980:SMH

[298] M. A. Melvin and T. R. Michalik. Symmetries and metrics of homogeneous cosmologies. Journal of Mathematical Physics, 21(7):1938–1951, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1938_s1.

Michalik:1980:SHN

[299] T. R. Michalik and M. A. Melvin. Spatially homogeneous neutrino cosmologies. Journal of Mathematical Physics, 21(7):1952–1964, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i7/p1952_s1.

Mizrahi:1980:EPS

[300] Maurice M. Mizrahi. Erratum: “Phase space path integrals, without limiting procedure”, [J. Math. Phys. 19 (1978), no. 1, 298–307; MR 56 #17627]. Journal of Mathematical Physics, 21(7):1965, July 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i7/p1965_

s1. See [3].

VandenBerghe:1980:SOTb

[301] G. Vanden Berghe and H. E. De Meyer. Shift operator techniques for the classification of multipole-phonon states. IV. Properties of shift operators in the G_{2} group. Journal of Mathematical Physics, 21(8):1967– 1972, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i8/p1967_s1.

DeMeyer:1980:SOTc

[302] H. E. De Meyer and G. Vanden Berghe. Shift operator techniques for the classification of multipole-phonon states. V. Properties of shift operators in the R(7) group. Journal of Mathematical Physics, 21(8):1973– 1976, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i8/p1973_s1.

Holman:1980:GBF

[303] Wayne J. Holman III. Generalized Bessel functions and the representation theory of U(2)σC^{2×2}. Journal of Mathematical Physics, 21(8):1977– 2010, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i8/p1977_s1.

Dirl:1980:SRT

[304] R. Dirl. Selection rules for type II Shubnikov space groups. Journal of Mathematical Physics, 21(8):2011–2018, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2011_s1.

Dirl:1980:SR

[305] R. Dirl. Selection rules for P n3′n. Journal of Mathematical Physics, 21(8):2019–2025, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2019_s1.

Lemire:1980:CNG

[306] F. Lemire and J. Patera. Congruence number, a generalization of SU(3) triality. Journal of Mathematical Physics, 21(8):2026–2027, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/
p2026_s1.

Jordan:1980:SDN

[307] Thomas F. Jordan. Simple derivation of the Newton–Wigner position operator. Journal of Mathematical Physics, 21(8):2028–2032, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2028_s1.

Iosifescu:1980:SDC

[308] Mircea Iosifescu and Horia Scutaru. On six-dimensional canonical realizations of the so(4, 2) algebra. Journal of Mathematical Physics, 21(8):2033–2045, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2033_s1.

Gonzalez-Gascon:1980:SDE

[309] F. González-Gascón. Symmetries of differential equations. II. Journal of Mathematical Physics, 21(8):2046–2051, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2046_s1.

Ramm:1980:VPR

[310] A. G. Ramm. A variational principle for resonances. Journal of Mathematical Physics, 21(8):2052–2053, August 1980. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2052_s1.

Berk:1980:WMS

[311] H. L. Berk and D. Pfirsch. WKB method for systems of integral equations. Journal of Mathematical Physics, 21(8):2054–2066, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/
p2054_s1.

Antoine:1980:PIPb

[312] J.-P. Antoine. Partial inner product spaces. IV. Topological considerations. Journal of Mathematical Physics, 21(8):2067–2079, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2067_s1.

Daubechies:1980:ITR

[313] I. Daubechies and A. Grossmann. An integral transform related to quantization. Journal of Mathematical Physics, 21(8):2080–2090, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2080_s1.

Bugajska:1980:NS

[314] Krystyna Bugajska. Nature of superspace. Journal of Mathematical Physics, 21(8):2091–2096, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i8/p2091_s1.

Bugajska:1980:GPS

[315] Krystyna Bugajska. On geometrical properties of spinor structure. Journal of Mathematical Physics, 21(8):2097–2101, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2097_s1.

Moore:1980:SFS

[316] Steven M. Moore. Stochastic fields from stochastic mechanics. Journal of Mathematical Physics, 21(8):2102–2110, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2102_s1.

Wiegel:1980:RCT

[317] F. W. Wiegel. On a remarkable class of two-dimensional random walks. Journal of Mathematical Physics, 21(8):2111–2113, August 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2111_
s1.

Rose:1980:TOM

[318] Conrad M. Rose and Ioannis M. Besieris. Nth-order multifrequency coherence functions: A functional path integral approach. II. Journal of Mathematical Physics, 21(8):2114–2120, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2114_s1.

Mortell:1980:SAD

[319] Michael P. Mortell and Brian R. Seymour. A simple approximate determination of stochastic transition for the standard mapping. Journal of Mathematical Physics, 21(8):2121–2123, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2121_s1.

Dominici:1980:PCI

[320] D. Dominici and J. Gomis. Poincaré–Cartan integral invariant and canonical transformations for singular Lagrangians. Journal of Mathematical Physics, 21(8):2124–2130, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2124_s1. See addendum

[934].

Wenzel:1980:SBW

[321] Alan R. Wenzel. Spectral broadening of waves propagating in a random medium. Journal of Mathematical Physics, 21(8):2131–2139, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2131_s1.

Terras:1980:IMC

[322] Riho Terras and Robert Swanson. Image methods for constructing Green’s functions and eigenfunctions for domains with plane boundaries. Journal of Mathematical Physics, 21(8):2140–2153, August 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2140_
s1.

Brinati:1980:QFN

[323] J. R. Brinati and S. S. Mizrahi. Quantum friction in the c-number picture: The damped harmonic oscillator. Journal of Mathematical Physics, 21(8):2154–2158, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2154_s1.

Landovitz:1980:TAT

[324] L. F. Landovitz, A. M. Levine, and W. M. Schreiber. Transition amplitudes for time-dependent harmonic oscillators. Journal of Mathematical Physics, 21(8):2159–2163, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i8/p2159_s1.

Brittin:1980:NOC

[325] Wesley E. Brittin and Arthur Y. Sakakura. Number operators for composite particles in nonrelativistic many-body theory. Journal of Mathematical Physics, 21(8):2164–2169, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2164_s1.

Bracken:1980:IHO

[326] A. J. Bracken and H. I. Leemon. The isotropic harmonic oscillator in an angular momentum basis: An algebraic formulation. Journal of Mathematical Physics, 21(8):2170–2181, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2170_s1.

Hill:1980:TLB

[327] Robert Nyden Hill. Tight lower bounds to eigenvalues of the Schrödinger equation. Journal of Mathematical Physics, 21(8):2182–2192, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2182_s1.

Newell:1980:ADE

[328] G. F. Newell. Asymptotic distribution of eigenvalues for the multidimensional Schrödinger equation. Journal of Mathematical Physics, 21(8):2193–2201, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2193_s1.

Goodmanson:1980:CSL

[329] David M. Goodmanson and John R. Taylor. Coulomb scattering as the limit of scattering off smoothly screened Coulomb potentials. Journal of Mathematical Physics, 21(8):2202–2207, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2202_s1.

Parravicini:1980:RST

[330] Guido Parravicini, Vittorio Gorini, and E. C. G. Sudarshan. Resonances, scattering theory, and rigged Hilbert spaces. Journal of Mathematical Physics, 21(8):2208–2226, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i8/p2208_s1.

Sakmar:1980:VPF

[331] I. A. Sakmar. A variational problem for the forward or backward scattering amplitude with three constraints. Journal of Mathematical Physics, 21(8):2227–2230, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2227_s1.

Kinnersley:1980:SSE

[332] William Kinnersley. Symmetries of the stationary Einstein–Maxwell field equations. VII. charging transformations. Journal of Mathematical Physics, 21(8):2231–2235, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i8/p2231_s1.

Harnad:1980:CCR

[333] J. Harnad, J. Tafel, and S. Shnider. Canonical connections on Riemannian symmetric spaces and solutions to the Einstein–Yang–Mills equations. Journal of Mathematical Physics, 21(8):2236–2240, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2236_s1.

Hoenselaers:1980:NSE

[334] C. Hoenselaers. On a new solution of Einstein’s equations. Journal of Mathematical Physics, 21(8):2241–2244, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2241_s1.

Lind:1980:NGC

[335] Robert W. Lind. A note on the good-cut equations. Journal of Mathematical Physics, 21(8):2245–2247, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2245_s1.

Das:1980:SEM

[336] A. Das and S. Kloster. On the stationary Einstein–Maxwell field equations. Journal of Mathematical Physics, 21(8):2248–2256, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/
p2248_s1.

Lun:1980:KLN

[337] Anthony W.-C. Lun. Kerr-like neutrino-gravitational solutions. Journal of Mathematical Physics, 21(8):2257–2260, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2257_s1.

Rao:1980:SEM

[338] K. N. Srinivasa Rao and A. V. Gopala Rao. Some electrovac models of homogeneous gravitational force fields in general relativity. Journal of Mathematical Physics, 21(8):2261–2268, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2261_s1.

Johri:1980:ESB

[339] V. B. Johri and G. K. Goswami. Exact solutions of Brans–Dicke theory for irrotational barotropic models with stiff matter. Journal of Mathematical Physics, 21(8):2269–2271, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i8/p2269_s1.

Chen:1980:ASB

[340] M. Chen. An alternative solution of the Bogoliubov’s equation for the two-particle distribution function in the kinetic theory of gases. Journal of Mathematical Physics, 21(8):2272–2277, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2272_s1.

Constantinescu:1980:ACC

[341] Florin Constantinescu. Analyticity in the coupling constant of the λP (φ) lattice theory. Journal of Mathematical Physics, 21(8):2278–2281, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2278_s1.

Kim:1980:TITa

[342] Shoon K. Kim. Theory of involutional transformations applied to the Dirac theory of the electron. I. Remarks on the Dirac plane waves. Journal of Mathematical Physics, 21(8):2282–2285, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2282_s1.

Kim:1980:TITb

[343] Shoon K. Kim. Theory of involutional transformations applied to the Dirac theory of the electron. II. Remarks on the Dirac–Coulomb waves. Journal of Mathematical Physics, 21(8):2286–2290, August 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2286_
s1.

Moncrief:1980:GEM

[344] Vincent Moncrief. Global existence of Maxwell–Klein–Gordon fields in (2 + 1)-dimensional space–time. Journal of Mathematical Physics, 21(8):2291–2296, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2291_s1.

Rosensteel:1980:EHF

[345] G. Rosensteel and Edwin Ihrig. Existence of Hartree–Fock solutions. Journal of Mathematical Physics, 21(8):2297–2301, August 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2297_
s1.

Peeters:1980:PSC

[346] F. Peeters and J. T. Devreese. Periodic solutions of the classical polaron and bipolaron systems. Journal of Mathematical Physics, 21(8):2302– 2306, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i8/p2302_s1.

Puga:1980:TEH

[347] M. W. Puga. Thermal excitations in Heisenberg-xy systems. Journal of Mathematical Physics, 21(8):2307–2312, August 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i8/p2307_s1.

Murray:1980:ESF

[348] J. C. Murray. Erratum: “On the sound field due to a moving source in a superfluid”, [J. Math. Phys. 20, 1409 (1979)]. Journal of Mathematical Physics, 21(8):2313, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2313_s1. See [11].

Pavelle:1980:EPI

[349] Richard Pavelle. Erratum: “The Planck integral cannot be evaluated in terms of a finite series of elementary functions”, [J. Math. Phys. 21 (1980), no. 1, 14; MR 80k:26001]. Journal of Mathematical Physics, 21(8):2313, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i8/p2313_s2. See [21].

Augusteijn:1980:EIN

[350] M. F. Augusteijn and E. Breitenberger. Erratum: “Integration of nearresonant systems in slow-fluctuation approximation”, [J. Math. Phys. 21, 462 (1980)]. Journal of Mathematical Physics, 21(8):2314, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i8/p2314_s1. See [88].

Murphy:1980:SOD

[351] T. Murphy. Shift operators for discrete representations of O(p, q). Journal of Mathematical Physics, 21(9):2317–2320, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2317_s1.

Stanley:1980:CG

[352] Richard P. Stanley. The character generator of SU(n). Journal of Mathematical Physics, 21(9):2321–2326, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i9/p2321_s1.

delAguila:1980:CTP

[353] F. del Aguila. Classification of three-particle states according to an orthonormal SU(3) ⊇ SO(3) basis. Journal of Mathematical Physics, 21 (9):2327–2334, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i9/p2327_s1.

Patera:1980:NRB

[354] J. Patera, R. T. Sharp, and R. Slansky. On a new relation between semisimple Lie algebras. Journal of Mathematical Physics, 21(9):2335– 2341, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2335_s1.

MartinezAlonso:1980:SSP

[355] L. Martínez Alonso. Schrödinger spectral problems with energydependent potentials as sources of nonlinear Hamiltonian evolution equations. Journal of Mathematical Physics, 21(9):2342–2349, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i9/p2342_s1.

Cawley:1980:GIH

[356] Robert Cawley. Gauge invariance and the Helmholtz conditions. Journal of Mathematical Physics, 21(9):2350–2359, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2350_s1.

Corum:1980:RCR

[357] James F. Corum. Relativistic covariance and rotational electrodynamics. Journal of Mathematical Physics, 21(9):2360–2364, September 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2360_
s1.

Girardeau:1980:FTR

[358] M. D. Girardeau. Fock–Tani representation for composite particles in a soluble model. Journal of Mathematical Physics, 21(9):2365–2375, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2365_s1.

Grossmann:1980:CES

[359] A. Grossmann, R. Høegh-Krohn, and M. Mebkhout. A class of explicitly soluble, local, many-center Hamiltonians for one-particle quantum mechanics in two and three dimensions. I. Journal of Mathematical Physics, 21(9):2376–2385, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i9/p2376_s1.

Mukunda:1980:RPP

[360] N. Mukunda, E. C. G. Sudarshan, J. K. Sharma, and C. L. Mehta. Representations and properties of para-Bose oscillator operators. I. Energy position and momentum eigenstates. Journal of Mathematical Physics, 21(9):2386–2394, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i9/p2386_s1.

Ramm:1980:NPA

[361] A. G. Ramm and B. A. Taylor. A new proof of absence of positive discrete spectrum of the Schrödinger operator. Journal of Mathematical Physics, 21(9):2395–2397, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i9/p2395_s1.

Darewych:1980:ULB

[362] J. W. Darewych and R. Pooran. Upper and lower bounds in nonrelativistic scattering theory. Journal of Mathematical Physics, 21(9):2398–2405, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2398_s1.

Devinatz:1980:EWO

[363] Allen Devinatz. The existence of wave operators for oscillating potentials. Journal of Mathematical Physics, 21(9):2406–2411, September 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2406_
s1.

Haggman:1980:PCS

[364] B. C. Haggman, G. W. Horndeski, and G. Mess. Properties of a covering space defined by Hawking. Journal of Mathematical Physics, 21(9):2412– 2416, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2412_s1.

Cosgrove:1980:RBG

[365] Christopher M. Cosgrove. Relationships between the group-theoretic and soliton-theoretic techniques for generating stationary axisymmetric gravitational solutions. Journal of Mathematical Physics, 21(9):2417– 2447, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2417_s1.

Larsen:1980:SEL

[366] Edward W. Larsen, G. C. Pomraning, and V. C. Badham. On the singular eigenfunctions for linear transport in an exponential atmosphere. Journal of Mathematical Physics, 21(9):2448–2454, September 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2448_
s1.

Bolsterli:1980:IIS

[367] M. Bolsterli. Interaction of isovector scalar mesons with simple sources. II. Isobars and meson scattering. Journal of Mathematical Physics, 21 (9):2455–2460, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i9/p2455_s1.

Hostler:1980:QFT

[368] Levere Hostler. Quantum field theory of particles of indefinite mass. I. Journal of Mathematical Physics, 21(9):2461–2467, September 1980. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2461_
s1.

Siewert:1980:CER

[369] C. E. Siewert. On computing eigenvalues in radiative transfer. Journal of Mathematical Physics, 21(9):2468–2470, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i9/p2468_s1.

Ben-Artzi:1980:ESS

[370] Matania Ben-Artzi and Allen Devinatz. Erratum: “Spectral and scattering theory for the adiabatic oscillator and related potentials”, [J. Math. Phys. 20, 594 (1979)]. Journal of Mathematical Physics, 21(9):2471, September 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i9/p2471_s1. See [8].

Antippa:1980:TSO

[371] Adel F. Antippa and Nguyen Ky Toan. Topological solution of ordinary and partial finite difference equations. Journal of Mathematical Physics, 21(10):2475–2480, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i10/p2475_s1.

Abresch:1980:FNS

[372] U. Abresch, A. Bovier, O. Lechtenfeld, M. Lüling, V. Rittenberg, and G. Weymans. Finite nonabelian subgroups of SU(n) with analytic expressions for the irreducible representations and the Clebsch–Gordan coefficients. Journal of Mathematical Physics, 21(10):2481–2486, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2481_s1.

vandenBroek:1980:EIR

[373] P. M. van den Broek and R. Dirl. Equivalence of induced representations. Journal of Mathematical Physics, 21(10):2487–2490, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2487_s1.

Beckers:1980:SFI

[374] J. Beckers, J. Harnad, and P. Jasselette. Spinor fields invariant under space–time transformations. Journal of Mathematical Physics, 21 (10):2491–2499, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i10/p2491_s1.

Soize:1980:OSS

[375] Christian Soize. Oscillators submitted to squared Gaussian processes. Journal of Mathematical Physics, 21(10):2500–2507, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i10/
p2500_s1.

Fordy:1980:FOM

[376] Allan P. Fordy and John Gibbons. Factorization of operators. I. Miura transformations. Journal of Mathematical Physics, 21(10):2508–2510, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i10/p2508_s1.

Onofri:1980:MBD

[377] E. Onofri. A modified Bars–Durgut equation with polynomial eigenfunctions. Journal of Mathematical Physics, 21(10):2511–2520, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2511_s1.

Klein:1980:VPP

[378] Abraham Klein, Ching Teh Li, and Moyez Vassanji. Variational principles for particles and fields in Heisenberg matrix mechanics. Journal of Mathematical Physics, 21(10):2521–2527, October 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i10/p2521_s1.

Schonfeld:1980:DSM

[379] Jonathan F. Schonfeld. Dynamical symmetry and magnetic charge. Journal of Mathematical Physics, 21(10):2528–2533, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i10/p2528_s1.

Nash:1980:FCP

[380] Patrick L. Nash. A factorization of M_{4}: construction of the principal bundle of orthogonal frames over M_{4} from \overline{O(3, 3)} spinors. Journal of Mathematical Physics, 21(10):2534–2537, October 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i10/p2534_s1.

Ludwig:1980:ABG

[381] Garry Ludwig. Asymptotic behavior of gravitational fields in a type II coordinate system. Journal of Mathematical Physics, 21(10):2538– 2542, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i10/p2538_s1.

Ludwig:1980:STG

[382] Garry Ludwig. Space–times with geodesic, shearfree, twistfree, nonexpanding rays. Journal of Mathematical Physics, 21(10):2543–2548, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2543_s1.

Rashid:1980:LGF

[383] M. A. Rashid. Lattice Green’s functions for cubic lattices. Journal of Mathematical Physics, 21(10):2549–2552, October 1980. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i10/p2549_s1.

Kerner:1980:COI

[384] Richard Kerner. Covariant objects and invariant equations on fiber bundles. Journal of Mathematical Physics, 21(10):2553–2559, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i10/
p2553_s1.

Palev:1980:MIF

[385] Tchavdar D. Palev. S-matrix for interacting A-fields. Journal of Mathematical Physics, 21(10):2560–2561, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i10/p2560_s1.

vanNieuwenhuizen:1980:USG

[386] P. van Nieuwenhuizen. Unitarity of supergravity and Z_{2} or Z_{2} × Z_{2} or Z_{2} × Z_{2} × Z_{2} gradings of gauge and ghost fields. Journal of Mathematical Physics, 21(10):2562–2566, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i10/p2562_s1.

Sommers:1980:SS

[387] Paul Sommers. Space spinors. Journal of Mathematical Physics, 21 (10):2567–2571, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i10/p2567_s1.

Leon:1980:DIS

[388] Jérôme Jp. Leon. The Dirac inverse spectral transform: Kinks and boomerons. Journal of Mathematical Physics, 21(10):2572–2578, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2572_s1.

Macfarlane:1980:PUM

[389] A. J. Macfarlane. Parametrizations of unitary matrices and related coset spaces. Journal of Mathematical Physics, 21(10):2579–2582, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i10/p2579_s1.

Beran:1980:FFT

[390] Mark J. Beran. Field fluctuations in a two-phase random medium. Journal of Mathematical Physics, 21(10):2583–2585, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i10/p2583_s1.

Bonilla:1980:LNP

[391] L. L. Bonilla and M. G. Velarde. Localized nonuniform patterns in a diffusion-reaction model with autocatalysis and the Langmuir– Hinshelwood saturation law. Journal of Mathematical Physics, 21 (10):2586–2589, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i10/p2586_s1.

Chew:1980:RNS

[392] W. C. Chew and J. A. Kong. Resonance of nonaxial symmetric modes in circular microstrip disk antenna. Journal of Mathematical Physics, 21(10):2590–2598, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i10/p2590_s1.

Collinson:1980:HMH

[393] C. D. Collinson. Homothetic motions and the Hauser metric. Journal of Mathematical Physics, 21(11):2601–2602, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i11/p2601_s1.

Devaney:1980:MST

[394] A. J. Devaney. Multiple scattering theory for discrete, elastic, random media. Journal of Mathematical Physics, 21(11):2603–2611, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i11/p2603_s1.

Gordon:1980:LBC

[395] Paul Gordon. Lateral boundary conditions for quasisteady atmospheric flows. Journal of Mathematical Physics, 21(11):2612–2627, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i11/p2612_s1.

Kumar:1980:TEO

[396] Sunil Kumar and C. L. Mehta. Time evolution operator for N interacting quantum harmonic oscillators. Journal of Mathematical Physics, 21 (11):2628–2631, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i11/p2628_s1.

VandenBerghe:1980:SOTc

[397] G. Vanden Berghe and H. E. De Meyer. Shift operator techniques for the classification of multipole-phonon states. VI. Properties of nonscalar R(3) product operators in the R(2λ+1) groups. Journal of Mathematical Physics, 21(11):2632–2634, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i11/p2632_s1.

DeMeyer:1980:SOTd

[398] H. E. De Meyer and G. Vanden Berghe. Shift operator techniques for the classification of multipole-phonon states. VII. self-consistent single step algorithm for R(5)O_{l}^{0} eigenstate and eigenvalue determination. Journal of Mathematical Physics, 21(11):2635–2639, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i11/p2635_s1.

Johnson:1980:CBR

[399] Bruce R. Johnson. On a connection between radial Schrödinger equations for different power-law potentials. Journal of Mathematical Physics, 21 (11):2640–2647, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i11/p2640_s1.

Prosser:1980:FSI

[400] Reese T. Prosser. Formal solutions of inverse scattering problems. III. Journal of Mathematical Physics, 21(11):2648–2653, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i11/
p2648_s1.

Kunzle:1980:UEC

[401] H. P. Künzle and J. R. Savage. On the uniqueness of the equilibrium configurations of slowly rotating relativistic fluids. Journal of Mathematical Physics, 21(11):2654–2661, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i11/p2654_s1.

Ryter:1980:PNI

[402] D. Ryter and U. Deker. Properties of the noise-induced “spurious” drift. I. Journal of Mathematical Physics, 21(11):2662–2665, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i11/
p2662_s1.

Deker:1980:PNI

[403] U. Deker and D. Ryter. Properties of the noise-induced (“spurious”) drift. II. Simplifications of Langevin equations. Journal of Mathematical Physics, 21(11):2666–2669, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i11/p2666_s1.

Garbaczewski:1980:CSA

[404] Piotr Garbaczewski. A concept of spin 1/2 approximation in the quantum theory of lattice Bose systems. Journal of Mathematical Physics, 21 (11):2670–2675, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i11/p2670_s1. See erratum [467].

Cahill:1980:NIS

[405] Kevin Cahill. Nonlinear internal symmetry. Journal of Mathematical Physics, 21(11):2676–2677, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i11/p2676_s1.

Bars:1980:IGF

[406] I. Bars. U(N) integral for the generating functional in lattice gauge theory. Journal of Mathematical Physics, 21(11):2678–2681, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i11/p2678_s1.

Gomis:1980:PWE

[407] J. Gomis, A. Poch, and J. M. Pons. Poincaré wave equations as Fourier transforms of Galilei wave equations. Journal of Mathematical Physics, 21(11):2682–2685, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i11/p2682_s1.

McGuire:1980:SIS

[408] D. McGuire. A solution of an inverse-source problem in coherence theory. Journal of Mathematical Physics, 21(11):2686–2689, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i11/
p2686_s1.

Zubairy:1980:QSM

[409] M. S. Zubairy. Quantum statistics of multimode m-photon absorption process. Journal of Mathematical Physics, 21(11):2690–2692, November 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i11/p2690_s1.

Samuel:1980:IWH

[410] Stuart Samuel. U(N) integrals, 1/N, and the De Wit–’t Hooft anomalies. Journal of Mathematical Physics, 21(12):2695–2703, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v21/i12/
p2695_s1.

Reiter:1980:LGS

[411] George Reiter. Lie groups, spin equations, and the geometrical interpretation of solitons. Journal of Mathematical Physics, 21(12):2704–2714, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i12/p2704_s1.

Taylor:1980:GGG

[412] J. G. Taylor. Generalized groups as global or local symmetries. Journal of Mathematical Physics, 21(12):2715–2718, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i12/p2715_s1.

Harnad:1980:GAP

[413] J. Harnad, S. Shnider, and Luc Vinet. Group actions on principal bundles and invariance conditions for gauge fields. Journal of Mathematical Physics, 21(12):2719–2724, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2719_s1.

Filter:1980:MRT

[414] Eckhard Filter and E. Otto Steinborn. A matrix representation of the translation operator with respect to a basis set of exponentially declining functions. Journal of Mathematical Physics, 21(12):2725–2736, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i12/p2725_s1.

McGuinness:1980:IPCa

[415] Mark J. McGuinness. Infinities of polynomial conserved densities for nonlinear evolution equations. Journal of Mathematical Physics, 21 (12):2737–2742, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i12/p2737_s1.

McGuinness:1980:IPCb

[416] Mark J. McGuinness. An infinity of polynomial conserved densities for a class of nonlinear evolution equations. Journal of Mathematical Physics, 21(12):2743–2748, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i12/p2743_s1.

Bruschi:1980:NTL

[417] M. Bruschi, S. V. Manakov, O. Ragnisco, and D. Levi. The nonabelian Toda lattice: Discrete analogue of the matrix Schrödinger spectral problem. Journal of Mathematical Physics, 21(12):2749–2753, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i12/p2749_s1.

Borghese:1980:ATV

[418] F. Borghese, P. Denti, G. Toscano, and O. I. Sindoni. An addition theorem for vector Helmholtz harmonics. Journal of Mathematical Physics, 21(12):2754–2755, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i12/p2754_s1.

Arnow:1980:PTF

[419] M. Arnow. Particle trajectories in 1/r fields. Journal of Mathematical Physics, 21(12):2756–2759, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2756_s1.

Siewert:1980:CAS

[420] C. E. Siewert, R. D. M. Garcia, and P. Grandjean. A concise and accurate solution for Poiseuille flow in a plane channel. Journal of Mathematical Physics, 21(12):2760–2763, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2760_s1.

Furutsu:1980:OMT

[421] K. Furutsu. Operator methods for time-dependent waves in random media with applications to the case of random particles. Journal of Mathematical Physics, 21(12):2764–2779, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i12/p2764_s1.

Bessis:1980:NSE

[422] N. Bessis and G. Bessis. A note on the Schrödinger equation for the x^{2} + λx^{2}/(1 + gx^{2}) potential. Journal of Mathematical Physics, 21(12):2780– 2785, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i12/p2780_s1.

King:1980:ESE

[423] Michael J. King and Fritz Rohrlich. Exactly solvable eigenvalue problem with hypergeometric eigenfunctions. Journal of Mathematical Physics, 21(12):2786–2788, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i12/p2786_s1.

Brill:1980:SSS

[424] Dieter R. Brill, John M. Cavallo, and James A. Isenberg. K-surfaces in the Schwarzschild space–time and the construction of lattice cosmologies. Journal of Mathematical Physics, 21(12):2789–2796, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i12/p2789_s1.

Ray:1980:GLC

[425] Dipankar Ray. Gödel-like cosmological solutions. Journal of Mathematical Physics, 21(12):2797–2798, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2797_s1.

Banerjee:1980:SGM

[426] A. Banerjee and S. B. Duttachoudhury. Static gravitational and Maxwell fields in the general scalar tensor theory. Journal of Mathematical Physics, 21(12):2799–2801, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2799_s1.

Wald:1980:DNH

[427] Robert M. Wald. Dynamics in nonglobally hyperbolic, static space– times. Journal of Mathematical Physics, 21(12):2802–2805, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i12/p2802_s1.

Samuel:1980:UAVa

[428] Stuart Samuel. The use of anticommuting variable integrals in statistical mechanics. I. The computation of partition functions. Journal of Mathematical Physics, 21(12):2806–2814, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v21/i12/p2806_s1.

Samuel:1980:UAVb

[429] Stuart Samuel. The use of anticommuting variable integrals in statistical mechanics. II. The computation of correlation functions. Journal of Mathematical Physics, 21(12):2815–2819, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i12/p2815_s1.

Samuel:1980:UAVc

[430] Stuart Samuel. The use of anticommuting variable integrals in statistical mechanics. III. Unsolved models. Journal of Mathematical Physics, 21 (12):2820–2833, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i12/p2820_s1.

Thierry-Mieg:1980:GRF

[431] Jean Thierry-Mieg. Geometrical reinterpretation of Faddeev–Popov ghost particles and BRS transformations. Journal of Mathematical Physics, 21(12):2834–2838, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2834_s1.

Bose:1980:GSB

[432] S. K. Bose. Gauge symmetry and its breakdown: The example of a BCS superconductor. Journal of Mathematical Physics, 21(12):2839– 2841, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i12/p2839_s1.

Stokes:1980:PVR

[433] Vijay K. Stokes. On the proper vectors of real third-order matrices. Journal of Mathematical Physics, 21(12):2842–2843, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v21/i12/p2842_s1.

Wannier:1980:BEM

[434] G. H. Wannier. Bloch electrons in a magnetic field — reduction to one dimension. Journal of Mathematical Physics, 21(12):2844–2846, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v21/i12/p2844_s1.

Jarem:1980:QFC

[435] John M. Jarem. Quality factor constrained, antenna pattern synthesis using a restricted class of aperture functions. Journal of Mathematical Physics, 21(12):2847–2851, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i12/p2847_s1.

CastilhoAlcaras:1980:ECS

[436] J. A. Castilho Alcarás and L. C. Biedenharn. Erratum: “On the coupling of self-conjugate systems with \overline{SL(3, R)} symmetry”, [J. Math. Phys. 20 (1979), no. 8, 1615–1621; MR 80i:81069]. Journal of Mathematical Physics, 21(12):2852, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v21/i12/p2852_s1. See [12].

Kesarwani:1980:EFT

[437] R. N. Kesarwani, Y. P. Varshni, Y. Ohnuki, S. Kamefuchi, Y. Ohnuki, and S. Kamefuchi. Erratum: “Five-term WKBJ approximation”, [J. Math. Phys. 21, 90 (1980)]. Journal of Mathematical Physics, 21(12): 2852, December 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i12/p2852_s2. See [35].

Chen:1981:IFI

[438] Jin Quan Chen. SU(mn) ⊇ SU(m) × SU(n) isoscalar factors and S(f_{1} + f_{2}) ⊇ S(f_{1})×S(f_{2}) isoscalar factors. Journal of Mathematical Physics, 22 (1):1–6, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i1/p1_s1.

Damhus:1981:ERC

[439] Ture Damhus. On the existence of real Clebsch–Gordan coefficients. Journal of Mathematical Physics, 22(1):7–14, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/p7_s1.

Gould:1981:MEG

[440] M. D. Gould. On the matrix elements of the U(n) generators. Journal of Mathematical Physics, 22(1):15–22, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i1/p15_s1.

Calogero:1981:RTM

[441] F. Calogero and A. Degasperis. Reduction technique for matrix nonlinear evolution equations solvable by the spectral transform. Journal of Mathematical Physics, 22(1):23–31, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i1/p23_s1.

Verbitskii:1981:OMS

[442] I. L. Verbitskiĭ. On one method of solving the Helmholtz equation. I. Journal of Mathematical Physics, 22(1):32–38, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/p32_s1.

Samaddar:1981:BWT

[443] S. N. Samaddar. A Bessel–Watson transform pair. Journal of Mathematical Physics, 22(1):39–41, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i1/p39_s1.

Kalnins:1981:NSC

[444] E. G. Kalnins and W. Miller, Jr. Nonorthogonal R-separable coordinates for four-dimensional complex Riemannian spaces. Journal of Mathematical Physics, 22(1):42–50, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i1/p42_s1.

Augusteijn:1981:SCA

[445] Marijke F. Augusteijn and Ernst Breitenberger. Stability of constantamplitude motions in slow-fluctuation approximation. Journal of Mathematical Physics, 22(1):51–66, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i1/p51_s1.

Blood:1981:IVQ

[446] F. A. Blood, Jr. Independent variables in quantum mechanics. Journal of Mathematical Physics, 22(1):67–77, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i1/p67_s1.

Sharma:1981:RPP

[447] J. K. Sharma, C. L. Mehta, N. Mukunda, and E. C. G. Sudarshan. Representation and properties of para-Bose oscillator operators. II. Coherent states and the minimum uncertainty states. Journal of Mathematical Physics, 22(1):78–90, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i1/p78_s1.

Ray:1981:ESV

[448] John R. Ray and James L. Reid. Ermakov systems, velocity-dependent potentials, and nonlinear superposition. Journal of Mathematical Physics, 22(1):91–95, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i1/p91_s1.

Ringwood:1981:RWQ

[449] G. A. Ringwood. Random walks and quantum currents in networks. Journal of Mathematical Physics, 22(1):96–101, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/p96_s1.

Mizrahi:1981:WAP

[450] Maurice M. Mizrahi. On the WKB approximation to the propagator for arbitrary Hamiltonians. Journal of Mathematical Physics, 22(1):102– 107, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i1/p102_s1.

vanHaeringen:1981:DPN

[451] H. van Haeringen and L. P. Kok. Degenerate perturbations in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 22(1):108– 115, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i1/p108_s1.

Henneberger:1981:ABS

[452] Walter C. Henneberger. Aharonov–Bohm scattering and the velocity operator. Journal of Mathematical Physics, 22(1):116–117, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/
p116_s1.

Cosenza:1981:SMA

[453] M. Cosenza, L. Herrera, M. Esculpi, and L. Witten. Some models of anisotropic spheres in general relativity. Journal of Mathematical Physics, 22(1):118–125, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i1/p118_s1.

Fennelly:1981:OFR

[454] A. J. Fennelly. Observer frame rotation rates and magnetic fields in spatially homogeneous universes. Journal of Mathematical Physics, 22(1):126–132, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i1/p126_s1.

Yamazaki:1981:HKX

[455] Masatoshi Yamazaki. On the Hoenselaers–Kinnersley–Xanthopoulos spinning mass fields. Journal of Mathematical Physics, 22(1):133–135, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i1/p133_s1.

Singh:1981:SAS

[456] T. Singh and L. N. Rai. On stationary axially symmetric Einstein– Maxwell scalar and Brans–Dicke–Maxwell fields. Journal of Mathematical Physics, 22(1):136–141, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i1/p136_s1.

Castagnino:1981:QEP

[457] Mario Castagnino and Ricardo Weder. The quantum equivalence principle and finite particle creation in expanding universes. Journal of Mathematical Physics, 22(1):142–151, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i1/p142_s1.

vandenBerg:1981:PTP

[458] J. van den Berg. Percolation theory on pairs of matching lattices. Journal of Mathematical Physics, 22(1):152–157, January 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i1/p152_s1.

Larsen:1981:SIP

[459] Edward W. Larsen. Solution of the inverse problem in multigroup transport theory. Journal of Mathematical Physics, 22(1):158–160, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i1/p158_s1.

Zeleznik:1981:TIE

[460] Frank J. Zeleznik. Thermodynamics. II. The extended thermodynamic system. Journal of Mathematical Physics, 22(1):161–178, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/
p161_s1.

Asorey:1981:SRC

[461] M. Asorey. Some remarks on the classical vacuum structure of gauge field theories. Journal of Mathematical Physics, 22(1):179–184, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v22/i1/p179_s1. See erratum [1720].

Kroger:1981:HSA

[462] H. Kröger. Hilbert space approach to the three- and four-nucleon systems. Journal of Mathematical Physics, 22(1):185–198, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/
p185_s1.

McCormick:1981:IPT

[463] N. J. McCormick and R. Sanchez. Inverse problem transport calculations for anisotropic scattering coefficients. Journal of Mathematical Physics, 22(1):199–208, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i1/p199_s1.

Kriegsmann:1981:SPE

[464] Gregory A. Kriegsmann. On the scattering of plane electromagnetic waves off spherically confined cold plasmas with overdense and steep densities. Journal of Mathematical Physics, 22(1):209–214, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i1/p209_s1.

Hiscock:1981:TCS

[465] William A. Hiscock. On the topology of charged spherical collapse. Journal of Mathematical Physics, 22(1):215–218, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i1/p215_s1.

Cornille:1981:ESN

[466] H. Cornille. Erratum: “Solutions of the nonlinear 3-wave equations in three spatial dimensions”, [J. Math. Phys. 20 (1979), no. 8, 1653–1666; MR 81d:35069]. Journal of Mathematical Physics, 22(1):219, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i1/p219_s1. See [13].

Garbaczewski:1981:ECS

[467] Piotr Garbaczewski. Erratum: “A concept of spin 1/2 approximation in the quantum theory of lattice Bose systems”, [J. Math. Phys. 21 (1980), no. 11, 2670–2675; MR 81j:82027]. Journal of Mathematical Physics, 22 (1):219, January 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i1/p219_s2. See [404].

Naudts:1981:GSA

[468] J. Naudts. Gaussian states on the algebra of the infinite classical system. Journal of Mathematical Physics, 22(2):221–225, February 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/p221_
s1.

Salingaros:1981:REC

[469] Nikos Salingaros. Realization, extension, and classification of certain physically important groups and algebras. Journal of Mathematical Physics, 22(2):226–232, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i2/p226_s1.

Newmarch:1981:RAGa

[470] J. D. Newmarch and R. M. Golding. The Racah algebra for groups with time reversal symmetry. Journal of Mathematical Physics, 22(2):233– 244, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i2/p233_s1.

Hughes:1981:RMR

[471] J. W. B. Hughes. Representations of Osp(2, 1) and the metaplectic representation. Journal of Mathematical Physics, 22(2):245–250, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p245_s1.

Miller:1981:SLG

[472] W. Miller, Jr., J. Patera, and P. Winternitz. Subgroups of Lie groups and separation of variables. Journal of Mathematical Physics, 22(2):251– 260, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i2/p251_s1.

Patera:1981:GFP

[473] J. Patera and R. T. Sharp. Generating functions for SU(2) plethysms with fixed exchange symmetry. Journal of Mathematical Physics, 22 (2):261–266, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i2/p261_s1.

Gould:1981:GRL

[474] M. D. Gould. General U(n) raising and lowering operators. Journal of Mathematical Physics, 22(2):267–270, February 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i2/p267_s1.

Rashid:1981:EFA

[475] M. A. Rashid. Expansion of a function about a displaced center. Journal of Mathematical Physics, 22(2):271–274, February 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i2/p271_s1.

Ramm:1981:SPS

[476] A. G. Ramm. On some properties of solutions of Helmholtz equation. Journal of Mathematical Physics, 22(2):275–276, February 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/p275_
s1.

Honerkamp:1981:GEE

[477] J. Honerkamp. Gauge equivalence of exactly integrable field theoretic models. Journal of Mathematical Physics, 22(2):277–281, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/
p277_s1.

Rockmore:1981:IFF

[478] R. Rockmore. Integrals over the Fermi function. Journal of Mathematical Physics, 22(2):282–283, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i2/p282_s1.

Ochs:1981:SAP

[479] Wilhelm Ochs. The set of all projective limits of a projective system of state operators. Journal of Mathematical Physics, 22(2):284–289, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p284_s1.

Sen:1981:DDB

[480] R. N. Sen. The domain of definition of bundle representations. Journal of Mathematical Physics, 22(2):290–292, February 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i2/p290_s1.

Schutz:1981:ASM

[481] John W. Schutz. An axiomatic system for Minkowski space–time. Journal of Mathematical Physics, 22(2):293–302, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/p293_s1.

Grabowski:1981:SFW

[482] Marian Grabowski. The spreading of free wave packets and the entropy of position. Journal of Mathematical Physics, 22(2):303–305, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p303_s1.

Blinder:1981:NCG

[483] S. M. Blinder. Nonrelativistic Coulomb Green’s function in parabolic coordinates. Journal of Mathematical Physics, 22(2):306–311, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p306_s1.

Gesztesy:1981:ASP

[484] F. Gesztesy and C. B. Lang. On the Abel summability of partial wave amplitudes for Coulomb-type interactions. Journal of Mathematical Physics, 22(2):312–319, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i2/p312_s1.

Dolzycki:1981:SEB

[485] Stanislaw J. Dolzycki. Source equivalency based on the energy specification of order l. Journal of Mathematical Physics, 22(2):320–324, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p320_s1.

Dolzycki:1981:LDA

[486] Stanislaw J. Dolzycki. On the Lorentz dipole approximation in static electrodynamics. Journal of Mathematical Physics, 22(2):325–330, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p325_s1.

Sparling:1981:ES

[487] G. A. J. Sparling and K. P. Tod. An example of an H-space. Journal of Mathematical Physics, 22(2):331–332, February 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i2/p331_s1.

Sobczyk:1981:STA

[488] G. Sobczyk. Space–time algebra approach to curvature. Journal of Mathematical Physics, 22(2):333–342, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i2/p333_s1.

Saraykar:1981:LSE

[489] R. V. Saraykar and N. E. Joshi. Linearization stability of Einstein equations coupled with self-gravitating scalar fields. Journal of Mathematical Physics, 22(2):343–347, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i2/p343_s1. See erratum [1162].

Cally:1981:SMG

[490] P. Cally and J. J. Monaghan. Statistical mechanics and the gravothermal catastrophe. Journal of Mathematical Physics, 22(2):348–351, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p348_s1.

Giannoni:1981:DTF

[491] M. J. Giannoni. A decomposition theorem for finite-dimensional Hermitian projections. Journal of Mathematical Physics, 22(2):352–356, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p352_s1.

Foussats:1981:MSF

[492] Adriana Foussats, Roberto Laurá, and Oscar Zandron. Massive spin-1 field in expanding universes and the quantum equivalence principle. Journal of Mathematical Physics, 22(2):357–370, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/p357_s1.

Baumann:1981:WPM

[493] Klaus Baumann. Wick polynomials with a mass gap are not infinitely divisible. Journal of Mathematical Physics, 22(2):371–373, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p371_s1.

Eichenherr:1981:RNS

[494] H. Eichenherr and J. Honerkamp. Reduction of the CP^{N} nonlinear sigma model. Journal of Mathematical Physics, 22(2):374–376, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p374_s1.

Talukdar:1981:QTS

[495] B. Talukdar, M. Sen, and D. Sen. Quantum tachyons in Schwarzschild space–time. Journal of Mathematical Physics, 22(2):377–379, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p377_s1.

Chaichian:1981:NEH

[496] Masud Chaichian, Jan Fischer, S. M. Molodensky, N. F. Nelipa, and Ivo Vrkoč. A new expression for the high-energy upper bound on the scattering amplitude. Journal of Mathematical Physics, 22(2):380–384, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p380_s1.

Johnson:1981:SUE

[497] Porter W. Johnson and Robert Lee Warnock. Solution of the unitarity equation with overlapping left and right cuts: A tool for study of the S^{∗} and similar systems. Journal of Mathematical Physics, 22(2):385– 399, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i2/p385_s1.

Zhu:1981:RMG

[498] Dong Pei Zhu. Representation matching and the gauge invariant energymomentum tensor. Journal of Mathematical Physics, 22(2):400–402, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i2/p400_s1.

Roman:1981:RQD

[499] P. Roman and J. Haavisto. Relativistic quantum dynamical group for hadrons. Journal of Mathematical Physics, 22(2):403–411, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p403_s1.

Quiros:1981:RPG

[500] M. Quirós and J. Ramírez Mittelbrunn. Representations of the Poincaré group in a K_{3}-basis and the infinite momentum limit. Journal of Mathematical Physics, 22(2):412–421, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i2/p412_s1.

Tucker:1981:ATG

[501] R. W. Tucker. Affine transformations and the geometry of superspace. Journal of Mathematical Physics, 22(2):422–429, February 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i2/p422_
s1.

Stacey:1981:TNIa

[502] Richard Stacey. Techniques for naturality isolation: I. with a Bosonic t channel. Journal of Mathematical Physics, 22(2):430–434, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p430_s1.

Stacey:1981:TNIb

[503] Richard Stacey. Techniques for naturality isolation. II. with a Fermionic t channel. Journal of Mathematical Physics, 22(2):435–441, February 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i2/p435_s1.

Rogers:1981:SEC

[504] Alice Rogers. Some examples of compact supermanifolds with non-Abelian fundamental group. Journal of Mathematical Physics, 22(3): 443–444, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p443_s1.

Abellanas:1981:CDN

[505] L. Abellanas and A. Galindo. Conserved densities for nonlinear evolution equations. II. Odd order case. Journal of Mathematical Physics, 22(3): 445–448, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p445_s1.

Kupershmidt:1981:NGT

[506] B. A. Kupershmidt. On the nature of the Gardner transformation. Journal of Mathematical Physics, 22(3):449–451, March 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i3/p449_s1.

Banuelos:1981:AEF

[507] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. Analytical expansions for Fermi–Dirac functions. Journal of Mathematical Physics, 22(3):452–455, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i3/p452_s1.

Kedem:1981:NND

[508] Benjamin Kedem. On nearest neighbor degeneracies of indistinguishable particles. Journal of Mathematical Physics, 22(3):456–461, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i3/p456_s1.

Masson:1981:MPA

[509] D. Masson. Moment problem approximants to the Chandrasekhar Hequation. Journal of Mathematical Physics, 22(3):462–464, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/
p462_s1.

Leach:1981:EIC

[510] P. G. L. Leach. An exact invariant for a class of time-dependent anharmonic oscillators with cubic anharmonicity. Journal of Mathematical Physics, 22(3):465–470, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p465_s1.

Nakayama:1981:SSW

[511] J. Nakayama, H. Ogura, and M. Sakata. Scattering of a scalar wave from a slightly random surface. Journal of Mathematical Physics, 22(3): 471–477, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p471_s1.

Cohn:1981:CED

[512] J. Cohn and N. Hong. Covariant electrodynamics of a dyon in a medium. Journal of Mathematical Physics, 22(3):478–483, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/p478_s1.

Frahm:1981:MFM

[513] Charles P. Frahm. A multipole formalism for material media. Journal of Mathematical Physics, 22(3):484–488, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i3/p484_s1.

Rosen:1981:EPN

[514] Gerald Rosen. Energy-particle-number inequality in nonlinear complexscalar field theory. Journal of Mathematical Physics, 22(3):489–490, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p489_s1.

Shucker:1981:MSM

[515] David S. Shucker. Measurement in stochastic mechanics. Journal of Mathematical Physics, 22(3):491–494, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i3/p491_s1.

Benioff:1981:QMH

[516] Paul Benioff. Quantum mechanical Hamiltonian models of discrete processes. Journal of Mathematical Physics, 22(3):495–507, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/
p495_s1.

Bhaumik:1981:LMC

[517] Debajyoti Bhaumik, Ajoy Choudhury, Mira De, and Binayak Dutta Roy. The Lipkin model and coherent states. Journal of Mathematical Physics, 22(3):508–510, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p508_s1.

Zinn-Justin:1981:EAI

[518] J. Zinn-Justin. Expansion around instantons in quantum mechanics. Journal of Mathematical Physics, 22(3):511–520, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/p511_s1.

Perelmuter:1981:DFS

[519] M. A. Perel′muter and Ya. A. Semenov. On decoupling of finite singularities in the scattering theory for the Schrödinger operator with a magnetic field. Journal of Mathematical Physics, 22(3):521–533, March 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/p521_
s1.

Arai:1981:SAS

[520] Asao Arai. Self-adjointness and spectrum of Hamiltonians in nonrelativistic quantum electrodynamics. Journal of Mathematical Physics, 22(3): 534–537, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p534_s1.

Maia:1981:CST

[521] M. D. Maia. Contacts of space–times. Journal of Mathematical Physics, 22(3):538–542, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p538_s1.

Szafron:1981:IIG

[522] D. A. Szafron. Intrinsic isometry groups in general relativity. Journal of Mathematical Physics, 22(3):543–549, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i3/p543_s1.

Ross:1981:FBS

[523] D. K. Ross. Fiber bundles, superspace, and the gravitational field. Journal of Mathematical Physics, 22(3):549–556, March 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i3/p549_s1.

Ross:1981:GGF

[524] D. K. Ross. Gauge groups in a fiber bundle approach to superspace and the gravitational field. Journal of Mathematical Physics, 22(3):557– 561, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p557_s1.

Weiss:1981:NLR

[525] George H. Weiss. A note on lattice random walks with an excluded point. Journal of Mathematical Physics, 22(3):562–563, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/p562_s1.

Bahurmuz:1981:LGF

[526] A. A. Bahurmuz and P. D. Loly. Lattice Green’s functions for a linear chain with next nearest neighbor interactions. Journal of Mathematical Physics, 22(3):564–568, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p564_s1.

Muraskin:1981:AFT

[527] M. Muraskin. Aesthetic field theory — extended particles in fourdimensional space–time. Journal of Mathematical Physics, 22(3):569– 571, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p569_s1.

Manoukian:1981:DTF

[528] Edward B. Manoukian. On the decoupling theorem in field theory. Journal of Mathematical Physics, 22(3):572–573, March 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i3/p572_s1.

Garbaczewski:1981:QSTa

[529] Piotr Garbaczewski. On quantum solitons and their classical relatives: Spin 1/2 approximation of the sine-Gordon system. Journal of Mathematical Physics, 22(3):574–580, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i3/p574_s1.

Hietarinta:1981:PMA

[530] Jarmo Hietarinta and Dennis Nemeschansky. Perturbation methods for the (almost) zero modes and the Green’s function in an instanton gas background. Journal of Mathematical Physics, 22(3):581–584, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i3/p581_s1.

Oh:1981:LSF

[531] C. H. Oh and Rosy Teh. Linearly superposable and finite-action classical Yang–Mills fields. Journal of Mathematical Physics, 22(3):585– 589, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i3/p585_s1.

Ghikas:1981:AFS

[532] Demetris P. K. Ghikas. An axiomatic framework for supersymmetries. Journal of Mathematical Physics, 22(3):590–599, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i3/p590_s1.

Sakmar:1981:UBS

[533] I. A. Sakmar. Upper bound on the spin-flip cross section from unitarity, total cross section and the forward slope. Journal of Mathematical Physics, 22(3):600–604, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p600_s1.

Chacon:1981:CMM

[534] E. Chacón, M. Moshinsky, and V. Vanagas. Confrontation of macroscopic and microscopic nuclear collective models. Journal of Mathematical Physics, 22(3):605–622, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i3/p605_s1.

Obregon:1981:BTI

[535] Octavio Obregón and Michael P. Ryan, Jr. Bianchi type IX cosmological models with homogeneous spinor fields. Journal of Mathematical Physics, 22(3):623–630, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i3/p623_s1.

Nash:1981:ETF

[536] Patrick L. Nash. Erratum: “Towards a factorization of M_{4}”, [J. Math. Phys. 21 (1980), no. 5, 1024–1031; MR 83a:81027]. Journal of Mathematical Physics, 22(3):631, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i3/p631_s1. See [171].

Newton:1981:EIS

[537] Roger G. Newton. Erratum: “Inverse scattering. II. Three dimensions”, [J. Math. Phys. 21 (1980), no. 7, 1698–1715; MR 81m:81066b]. Journal of Mathematical Physics, 22(3):631, March 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i3/p631_s2. See [268, 1005].

Wongtawatnugool:1981:EPI

[538] C. Wongtawatnugool, S. Y. Wu, and C. C. Shih. Eigenvalue problem for an infinite tridiagonal matrix. Journal of Mathematical Physics, 22(4): 633–636, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p633_s1.

Fischler:1981:YTM

[539] Mark Fischler. Young-tableau methods for Kronecker products of representations of the classical groups. Journal of Mathematical Physics, 22(4):637–648, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p637_s1.

Brenig:1981:AAI

[540] L. Brenig and V. Fairén. Analytical approach to initial-value problems in nonlinear systems. Journal of Mathematical Physics, 22(4):649–652, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i4/p649_s1.

Glasser:1981:LTA

[541] M. L. Glasser. Laplace transforms and asymptotic expansions of orthogonal polynomials. Journal of Mathematical Physics, 22(4):653–655, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i4/p653_s1.

Behroozi:1981:PEB

[542] F. Behroozi and R. Gagnon. Perimeter expansion in the n-bug system and its relationship to stability. Journal of Mathematical Physics, 22(4):

656–662, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p656_s1.

Krisch:1981:KSE

[543] J. P. Krisch. On the Killing surface-event horizon relation. Journal of Mathematical Physics, 22(4):663–666, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p663_s1.

Finley:1981:AAD

[544] J. D. Finley III and J. F. Plebański. All algebraically degenerate H spaces, via HH spaces. Journal of Mathematical Physics, 22(4):667– 674, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p667_s1.

Gowdy:1981:WEA

[545] Robert H. Gowdy. The wave equation in asymptotically retarded time coordinates: Waves as simple, regular functions on a compact manifold. Journal of Mathematical Physics, 22(4):675–678, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p675_s1.

Leach:1981:FNH

[546] P. G. L. Leach. A further note on the Hénon–Heiles problem. Journal of Mathematical Physics, 22(4):679–682, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p679_s1.

Sharma:1981:FCB

[547] V. D. Sharma and V. V. Menon. Further comments on the behavior of acceleration waves of arbitrary shape. Journal of Mathematical Physics, 22(4):683–684, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p683_s1.

Willemsen:1981:FMR

[548] Jorge F. Willemsen. Functional methods in random classical field theory. Journal of Mathematical Physics, 22(4):685–689, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p685_s1.

Gessner:1981:MLS

[549] W. Gessner. The maximality of Lorentz spaces and their importance with respect to classical and quantum electrodynamics. Journal of Mathematical Physics, 22(4):690–704, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i4/p690_s1.

Kaiser:1981:PSA

[550] Gerald Kaiser. Phase-space approach to relativistic quantum mechanics. III. Quantization, relativity, localization and gauge freedom. Journal of Mathematical Physics, 22(4):705–714, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p705_s1.

Bernido:1981:PIP

[551] Christopher C. Bernido and Akira Inomata. Path integrals with a periodic constraint: The Aharonov–Bohm effect. Journal of Mathematical Physics, 22(4):715–718, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p715_s1.

Bracken:1981:NSC

[552] A. J. Bracken and H. I. Leemon. A new set of coherent states for the isotropic harmonic oscillator: coherent, angular momentum states. Journal of Mathematical Physics, 22(4):719–732, April 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i4/p719_s1.

Muller:1981:IST

[553] R. Müller and H. J. W. Müller-Kirsten. Iteration of single- and twochannel Schrödinger equations. Journal of Mathematical Physics, 22(4): 733–749, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p733_s1.

Wiesler:1981:DBH

[554] Andreas Wiesler. Derivation of “Bethe’s hypothesis” from the quantum inverse scattering transformation for the nonlinear Schrödinger equation. Journal of Mathematical Physics, 22(4):750–753, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p750_s1.

Fernandez:1981:HRO

[555] Francisco M. Fernández and Eduardo A. Castro. Hypervirial relations and orthogonalization conditions. Journal of Mathematical Physics, 22(4):754–756, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p754_s1.

Kuratsuji:1981:STP

[556] Hiroshi Kuratsuji and Yutaka Mizobuchi. A semiclassical treatment of path integrals for the spin system. Journal of Mathematical Physics, 22(4):757–764, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p757_s1.

Moore:1981:SAP

[557] Steven M. Moore. Semiclassical approximations at positive temperatures in stochastic physics. Journal of Mathematical Physics, 22(4): 765–769, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p765_s1.

Battle:1981:IVL

[558] Guy A. Battle and Lon Rosen. On the infinite volume limit of the strongly coupled Yukawa_{2} model. Journal of Mathematical Physics, 22(4):770– 776, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p770_s1.

Swanson:1981:PIF

[559] Mark S. Swanson. The path integral formulation of the dynamical map. Journal of Mathematical Physics, 22(4):777–784, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p777_s1.

Tang:1981:BET

[560] Henry H. K. Tang and Jan S. Vaagen. Bipolar expansion of tensor fields: Recurrence relations and analytic construction of form factors. Journal of Mathematical Physics, 22(4):785–793, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p785_s1.

Fanchi:1981:SFF

[561] John R. Fanchi. 4-space formulation of field equations for multicomponent eigenfunctions. Journal of Mathematical Physics, 22(4):794–797, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i4/p794_s1.

Polyzou:1981:IDU

[562] Wayne Polyzou. Interior dynamics in the unified connected reaction theory of Polyzou and Redish. Journal of Mathematical Physics, 22(4): 798–802, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p798_s1.

Geroch:1981:LGR

[563] Robert Geroch and Jeffrey Winicour. Linkages in general relativity. Journal of Mathematical Physics, 22(4):803–812, April 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i4/p803_s1.

Frankel:1981:EDS

[564] Theodore Frankel and Gregory J. Galloway. Energy density and spatial curvature in general relativity. Journal of Mathematical Physics, 22(4): 813–817, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p813_s1.

Ludvigsen:1981:AFS

[565] M. Ludvigsen, E. T. Newman, and K. P. Tod. Asymptotically flat Hspaces. Journal of Mathematical Physics, 22(4):818–823, April 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p818_
s1.

Banerjee:1981:SCP

[566] A. Banerjee and N. O. Santos. Static charged perfect fluid in a conformally flat spacetime. Journal of Mathematical Physics, 22(4):824– 826, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p824_s1.

Letelier:1981:SGG

[567] P. S. Letelier and R. Machado. Self-graviting [-gravitating] anisotropic fluids with plane symmetry. Journal of Mathematical Physics, 22(4): 827–829, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p827_s1.

Moschetti:1981:HSE

[568] Gaetano Moschetti. Homothetic solutions of Einstein’s equations and shock waves. Journal of Mathematical Physics, 22(4):830–834, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i4/p830_s1.

Marmo:1981:LDP

[569] G. Marmo, E. J. Saletan, A. Simoni, and F. Zaccaria. Liouville dynamics and Poisson brackets. Journal of Mathematical Physics, 22(4): 835–842, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p835_s1.

Daniels:1981:PEC

[570] H. A. M. Daniëls. Passivity and equilibrium for classical Hamiltonian systems. Journal of Mathematical Physics, 22(4):843–846, April 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/p843_
s1.

Sanchez:1981:GSI

[571] R. Sanchez and N. J. McCormick. General solutions to inverse transport problems. Journal of Mathematical Physics, 22(4):847–855, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i4/p847_s1.

Larsen:1981:LTE

[572] Edward W. Larsen and Thomas W. Mullikin. Linear transport in an exponential atmosphere. Journal of Mathematical Physics, 22(4):856– 861, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p856_s1.

McKellar:1981:LGF

[573] R. J. McKellar. Local gauge field theory formalism. Journal of Mathematical Physics, 22(4):862–869, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i4/p862_s1.

Cant:1981:RNLa

[574] A. Cant. The role of noncompact Lie algebras in relativistic wave equations. I. Journal of Mathematical Physics, 22(4):870–877, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/
p870_s1.

Cant:1981:RNLb

[575] A. Cant. The role of noncompact Lie algebras in relativistic wave equations. II. Journal of Mathematical Physics, 22(4):878–882, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i4/
p878_s1.

Bolle:1981:SRC

[576] D. Bollé and T. A. Osborn. Sum rules in classical scattering. Journal of Mathematical Physics, 22(4):883–892, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p883_s1.

Singh:1981:CLB

[577] S. R. Singh. Converging lower bounds to atomic binding energies. Journal of Mathematical Physics, 22(4):893–896, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i4/p893_s1.

Hillion:1981:TBC

[578] P. Hillion and S. Quinnez. Thermal blooming calculations with analytical diffraction approximated expressions. Journal of Mathematical Physics, 22(4):897–907, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p897_s1.

Beyer:1981:ANI

[579] W. A. Beyer, G. M. Wing, and A. Migliori. Analysis of a nonlinear integral equation arising in the study of the magnetic field in the critical state model of superconductivity. Journal of Mathematical Physics, 22(4): 908–913, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p908_s1.

Jedrzejewski:1981:SMO

[580] J. Jedrzejewski and A. Verbeure. On the spontaneous magnetic order in the Hubbard model. Journal of Mathematical Physics, 22(4):914– 915, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i4/p914_s1.

Kim:1981:ENM

[581] Shoon K. Kim. Erratum: “A new method of matrix transformations. II. General theory of matrix diagonalizations via reduced characteristic equations and its application to angular momentum coupling”, [J. Math. Phys. 20 (1979), no. 11, 2159–2167; MR 81j:15015b]. Journal of Mathematical Physics, 22(4):916, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i4/p916_s1. See [16].

Calogero:1981:MDO

[582] F. Calogero. Matrices, differential operators, and polynomials. Journal of Mathematical Physics, 22(5):919–934, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i5/p919_s1.

Wong:1981:EER

[583] M. K. F. Wong and Hsin Yang Yeh. Explicit evaluation of the representation functions of IU(n). Journal of Mathematical Physics, 22(5): 935–938, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p935_s1.

Rogers:1981:SLG

[584] Alice Rogers. Super Lie groups: global topology and local structure. Journal of Mathematical Physics, 22(5):939–945, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p939_s1.

Basu:1981:LGO

[585] Debabrata Basu and D. Mitra. The Lorentz group in the oscillator realization. III. The group SO(3, 1). Journal of Mathematical Physics, 22(5):

946–953, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p946_s1.

Beals:1981:PRC

[586] Richard Beals. Partial-range completeness and existence of solutions to two-way diffusion equations. Journal of Mathematical Physics, 22(5): 954–960, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p954_s1. See erratum [1566].

Omote:1981:BTE

[587] M. Omote and M. Wadati. Bäcklund transformations for the Ernst equation. Journal of Mathematical Physics, 22(5):961–964, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p961_
s1.

Leon:1981:ISC

[588] Jérôme Jp. Leon. Inverse scattering connections. Journal of Mathematical Physics, 22(5):965–969, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i5/p965_s1.

Steeb:1981:SDE

[589] W.-H. Steeb, W. Erig, and W. Strampp. Symmetries and the Dirac equation. Journal of Mathematical Physics, 22(5):970–973, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p970_
s1.

Epifanio:1981:SSP

[590] G. Epifanio and C. Trapani. Some spectral properties in algebras of unbounded operators. Journal of Mathematical Physics, 22(5):974–978, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i5/p974_s1.

Matsushita:1981:ECC

[591] Yasuo Matsushita. On Euler characteristics of compact Einstein 4manifolds of metric signature (+ + −−). Journal of Mathematical Physics, 22(5):979–982, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i5/p979_s1. See erratum [1129].

Nash:1981:SCO

[592] Patrick L. Nash. On the substructure of the classical observables x^{α}, p_{α}, and J^{αβ}. Journal of Mathematical Physics, 22(5):983–987, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i5/p983_s1.

Gowdy:1981:GSP

[593] Robert H. Gowdy. Geometrical spacetime perturbation theory: Regular higher order structures. Journal of Mathematical Physics, 22(5): 988–994, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p988_s1.

Ishikawa:1981:NFR

[594] Hirohisa Ishikawa. Note on Finslerian relativity. Journal of Mathematical Physics, 22(5):995–1004, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i5/p995_s1.

Strauss:1981:ELS

[595] Walter A. Strauss and Luis Vázquez. Existence of localized solutions for certain model field theories. Journal of Mathematical Physics, 22(5): 1005–1009, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p1005_s1.

Yasue:1981:QMS

[596] Kunio Yasue. Quantum mechanics and stochastic control theory. Journal of Mathematical Physics, 22(5):1010–1020, May 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1010_s1.

Phares:1981:MECa

[597] Alain J. Phares and Robert J. Meier, Jr. A multidimensional extension of the combinatorics function technique. I. Linear and homogeneous partial difference equations. Journal of Mathematical Physics, 22(5): 1021–1029, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i5/p1021_s1.

Phares:1981:MECb

[598] Alain J. Phares. A multidimensional extension of the combinatorics function technique. II. Linear and inhomogeneous partial difference equations. Journal of Mathematical Physics, 22(5):1030–1032, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1030_
s1.

Kovalenko:1981:SOP

[599] V. F. Kovalenko, M. A. Perel′muter, and Ya. A. Semenov. Schrödinger operators with L_{w}^{l/2}(R^{l})-potentials. Journal of Mathematical Physics, 22(5):1033–1044, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i5/p1033_s1.

Das:1981:VPA

[600] U. Das. Variable-phase approach to off-shell scattering by nonlocal potentials. Journal of Mathematical Physics, 22(5):1045–1046, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1045_
s1.

Sakmar:1981:BDP

[601] I. A. Sakmar. Bound on the diffraction peak for the spin 0–spin 1/2 case. Journal of Mathematical Physics, 22(5):1047–1050, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1047_s1.

Hauser:1981:PGC

[602] Isidore Hauser and Frederick J. Ernst. Proof of a Geroch conjecture. Journal of Mathematical Physics, 22(5):1051–1063, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1051_s1.

Pereira:1981:NPT

[603] C. Martin Pereira. Nonglobal proof of the thin-sandwich conjecture. Journal of Mathematical Physics, 22(5):1064–1074, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1064_s1.

Banerjee:1981:CFS

[604] A. Banerjee and N. O. Santos. Conformally flat static space–time in Brans–Dicke theory. Journal of Mathematical Physics, 22(5):1075–1080, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i5/p1075_s1.

Anderson:1981:BCZ

[605] James L. Anderson and Ronald E. Kates. Boundary conditions for zerorest-mass fields in curved space–time: A regularity theorem. Journal of Mathematical Physics, 22(5):1081–1083, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i5/p1081_s1.

Creamer:1981:SME

[606] Dennis B. Creamer, H. B. Thacker, and David Wilkinson. Statistical mechanics of an exactly integrable system. Journal of Mathematical Physics, 22(5):1084–1089, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i5/p1084_s1.

Boon:1981:ANL

[607] M. Boon and J. Zak. Amplitudes on von Neumann lattices. Journal of Mathematical Physics, 22(5):1090–1099, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i5/p1090_s1.

Flato:1981:QFT

[608] M. Flato and C. Fronsdal. Quantum field theory of singletons. The Rac. Journal of Mathematical Physics, 22(5):1100–1105, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1100_s1.

Garczynski:1981:QAF

[609] Włodzimierz Garczyński and Jerzy Stelmach. Quasicontinual approach to a field theory on a lattice: General theory. Journal of Mathematical Physics, 22(5):1106–1113, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i5/p1106_s1.

Laursen:1981:BDC

[610] M. L. Laursen and M. A. Samuel. The n-bubble diagram contribution to g−2. Journal of Mathematical Physics, 22(5):1114–1126, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1114_
s1.

Coen:1981:ISP

[611] Shimon Coen. Inverse scattering of the permittivity and permeability profiles of a plane stratified medium. Journal of Mathematical Physics, 22(5):1127–1129, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i5/p1127_s1.

Tur:1981:PFO

[612] Moshe Tur and Mark J. Beran. Propagation of the fourth-order coherence function in an anisotropic random medium. Journal of Mathematical Physics, 22(5):1130–1136, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i5/p1130_s1.

Antoine:1981:EPI

[613] J.-P. Antoine. Erratum: “Partial inner product spaces. III. Compatibility relations revisited”, [J. Math. Phys. 21 (1980), no. 2, 268–279; MR 81g:46032]. Journal of Mathematical Physics, 22(5):1137, May 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i5/p1137_

s1. See [62].

Aviram:1981:EMC

[614] I. Aviram and S. Goshen. Erratum: “A multigroup criticality condition for a space-independent multiplying assembly via the Chapman– Kolmogorov equation”, [J. Math. Phys. 21, 1897 (1980)]. Journal of Mathematical Physics, 22(5):1137, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i5/p1137_s2. See [291].

Dirl:1981:IPC

[615] R. Dirl. On the irreducibility of projective corepresentations — application to double Shubnikov point groups. Journal of Mathematical Physics, 22(6):1139–1143, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i6/p1139_s1.

Soto:1981:OSG

[616] M. F. Soto, Jr. and R. Mirman. Objects for the symmetric group. Journal of Mathematical Physics, 22(6):1144–1148, June 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1144_s1. See ad-

dendum [895].

Balantekin:1981:DCF

[617] A. Baha Balantekin and Itzhak Bars. Dimension and character formulas for Lie supergroups. Journal of Mathematical Physics, 22(6):1149– 1162, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1149_s1.

Hermann:1981:IHF

[618] Robert Hermann. Infeld–Hull factorization, Galois–Picard–Vessiot theory for differential operators. Journal of Mathematical Physics, 22(6): 1163–1167, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1163_s1.

Hermann:1981:ICZ

[619] Robert Hermann. Isomonodromy as a curvature-zero condition. Journal of Mathematical Physics, 22(6):1168–1169, June 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1168_s1.

Fordy:1981:FOI

[620] Allan P. Fordy and John Gibbons. Factorization of operators. II. Journal of Mathematical Physics, 22(6):1170–1175, June 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1170_s1.

Kaup:1981:LSB

[621] D. J. Kaup. The lump solutions and the Bäcklund transformation for the three-dimensional three-wave resonant interaction. Journal of Mathematical Physics, 22(6):1176–1181, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i6/p1176_s1.

Heiss:1981:PSE

[622] W. D. Heiss. Power series expansion for a crossing-symmetric vertex function. Journal of Mathematical Physics, 22(6):1182–1185, June 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1182_
s1.

Morikawa:1981:COV

[623] Tetsuo Morikawa and Y. J. I’Haya. Construction of orthonormal vector sets for atoms and molecules by means of recursive variation. Journal of Mathematical Physics, 22(6):1186–1189, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1186_s1.

Froman:1981:CHO

[624] Nanny Fröman and Per Olof Fröman. Comments on: “Higher order modified potentials for the effective phase integral approximation”, [J. Math. Phys. 20 (1979), no. 1, 83–85; MR 80f:81031 ] by Edward R. Floyd. Journal of Mathematical Physics, 22(6):1190–1191, June 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1190_s1. See [6].

Nemeth:1981:TSP

[625] G. Németh, Á. Ág, and Gy. Páris. Two-sided Padé approximations for the plasma dispersion function. Journal of Mathematical Physics, 22(6): 1192–1195, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1192_s1.

Breitenberger:1981:EPN

[626] Ernst Breitenberger and Robert D. Mueller. The elastic pendulum: A nonlinear paradigm. Journal of Mathematical Physics, 22(6):1196–1210, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i6/p1196_s1.

Ikeda:1981:CLT

[627] Satoshi Ikeda. On the conservation laws in the theory of fields in Finsler spaces. Journal of Mathematical Physics, 22(6):1211–1214, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i6/
p1211_s1.

Ikeda:1981:TFF

[628] Satoshi Ikeda. On the theory of fields in Finsler spaces. Journal of Mathematical Physics, 22(6):1215–1218, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1215_s1.

Munier:1981:SET

[629] A. Munier, J. R. Burgan, M. Feix, and E. Fijalkow. Schrödinger equation with time-dependent boundary conditions. Journal of Mathematical Physics, 22(6):1219–1223, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i6/p1219_s1.

Tikochinsky:1981:EPM

[630] Yoel Tikochinsky. Exact propagators for multidimensional quadratic Hamiltonians — initial-value approach. Journal of Mathematical Physics, 22(6):1224–1228, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i6/p1224_s1.

Brander:1981:HEB

[631] O. Brander. High-energy behavior of phase shifts for scattering from singular potentials. Journal of Mathematical Physics, 22(6):1229–1235, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i6/p1229_s1.

Kundu:1981:MES

[632] P. Kundu. Multipole expansion of stationary asymptotically flat vacuum metrics in general relativity. Journal of Mathematical Physics, 22(6): 1236–1242, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1236_s1.

Akolia:1981:AC

[633] G. M. Akolia, P. S. Joshi, and U. D. Vyas. On almost causality. Journal of Mathematical Physics, 22(6):1243–1247, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1243_s1.

Lukacs:1981:NKV

[634] B. Lukács, Z. Perjés, and Á. Sebestyén. Null Killing vectors. Journal of Mathematical Physics, 22(6):1248–1253, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1248_s1.

Xanthopoulos:1981:ESR

[635] Basilis C. Xanthopoulos. Exterior spacetimes for rotating stars. Journal of Mathematical Physics, 22(6):1254–1259, June 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1254_s1.

McQuistan:1981:ECN

[636] R. B. McQuistan. Exact composite nearest neighbor degeneracies for one-dimensional lattice spaces. Journal of Mathematical Physics, 22(6): 1260–1263, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1260_s1.

Baker:1981:ADC

[637] George A. Baker, Jr. and Daniel Bessis. Asymptotic decay of correlations in the spin-1/2 Ising model. Journal of Mathematical Physics, 22(6): 1264–1266, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1264_s1.

Mintchev:1981:RPH

[638] Mihail Mintchev and Emilio d’Emilio. On the reconstruction of the physical Hilbert spaces for quantum field theories with indefinite metric. Journal of Mathematical Physics, 22(6):1267–1271, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i6/p1267_s1.

Garbaczewski:1981:QSTb

[639] Piotr Garbaczewski. On quantum solitons and their classical relatives. II. “Fermion-boson reciprocity” and classical vs. quantum problem for the sine-Gordon system. Journal of Mathematical Physics, 22(6):1272– 1279, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1272_s1.

Ahner:1981:DVI

[640] Henry F. Ahner. Dynamically variable internal symmetries. Journal of Mathematical Physics, 22(6):1280–1282, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1280_s1.

Rodriguez:1981:BWE

[641] M. A. Rodríguez and M. Lorente. Bargmann–Wigner equations: Symmetries of multispinors and equations of motion. Journal of Mathematical Physics, 22(6):1283–1288, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i6/p1283_s1.

Szczyrba:1981:MOM

[642] Igor Szczyrba. On the mass operators for multiquark systems. Journal of Mathematical Physics, 22(6):1289–1298, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1289_s1.

Combes:1981:AAG

[643] J. M. Combes, M. Hoffmann-Ostenhof, and T. Hoffmann-Ostenhof. Asymptotics of atomic ground states: The relation between the ground state of helium and the ground state of He^{+}. Journal of Mathematical Physics, 22(6):1299–1305, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i6/p1299_s1.

Ramirez:1981:ESD

[644] Ricardo Ramirez and Miguel Orszag. Exact solution of the Dicke model for the interaction of one atom with the electromagnetic field. Journal of Mathematical Physics, 22(6):1306–1310, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1306_s1.

Morita:1981:ECF

[645] Nagayoshi Morita and Nobuaki Kumagai. The expansion of cylindrical fiber modes in a spherical coordinate. Journal of Mathematical Physics, 22(6):1311–1316, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i6/p1311_s1.

Leuchtag:1981:FDE

[646] H. Richard Leuchtag. A family of differential equations arising from multi-ion electrodiffusion. Journal of Mathematical Physics, 22(6):1317– 1320, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i6/p1317_s1.

King:1981:IIO

[647] Frederick W. King. Integral inequalities for optical constants. Journal of Mathematical Physics, 22(6):1321–1323, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i6/p1321_s1.

Lindblom:1981:SPS

[648] Lee Lindblom. Some properties of static general relativistic stellar models. II. Journal of Mathematical Physics, 22(6):1324–1326, June 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i6/
p1324_s1.

Whitehead:1981:CPC

[649] R. R. Whitehead. Comments on a paper: “Continuum calculus. IV. The Laplace transform method in the evaluation of the Feynman path integrals with a Gaussian measure and applications in quantum mechanics”, [J. Math. Phys. 21 (1980), no. 5, 1055–1069; MR 81f:81018 ] by L. L. Lee. Journal of Mathematical Physics, 22(7):1329, July 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1329_

s1. See [175].

Kennedy:1981:CAD

[650] A. D. Kennedy. Clifford algebras in 2ω dimensions. Journal of Mathematical Physics, 22(7):1330–1337, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i7/p1330_s1.

Moshinsky:1981:SPS

[651] M. Moshinsky and T. H. Seligman. On the structure of phase space. Journal of Mathematical Physics, 22(7):1338–1344, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1338_s1.

Bayen:1981:QS

[652] F. Bayen and C. Fronsdal. Quantization on the sphere. Journal of Mathematical Physics, 22(7):1345–1349, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1345_s1.

Garola:1981:ILA

[653] C. Garola and L. Solombrino. Irreducible linear-antilinear representations and internal symmetries. Journal of Mathematical Physics, 22(7): 1350–1358, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i7/p1350_s1.

Geronimo:1981:MOP

[654] Jeffrey S. Geronimo. Matrix orthogonal polynomials on the unit circle. Journal of Mathematical Physics, 22(7):1359–1365, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1359_s1.

Kerner:1981:UFN

[655] Edward H. Kerner. Universal formats for nonlinear ordinary differential systems. Journal of Mathematical Physics, 22(7):1366–1371, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1366_s1.

Blanchard:1981:TSS

[656] Ph. Blanchard and M. Sirugue. Treatment of some singular potentials by change of variables in Wiener integrals. Journal of Mathematical Physics, 22(7):1372–1376, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i7/p1372_s1.

Costen:1981:JCF

[657] Robert C. Costen. Jump conditions for fields that have infinite, integrable singularities at an interface. Journal of Mathematical Physics, 22(7): 1377–1385, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i7/p1377_s1.

Mathot:1981:DRA

[658] Françoise Mathot. Distributionlike representations of ∗-algebras. Journal of Mathematical Physics, 22(7):1386–1389, July 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1386_s1.

Ray:1981:CSC

[659] Dipankar Ray. A class of solutions of Calapso–Guichard equations. Journal of Mathematical Physics, 22(7):1390–1392, July 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1390_s1.

Gurses:1981:NNE

[660] Metin Gürses and Yavuz Nutku. New nonlinear evolution equations from surface theory. Journal of Mathematical Physics, 22(7):1393–1398, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1393_s1.

Lehman:1981:NEI

[661] D. R. Lehman, William C. Parke, and L. C. Maximon. Numerical evaluation of integrals containing a spherical Bessel function by product integration. Journal of Mathematical Physics, 22(7):1399–1413, July 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1399_
s1.

Hojman:1981:ELM

[662] S. Hojman and H. Harleston. Equivalent Lagrangians: Multidimensional case. Journal of Mathematical Physics, 22(7):1414–1419, July 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1414_
s1.

VilelaMendes:1981:DVF

[663] R. Vilela Mendes and J. Taborda Duarte. Decomposition of vector fields and mixed dynamics. Journal of Mathematical Physics, 22(7):1420–1422, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1420_s1.

DeMeyer:1981:ADT

[664] H. E. De Meyer and G. Vanden Berghe. An alternative derivation of transition amplitudes for time-dependent harmonic oscillators. Journal of Mathematical Physics, 22(7):1423–1424, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1423_s1.

Aldaya:1981:SSB

[665] V. Aldaya and J. A. de Azcárraga. Symmetries of the Schrödinger bundle. Journal of Mathematical Physics, 22(7):1425–1427, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1425_s1.

Au:1981:BSP

[666] C. K. Au and Y. Aharonov. Bound state perturbation theory for the one-space and one-time dimension Klein–Gordon equation. Journal of Mathematical Physics, 22(7):1428–1432, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1428_s1.

Kalikstein:1981:FVP

[667] Kalman Kalikstein. Formulation of variational principles via Lagrange multipliers. Journal of Mathematical Physics, 22(7):1433–1437, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1433_s1.

Sakmar:1981:IRP

[668] I. A. Sakmar. Integrated rotational parameter as a constraint for the variational lower bound of the diffraction peak. Journal of Mathematical Physics, 22(7):1438–1442, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i7/p1438_s1.

Williams:1981:DRC

[669] Gareth Williams. A discussion of the relation “closer than” within simple regions of curved space–time. Journal of Mathematical Physics, 22(7): 1443–1444, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i7/p1443_s1.

Agacy:1981:AGF

[670] R. L. Agacy and L. M. Clendenning. On the analysis of the gravitational field by reduction of the Weyl tensor to canonical form: Algebraically general case. Journal of Mathematical Physics, 22(7):1445–1451, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1445_s1.

Miller:1981:VCC

[671] Bonnie D. Miller. Vacuum counterexamples to the cosmic censorship hypothesis. Journal of Mathematical Physics, 22(7):1452–1456, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i7/p1452_s1.

Couch:1981:SWE

[672] W. E. Couch. Solutions to wave equations on black hole geometries. Journal of Mathematical Physics, 22(7):1457–1462, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1457_s1.

Larsen:1981:LTN

[673] Edward W. Larsen. Linear transport in nonhomogeneous media. I. Journal of Mathematical Physics, 22(7):1463–1466, July 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1463_s1.

Cyranski:1981:ECQ

[674] John F. Cyranski. Entropy in classical and quantum physics. Journal of Mathematical Physics, 22(7):1467–1478, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1467_s1.

Brihaye:1981:YMC

[675] Yves Brihaye. Yang–Mills classical solutions with a remaining U(1) invariance. Journal of Mathematical Physics, 22(7):1479–1481, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i7/
p1479_s1.

Quesne:1981:MEO

[676] C. Quesne. Matrix elements of operators in symmetrical U(6) ⊇ U(3) ⊇ U(2) ⊇ U(1) and U(6) ⊇ SU(3) ⊇ SO(3) ⊇ SO(2) basis. Journal of Mathematical Physics, 22(7):1482–1496, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1482_s1.

Small:1981:PST

[677] Robert D. Small. Paraxial self-trapped beams in nonlinear optics. Journal of Mathematical Physics, 22(7):1497–1503, July 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i7/p1497_s1.

Sinanoglu:1981:TRN

[678] Oktay Sinanoğlu. 1- and 2-topology of reaction networks. Journal of Mathematical Physics, 22(7):1504–1512, July 1981. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i7/p1504_s1.

Pechenick:1981:ISE

[679] Kay R. Pechenick and Jeffrey M. Cohen. Inverse scattering — exact solution of the Gel’fand–Levitan equation. Journal of Mathematical Physics, 22(7):1513–1516, July 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i7/p1513_s1.

Sharma:1981:SPO

[680] C. S. Sharma and U. Egele. Some properties of the operator algebra generated by Hodge’s star and the exterior derivative. Journal of Mathematical Physics, 22(8):1519–1520, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i8/p1519_s1.

Valle:1981:FCS

[681] J. W. F. Valle. Fermionic coherent states in a Fock superspace. Journal of Mathematical Physics, 22(8):1521–1525, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1521_s1.

Moshinsky:1981:CTR

[682] M. Moshinsky and T. H. Seligman. Canonical transformations relating the oscillator and Coulomb problems and their relevance for collective motions. Journal of Mathematical Physics, 22(8):1526–1535, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1526_s1.

Bovier:1981:RCG

[683] A. Bovier, M. Lüling, and D. Wyler. Representations and Clebsch– Gordan coefficients of Z-metacyclic groups. Journal of Mathematical Physics, 22(8):1536–1542, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1536_s1.

Bovier:1981:FS

[684] A. Bovier, M. Lüling, and D. Wyler. Finite subgroups of SU(3). Journal of Mathematical Physics, 22(8):1543–1547, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1543_s1. See er-

ratum [1194].

Carinena:1981:SPR

[685] José F. Cariñena and Mariano Santander. Semiunitary projective representations of the complete Galilei group. Journal of Mathematical Physics, 22(8):1548–1558, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1548_s1.

Wong:1981:UTR

[686] M. K. F. Wong and Hsin Yang Yeh. A unified treatment of the representation functions of SO(n, 1), SO(n + 1), and ISO(n). Journal of Mathematical Physics, 22(8):1559–1565, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1559_s1.

Heidenreich:1981:TPP

[687] W. Heidenreich. Tensor products of positive energy representations of S̃O(3, 2) and S̃O(4, 2). Journal of Mathematical Physics, 22(8):1566– 1574, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1566_s1.

Boyd:1981:SLE

[688] John P. Boyd. Sturm–Liouville eigenproblems with an interior pole. Journal of Mathematical Physics, 22(8):1575–1590, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1575_s1.

Luning:1981:CBI

[689] C. D. Luning and W. L. Perry. Convergence of Berryman’s iterative method for some Emden–Fowler equations. Journal of Mathematical Physics, 22(8):1591–1595, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1591_s1.

Coz:1981:DSM

[690] Marcel Coz. A direct study of a Marchenko fundamental equation with centripetal potential. Journal of Mathematical Physics, 22(8):1596– 1607, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1596_s1.

Nakamura:1981:CBT

[691] Akira Nakamura. Chain of the Bäcklund transformation for the KdV equation. Journal of Mathematical Physics, 22(8):1608–1613, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1608_s1.

Fleming:1981:LVB

[692] Gordon N. Fleming. A limit on the variation of bounded positive operators. Journal of Mathematical Physics, 22(8):1614–1618, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/
p1614_s1.

Pfeifer:1981:AND

[693] Peter Pfeifer. Adjoints of nondensely defined Hilbert space operators. Journal of Mathematical Physics, 22(8):1619–1622, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1619_

s1. See erratum [1084].

Pomraning:1981:CPE

[694] G. C. Pomraning. The criticality problem for an exponential atmosphere. Journal of Mathematical Physics, 22(8):1623–1625, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1623_
s1.

Lutzky:1981:CLD

[695] M. Lutzky. Conservation laws and discrete symmetries in classical mechanics. Journal of Mathematical Physics, 22(8):1626–1627, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1626_s1.

Iwai:1981:RFD

[696] Toshihiro Iwai. On reduction of the four-dimensional harmonic oscillator. Journal of Mathematical Physics, 22(8):1628–1632, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1628_
s1.

Iwai:1981:CKP

[697] Toshihiro Iwai. On a “conformal” Kepler problem and its reduction. Journal of Mathematical Physics, 22(8):1633–1639, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1633_
s1.

Fahnline:1981:COF

[698] Donald E. Fahnline. Classification of orbits of Fokker’s time-asymmetric relativistic two-body problem. Journal of Mathematical Physics, 22 (8):1640–1646, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i8/p1640_s1.

Bojarski:1981:ISI

[699] Norbert N. Bojarski. Inverse scattering inverse source theory. Journal of Mathematical Physics, 22(8):1647–1650, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1647_s1.

Gadella:1981:GWC

[700] M. Gadella and E. Santos. A generalized Weyl correspondence: Applications. Journal of Mathematical Physics, 22(8):1651–1659, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1651_s1.

Steinberg:1981:IIP

[701] Stanly Steinberg and Kurt Bernardo Wolf. Invariant inner products on spaces of solutions of the Klein–Gordon and Helmholtz equations. Journal of Mathematical Physics, 22(8):1660–1663, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1660_s1.

Goldin:1981:RLC

[702] G. A. Goldin, R. Menikoff, and D. H. Sharp. Representations of a local current algebra in nonsimply connected space and the Aharonov– Bohm effect. Journal of Mathematical Physics, 22(8):1664–1668, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1664_s1.

Fernandez:1981:HCE

[703] Francisco M. Fernández and Eduardo A. Castro. Hypervirial calculation of energy eigenvalues of a bounded centrally located harmonic oscillator. Journal of Mathematical Physics, 22(8):1669–1671, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1669_
s1.

Evans:1981:MSA

[704] J. W. Evans. The mathematical structure of arrangement channel quantum mechanics. Journal of Mathematical Physics, 22(8):1672–1686, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1672_s1.

Haba:1981:FRF

[705] Z. Haba. Fluctuations of P (φ) random fields. Journal of Mathematical Physics, 22(8):1687–1691, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1687_s1.

Mattis:1981:SSF

[706] Daniel C. Mattis and Bill Sutherland. Strange solutions to field theories in one spatial dimension. Journal of Mathematical Physics, 22(8):1692– 1695, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1692_s1.

Schendel:1981:RRR

[707] J. Schendel and J. Winicour. Restrictions on relativistically rotating fluids. Journal of Mathematical Physics, 22(8):1696–1702, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/
p1696_s1.

Giachetti:1981:HFG

[708] Riccardo Giachetti, Riccardo Ricci, and Emanuele Sorace. Hamiltonian formulation for the gauge theory of the gravitational coupling. Journal of Mathematical Physics, 22(8):1703–1710, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1703_s1.

Debever:1981:OTI

[709] R. Debever and R. G. McLenaghan. Orthogonal transitivity, invertibility, and null geodesic separability in type D electrovac solutions of Einstein’s field equations with cosmological constant. Journal of Mathematical Physics, 22(8):1711–1726, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i8/p1711_s1.

Moody:1981:AEC

[710] Michael C. Moody and John R. Ray. On the application of energy conditions and algebraic classification in general relativity. Journal of Mathematical Physics, 22(8):1727–1735, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i8/p1727_s1.

Gegenberg:1981:ESS

[711] J. D. Gegenberg and A. Das. An exact stationary solution of the combined Einstein–Maxwell–Klein–Gordon equations. Journal of Mathematical Physics, 22(8):1736–1739, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i8/p1736_s1.

Marchioro:1981:DIA

[712] C. Marchioro, A. Pellegrinotti, and M. Pulvirenti. On the dynamics of infinite anharmonic systems. Journal of Mathematical Physics, 22(8):1740– 1745, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1740_s1.

Maltz:1981:PSO

[713] A. Maltz and E. E. Mola. Pair site occupation probability for onedimensional array of dumbbells. Journal of Mathematical Physics, 22 (8):1746–1748, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i8/p1746_s1.

Hampton:1981:DRL

[714] Richard R. Hampton. Distributional and regulator limits for circuit breaker regularized Feynman amplitudes. Journal of Mathematical Physics, 22(8):1749–1758, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1749_s1.

Krause:1981:RGC

[715] Jorge Krause. Radiation gauge covariance. Journal of Mathematical Physics, 22(8):1759–1766, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1759_s1.

Quiros:1981:GSF

[716] M. Quirós, F. J. de Urríes, J. Hoyos, M. L. Mazón, and E. Rodríguez. Geometrical structure of Faddeev–Popov fields and invariance properties of gauge theories. Journal of Mathematical Physics, 22(8):1767– 1774, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1767_s1.

Hlavaty:1981:FEI

[717] L. Hlavatý and J. Niederle. Field equations invariant under indecomposable representations of the Lorentz group. Journal of Mathematical Physics, 22(8):1775–1780, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i8/p1775_s1.

Sen:1981:ENZ

[718] Amitabha Sen. On the existence of neutrino “zero-modes” in vacuum spacetimes. Journal of Mathematical Physics, 22(8):1781–1786, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i8/p1781_s1.

Gilain:1981:NIE

[719] C. Gilain and D. Levy. A new integral equation for summing Feynman graph series (φ^{3} ladder graph case). Journal of Mathematical Physics, 22(8):1787–1809, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i8/p1787_s1.

Balantekin:1981:RS

[720] A. Baha Balantekin and Itzhak Bars. Representations of supergroups. Journal of Mathematical Physics, 22(8):1810–1818, August 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1810_
s1.

Rosen:1981:IFT

[721] Gerald Rosen. Incompressible fluid turbulence at large Reynolds numbers: Theoretical basis for the t^{−1} decay law and the form of the longitudinal correlation function. Journal of Mathematical Physics, 22(8):1819– 1823, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1819_s1.

Ruggeri:1981:CCE

[722] Tommaso Ruggeri and Alberto Strumia. Convex covariant entropy density, symmetric conservative form, and shock waves in relativistic magnetohydrodynamics. Journal of Mathematical Physics, 22(8):1824– 1827, August 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i8/p1824_s1.

Hiscock:1981:BHM

[723] William A. Hiscock. On black holes in magnetic universes. Journal of Mathematical Physics, 22(8):1828–1833, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i8/p1828_s1.

Batalin:1981:QCF

[724] I. A. Batalin. Quasigroup construction and first class constraints. Journal of Mathematical Physics, 22(9):1837–1850, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1837_s1.

Berdjis:1981:CCC

[725] F. Berdjis. A criterion for completeness of Casimir operators. Journal of Mathematical Physics, 22(9):1851–1856, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1851_s1.

Berdjis:1981:CO

[726] F. Berdjis and E. Beslmüller. Casimir operators for F_{4}, E_{6}, E_{7}, and E_{8}. Journal of Mathematical Physics, 22(9):1857–1860, September 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1857_
s1.

Daskaloyannis:1981:APS

[727] C. Daskaloyannis. The analytical properties of the off-shell T matrix for infinite rank nonlocal separable potentials. Journal of Mathematical Physics, 22(9):1861–1866, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i9/p1861_s1.

Bountis:1981:SPO

[728] Tassos Bountis and Robert H. G. Helleman. On the stability of periodic orbits of two-dimensional mappings. Journal of Mathematical Physics, 22(9):1867–1877, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i9/p1867_s1.

Katzin:1981:GFI

[729] Gerald H. Katzin and Jack Levine. Geodesic first integrals with explicit path-parameter dependence in Riemannian space–times. Journal of Mathematical Physics, 22(9):1878–1891, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1878_s1.

Percacci:1981:GDN

[730] R. Percacci. Global definition of nonlinear sigma model and some consequences. Journal of Mathematical Physics, 22(9):1892–1895, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p1892_s1.

Hojman:1981:IPC

[731] Sergio Hojman and Luis F. Urrutia. On the inverse problem of the calculus of variations. Journal of Mathematical Physics, 22(9):1896–1903, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p1896_s1.

Kaushal:1981:DNI

[732] R. S. Kaushal and H. J. Korsch. Dynamical Noether invariants for time-dependent nonlinear systems. Journal of Mathematical Physics, 22(9):1904–1908, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i9/p1904_s1.

Weaver:1981:DRL

[733] Richard L. Weaver and Yih Hsing Pao. Dispersion relations for linear wave propagation in homogeneous and inhomogeneous media. Journal of Mathematical Physics, 22(9):1909–1918, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1909_s1.

Salingaros:1981:EHP

[734] Nikos Salingaros. Electromagnetism and the holomorphic properties of spacetime. Journal of Mathematical Physics, 22(9):1919–1925, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p1919_s1.

Szczyrba:1981:HSI

[735] Wiktor Szczyrba. A Hamiltonian structure of the interacting gravitational and matter fields. Journal of Mathematical Physics, 22(9):1926– 1943, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p1926_s1.

Nelson:1981:PIT

[736] Bruce Nelson and Benny Sheeks. Path integration for time-dependent metrics. Journal of Mathematical Physics, 22(9):1944–1947, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p1944_s1.

Mitchell:1981:DSS

[737] Thomas P. Mitchell. Dynamical symmetries of the Schrödinger equation. Journal of Mathematical Physics, 22(9):1948–1951, September 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1948_
s1.

Maioli:1981:EPH

[738] Marco Maioli. Exponential perturbations of the harmonic oscillator. Journal of Mathematical Physics, 22(9):1952–1958, September 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p1952_
s1.

Truax:1981:STD

[739] D. Rodney Truax. Symmetry of time-dependent Schrödinger equations. I. A classification of time-dependent potentials by their maximal kinematical algebras. Journal of Mathematical Physics, 22(9):1959–1964, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p1959_s1.

Kowalski:1981:PCM

[740] K. L. Kowalski, Wayne N. Polyzou, and Edward F. Redish. Partition combinatorics and multiparticle scattering theory. Journal of Mathematical Physics, 22(9):1965–1982, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i9/p1965_s1.

Kesarwani:1981:EAO

[741] R. N. Kesarwani and Y. P. Varshni. Eigenvalues of an anharmonic oscillator. Journal of Mathematical Physics, 22(9):1983–1989, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p1983_s1.

Mashford:1981:NTS

[742] J. S. Mashford. A nonmanifold theory of space–time. Journal of Mathematical Physics, 22(9):1990–1993, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i9/p1990_s1.

Monroe:1981:LTT

[743] David K. Monroe. Local transverse-traceless tensor operators in general relativity. Journal of Mathematical Physics, 22(9):1994–2000, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p1994_s1.

Oliveira:1981:WTA

[744] Colber G. Oliveira. Weyl tensors for asymmetric complex curvatures. Journal of Mathematical Physics, 22(9):2001–2005, September 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p2001_
s1.

Kundu:1981:ASG

[745] P. Kundu. On the analyticity of stationary gravitational fields at spatial infinity. Journal of Mathematical Physics, 22(9):2006–2011, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p2006_s1.

Magnon-Ashtekar:1981:TFQ

[746] Anne Magnon-Ashtekar. Thermal fluctuations in quantum gravity: A semiclassical description. Journal of Mathematical Physics, 22(9):2012– 2016, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p2012_s1.

Fung:1981:VBY

[747] M. K. Fung. Validity of the Bethe–Yang hypothesis in the delta-function interaction problem. Journal of Mathematical Physics, 22(9):2017–2019, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p2017_s1.

Malyshev:1981:RGC

[748] V. A. Malyshev and B. Tirozzi. Renormalization group convergence for small perturbations of Gaussian random fields with slowly decaying correlations. Journal of Mathematical Physics, 22(9):2020–2025, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p2020_s1.

Ktorides:1981:DLS

[749] C. N. Ktorides. Diagonalizability of the longitudinal sector of the functional integral in massive SU(2) Yang–Mills theories via topological contributions. Journal of Mathematical Physics, 22(9):2026–2033, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p2026_s1.

Mielke:1981:TES

[750] Eckehard W. Mielke. Toward exact solutions of the nonlinear Heisenberg– Pauli–Weyl spinor equation. Journal of Mathematical Physics, 22(9): 2034–2039, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i9/p2034_s1.

Fulp:1981:ASY

[751] R. O. Fulp, Paul Sommers, and L. K. Norris. Algebraically special Yang– Mills solutions without sources. Journal of Mathematical Physics, 22 (9):2040–2054, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i9/p2040_s1. See erratum [893].

Arisue:1981:GSR

[752] Hiroaki Arisue, Toshiaki Fujiwara, Takuo Inoue, and Kaku Ogawa. Generalized Schrödinger representation and its application to gauge field theories. Journal of Mathematical Physics, 22(9):2055–2059, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i9/p2055_s1.

Ogielski:1981:PIQ

[753] A. T. Ogielski and J. Sobczyk. Path integral quantization of the supersymmetric model of the Dirac particle. Journal of Mathematical Physics, 22(9):2060–2065, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i9/p2060_s1.

Hieu:1981:DNR

[754] Nguyen Văn Hiêu. Decomposition of the natural representation of the U(2) color charge algebra in the three-quark sector. Journal of Mathematical Physics, 22(9):2066–2068, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i9/p2066_s1.

Mitra:1981:HMA

[755] P. Mitra. Higgs multiplets in the adjoint representation of SU(N). Journal of Mathematical Physics, 22(9):2069–2072, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i9/p2069_s1.

Ott:1981:PTD

[756] R. H. Ott. Propagation in a tropospheric duct with a single-step discontinuity in the refractive index in the direction of propagation. Journal of Mathematical Physics, 22(9):2073–2079, September 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i9/p2073_s1.

Hameiri:1981:SES

[757] Eliezer Hameiri. Spectral estimates, stability conditions, and the rotating screw-pinch. Journal of Mathematical Physics, 22(9):2080–2088, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i9/p2080_s1.

Ilamed:1981:ATA

[758] Yehiel Ilamed and Nikos Salingaros. Algebras with three anticommuting elements. I. Spinors and quaternions. Journal of Mathematical Physics, 22(10):2091–2095, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2091_s1.

Salingaros:1981:ATA

[759] Nikos Salingaros. Algebras with three anticommuting elements. II. Two algebras over a singular field. Journal of Mathematical Physics, 22 (10):2096–2100, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2096_s1.

Kim:1981:UTP

[760] Shoon K. Kim. A unified theory of the point groups and their general irreducible representations. Journal of Mathematical Physics, 22 (10):2101–2107, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2101_s1.

Bovier:1981:PCF

[761] Anton Bovier and Daniel Wyler. On the parametrization of certain finite groups and their representations. Journal of Mathematical Physics, 22(10):2108–2112, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2108_s1.

Newmarch:1981:RAGb

[762] J. D. Newmarch and R. M. Golding. The Racah algebra for groups with time reversal symmetry. II. Journal of Mathematical Physics, 22 (10):2113–2119, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2113_s1.

Bretin:1981:GBS

[763] C. Bretin and J. P. Gazeau. Generalized Bessel series and multiplicity problem in complex semisimple Lie algebra theory. Journal of Mathematical Physics, 22(10):2120–2126, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2120_s1.

Palev:1981:CRL

[764] Tchavdar D. Palev. Canonical realizations of Lie superalgebras: Ladder representations of the Lie superalgebra A(m, n). Journal of Mathematical Physics, 22(10):2127–2131, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i10/p2127_s1.

Nelson:1981:FDA

[765] Bruce Nelson and Benny Sheeks. Fredholm determinants associated with Wiener integrals. Journal of Mathematical Physics, 22(10):2132– 2136, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2132_s1.

Blenski:1981:MSP

[766] Tomasz Błeński. Modified singular perturbation method for a stiff system of linear evolution equations. Journal of Mathematical Physics, 22 (10):2137–2143, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2137_s1.

Elworthy:1981:CMD

[767] David Elworthy and Aubrey Truman. Classical mechanics, the diffusion (heat) equation and the Schrödinger equation on a Riemannian manifold. Journal of Mathematical Physics, 22(10):2144–2166, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i10/
p2144_s1.

Rao:1981:RPE

[768] K. N. Srinivasa Rao, D. Saroja, and A. V. Gopala Rao. On rotations in a pseudo-Euclidean space and proper Lorentz transformations. Journal of Mathematical Physics, 22(10):2167–2179, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2167_s1.

Gupta:1981:EIS

[769] M. R. Gupta and Jayanti Ray. Extension of inverse scattering method to nonlinear evolution equation in nonuniform medium. Journal of Mathematical Physics, 22(10):2180–2183, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2180_s1.

Plohr:1981:BII

[770] Bradley Plohr. The behavior at infinity of isotropic vortices and monopoles. Journal of Mathematical Physics, 22(10):2184–2190, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i10/p2184_s1.

Newton:1981:ISI

[771] Roger G. Newton. Inverse scattering. III. Three dimensions, continued. Journal of Mathematical Physics, 22(10):2191–2200, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i10/

p2191_s1. See erratum [1004].

Furukawa:1981:NDP

[772] Masao Furukawa. A new direct proof of the expansion of a Coulombdistorted plane wave in Coulomb-distorted spherical waves. Journal of Mathematical Physics, 22(10):2201–2202, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2201_s1.

Hermann:1981:AGS

[773] Robert Hermann. Algebraic and geometric structure of linear filters and scattering systems. I. Journal of Mathematical Physics, 22(10):2203– 2207, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2203_s1.

Semenoff:1981:PLD

[774] G. Semenoff, H. Matsumoto, and H. Umezawa. A perturbative look at the dynamics of extended systems in quantum field theory. Journal of Mathematical Physics, 22(10):2208–2222, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2208_s1.

Ferrando:1981:IRC

[775] J. J. Ferrando and J. Olivert. Incompressibility in relativistic continuous media. Journal of Mathematical Physics, 22(10):2223–2227, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i10/p2223_s1.

Finkelstein:1981:CSE

[776] Robert J. Finkelstein and Jaegu Kim. Classical solutions of the equations of supergravity. Journal of Mathematical Physics, 22(10):2228– 2234, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2228_s1.

Krori:1981:CDM

[777] K. D. Krori, T. Chaudhury, and R. Bhattacharjee. Charged Demianski metric. Journal of Mathematical Physics, 22(10):2235–2236, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i10/p2235_s1.

Raychaudhuri:1981:VFI

[778] A. K. Raychaudhuri and S. K. Saha. Viscous fluid interpretation of electromagnetic fields. I. Journal of Mathematical Physics, 22(10):2237– 2239, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2237_s1.

Ludwig:1981:STG

[779] Garry Ludwig. Space–times with geodesic, shear-free, twist-free, nonexpanding rays. II. Journal of Mathematical Physics, 22(10):2240–2244, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2240_s1.

VandenBergh:1981:PVP

[780] N. Van den Bergh. The Palatini variational principle for the general Bergmann–Wagoner–Nordtvedt theory of gravitation. Journal of Mathematical Physics, 22(10):2245–2248, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2245_s1.

Takimoto:1981:CFM

[781] Noboru Takimoto. Coherent frequency modulation in Gaussian stochastic processes. Journal of Mathematical Physics, 22(10):2249–2257, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i10/p2249_s1.

Manoukian:1981:GDT

[782] Edward B. Manoukian. Generalized decoupling theorem in quantum field theory. Journal of Mathematical Physics, 22(10):2258–2262, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i10/p2258_s1.

Barr:1981:ECV

[783] S. Barr and A. Zee. Electric charge is vectorlike. Journal of Mathematical Physics, 22(10):2263–2268, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i10/p2263_s1.

Bonora:1981:SAP

[784] L. Bonora, A. de Pantz, and P. Pasti. Some applications of Postnikov systems in gauge theories. Journal of Mathematical Physics, 22(10):2269– 2271, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2269_s1.

Edelen:1981:LIE

[785] Dominic G. B. Edelen. The linear integral equations associated with any antiexact solution of the Yang–Mills free field equations. Journal of Mathematical Physics, 22(10):2272–2275, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2272_s1.

Eriksson:1981:IGI

[786] K. E. Eriksson, N. Svartholm, and B. S. Skagerstam. On invariant group integrals in lattice QCD. Journal of Mathematical Physics, 22(10):2276– 2278, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2276_s1.

Kyriakopoulos:1981:MSY

[787] E. Kyriakopoulos. Minkowski space Yang–Mills fields from solutions of equations in the three-dimensional Euclidean space. Journal of Mathematical Physics, 22(10):2279–2282, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2279_s1.

Cant:1981:ICM

[788] A. Cant. Invariant connections and magnetic monopoles. Journal of Mathematical Physics, 22(10):2283–2288, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2283_s1.

Kim:1981:SFR

[789] Y. S. Kim and Marilyn E. Noz. Symplectic formulation of relativistic quantum mechanics. Journal of Mathematical Physics, 22(10):2289– 2293, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i10/p2289_s1.

Jagannathan:1981:BPB

[790] R. Jagannathan and R. Vasudevan. Boson, para-boson, and bosonfermion representations of some graded Lie algebras. Journal of Mathematical Physics, 22(10):2294–2299, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2294_s1.

Ross:1981:PWT

[791] D. K. Ross. Plane waves and topology. Journal of Mathematical Physics, 22(10):2300–2302, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2300_s1.

Kim:1981:NMC

[792] Shoon K. Kim. A new method of constructing the symmetry coordinates of molecular vibrations based on the correspondence theorem. Journal of Mathematical Physics, 22(10):2303–2306, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2303_s1.

Hostler:1981:QFT

[793] Levere Hostler. Quantum field theory of particles of indefinite mass. II. An electromagnetic model. Journal of Mathematical Physics, 22 (10):2307–2316, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2307_s1.

Hillion:1981:ASG

[794] P. Hillion and S. Quinnez. Analytical solutions of geometric optics with an approximation of diffraction. Journal of Mathematical Physics, 22 (10):2317–2325, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i10/p2317_s1.

VandenBerghe:1981:SOTa

[795] G. Vanden Berghe and H. E. De Meyer. Shift operator techniques for the classification of multipole-phonon states. VIII. R(5) ↓ SU(2) ⊗ SU(2) reduction for obtaining quadrupole O_{l}^{0} eigenvalues. Journal of Mathematical Physics, 22(10):2326–2331, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2326_s1.

VandenBerghe:1981:SOTb

[796] G. Vanden Berghe, H. E. De Meyer, and J. Van der Jeugt. Shift operator techniques for the classification of multipole-phonon states. IX. Properties of nonscalar R(3) product operators in the G_{2} group. Journal of Mathematical Physics, 22(10):2332–2337, October 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i10/p2332_s1.

Coen:1981:IPS

[797] Shimon Coen. The inverse problem of the shear modulus and density profiles of a layered earth — torsional vibration data. Journal of Mathematical Physics, 22(10):2338–2341, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2338_s1.

DeFacio:1981:EGL

[798] B. DeFacio and H. E. Moses. Erratum: “The Gel’fand–Levitan equation can give simple examples of non-self-adjoint operators with complete eigenfunctions and spectral representations. I. Ghosts and resonances”, [J. Math. Phys. 21 (1980), no. 7, 1716–1723; MR 81m:81053]. Journal of Mathematical Physics, 22(10):2342, October 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i10/p2342_s1. See [269].

Alberti:1981:NSD

[799] Peter M. Alberti and Armin Uhlmann. A note on stochastic dynamics in the state space of a commutative C^{∗}-algebra. Journal of Mathematical Physics, 22(11):2345–2351, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2345_s1.

Patera:1981:TAN

[800] J. Patera and R. T. Sharp. On the triangle anomaly number of SU(n) representations. Journal of Mathematical Physics, 22(11):2352–2356, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2352_s1.

Earp:1981:WSP

[801] C. Earp. Wigner 9j symbols and the product group. Journal of Mathematical Physics, 22(11):2357–2359, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2357_s1.

DeMeyer:1981:NESa

[802] H. E. De Meyer, G. Vanden Berghe, and J. W. B. Hughes. Nonscalar extension of shift operator techniques for SU(3) in an O(3) basis. I. Theory. Journal of Mathematical Physics, 22(11):2360–2365, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/
p2360_s1.

DeMeyer:1981:NESb

[803] H. E. De Meyer, G. Vanden Berghe, and J. W. B. Hughes. Nonscalar extension of shift operator techniques for SU(3) in an O(3) basis. II. Applications. Journal of Mathematical Physics, 22(11):2366–2371, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2366_s1.

Adams:1981:LAL

[804] Mitchell D. Adams. Lorentz and SU(2)⊗U(N) analysis of the Lie algebra sp(4N; r)[gl(4N; r)] for any integer N ≥ 1. Journal of Mathematical Physics, 22(11):2372–2375, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2372_s1.

Gould:1981:WCS

[805] M. D. Gould. Wigner coefficients for a semisimple Lie group and the matrix elements of the O(n) generators. Journal of Mathematical Physics, 22(11):2376–2388, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2376_s1.

Giovannini:1981:CQS

[806] N. Giovannini. Classical and quantal systems of imprimitivity. Journal of Mathematical Physics, 22(11):2389–2396, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2389_s1.

Giovannini:1981:SSC

[807] N. Giovannini. State spaces for classical and quantal, relativistic and nonrelativistic elementary particles. Journal of Mathematical Physics, 22(11):2397–2403, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2397_s1.

Kasperkovitz:1981:NRT

[808] P. Kasperkovitz and G. Kahl. Noncomplex representations and their relation to antiunitary symmetry. Journal of Mathematical Physics, 22 (11):2404–2416, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2404_s1.

Kasperkovitz:1981:RTC

[809] P. Kasperkovitz. Representation theory of compact groups over fields of characteristic zero. Journal of Mathematical Physics, 22(11):2417– 2426, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2417_s1.

Demiralp:1981:AEC

[810] Metin Demiralp and N. Abdülbaki Baykara. Analytic evaluation of certain zeroth order Coulombic hyperangular interaction integrals. Journal of Mathematical Physics, 22(11):2427–2432, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2427_s1.

Weinert:1981:SPE

[811] M. Weinert. Solution of Poisson’s equation: Beyond Ewald-type methods. Journal of Mathematical Physics, 22(11):2433–2439, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2433_s1.

Derrick:1981:LEI

[812] G. H. Derrick. Linear equations invariant under arbitrary coordinate changes. Journal of Mathematical Physics, 22(11):2440–2446, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2440_s1.

Kaliappan:1981:SSE

[813] P. Kaliappan and M. Lakshmanan. Similarity solutions for the Ernst equations with electromagnetic fields. Journal of Mathematical Physics, 22(11):2447–2451, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2447_s1.

vandenBerg:1981:BGF

[814] M. van den Berg. Bounds on Green’s functions of second-order differential equations. Journal of Mathematical Physics, 22(11):2452–2455, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2452_s1.

Nakamura:1981:DMS

[815] Akira Nakamura. Decay mode solution of the two-dimensional KdV equation and the generalized Bäcklund transformation. Journal of Mathematical Physics, 22(11):2456–2462, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2456_s1.

Bruschi:1981:EEA

[816] M. Bruschi, O. Ragnisco, and D. Levi. Evolution equations associated with the discrete analog of the matrix Schrödinger spectral problem solvable by the inverse spectral transform. Journal of Mathematical Physics, 22(11):2463–2471, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2463_s1.

Auberson:1981:SPB

[817] G. Auberson and G. Mennessier. Some properties of Borel summable functions. Journal of Mathematical Physics, 22(11):2472–2481, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2472_s1.

vanHaeringen:1981:CUR

[818] H. van Haeringen and L. P. Kok. The Coulomb unitarity relation and some series of products of three Legendre functions. Journal of Mathematical Physics, 22(11):2482–2483, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2482_s1.

Krivine:1981:EMP

[819] H. Krivine and J. Treiner. An expansion of the moments of powers of a Fermi distribution. Journal of Mathematical Physics, 22(11):2484– 2485, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2484_s1.

Joshi:1981:IRC

[820] K. N. Joshi and B. S. Rajput. Irreducible representations of the complex rotation group in terms of the axis and angles of rotations. Journal of Mathematical Physics, 22(11):2486–2488, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2486_s1.

Antoine:1981:ITR

[821] J.-P. Antoine and W. Karwowski. Interpolation theory and refinement of nested Hilbert spaces. Journal of Mathematical Physics, 22(11):2489– 2496, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2489_s1. See erratum [1160].

MartinezAlonso:1981:MHS

[822] L. Martínez Alonso and F. Guil Guerrero. Modified Hamiltonian systems and canonical transformations arising from the relationship between generalized Zakharov–Shabat and energy-dependent Schrödinger operators. Journal of Mathematical Physics, 22(11):2497–2503, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i11/p2497_s1.

Sarlet:1981:CST

[823] W. Sarlet and J. R. Ray. Classification scheme for two-dimensional Ermakov-type systems and generalizations. Journal of Mathematical Physics, 22(11):2504–2511, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2504_s1.

Collas:1981:EPC

[824] Peter Collas. Equivalent potentials in classical mechanics. Journal of Mathematical Physics, 22(11):2512–2517, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2512_s1.

Chudnovsky:1981:BTC

[825] D. V. Chudnovsky and G. V. Chudnovsky. Bäcklund transformations connecting different isospectral deformation equations. Journal of Mathematical Physics, 22(11):2518–2522, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2518_s1.

Weston:1981:IPR

[826] V. H. Weston. Inverse problem for the reduced wave equation with fixed incident wave. II. Journal of Mathematical Physics, 22(11):2523–2529, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2523_s1.

Bialynicki-Birula:1981:NH

[827] Iwo Białynicki-Birula, E. T. Newman, J. Porter, J. Winicour, B. Lukács, Z. Perjés, and Á. Sebestyén. A note on helicity. Journal of Mathematical Physics, 22(11):2530–2532, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2530_s1.

Ranada:1981:BSC

[828] Antonio F. Rañada and Juan M. Usón. Bound states of a classical charged nonlinear Dirac field in a Coulomb potential. Journal of Mathematical Physics, 22(11):2533–2538, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2533_s1.

Arai:1981:MHOa

[829] Asao Arai. On a model of a harmonic oscillator coupled to a quantized, massless, scalar field. I. Journal of Mathematical Physics, 22(11):2539– 2548, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2539_s1.

Arai:1981:MHOb

[830] Asao Arai. On a model of a harmonic oscillator coupled to a quantized, massless, scalar field. II. Journal of Mathematical Physics, 22(11):2549– 2552, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2549_s1.

Gudder:1981:GMC

[831] S. Gudder and J. Zerbe. Generalized monotone convergence and Radon– Nikodým theorems. Journal of Mathematical Physics, 22(11):2553–2561, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2553_s1.

Schroeck:1981:MQM

[832] Franklin E. Schroeck, Jr. A model of a quantum mechanical treatment of measurement with a physical interpretation. Journal of Mathematical Physics, 22(11):2562–2572, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2562_s1.

Johnson:1981:EPL

[833] Bruce R. Johnson. An expansion for power-law wavefunctions. Journal of Mathematical Physics, 22(11):2573–2580, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2573_s1.

Jimbo:1981:MPD

[834] M. Jimbo, M. Kashiwara, and T. Miwa. Monodromy preserving deformation of ramified solutions to (Δ − m^{2})u = 0. Journal of Mathematical Physics, 22(11):2581–2587, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i11/p2581_s1.

Davidson:1981:SQE

[835] Mark Davidson. Stochastic quantization of the electromagnetic field. Journal of Mathematical Physics, 22(11):2588–2593, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/
p2588_s1.

Sanz:1981:TMP

[836] P. Sanz, J. Sañudo, and J. Sesma. Trajectories of the S-matrix poles in the complex wavenumber plane. Journal of Mathematical Physics, 22 (11):2594–2597, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2594_s1.

Coleman:1981:SST

[837] Robert Alan Coleman and Herbert Korte. Spacetime G structures and their prolongations. Journal of Mathematical Physics, 22(11):2598–2611, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2598_s1.

Ramaswamy:1981:DMG

[838] Sriram Ramaswamy and Amitabha Sen. Dual-mass in general relativity. Journal of Mathematical Physics, 22(11):2612–2619, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/
p2612_s1.

McIntosh:1981:CRP

[839] C. B. G. McIntosh, J. M. Foyster, and A. W.-C. Lun. The classification of the Ricci and Plebański tensors in general relativity using Newman– Penrose formalism. Journal of Mathematical Physics, 22(11):2620–2623, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2620_s1.

Cosgrove:1981:BTH

[840] Christopher M. Cosgrove. Bäcklund transformations in the Hauser–Ernst formalism for stationary axisymmetric spacetimes. Journal of Mathematical Physics, 22(11):2624–2639, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i11/p2624_s1.

Kuchar:1981:GRD

[841] Karel Kuchař. General relativity: Dynamics without symmetry. Journal of Mathematical Physics, 22(11):2640–2654, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2640_s1.

GarciaDiaz:1981:ANC

[842] A. García Díaz and J. F. Plebański. All nontwisting N’s with cosmological constant. Journal of Mathematical Physics, 22(11):2655–2658, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2655_s1.

Wald:1981:ESS

[843] Robert M. Wald and Ping Yip. On the existence of simultaneous synchronous coordinates in spacetimes with spacelike singularities. Journal of Mathematical Physics, 22(11):2659–2665, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/

p2659_s1. See erratum [1006].

Tupper:1981:EEF

[844] B. O. J. Tupper. The equivalence of electromagnetic fields and viscous fluids in general relativity. Journal of Mathematical Physics, 22 (11):2666–2673, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2666_s1.

Schaer:1981:GCS

[845] Carl Schaer. A generalization of the concept of spacetime. Journal of Mathematical Physics, 22(11):2674–2677, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2674_s1.

Lim:1981:TMS

[846] Chee-Seng Lim. Transformation modified singular spherical integrals for an arbitrarily time-dependent charged particle in a crystal. Journal of Mathematical Physics, 22(11):2678–2688, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/p2678_s1.

Gonzalez-Diaz:1981:ASC

[847] P. F. González-Díaz. Analytical solution for the change of Raman transitions quantum statistics. Journal of Mathematical Physics, 22(11):2689– 2691, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2689_s1.

Kriegsmann:1981:LWC

[848] Gregory A. Kriegsmann. Linear wave conversion in a warm, unmagnetized, collisionless plasma. Journal of Mathematical Physics, 22 (11):2692–2698, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2692_s1. See erratum [1034].

Reboucas:1981:TDF

[849] Marcelo J. Rebouças and J. Ademir S. de Lima. Time-dependent, finite, rotating universes. Journal of Mathematical Physics, 22(11):2699– 2703, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i11/p2699_s1. See erratum [1130].

Atkinson:1981:NCQ

[850] D. Atkinson, J. K. Drohm, P. W. Johnson, and K. Stam. Nonperturbative confinement in quantum chromodynamics. I. Study of an approximate equation of Mandelstam. Journal of Mathematical Physics, 22 (11):2704–2712, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2704_s1.

Iyanaga:1981:GST

[851] M. Iyanaga. Generalization of Stokes theorem and pure gauge fields. Journal of Mathematical Physics, 22(11):2713–2715, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i11/
p2713_s1.

Michielsen:1981:TDR

[852] B. L. Michielsen, G. C. Herman, A. T. de Hoop, and D. de Zutter. Threedimensional relativistic scattering of electromagnetic waves by an object in uniform translational motion. Journal of Mathematical Physics, 22 (11):2716–2722, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i11/p2716_s1.

Chen:1981:NRI

[853] H. H. Chen and Felix Lee. Numerical representation and identification of graphs. Journal of Mathematical Physics, 22(12):2727–2731, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2727_s1.

Gaskell:1981:SB

[854] R. Gaskell, G. Rosensteel, and R. T. Sharp. Sp(6) states in an SU(3) × U(1) basis. Journal of Mathematical Physics, 22(12):2732–2735, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2732_s1.

Gaskell:1981:GFC

[855] R. Gaskell and R. T. Sharp. Generating functions for G_{2} characters and subgroup branching rules. Journal of Mathematical Physics, 22 (12):2736–2739, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2736_s1.

Marcu:1981:GSSa

[856] M. Marcu, A. Regev, and V. Rittenberg. The global symmetries of spin systems defined on abelian groups. I. Journal of Mathematical Physics, 22(12):2740–2752, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2740_s1.

Marcu:1981:GSSb

[857] M. Marcu and V. Rittenberg. The global symmetries of spin systems defined on abelian groups. II. Journal of Mathematical Physics, 22 (12):2753–2758, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2753_s1.

Cianci:1981:GTT

[858] Roberto Cianci. On generalized torsion tensor fields and the reduced fiber bundle. Journal of Mathematical Physics, 22(12):2759–2761, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2759_s1.

Gruber:1981:RGB

[859] Bruno Gruber and Anatoli U. Klimyk. Representations of the groups GL(n, R) and SU(n) in an SO(n) basis. Journal of Mathematical Physics, 22(12):2762–2769, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i12/p2762_s1.

McKay:1981:SFI

[860] J. McKay, J. Patera, and R. T. Sharp. Second and fourth indices of plethysms. Journal of Mathematical Physics, 22(12):2770–2774, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2770_s1.

Hughes:1981:IRC

[861] J. W. B. Hughes. Irreducible representations of the central extension of Sl(2) ∧ T_{2}. Journal of Mathematical Physics, 22(12):2775–2779, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2775_s1.

Asano:1981:NSA

[862] Naruyoshi Asano and Yusuke Kato. Non-self-adjoint Zakharov–Shabat operator with a potential of the finite asymptotic values. I. Direct spectral and scattering problems. Journal of Mathematical Physics, 22(12):2780– 2793, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2780_s1.

Cacuci:1981:STNa

[863] Dan G. Cacuci. Sensitivity theory for nonlinear systems. I. Nonlinear functional analysis approach. Journal of Mathematical Physics, 22 (12):2794–2802, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2794_s1.

Cacuci:1981:STNb

[864] Dan G. Cacuci. Sensitivity theory for nonlinear systems. II. Extensions to additional classes of responses. Journal of Mathematical Physics, 22 (12):2803–2812, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2803_s1.

Bohm:1981:RPG

[865] A. Böhm. Resonance poles and Gamow vectors in the rigged Hilbert space formulation of quantum mechanics. Journal of Mathematical Physics, 22(12):2813–2823, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i12/p2813_s1.

Case:1981:PHS

[866] K. M. Case and A. M. Roos. Perturbed Hamiltonian systems. Journal of Mathematical Physics, 22(12):2824–2830, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/p2824_s1.

Ikeda:1981:CDS

[867] Satoshi Ikeda. On the covariant differential of spin direction in the Finslerian deformation theory of ferromagnetic substances. Journal of Mathematical Physics, 22(12):2831–2834, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i12/p2831_s1.

AlarconGutierrez:1981:SDG

[868] Sabás Alarcón Gutiérrez, Alan L. Dudley, and Jerzy F. Plebañski. Signals and discontinuities in general relativistic nonlinear electrodynamics. Journal of Mathematical Physics, 22(12):2835–2848, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/
p2835_s1.

Abdel-Raouf:1981:PLB

[869] M. A. Abdel-Raouf. Partitioning lower bounds for Bubnov–Galerkin’s eigenvalues. Journal of Mathematical Physics, 22(12):2849–2853, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2849_s1.

Polyzou:1981:CCG

[870] Wayne Polyzou. Construction of (JJ^{∗})^{−1} in the Chandler–Gibson reaction theory. Journal of Mathematical Physics, 22(12):2854–2857, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2854_s1.

Evans:1981:FED

[871] J. W. Evans and D. K. Hoffman. Faddeev’s equations in differential form: Completeness of physical and spurious solutions and spectral properties. Journal of Mathematical Physics, 22(12):2858–2871, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2858_s1.

Saenz:1981:QMS

[872] A. W. Sáenz. Quantum-mechanical scattering by impenetrable periodic surfaces. Journal of Mathematical Physics, 22(12):2872–2884, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2872_s1.

Ashtekar:1981:RDF

[873] Abhay Ashtekar. Radiative degrees of freedom of the gravitational field in exact general relativity. Journal of Mathematical Physics, 22(12):2885– 2895, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2885_s1.

Derrick:1981:CSC

[874] G. H. Derrick. On a completely symmetric choice of space–time coordinates. Journal of Mathematical Physics, 22(12):2896–2902, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2896_s1.

Nishigori:1981:MFA

[875] Takeo Nishigori. Memory function approach to nonlinear deterministic systems: An exact linear equation. Journal of Mathematical Physics, 22 (12):2903–2909, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2903_s1.

Messer:1981:GPT

[876] J. Messer. On the gravitational phase transition in the Thomas–Fermi model. Journal of Mathematical Physics, 22(12):2910–2917, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2910_s1.

deSmedt:1981:KCR

[877] Ph. de Smedt and G. Stragier. The KMS condition and regularity for the ideal Bose gas. Journal of Mathematical Physics, 22(12):2918–2920, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2918_s1.

Majewski:1981:TDB

[878] W. A. Majewski. Time development of Bose systems. Journal of Mathematical Physics, 22(12):2921–2925, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v22/i12/p2921_s1.

Lavenda:1981:KAB

[879] B. H. Lavenda and E. Santamato. The kinetic analog of Boltzmann’s principle. Journal of Mathematical Physics, 22(12):2926–2933, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2926_s1.

McKellar:1981:UGP

[880] R. J. McKellar. The uniqueness of gravity as a Poincaré or Lorentz gauge theory. Journal of Mathematical Physics, 22(12):2934–2942, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2934_s1.

Doria:1981:QAF

[881] Francisco Antonio Doria. Quasi-abelian and fully non-abelian gauge field copies: A classification. Journal of Mathematical Physics, 22(12):2943– 2951, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2943_s1.

Lopuszanski:1981:DIS

[882] Jan T. Łopuszański. Democracy of internal symmetries in supersymmetrical quantum field theory. Journal of Mathematical Physics, 22 (12):2952–2956, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2952_s1.

Nash:1981:NBF

[883] Patrick L. Nash. On a new Bose–Fermi mixing symmetry. Journal of Mathematical Physics, 22(12):2957–2960, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/p2957_s1.

Halemane:1981:SAP

[884] Thirumala Raya Halemane. A statistical approach to perturbation theory and inverse-energy-weighted sum rules. Journal of Mathematical Physics, 22(12):2961–2967, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i12/p2961_s1.

Sen:1981:CCF

[885] Siddhartha Sen. A calculation of the Casimir force on a circular boundary. Journal of Mathematical Physics, 22(12):2968–2973, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/
p2968_s1.

Sutton:1981:ASW

[886] G. R. Sutton. Application of a stochastic waveguide propagation model to ocean acoustics. Journal of Mathematical Physics, 22(12):2974–2976, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2974_s1.

Vatsya:1981:ASC

[887] S. R. Vatsya. On approximating the solutions and critical points of some nonlinear phenomena. Journal of Mathematical Physics, 22(12):2977– 2982, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p2977_s1.

Rosen:1981:IIH

[888] Gerald Rosen. Inequalities in isotropic homogeneous turbulence theory. Journal of Mathematical Physics, 22(12):2983–2985, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/
p2983_s1.

Peyrard:1981:DPD

[889] M. Peyrard. Dynamic polarizability of the double quadratic and multiquadratic chains with anharmonic coupling. Journal of Mathematical Physics, 22(12):2986–2995, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i12/p2986_s1.

Hall:1981:EST

[890] George L. Hall. Electrostatic structural transitions in a Gaussian Wigner solid. Journal of Mathematical Physics, 22(12):2996–3001, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p2996_s1.

Roepstorff:1981:PGA

[891] Gert Roepstorff. The Peierls–Griffiths argument for disordered Ising systems. Journal of Mathematical Physics, 22(12):3002–3005, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v22/i12/p3002_s1.

Bose:1981:NTA

[892] S. K. Bose and E. Esteban. A null tetrad analysis of the Ernst metric. Journal of Mathematical Physics, 22(12):3006–3009, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/
p3006_s1.

Fulp:1981:EAS

[893] R. O. Fulp, Paul Sommers, and L. K. Norris. Erratum: “Algebraically special Yang–Mills solutions without sources”, [J. Math. Phys. 22, 2040 (1981)]. Journal of Mathematical Physics, 22(12):3010, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v22/i12/

p3010_s2. See [751].

Moses:1981:EKG

[894] H. E. Moses. Erratum: “A kernel of Gel’fand–Levitan type for the threedimensional Schrödinger equation”, [J. Math. Phys. 21 (1980), no. 1, 83–89; MR 81a:81023]. Journal of Mathematical Physics, 22(12):3010, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v22/i12/p3010_s3. See [34].

Soto:1981:AOS

[895] M. F. Soto, Jr. and R. Mirman. Addendum: “Objects for the symmetric group”, [J. Math. Phys. 22, 1144 (1981)]. Journal of Mathematical Physics, 22(12):3010, December 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v22/i12/p3010_s1. See [616].

Salingaros:1982:CCA

[896] Nikos Salingaros. On the classification of Clifford algebras and their relation to spinors in n dimensions. Journal of Mathematical Physics, 23 (1):1–7, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i1/p1_s1. See erratum [1085].

Okubo:1982:MFO

[897] Susumu Okubo. Modified fourth-order Casimir invariants and indices for simple Lie algebras. Journal of Mathematical Physics, 23(1):8–20, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i1/p8_s1.

Malec:1982:GAB

[898] Edward Malec. Gribov ambiguities from the bifurcation theory viewpoint. Journal of Mathematical Physics, 23(1):21–23, January 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/p21_s1.

Iosilevskii:1982:LSE

[899] Ya. A. Iosilevskii. Lippmann–Schwinger equations for a scalar macroscopic field in an anisotropic stratified medium. Journal of Mathematical Physics, 23(1):24–38, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p24_s1.

Epifanio:1982:STP

[900] G. Epifanio and C. Trapani. Some topics in pre-Hilbert space. Journal of Mathematical Physics, 23(1):39–42, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i1/p39_s1.

Truax:1982:STD

[901] D. Rodney Truax. Symmetry of time-dependent Schrödinger equations. II. exact solutions for the equation {∂_{xx} + 2i∂_{t} − 2g_{2}(t)x^{2} − 2g_{1}(t)x − 2g_{0}(t)}Ψ(x, t) = 0. Journal of Mathematical Physics, 23(1):43–54, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i1/p43_s1.

Robbins:1982:SNO

[902] Steven Robbins. Statistics with number operator CC^{∗}. Journal of Mathematical Physics, 23(1):55–63, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i1/p55_s1.

Crandall:1982:GSE

[903] R. E. Crandall and Mary Hall Reno. Ground state energy bounds for potentials |x|^{ν}. Journal of Mathematical Physics, 23(1):64–70, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v23/i1/p64_s1. See erratum [1161].

Majumdar:1982:TMB

[904] Chanchal K. Majumdar and S. Banerjee. The three magnon bound-state equation in one dimension. Journal of Mathematical Physics, 23(1):71– 73, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i1/p71_s1.

Gesztesy:1982:SCT

[905] F. Gesztesy. On the structure of Coulomb-type scattering amplitudes. Journal of Mathematical Physics, 23(1):74–82, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/p74_s1.

Yamani:1982:EFM

[906] Hashim A. Yamani. The equivalence of the Feshbach and J-matrix methods. Journal of Mathematical Physics, 23(1):83–86, January 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/p83_s1.

Terry:1982:AFC

[907] Paul Terry. The asymptotic form of the continuum wavefunctions and redundant poles in the Heisenberg condition. Journal of Mathematical Physics, 23(1):87–91, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p87_s1.

Kesarwani:1982:QPW

[908] R. N. Kesarwani and Y. P. Varshni. The quantum pendulum in the WKBJ approximation. Journal of Mathematical Physics, 23(1):92–95, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i1/p92_s1.

Retzloff:1982:NMFa

[909] D. G. Retzloff, B. DeFacio, and P. W. Dennis. A new mathematical formulation of accelerated observers in general relativity. I. Journal of Mathematical Physics, 23(1):96–104, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i1/p96_s1.

Retzloff:1982:NMFb

[910] D. G. Retzloff, B. DeFacio, and P. W. Dennis. A new mathematical formulation of accelerated observers in general relativity. II. Journal of Mathematical Physics, 23(1):105–108, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i1/p105_s1.

Friedman:1982:VHC

[911] John L. Friedman and Steven Mayer. Vacuum handles carrying angular momentum; electrovac handles carrying net charge. Journal of Mathematical Physics, 23(1):109–115, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i1/p109_s1.

Sonnleitner:1982:FKV

[912] Stephanie A. Sonnleitner and J. D. Finley III. The form of Killing vectors in expanding HH spaces. Journal of Mathematical Physics, 23(1):116– 122, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i1/p116_s1.

GarciaD:1982:ETS

[913] Alberto García D. and Jerzy F. Plebański. An exceptional type D shearing twisting electrovac with λ. Journal of Mathematical Physics, 23(1):123–125, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p123_s1.

Krori:1982:SLI

[914] K. D. Krori and B. B. Paul. A Schwarzschild-like interior solution for charged spheres. Journal of Mathematical Physics, 23(1):126–127, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i1/p126_s1.

Bradfield:1982:JKK

[915] T. Bradfield and R. Kantowski. Jordan–Kaluza–Klein type unified theories of gauge and gravity fields. Journal of Mathematical Physics, 23(1):128–131, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p128_s1.

Davidson:1982:SQL

[916] Mark Davidson. Stochastic quantization of the linearized gravitational field. Journal of Mathematical Physics, 23(1):132–137, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/
p132_s1.

Isaacson:1982:CLC

[917] D. Isaacson. The continuum limit of a classical 3-component, onedimensional Heisenberg model is Brownian motion on the surface of a sphere. Journal of Mathematical Physics, 23(1):138–140, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/
p138_s1.

Kowall:1982:CEF

[918] James Kowall and H. M. Fried. A cluster expansion in field theory. Journal of Mathematical Physics, 23(1):141–144, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i1/p141_s1.

Steeb:1982:SSN

[919] W.-H. Steeb, W. Erig, and W. Strampp. Similarity solutions of nonlinear Dirac equations and conserved currents. Journal of Mathematical Physics, 23(1):145–153, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i1/p145_s1.

Talukdar:1982:MCN

[920] B. Talukdar, N. Mallick, M. Kundu, and R. N. Panigrahi. T matrix for Coulomb-nuclear admixtures. Journal of Mathematical Physics, 23(1):154–155, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p154_s1.

Gazeau:1982:FEC

[921] Jean-Pierre Gazeau. Four Euclidean conformal group approach to the multiphoton processes in the H-atom. Journal of Mathematical Physics, 23(1):156–164, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p156_s1.

Lewis:1982:EIC

[922] H. Ralph Lewis and P. G. L. Leach. Exact invariants for a class of time-dependent nonlinear Hamiltonian systems. Journal of Mathematical Physics, 23(1):165–175, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i1/p165_s1.

Larsson:1982:CRFa

[923] Jonas Larsson. Current responses of first and second order in a collisionless plasma. I. Stationary plasma. Journal of Mathematical Physics, 23(1):176–182, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i1/p176_s1.

Larsson:1982:CRFb

[924] Jonas Larsson. Current responses of first and second order in a collisionless plasma. II. Three-wave interaction. Journal of Mathematical Physics, 23(1):183–187, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basu:1982:UIR

[925] Debabrata Basu and Kurt Bernardo Wolf. The unitary irreducible representations of SL(2, R) in all subgroup reductions. Journal of Mathematical Physics, 23(2):189–205, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i2/p189_s1.

Sitaram:1982:GQR

[926] B. R. Sitaram and K. C. Tripathy. Geometric quantization and representations of semisimple Lie groups: Spin(2, 1) and Spin(2, 2). Journal of Mathematical Physics, 23(2):206–210, February 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i2/p206_s1.

GuilGuerrero:1982:REH

[927] F. Guil Guerrero. The Riccati equation and Hamiltonian systems. Journal of Mathematical Physics, 23(2):211–216, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p211_s1.

GuilGuerrero:1982:FOC

[928] F. Guil Guerrero. Factorization of operators and completely integrable Hamiltonian systems. Journal of Mathematical Physics, 23(2):217–226, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i2/p217_s1.

Urani:1982:GIA

[929] John R. Urani and Marilyn H. Kemp. A global isometry approach to accelerating observers in flat space–time. Journal of Mathematical Physics, 23(2):227–230, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i2/p227_s1.

Ablamowicz:1982:CAA

[930] R. Abłamowicz, Z. Oziewicz, and J. Rzewuski. Clifford algebra approach to twistors. Journal of Mathematical Physics, 23(2):231–242, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i2/p231_s1.

Adamowski:1982:EEC

[931] Janusz Adamowski, Bernd Gerlach, and Hajo Leschke. Explicit evaluation of certain Gaussian functional integrals arising in problems of statistical physics. Journal of Mathematical Physics, 23(2):243–249, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i2/p243_s1.

Rubin:1982:RWL

[932] Robert J. Rubin and George H. Weiss. Random walks on lattices. The problem of visits to a set of points revisited. Journal of Mathematical Physics, 23(2):250–253, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i2/p250_s1.

Fernandez:1982:HCI

[933] F. M. Fernández, A. Mesón, and E. A. Castro. Hypervirial calculation of integrals involving Bessel functions. Journal of Mathematical Physics, 23(2):254–255, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i2/p254_s1.

Dominici:1982:PCI

[934] D. Dominici and J. Gomis. “Poincaré–Cartan integral invariant and canonical transformations for singular Lagrangians”: An addendum, [J. Math. Phys. 21 (1980), no. 8, 2124–2130; MR 81k:70013]. Journal of Mathematical Physics, 23(2):256–257, February 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i2/p256_s1. See [320].

Jaulent:1982:SSI

[935] M. Jaulent and C. Jean. Solution of a Schrödinger inverse scattering problem with a polynomial spectral dependence in the potential. Journal of Mathematical Physics, 23(2):258–266, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p258_s1.

Rosendorff:1982:GSO

[936] S. Rosendorff. Generalized second-order Coulomb phase shift functions. Journal of Mathematical Physics, 23(2):267–271, February 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p267_
s1.

Kloster:1982:PAM

[937] S. Kloster and A. Das. Pre-atlas manifolds and coordinate-free general relativity. Journal of Mathematical Physics, 23(2):272–275, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i2/p272_s1.

Lugo:1982:SAT

[938] Gabriel Lugo. Structure of asymptotic twistor space. Journal of Mathematical Physics, 23(2):276–282, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i2/p276_s1.

Persides:1982:TI

[939] S. Persides. Timelike infinity. Journal of Mathematical Physics, 23 (2):283–288, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i2/p283_s1.

Persides:1982:UFT

[940] S. Persides. A unified formulation of timelike, null and spatial infinity. Journal of Mathematical Physics, 23(2):289–292, February 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p289_
s1.

Nevin:1982:EET

[941] J. M. Nevin. Electromagnetic energy tensors and the Lorentz equation of motion for fields with electric and magnetic charge distributions. Journal of Mathematical Physics, 23(2):293–295, February 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i2/p293_s1.

Steiger:1982:CTR

[942] Ulrich Steiger and Ronald F. Fox. Coupled translational and rotational diffusion in liquids. Journal of Mathematical Physics, 23(2):296–309, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i2/p296_s1.

Gmitro:1982:TVR

[943] M. Gmitro and A. A. Ovchinnikova. Two variable relativistic tensor harmonics. Journal of Mathematical Physics, 23(2):310–314, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i2/p310_s1.

Swanson:1982:VGS

[944] D. G. Swanson. A 4-vector generalization of the sine-Gordon equation. Journal of Mathematical Physics, 23(2):315–317, February 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p315_
s1.

Charbonneau:1982:LRT

[945] M. Charbonneau, K. M. van Vliet, and P. Vasilopoulos. Linear response theory revisited. III. one-body response formulas and generalized Boltzmann equations. Journal of Mathematical Physics, 23(2):318–336, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i2/p318_s1.

Litvin:1982:TFC

[946] D. B. Litvin. Tensor fields on crystals. Journal of Mathematical Physics, 23(2):337–344, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i2/p337_s1.

Coleman:1982:EJB

[947] R. A. Coleman and H. Korte. Erratum: “Jet bundles and path structures”, [J. Math. Phys. 21 (1980), no. 6, 1340–1351; MR 82b:83050]. Journal of Mathematical Physics, 23(2):345, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i2/p345_s1. See

[215].

Bincer:1982:MLO

[948] Adam M. Bincer. Mickelsson lowering operators for the symplectic group. Journal of Mathematical Physics, 23(3):347–349, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p347_s1.

Blank:1982:BFR

[949] Jiří Blank, Miloslav Havlíček, Pavel Exner, and Wolfgang Lassner. Boson-fermion representations of Lie superalgebras: the example of osp(1, 2). Journal of Mathematical Physics, 23(3):350–353, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/
p350_s1.

Cant:1982:LAP

[950] A. Cant. Lie-algebraic properties of infinite-dimensional wave equations. Journal of Mathematical Physics, 23(3):354–356, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p354_s1.

Srivastava:1982:RCD

[951] H. M. Srivastava. Remarks on certain dual series equations involving the Konhauser biorthogonal polynomials. Journal of Mathematical Physics, 23(3):357, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i3/p357_s1.

Fuchssteiner:1982:BHS

[952] Benno Fuchssteiner and Walter Oevel. The bi-Hamiltonian structure of some nonlinear fifth- and seventh-order differential equations and recursion formulas for their symmetries and conserved covariants. Journal of Mathematical Physics, 23(3):358–363, March 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i3/p358_s1.

Kupershmidt:1982:SSI

[953] B. A. Kupershmidt. Singular symmetries of integrable curves and surfaces. Journal of Mathematical Physics, 23(3):364–366, March 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p364_
s1.

Muller-Kirsten:1982:RRC

[954] H. J. W. Müller-Kirsten and L. K. Sharma. Recurrence relations for the coefficients of perturbation expansions. Journal of Mathematical Physics, 23(3):367–372, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i3/p367_s1.

Nieh:1982:IRC

[955] H. T. Nieh and M. L. Yan. An identity in Riemann–Cartan geometry. Journal of Mathematical Physics, 23(3):373–374, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p373_s1.

Boyd:1982:TFG

[956] John P. Boyd. Theta functions, Gaussian series, and spatially periodic solutions of the Korteweg–de Vries equation. Journal of Mathematical Physics, 23(3):375–387, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i3/p375_s1.

Mitra:1982:SMB

[957] P. Mitra. String mechanics based on 2-forms. Journal of Mathematical Physics, 23(3):388–391, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i3/p388_s1.

Case:1982:SGM

[958] K. M. Case and A. M. Roos. Sine–Gordon and modified Korteweg–de Vries charges. Journal of Mathematical Physics, 23(3):392–395, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i3/p392_s1.

Krueger:1982:IPN

[959] Robert J. Krueger. Inverse problems for nonabsorbing media with discontinuous material properties. Journal of Mathematical Physics, 23(3): 396–404, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p396_s1.

Combe:1982:FPI

[960] Ph. Combe, R. Høegh-Krohn, R. Rodriguez, M. Sirugue, and M. Sirugue-Collin. Feynman path integral and Poisson processes with piecewise classical paths. Journal of Mathematical Physics, 23(3):405–411, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i3/p405_s1.

Chen:1982:TBC

[961] Augustine C. Chen. Theoretical basis for Coulomb matrix elements in the oscillator representation. Journal of Mathematical Physics, 23(3): 412–416, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p412_s1.

Nakamura:1982:SME

[962] Akira Nakamura. Simple multiple explode-decay mode solutions of a twodimensional nonlinear Schrödinger equation. Journal of Mathematical Physics, 23(3):417–420, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i3/p417_s1.

Karlsson:1982:CRR

[963] Bengt R. Karlsson. A combinatoric result related to the N-body problem. Journal of Mathematical Physics, 23(3):421–422, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p421_s1.

Urani:1982:GDE

[964] John R. Urani and Marilyn H. Kemp. A generalization of the Dirac equation to accelerating reference frames. Journal of Mathematical Physics, 23(3):423–424, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i3/p423_s1.

Letaw:1982:SCS

[965] John R. Letaw and Jonathan D. Pfautsch. The stationary coordinate systems in flat spacetime. Journal of Mathematical Physics, 23(3):425– 431, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p425_s1.

Geroch:1982:SBS

[966] Robert Geroch, Can Bin Liang, and Robert M. Wald. Singular boundaries of space–times. Journal of Mathematical Physics, 23(3):432– 435, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p432_s1.

McIntosh:1982:RTM

[967] C. B. G. McIntosh and W. D. Halford. The Riemann tensor, the metric tensor, and curvature collineations in general relativity. Journal of Mathematical Physics, 23(3):436–441, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i3/p436_s1.

Garbaczewski:1982:QSF

[968] Piotr Garbaczewski. Quantization of spinor fields. II. Meaning of “bosonization” in 1 + 1 and 1 + 3 dimensions. Journal of Mathematical Physics, 23(3):442–450, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i3/p442_s1.

Nilsson:1982:RSP

[969] Bengt E. W. Nilsson. Reduction of the super phase space for a massless Dirac particle. Journal of Mathematical Physics, 23(3):451–455, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i3/p451_s1.

Kaliappan:1982:CBI

[970] P. Kaliappan and M. Lakshmanan. Connection between the infinite sequence of Lie–Bäcklund symmetries of the Korteweg–de Vries and sine-Gordon equations. Journal of Mathematical Physics, 23(3):456– 459, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p456_s1.

Laedke:1982:CEP

[971] E. W. Laedke and K. H. Spatschek. Complementary energy principles in dissipative fluids. Journal of Mathematical Physics, 23(3):460– 463, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i3/p460_s1.

Lee:1982:LRO

[972] M. Howard Lee. Long-range order in the spin van der Waals model. Journal of Mathematical Physics, 23(3):464–471, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i3/p464_s1.

Antillon:1982:EED

[973] A. Antillón and T. H. Seligman. Explicit expressions for double Gel’fand states. Journal of Mathematical Physics, 23(4):473–480, April 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p473_
s1.

Sitaram:1982:GQUa

[974] B. R. Sitaram and K. C. Tripathy. Geometric quantization and UIRs of semisimple Lie groups. II. Discrete series. Journal of Mathematical Physics, 23(4):481–483, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p481_s1.

Sitaram:1982:GQUb

[975] B. R. Sitaram and K. C. Tripathy. Geometric quantization and UIRs of semisimple Lie groups. III. Principal and supplementary series. Journal of Mathematical Physics, 23(4):484–485, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i4/p484_s1.

Balantekin:1982:AEC

[976] A. Baha Balantekin. Anomalies and eigenvalues of Casimir operators for Lie groups and supergroups. Journal of Mathematical Physics, 23(4): 486–489, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i4/p486_s1.

Patera:1982:DOS

[977] J. Patera, C. Rousseau, and D. Schlomiuk. Dimensions of orbits and strata in complex and real classical Lie algebras. Journal of Mathematical Physics, 23(4):490–494, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p490_s1.

Fliess:1982:ANF

[978] Michel Fliess and Françoise Lamnabhi-Lagarrigue. Application of a new functional expansion to the cubic anharmonic oscillator. Journal of Mathematical Physics, 23(4):495–502, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p495_s1.

Reid:1982:NSH

[979] James L. Reid and John R. Ray. Nonlinear superposition, higher-order nonlinear equations, and classical linear invariants. Journal of Mathematical Physics, 23(4):503–509, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p503_s1.

Verbitskii:1982:OMS

[980] I. L. Verbitskiĭ. On one method of solving the Helmholtz equation. II. Journal of Mathematical Physics, 23(4):510–516, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p510_s1.

Harnad:1982:PLS

[981] J. Harnad and P. Winternitz. Pseudopotentials and Lie symmetries for the generalized nonlinear Schrödinger equation. Journal of Mathematical Physics, 23(4):517–525, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p517_s1.

Mola:1982:ERR

[982] E. E. Mola. Exact recursion relation for pseudobidimensional arrays of dumbbells. Journal of Mathematical Physics, 23(4):526–530, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i4/p526_s1.

Chang:1982:ASH

[983] Y. F. Chang, M. Tabor, and J. Weiss. Analytic structure of the Hénon– Heiles Hamiltonian in integrable and nonintegrable regimes. Journal of Mathematical Physics, 23(4):531–538, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i4/p531_s1.

Elskens:1982:CMNa

[984] Yves Elskens and David Speiser. Classical mechanics of nonspherical bodies. I. Binary collisions in two dimensions. Journal of Mathematical Physics, 23(4):539–546, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p539_s1.

Elskens:1982:CMNb

[985] Yves Elskens and David Speiser. Classical mechanics of nonspherical bodies. II. Boltzmann equation and H-theorem in two dimensions. Journal of Mathematical Physics, 23(4):547–551, April 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i4/p547_s1.

Katzin:1982:TDQ

[986] Gerald H. Katzin and Jack Levine. Time-dependent quadratic constants of motion, symmetries, and orbit equations for classical particle dynamical systems with time-dependent Kepler potentials. Journal of Mathematical Physics, 23(4):552–563, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p552_s1.

Kodama:1982:DIS

[987] Y. Kodama, M. J. Ablowitz, and J. Satsuma. Direct and inverse scattering problems of the nonlinear intermediate long wave equation. Journal of Mathematical Physics, 23(4):564–576, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i4/p564_s1.

Corones:1982:PAT

[988] J. Corones, B. DeFacio, and R. J. Krueger. Parabolic approximations to the time-independent elastic wave equation. Journal of Mathematical Physics, 23(4):577–586, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p577_s1.

Strand:1982:GRL

[989] Michael P. Strand and R. Stephen Berry. Group representations in the Liouville representation and the algebraic approach. Journal of Mathematical Physics, 23(4):587–593, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p587_s1.

Newton:1982:ISI

[990] Roger G. Newton. Inverse scattering. IV. three dimensions: generalized Marchenko construction with bound states, and generalized Gel’fand– Levitan equations. Journal of Mathematical Physics, 23(4):594–604, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i4/p594_s1.

Sakmar:1982:MSFa

[991] I. A. Sakmar. Maximum of the spin-flip cross section from unitarity and three constraints. Journal of Mathematical Physics, 23(4):605–610, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i4/p605_s1.

Geroch:1982:BTG

[992] Robert Geroch. Building things in general relativity. Journal of Mathematical Physics, 23(4):611–614, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p611_s1.

Cosgrove:1982:RBI

[993] Christopher M. Cosgrove. Relationship between the inverse scattering techniques of Belinskiĭ–Zakharov and Hauser–Ernst in general relativity. Journal of Mathematical Physics, 23(4):615–633, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p615_s1.

McDow:1982:CCI

[994] J. C. McDow and J. W. Moffat. Consistency of the Cauchy initial value problem in a nonsymmetric theory of gravitation. Journal of Mathematical Physics, 23(4):634–636, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p634_s1.

Krori:1982:KLM

[995] K. D. Krori and D. R. Bhattacharjee. Kerr-like metric in Brans–Dicke theory. Journal of Mathematical Physics, 23(4):637–638, April 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p637_
s1.

Turchetti:1982:SGT

[996] G. Turchetti. On the structure of the generalized Tjon–Wu equations. Journal of Mathematical Physics, 23(4):639–645, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p639_s1.

Courbage:1982:MPIa

[997] M. Courbage. Mathematical problems of irreversible statistical mechanics for quantum systems. I: Analytic continuation of the collision and destruction operators by spectral deformation method. Journal of Mathematical Physics, 23(4):646–651, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p646_s1.

Courbage:1982:MPIb

[998] M. Courbage. Mathematical problems of irreversible statistical mechanics for quantum systems. II: On the singularities of (Ψ(z) − z)^{−1} and the pseudo-Markovian equation. Application to Lee model. Journal of Mathematical Physics, 23(4):652–658, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p652_s1.

Cattaneo:1982:DCO

[999] U. Cattaneo. Densities of covariant observables. Journal of Mathematical Physics, 23(4):659–664, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p659_s1.

Datta:1982:RTC

[1000] D. P. Datta and S. Mukherjee. Regge trajectories in confining potentials. Journal of Mathematical Physics, 23(4):665–668, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i4/p665_s1. See

erratum [1428].

Holmes:1982:MMA

[1001] Philip J. Holmes and Jerrold E. Marsden. Melnikov’s method and Arnol′d diffusion for perturbations of integrable Hamiltonian systems. Journal of Mathematical Physics, 23(4):669–675, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i4/p669_s1.

Rosen:1982:GTB

[1002] Gerald Rosen. Global theorems for boundary-free viscous incompressible fluid flows of finite energy. Journal of Mathematical Physics, 23(4):676– 679, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i4/p676_s1.

Geroch:1982:DBH

[1003] Robert Geroch and James B. Hartle. Distorted black holes. Journal of Mathematical Physics, 23(4):680–692, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i4/p680_s1.

Newton:1982:EISa

[1004] Roger G. Newton. Erratum: “Inverse scattering. II. Three dimensions”, [J. Math. Phys. 21 (1980), no. 7, 1698–1715; MR 81m:81066b]. Journal of Mathematical Physics, 23(4):693, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i4/p693_s1. See [268].

Newton:1982:EISb

[1005] Roger G. Newton. Erratum: “Inverse Scattering. III. Three dimensions, continued”, [J. Math. Phys. 22, 2191 (1981)]. Journal of Mathematical Physics, 23(4):693, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i4/p693_s2. See [537].

Wald:1982:EES

[1006] Robert M. Wald and Ping Yip. Erratum: “On the existence of simultaneous synchronous coordinates in spacetimes with spacelike singularities”, [J. Math. Phys. 22, 2659 (1981)]. Journal of Mathematical Physics, 23 (4):693, April 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i4/p693_s3. See [843].

Newmarch:1982:CTC

[1007] J. D. Newmarch and R. M. Golding. The character table for the corepresentations of magnetic groups. Journal of Mathematical Physics, 23(5): 695–704, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p695_s1.

Patera:1982:COS

[1008] J. Patera and C. Rousseau. Complex orthogonal and symplectic matrices depending on parameters. Journal of Mathematical Physics, 23(5): 705–714, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p705_s1.

Kroger:1982:TPM

[1009] Helmut Kröger and Rainer Perne. A theorem on A-proper mappings and its application in scattering theory. Journal of Mathematical Physics, 23(5):715–719, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p715_s1.

Janssen:1982:BTZ

[1010] A. J. E. M. Janssen. Bargmann transform, Zak transform, and coherent states. Journal of Mathematical Physics, 23(5):720–731, May 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p720_
s1.

Rowe:1982:VFL

[1011] D. J. Rowe and A. Ryman. Valleys and fall lines on a Riemannian manifold. Journal of Mathematical Physics, 23(5):732–735, May 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p732_
s1.

Capri:1982:MMC

[1012] A. Z. Capri and M. Kobayashi. A mechanical model with constraints. Journal of Mathematical Physics, 23(5):736–741, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p736_s1.

Littlejohn:1982:HPT

[1013] Robert G. Littlejohn. Hamiltonian perturbation theory in noncanonical coordinates. Journal of Mathematical Physics, 23(5):742–747, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i5/p742_s1.

Weir:1982:EMP

[1014] G. J. Weir. Electromagnetic multipole propagation in a homogeneous conducting wholespace. Journal of Mathematical Physics, 23(5):748–752, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i5/p748_s1.

Claverie:1982:NSP

[1015] Pierre Claverie and Francisco Soto. Nonrecurrence of the stochastic process for the hydrogen atom problem in stochastic electrodynamics. Journal of Mathematical Physics, 23(5):753–759, May 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i5/p753_s1.

Battle:1982:PIC

[1016] Guy A. Battle III. Pseudoscalar interaction of coupled quantummechanical oscillators with independent Fermi systems. Journal of Mathematical Physics, 23(5):760–779, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i5/p760_s1.

Sharma:1982:SWE

[1017] L. K. Sharma, V. P. Iyer, Sheilly Mehta, S. V. Kulkarni, and Jyoti Choubey. Solutions of the wave equation for superposed potentials with application to charmonium spectroscopy. Journal of Mathematical Physics, 23(5):780–789, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p780_s1.

Bourgeois:1982:QMS

[1018] Brian Bourgeois. Quantum mechanical scattering theory for potentials of the form V (r) = (sin r)/r^{β}, 1/3 < β ≤ 1/2. Journal of Mathematical Physics, 23(5):790–797, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p790_s1.

Zohni:1982:GCR

[1019] O. Zohni. Generalized completeness relations in the theory of resonant scattering. Journal of Mathematical Physics, 23(5):798–802, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i5/p798_s1.

Kesarwani:1982:EAO

[1020] R. N. Kesarwani and Y. P. Varshni. Eigenvalues of an anharmonic oscillator. II. Journal of Mathematical Physics, 23(5):803–809, May 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p803_
s1.

Reula:1982:ETS

[1021] Oscar Reula. Existence theorem for solutions of Witten’s equation and nonnegativity of total mass. Journal of Mathematical Physics, 23(5): 810–814, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p810_s1.

Larsen:1982:LTN

[1022] Edward W. Larsen, G. C. Pomraning, and V. C. Badham. Linear transport in nonhomogeneous media. II. Journal of Mathematical Physics, 23(5):815–821, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p815_s1.

Yoshimura:1982:USD

[1023] Tetz Yoshimura. Unrenormalized Schwinger–Dyson equations and dynamical mass generation. Journal of Mathematical Physics, 23(5):822– 829, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p822_s1.

Henneaux:1982:RSS

[1024] Marc Henneaux. Remarks on spacetime symmetries and nonabelian gauge fields. Journal of Mathematical Physics, 23(5):830–833, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/
p830_s1.

Kovacs:1982:WS

[1025] Eve Kovacs. SU(3) wave solutions. Journal of Mathematical Physics, 23(5):834–838, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p834_s1.

Bonora:1982:SBT

[1026] L. Bonora, P. Pasti, and M. Tonin. Supermanifolds and BRS transformations. Journal of Mathematical Physics, 23(5):839–845, May 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p839_
s1.

Cooper:1982:HLT

[1027] Alan Cooper, Joel Feldman, and Lon Rosen. Higher Legendre transforms and their relationship to Bethe–Salpeter kernels and r-field projectors. Journal of Mathematical Physics, 23(5):846–868, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p846_s1.

Ross:1982:HGS

[1028] D. K. Ross. The holonomy group and spontaneous symmetry breaking. Journal of Mathematical Physics, 23(5):869–872, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i5/p869_s1.

Quiros:1982:TMM

[1029] M. Quirós, J. Ramírez Mittelbrunn, and E. Rodríguez. On the topological meaning of magnetic charge quantization. Journal of Mathematical Physics, 23(5):873–877, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p873_s1.

Deenen:1982:DGMa

[1030] J. Deenen and C. Quesne. Dynamical group of microscopic collective states. I. One-dimensional case. Journal of Mathematical Physics, 23(5): 878–889, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p878_s1.

Bruss:1982:EES

[1031] Anna R. Bruss. The eikonal equation: some results applicable to computer vision. Journal of Mathematical Physics, 23(5):890–896, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i5/p890_s1.

Grimes:1982:BRA

[1032] Dale M. Grimes. Biconical receiving antenna. Journal of Mathematical Physics, 23(5):897–914, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p897_s1.

Bleistein:1982:ENI

[1033] Norman Bleistein and Jack K. Cohen. Erratum: “Nonuniqueness in the inverse source problem in acoustics and electromagnetics”, [J. Math. Phys. 18 (1977), no. 2, 194–201; MR 55 #11812]. Journal of Mathematical Physics, 23(5):915, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i5/p915_s2. See [1].

Kriegsmann:1982:ELW

[1034] G. A. Kriegsmann. Erratum: “Linear wave conversion in a warm, unmagnetized, collisionless plasma”, [J. Math. Phys. 22, 2692 (1981)]. Journal of Mathematical Physics, 23(5):915, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i5/p915_s1. See [848].

Geronimo:1982:UBN

[1035] Jeffrey S. Geronimo. An upper bound on the number of eigenvalues of an infinite-dimensional Jacobi matrix. Journal of Mathematical Physics, 23(6):917–921, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p917_s1.

Chatterjee:1982:PPG

[1036] R. Chatterjee and T. Lulek. Permutational properties of the generalized Clebsch–Gordan coefficients. Journal of Mathematical Physics, 23(6): 922–927, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i6/p922_s1.

Chen:1982:NAP

[1037] Jin Quan Chen and Mei Juan Gao. A new approach to permutation group representation. Journal of Mathematical Physics, 23(6):928–943, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p928_s1.

Baake:1982:SRS

[1038] M. Baake, B. Gemünden, and R. Oedingen. Structure and representations of the symmetry group of the four-dimensional cube. Journal of Mathematical Physics, 23(6):944–953, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p944_s1. See erra-

tum [1281].

Inaba:1982:IDR

[1039] I. Inaba, T. Maekawa, and T. Yamamoto. Infinite-dimensional representations of the graded Lie algebra (Sp(4): 4). Representation of the para-Bose operators with real order of quantization. Journal of Mathematical Physics, 23(6):954–963, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i6/p954_s1.

vanHaeringen:1982:FSP

[1040] H. van Haeringen. A family of sums of products of Legendre functions. Journal of Mathematical Physics, 23(6):964–966, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p964_s1.

Csavinszky:1982:VCM

[1041] P. Csavinszky. Variational calculation of the multipole potential from an arbitrary localized charge distribution. Journal of Mathematical Physics, 23(6):967–968, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p967_s1.

Pecora:1982:NID

[1042] Louis M. Pecora. A nonstandard infinite-dimensional vector space approach to Gaussian functional measures. Journal of Mathematical Physics, 23(6):969–982, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p969_s1.

Berryman:1982:ABN

[1043] James G. Berryman and Charles J. Holland. Asymptotic behavior of the nonlinear diffusion equation n_{t} = (n^{−1}n_{x})_{x}. Journal of Mathematical Physics, 23(6):983–987, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p983_s1.

Berk:1982:NSL

[1044] H. L. Berk, William McCay Nevins, and K. V. Roberts. New Stokes’ line in WKB theory. Journal of Mathematical Physics, 23(6):988–1002, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p988_s1.

Behnam:1982:SMD

[1045] E. Behnam. Some more developments on operators in Kreĭn space. The exponential map. Journal of Mathematical Physics, 23(6):1003–1010, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1003_s1.

Luttinger:1982:NMA

[1046] J. M. Luttinger. A new method for the asymptotic evaluation of a class of path integrals. Journal of Mathematical Physics, 23(6):1011–1016, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1011_s1.

Papiez:1982:SOC

[1047] Lech Papież. Stochastic optimal control and quantum mechanics. Journal of Mathematical Physics, 23(6):1017–1019, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1017_s1.

Morato:1982:DDR

[1048] L. M. Morato. On the dynamics of diffusions and the related general electromagnetic potentials. Journal of Mathematical Physics, 23(6):1020– 1024, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i6/p1020_s1.

Russ:1982:PTDa

[1049] Muchere C. Russ and C. A. Hollingsworth. Properties of threedimensional Cartesian tensors. I. Some properties of irreducible tensors. Journal of Mathematical Physics, 23(6):1025–1027, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1025_s1.

Russ:1982:PTDb

[1050] Muchere C. Russ and C. A. Hollingsworth. Properties of threedimensional Cartesian tensors. II. Arbitrarily oriented tensors expressed in generalized spherical functions. Ranks through 4. Journal of Mathematical Physics, 23(6):1028–1030, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i6/p1028_s1.

Hollingsworth:1982:PTD

[1051] C. A. Hollingsworth and Muchere C. Russ. Properties of threedimensional Cartesian tensors. III. Concerning tensor averages. Journal of Mathematical Physics, 23(6):1031–1036, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1031_s1.

Holt:1982:CNI

[1052] Craig R. Holt. Construction of new integrable Hamiltonians in two degrees of freedom. Journal of Mathematical Physics, 23(6):1037–1046, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1037_s1.

Marx:1982:DGF

[1053] Egon Marx and Daniel Maystre. Dyadic Green functions for the timedependent wave equation. Journal of Mathematical Physics, 23(6):1047– 1056, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i6/p1047_s1.

Marx:1982:SIE

[1054] Egon Marx. Single integral equation for wave scattering. Journal of Mathematical Physics, 23(6):1057–1065, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i6/p1057_s1.

Fokas:1982:UIE

[1055] A. S. Fokas and R. L. Anderson. On the use of isospectral eigenvalue problems for obtaining hereditary symmetries for Hamiltonian systems. Journal of Mathematical Physics, 23(6):1066–1073, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1066_s1.

Cyranski:1982:QTP

[1056] John F. Cyranski. Quantum theory as a probability theory on Hilbert space. Journal of Mathematical Physics, 23(6):1074–1077, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/
p1074_s1.

Bach:1982:PFQ

[1057] Alexander Bach and Thomas Wenning. A probabilistic formulation of quantum theory. II. Journal of Mathematical Physics, 23(6):1078–1081, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1078_s1.

Huynh:1982:IQA

[1058] T. V. Huynh. Invariant *-quantization associated with the affine group. Journal of Mathematical Physics, 23(6):1082–1087, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1082_s1.

Iwai:1982:SGH

[1059] Toshihiro Iwai. The symmetry group of the harmonic oscillator and its reduction. Journal of Mathematical Physics, 23(6):1088–1092, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1088_s1.

Iwai:1982:QCK

[1060] Toshihiro Iwai. Quantization of the conformal Kepler problem and its application to the hydrogen atom. Journal of Mathematical Physics, 23(6):

1093–1099, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i6/p1093_s1.

Palev:1982:PBP

[1061] Tchavdar D. Palev. Para-Bose and para-Fermi operators as generators of orthosymplectic Lie superalgebras. Journal of Mathematical Physics, 23(6):1100–1102, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i6/p1100_s1.

Fernandez:1982:HPT

[1062] Francisco M. Fernández and Eduardo A. Castro. Hypervirialperturbational treatment of the bounded hydrogen atom. Journal of Mathematical Physics, 23(6):1103–1111, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i6/p1103_s1.

Ramm:1982:VPR

[1063] A. G. Ramm. Variational principles for resonances. II. Journal of Mathematical Physics, 23(6):1112–1114, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i6/p1112_s1.

Saenz:1982:SPS

[1064] A. W. Sáenz. Scattering by periodic surfaces. Journal of Mathematical Physics, 23(6):1115–1122, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i6/p1115_s1.

Ramm:1982:CMAa

[1065] A. G. Ramm. Convergence of the T -matrix approach to scattering theory. Journal of Mathematical Physics, 23(6):1123–1125, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1123_s1.

Boyer:1982:KVS

[1066] C. P. Boyer and J. D. Finley III. Killing vectors in self-dual, Euclidean Einstein spaces. Journal of Mathematical Physics, 23(6):1126–1130, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1126_s1.

Dodd:1982:LDP

[1067] R. K. Dodd and H. C. Morris. Linear deformation problems for the Ernst equation. Journal of Mathematical Physics, 23(6):1131–1136, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1131_s1.

Jantzen:1982:SSC

[1068] Robert T. Jantzen. Spinor sources in cosmology. Journal of Mathematical Physics, 23(6):1137–1146, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i6/p1137_s1.

Tod:1982:SDK

[1069] K. P. Tod. Self-dual Kerr–Schild metrics and null Maxwell fields. Journal of Mathematical Physics, 23(6):1147–1148, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1147_s1.

McIntosh:1982:SAV

[1070] C. B. G. McIntosh and E. H. van Leeuwen. Spacetimes admitting a vector field whose inner product with the Riemann tensor is zero. Journal of Mathematical Physics, 23(6):1149–1152, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i6/p1149_s1.

Birss:1982:MDF

[1071] R. R. Birss. The metric dependence of four-dimensional formulations of electromagnetism. Journal of Mathematical Physics, 23(6):1153–1154, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1153_s1.

Garrido:1982:CES

[1072] L. Garrido and J. Masoliver. On a class of exact solutions to the Fokker– Planck equations. Journal of Mathematical Physics, 23(6):1155–1158, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1155_s1.

vandenBerg:1982:BCI

[1073] M. van den Berg. On boson condensation into an infinite number of lowlying levels. Journal of Mathematical Physics, 23(6):1159–1161, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1159_s1.

Percus:1982:HSO

[1074] J. K. Percus. Higher spin one-dimensional Ising lattice in arbitrary external field. Journal of Mathematical Physics, 23(6):1162–1167, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1162_s1.

Garczynski:1982:QRF

[1075] Włodzimierz Garczyński and Jerzy Stelmach. Quasicontinual representation of a field given on a finite lattice. Interpolation properties. Journal of Mathematical Physics, 23(6):1168–1174, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1168_s1.

Letelier:1982:SSS

[1076] Patrício S. Letelier. Soliton solutions for self-dual SU(2) gauge fields on Euclidean space. Journal of Mathematical Physics, 23(6):1175–1178, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i6/p1175_s1.

Hostler:1982:IRD

[1077] Levere C. Hostler. An SL(2, C)-invariant representation of the Dirac equation. Journal of Mathematical Physics, 23(6):1179–1184, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/
p1179_s1.

Sharma:1982:HEA

[1078] L. K. Sharma and V. P. Iyer. High-energy asymptotic expansion and Regge trajectories for the general even power potential in the Klein– Gordon equation with applications. Journal of Mathematical Physics, 23(6):1185–1189, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p1185_s1.

SabbaStefanescu:1982:DPS

[1079] I. Sabba Stefănescu. On the determination of πN phase shifts from isospin constraints and fixed t analyticity. Journal of Mathematical Physics, 23(6):1190–1209, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i6/p1190_s1.

Koh:1982:CSM

[1080] I. G. Koh, Y. D. Kim, Y. J. Park, C. H. Kim, and Y. S. Kim. Complete set of SU(5) monopole solutions. Journal of Mathematical Physics, 23(6): 1210–1212, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i6/p1210_s1.

Fante:1982:SEW

[1081] Ronald L. Fante. Scattering of electromagnetic waves from random media with multiple scattering included. Journal of Mathematical Physics, 23(6):1213–1222, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p1213_s1.

DeMeyer:1982:SOT

[1082] H. E. De Meyer, G. Vanden Berghe, and J. Van der Jeugt. Shift operator techniques for the classification of multipole-photon states. X. P_{l}^{0} eigenstate and eigenvalue determination in G_{2}. Journal of Mathematical Physics, 23(6):1223–1227, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i6/p1223_s1.

Nakamura:1982:BEC

[1083] Mitsunobu Nakamura and Makoto Mizuno. Bounds for effective conductivity of heterogeneous materials by the finite element method. Journal of Mathematical Physics, 23(6):1228–1230, June 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1228_s1.

Pfeifer:1982:EAN

[1084] Peter Pfeifer. Erratum: “Adjoints of nondensely defined Hilbert space operators”, [J. Math. Phys. 22, 1619 (1981)]. Journal of Mathematical Physics, 23(6):1231, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i6/p1231_s2. See [693].

Salingaros:1982:ECC

[1085] Nikos Salingaros. Erratum: “On the classification of Clifford algebras and their relation to spinors in n dimensions”, [J. Math. Phys. 23, 1 (1982)]. Journal of Mathematical Physics, 23(6):1231, June 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i6/p1231_

s1. See [896].

Sarma:1982:DBI

[1086] C. R. Sarma. Determination of basis for the irreducible representations of the unitary group for U(p+q) ↓ U(p)⊗U(q). Journal of Mathematical Physics, 23(7):1235–1238, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1235_s1.

Balantekin:1982:BRS

[1087] A. Baha Balantekin and Itzhak Bars. Branching rules for the supergroup SU(N/M) from those of SU(N + M). Journal of Mathematical Physics, 23(7):1239–1247, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1239_s1.

Forest:1982:STP

[1088] M. Gregory Forest and David W. McLaughlin. Spectral theory for the periodic sine-Gordon equation: A concrete viewpoint. Journal of Mathematical Physics, 23(7):1248–1277, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i7/p1248_s1.

Newberger:1982:NSR

[1089] Barry S. Newberger. New sum rule for products of Bessel functions with application to plasma physics. Journal of Mathematical Physics, 23(7): 1278–1281, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v23/i7/p1278_s1. See erratum [1608] and comments [1886].

Hermann:1982:SAG

[1090] Robert Hermann. Some algebraic, geometric, and system-theoretic properties of the Special Functions of mathematical physics. Journal of Mathematical Physics, 23(7):1282–1294, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i7/p1282_s1.

Fischer:1982:NSS

[1091] Elliot Fischer. New similarity solutions for the Ernst equations with electromagnetic fields. Journal of Mathematical Physics, 23(7):1295– 1296, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1295_s1.

Aldaya:1982:QCS

[1092] Víctor Aldaya and José A. de Azcárraga. Quantization as a consequence of the symmetry group: An approach to geometric quantization. Journal of Mathematical Physics, 23(7):1297–1305, July 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i7/p1297_s1.

Fine:1982:JDQ

[1093] Arthur Fine. Joint distributions, quantum correlations, and commuting observables. Journal of Mathematical Physics, 23(7):1306–1310, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i7/p1306_s1.

Boiteux:1982:TNC

[1094] M. Boiteux. Theory of nonbijective canonical transformations in mechanics: Application to the Coulomb problem. Journal of Mathematical Physics, 23(7):1311–1314, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1311_s1.

Pesquera:1982:QAO

[1095] L. Pesquera and P. Claverie. The quartic anharmonic oscillator in stochastic electrodynamics. Journal of Mathematical Physics, 23(7): 1315–1322, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1315_s1.

Singh:1982:CFT

[1096] Virendra Singh, S. N. Biswas, and K. Datta. Continued fraction theory of the rotating harmonic oscillator. Journal of Mathematical Physics, 23(7): 1323–1326, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1323_s1.

Truini:1982:IQM

[1097] P. Truini and L. C. Biedenharn. An E_{6} ⊗ U(1) invariant quantum mechanics for a Jordan pair. Journal of Mathematical Physics, 23(7):1327– 1345, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1327_s1.

Vaidyanathan:1982:LIT

[1098] R. Vaidyanathan. Linear invariants of a time-dependent quantal oscillator. Journal of Mathematical Physics, 23(7):1346–1348, July 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i7/p1346_
s1.

Plebanski:1982:SAD

[1099] J. F. Plebański and G. F. Torres del Castillo. HH spaces with an algebraically degenerate right side. Journal of Mathematical Physics, 23(7): 1349–1352, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1349_s1.

Bogoyavlensky:1982:ISS

[1100] O. I. Bogoyavlensky and G. Moschetti. The investigation of some self-similar solutions of Einstein’s equations. Journal of Mathematical Physics, 23(7):1353–1358, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1353_s1.

Guo:1982:EGN

[1101] Dong Sheng Guo and Frederick J. Ernst. Electrovac generalization of Neugebauer’s N = 2 solution of the Einstein vacuum field equations. Journal of Mathematical Physics, 23(7):1359–1363, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i7/p1359_s1.

Ibanez:1982:NES

[1102] J. Ibañez and J. L. Sanz. New exact static solutions to Einstein’s equations for spherically symmetric perfect fluid distributions. Journal of Mathematical Physics, 23(7):1364–1365, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i7/p1364_s1.

Choudhury:1982:SGF

[1103] S. B. Dutta Choudhury and D. Bhattacharya. Static gravitational fields in a general class of scalar-tensor theories. Journal of Mathematical Physics, 23(7):1366–1369, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1366_s1.

Alicki:1982:PIS

[1104] Robert Alicki. Path integrals and stationary phase approximation for quantum dynamical semigroups. Quadratic systems. Journal of Mathematical Physics, 23(7):1370–1375, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i7/p1370_s1.

Siewert:1982:IPP

[1105] C. E. Siewert and W. L. Dunn. On inverse problems for plane-parallel media with nonuniform surface illumination. Journal of Mathematical Physics, 23(7):1376–1378, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i7/p1376_s1.

Schlafly:1982:CNG

[1106] R. Schlafly. A Chern number for gauge fields on R^{4}. Journal of Mathematical Physics, 23(7):1379–1384, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i7/p1379_s1.

Kamin:1982:NTE

[1107] Shoshana Kamin and Philip Rosenau. Nonlinear thermal evolution in an inhomogeneous medium. Journal of Mathematical Physics, 23(7):1385– 1390, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i7/p1385_s1.

Desmier:1982:ASF

[1108] P. E. Desmier, R. T. Sharp, and J. Patera. Analytic SU(3) states in a finite subgroup basis. Journal of Mathematical Physics, 23(8):1393– 1398, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1393_s1.

Klimyk:1982:SME

[1109] Anatoli U. Klimyk and Bruno Gruber. Structure and matrix elements of the degenerate series representations of U(p + q) and U(p, q) in a U(p) × U(q) basis. Journal of Mathematical Physics, 23(8):1399– 1408, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1399_s1.

Lemire:1982:CCF

[1110] F. W. Lemire and J. Patera. Congruence classes of finite representations of simple Lie superalgebra. Journal of Mathematical Physics, 23(8):1409– 1414, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1409_s1.

Levine:1982:SLA

[1111] R. Y. Levine and Y. Tomozawa. Supersymmetry and Lie algebras. Journal of Mathematical Physics, 23(8):1415–1421, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1415_s1.

Nakamura:1982:GSS

[1112] Akira Nakamura. General superposition of solitons and various ripplons of a two-dimensional nonlinear Schrödinger equation. Journal of Mathematical Physics, 23(8):1422–1426, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i8/p1422_s1.

Kupershmidt:1982:KVS

[1113] B. A. Kupershmidt. Korteweg–de Vries surfaces and Bäcklund curves. Journal of Mathematical Physics, 23(8):1427–1432, August 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1427_
s1.

Jain:1982:IES

[1114] D. L. Jain and R. P. Kanwal. Interior and exterior solutions for boundary value problems in composite elastic media. I. Two-dimensional problems. Journal of Mathematical Physics, 23(8):1433–1443, August 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1433_
s1.

Bostrom:1982:TDS

[1115] Anders Boström. Time-dependent scattering by a bounded obstacle in three dimensions. Journal of Mathematical Physics, 23(8):1444– 1450, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1444_s1.

Asorey:1982:QEP

[1116] M. Asorey, J. F. Cariñena, and M. Paramio. Quantum evolution as a parallel transport. Journal of Mathematical Physics, 23(8):1451–1458, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i8/p1451_s1.

Parthasarathy:1982:FPI

[1117] K. R. Parthasarathy and K. B. Sinha. Feynman path integrals of operator-valued maps. Journal of Mathematical Physics, 23(8):1459– 1462, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1459_s1.

GarciaD:1982:STP

[1118] Alberto García D. and Jerzy F. Plebański. Solutions of type D possessing a group with null orbits as contractions of the seven-parameter solution. Journal of Mathematical Physics, 23(8):1463–1465, August 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1463_
s1.

Cervero:1982:ESC

[1119] José M. Cerveró. Explicit solutions of the conformal scalar equations in arbitrary dimensions. Journal of Mathematical Physics, 23(8):1466– 1470, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1466_s1.

Luehr:1982:TBG

[1120] C. P. Luehr and M. P. Rosenbaum. Twistor bundles and gauge action principles of gravitation. Journal of Mathematical Physics, 23(8):1471– 1488, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1471_s1.

Mohanty:1982:PMC

[1121] Udayan Mohanty. Properties of Mayer cluster expansion. Journal of Mathematical Physics, 23(8):1489–1492, August 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1489_s1.

Haba:1982:CSP

[1122] Z. Haba. Continuity of sample paths and weak compactness of measures in Euclidean field theory. Journal of Mathematical Physics, 23(8):1493– 1500, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i8/p1493_s1.

Kalinowski:1982:CCH

[1123] M. W. Kalinowski, M. Seweryński, and L. Szymanowski. Comment on calculations of higher-order tadpoles and triangle vertex integral. Journal of Mathematical Physics, 23(8):1501–1503, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1501_s1.

Hoyos:1982:SBS

[1124] J. Hoyos, M. Quirós, J. Ramírez Mittelbrunn, and F. J. de Urríes. Superfiber bundle structure of gauge theories with Faddeev–Popov fields. Journal of Mathematical Physics, 23(8):1504–1510, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1504_s1.

Binegar:1982:RFT

[1125] Birne Binegar. Relativistic field theories in three dimensions. Journal of Mathematical Physics, 23(8):1511–1517, August 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1511_s1.

MartinezAlonso:1982:NSE

[1126] L. Martínez Alonso. The nonlinear Schrödinger equation as a Galileaninvariant dynamical system. Journal of Mathematical Physics, 23 (8):1518–1523, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i8/p1518_s1.

Faber:1982:ATT

[1127] V. Faber, A. B. White, Jr., and G. M. Wing. An analysis of Taylor’s theory of toroidal plasma relaxation. Journal of Mathematical Physics, 23(8):1524–1537, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i8/p1524_s1.

Folland:1982:FSA

[1128] N. O. Folland and Junaidah Osman. Finite sum approximations to Brillouin zone integrals with symmetrized plane waves. Journal of Mathematical Physics, 23(8):1538–1546, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i8/p1538_s1.

Matsushita:1982:EEC

[1129] Yasuo Matsushita. Erratum: “On Euler characteristics of compact Einstein 4-manifolds of metric signature (+ + −−)”, [J. Math. Phys. 22 (1981), no. 5, 979–982; MR 82j:53106]. Journal of Mathematical Physics, 23(8):1547, August 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i8/p1547_s1. See [591].

Reboucas:1982:ETD

[1130] Marcelo J. Rebouças and J. Ademir S. de Lima. Erratum: “Timedependent, finite, rotating universes”, [J. Math. Phys. 22, 2699 (1981)]. Journal of Mathematical Physics, 23(8):1547, August 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i8/p1547_

s2. See [849].

Berenson:1982:MGF

[1131] Rhoda Berenson. Molien generating functions, invariants and covariants of magnetic point groups. Journal of Mathematical Physics, 23(9):1551– 1559, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1551_s1.

Bystricky:1982:GSH

[1132] J. Bystricky, R. Gaskell, J. Patera, and R. T. Sharp. Generalized SU(2) spherical harmonics. Journal of Mathematical Physics, 23(9):1560–1565, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1560_s1.

Leibbrandt:1982:BGS

[1133] George Leibbrandt, Shein Shion Wang, and Nader Zamani. Bäcklund generated solutions of Liouville’s equation and their graphical representations in three spatial dimensions. Journal of Mathematical Physics, 23(9):1566–1572, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i9/p1566_s1.

Rzewuski:1982:PRS

[1134] Jan Rzewuski. On the projections of representation spaces of the symmetry group on the Minkowski space. Journal of Mathematical Physics, 23(9):1573–1576, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i9/p1573_s1.

Yasue:1982:CCT

[1135] Kunio Yasue. On a certain class of two-sided continuous local semimartingales: Toward a sample-wise characterization of the Nelson process. Journal of Mathematical Physics, 23(9):1577–1583, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1577_s1.

Yamada:1982:JDM

[1136] K. Yamada and S. Shieh. On Jacobi’s decomposition of the motion of a heavy symmetrical top into the motions of two free triaxial tops. Journal of Mathematical Physics, 23(9):1584–1588, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1584_s1.

Cantrijn:1982:VFG

[1137] F. Cantrijn. Vector fields generating invariants for classical dissipative systems. Journal of Mathematical Physics, 23(9):1589–1595, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1589_s1.

Augusteijn:1982:RPP

[1138] Marijke F. Augusteijn and Ernst Breitenberger. Root parities and phase behavior in the slow-fluctuation technique. Journal of Mathematical Physics, 23(9):1596–1609, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i9/p1596_s1.

Gallone:1982:SRS

[1139] Franco Gallone and Antonio Sparzani. Structure results for the Segal quantization of Fermi systems. Journal of Mathematical Physics, 23 (9):1610–1616, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i9/p1610_s1.

Faulkner:1982:MSJ

[1140] John R. Faulkner. Measurement systems and Jordan algebras. Journal of Mathematical Physics, 23(9):1617–1621, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1617_s1.

Demiralp:1982:ULB

[1141] Metin Demiralp and N. Abdülbaki Baykara. Upper and lower bounds to zeroth order Coulombic hyperangular interaction integrals. Journal of Mathematical Physics, 23(9):1622–1627, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1622_s1.

Roychoudhury:1982:MDS

[1142] R. K. Roychoudhury. Mass dependence of Schrödinger wavefunctions for an exponential potential. Journal of Mathematical Physics, 23(9):1628– 1629, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1628_s1.

Kolassis:1982:EST

[1143] Charalampos A. Kolassis. On the effect of space–time isometries on the neutrino field. Journal of Mathematical Physics, 23(9):1630–1638, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1630_s1.

Bonnor:1982:MTS

[1144] W. B. Bonnor and M. A. H. MacCallum. The Melnick–Tabensky solutions have high symmetry. Journal of Mathematical Physics, 23(9):1639– 1640, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1639_s1.

Lim:1982:MCE

[1145] P. H. Lim. Manifestly covariant equations of motion for a particle in an external field. Journal of Mathematical Physics, 23(9):1641–1646, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1641_s1.

Kuchar:1982:CSP

[1146] Karel Kuchař. Conditional symmetries in parametrized field theories. Journal of Mathematical Physics, 23(9):1647–1661, September 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1647_
s1.

Chyba:1982:TDE

[1147] Christopher F. Chyba. Time-dependent embeddings for Schwarzschildlike solutions to the gravitational field equations. Journal of Mathematical Physics, 23(9):1662–1667, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i9/p1662_s1.

Isham:1982:STY

[1148] C. J. Isham and G. Kunstatter. Spatial topology and Yang–Mills vacua. Journal of Mathematical Physics, 23(9):1668–1677, September 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1668_
s1.

Steiger:1982:BOR

[1149] Ulrich R. Steiger and Ronald F. Fox. Boson operator representation of Brownian motion. Journal of Mathematical Physics, 23(9):1678–1687, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1678_s1.

Hughes:1982:LDR

[1150] Barry D. Hughes and Michael F. Shlesinger. Lattice dynamics, random walks, and nonintegral effective dimensionality. Journal of Mathematical Physics, 23(9):1688–1692, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i9/p1688_s1.

Jacquot:1982:LIE

[1151] J. L. Jacquot and M. Umezawa. Lorentz invariance of the extended object. Journal of Mathematical Physics, 23(9):1693–1696, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1693_s1.

VanAlstine:1982:SIS

[1152] Peter Van Alstine and Horace Crater. Scalar interactions of supersymmetric relativistic spinning particles. Journal of Mathematical Physics, 23(9):1697–1699, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i9/p1697_s1.

Talukdar:1982:CMN

[1153] B. Talukdar, D. K. Ghosh, and T. Sasakawa. Coulomb-modified nuclear scattering. I. Journal of Mathematical Physics, 23(9):1700–1707, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1700_s1.

Przezdziecki:1982:REF

[1154] S. Przeździecki and W. Laprus. On the representation of electromagnetic fields in gyrotropic media in terms of scalar Hertz potentials. Journal of Mathematical Physics, 23(9):1708–1712, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1708_s1.

Jarem:1982:PSF

[1155] John M. Jarem. The pattern space factor and quality factor of cylindrical source antennas. Journal of Mathematical Physics, 23(9):1713–1716, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i9/p1713_s1.

Chen:1982:PTO

[1156] M. Chen. On the phase transition of the one-dimensional Percus–Yevick equation for an arbitrary potential of finite range. Journal of Mathematical Physics, 23(9):1717–1722, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i9/p1717_s1.

Akyildiz:1982:CLS

[1157] Yılmaz Akyıldız. Conservation laws for shallow water waves on a sloping beach. Journal of Mathematical Physics, 23(9):1723–1727, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1723_s1.

Vatsya:1982:ASC

[1158] S. R. Vatsya. On approximating the solutions of the Chandrasekhar Hequation. Journal of Mathematical Physics, 23(9):1728–1731, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1728_s1.

Sanz:1982:DIV

[1159] J. L. Sanz. Dynamical importance of vorticity and shear in the universe. Journal of Mathematical Physics, 23(9):1732–1736, September 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1732_
s1.

Antoine:1982:EIT

[1160] J.-P. Antoine and W. Karwowski. Erratum: “Interpolation theory and refinement of nested Hilbert spaces”, [J. Math. Phys. 22, 2489 (1981)]. Journal of Mathematical Physics, 23(9):1737, September 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i9/p1737_

s1. See [821].

Crandall:1982:EGS

[1161] R. E. Crandall and Mary Hall Reno. Erratum: “Ground state energy bounds for potentials |x|^{ν}”, [J. Math. Phys. 23 (1982), no. 1, 64–70; MR 82m:81021]. Journal of Mathematical Physics, 23(9):1737, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i9/p1737_s2. See [903].

Saraykar:1982:ELS

[1162] R. V. Saraykar and N. E. Joshi. Erratum: “Linearization stability of Einstein equations coupled with self-gravitating scalar fields”, [J. Math. Phys. 22 (1981), no. 2, 343–347; MR 82f:83010]. Journal of Mathematical Physics, 23(9):1738, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v23/i9/p1738_s1. See [489].

Jantzen:1982:GQS

[1163] Robert T. Jantzen. Generalized quaternions and spacetime symmetries. Journal of Mathematical Physics, 23(10):1741–1746, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/
p1741_s1.

Fairbairn:1982:SCF

[1164] W. M. Fairbairn and T. Fulton. Some comments on finite subgroups of SU(3). Journal of Mathematical Physics, 23(10):1747–1748, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1747_s1.

Puta:1982:SSP

[1165] Mircea Puta. Some spectral properties of the Kostant complex. Journal of Mathematical Physics, 23(10):1749–1751, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1749_s1.

Ciulli:1982:ACD

[1166] S. Ciulli and T. D. Spearman. Analytic continuation from data points with unequal errors. Journal of Mathematical Physics, 23(10):1752– 1764, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1752_s1.

Wilhelm:1982:IBV

[1167] H. E. Wilhelm. Initial-boundary-value problem for diffusion of magnetic fields into conductors with external electromagnetic transients. Journal of Mathematical Physics, 23(10):1765–1768, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1765_s1.

Fernandez:1982:SCA

[1168] Francisco M. Fernández and Eduardo A. Castro. Some comments about the tensor virial theorem and orthogonal linear transformations. Journal of Mathematical Physics, 23(10):1769–1770, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1769_s1.

Arthur:1982:GTC

[1169] M. D. Arthur and K. M. Case. Gradient theorem for completely integrable Hamiltonian systems. Journal of Mathematical Physics, 23 (10):1771–1777, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i10/p1771_s1.

Palev:1982:WAQ

[1170] Tchavdar D. Palev. Wigner approach to quantization. Noncanonical quantization of two particles interacting via a harmonic potential. Journal of Mathematical Physics, 23(10):1778–1784, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1778_s1.

Emch:1982:QCM

[1171] Gérard G. Emch. Quantum and classical mechanics on homogeneous Riemannian manifolds. Journal of Mathematical Physics, 23(10):1785– 1791, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1785_s1.

Ghirardi:1982:QEP

[1172] G. C. Ghirardi, A. Rimini, and T. Weber. Quantum evolution in the presence of additive conservation laws and the quantum theory of measurement. Journal of Mathematical Physics, 23(10):1792–1796, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1792_s1.

Klauder:1982:CMS

[1173] John R. Klauder. Constructing measures for spin-variable path integrals. Journal of Mathematical Physics, 23(10):1797–1805, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/
p1797_s1.

Daubechies:1982:CMP

[1174] Ingrid Daubechies and John R. Klauder. Constructing measures for path integrals. Journal of Mathematical Physics, 23(10):1806–1822, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1806_s1.

Benoist-Gueutal:1982:PSB

[1175] P. Benoist-Gueutal and M. L’Huillier. Properties of solutions for Nbody Yakubovskiĭ-Faddeev equations. Journal of Mathematical Physics, 23(10):1823–1834, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i10/p1823_s1.

Mansour:1982:PTA

[1176] H. M. M. Mansour and H. J. W. Müller-Kirsten. Perturbative technique as an alternative to the WKB method applied to the double-well potential. Journal of Mathematical Physics, 23(10):1835–1845, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1835_s1.

Krori:1982:DTM

[1177] K. D. Krori and D. R. Bhattacharjee. Demiański-type metric in Brans– Dicke theory. Journal of Mathematical Physics, 23(10):1846–1847, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1846_s1.

Maassen:1982:IMM

[1178] Hans Maassen. On the invertibility of Møller morphisms. Journal of Mathematical Physics, 23(10):1848–1851, October 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1848_s1.

Haber:1982:RBE

[1179] Howard E. Haber and H. Arthur Weldon. On the relativistic Bose– Einstein integrals. Journal of Mathematical Physics, 23(10):1852–1858, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1852_s1.

Boffi:1982:RIS

[1180] V. C. Boffi and G. Spiga. Rigorous iterated solutions to a nonlinear integral evolution problem in particle transport theory. Journal of Mathematical Physics, 23(10):1859–1864, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i10/p1859_s1.

Ceccatto:1982:IPC

[1181] H. Ceccatto, A. Foussats, H. Giacomini, and O. Zandron. On the inconsistency of a photon creation mechanism in an expanding universe. Journal of Mathematical Physics, 23(10):1865–1872, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1865_s1.

Yao:1982:MAA

[1182] Te Hai Yao. On mapping approaches in axiomatic quantum field theory. Journal of Mathematical Physics, 23(10):1873–1880, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/
p1873_s1.

Bongaarts:1982:MEA

[1183] P. J. M. Bongaarts. Maxwell’s equations in axiomatic quantum field theory. II. Covariant and noncovariant gauges. Journal of Mathematical Physics, 23(10):1881–1898, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i10/p1881_s1.

Cooper:1982:SLT

[1184] Alan Cooper, Joel Feldman, and Lon Rosen. The second Legendre transform for the weakly coupled P (φ)_{2} model. Journal of Mathematical Physics, 23(10):1899–1916, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i10/p1899_s1.

Atkinson:1982:NCQ

[1185] D. Atkinson, P. W. Johnson, and K. Stam. Nonperturbative confinement in quantum chromodynamics. II. Mandelstam’s gluon propagator. Journal of Mathematical Physics, 23(10):1917–1924, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1917_s1.

Bracken:1982:LCIa

[1186] A. J. Bracken and Barry Jessup. Local conformal-invariance of the wave equation for finite-component fields. I. The conditions for invariance, and fully-reducible fields. Journal of Mathematical Physics, 23(10):1925– 1946, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1925_s1.

Bracken:1982:LCIb

[1187] A. J. Bracken and Barry Jessup. Local conformal-invariance of the wave equation for finite-component fields. II. Classification of relevant indecomposable fields. Journal of Mathematical Physics, 23(10):1947– 1952, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1947_s1.

DeWitt:1982:ECE

[1188] Bryce S. DeWitt and Peter van Nieuwenhuizen. Explicit construction of the exceptional superalgebras F(4) and G(3). Journal of Mathematical Physics, 23(10):1953–1963, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i10/p1953_s1.

Negi:1982:SSE

[1189] O. P. S. Negi and B. S. Rajput. Study of superluminal electromagnetic fields. Journal of Mathematical Physics, 23(10):1964–1967, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i10/p1964_s1.

Ogata:1982:FSE

[1190] H. Ogata and J. Asai. Finite-size effects on the relativistic electrons in a Coulomb field. I. Journal of Mathematical Physics, 23(10):1968– 1978, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1968_s1.

Kerrighan:1982:UEM

[1191] D. B. Kerrighan. On the uniqueness of the energy-momentum tensor for electromagnetism. Journal of Mathematical Physics, 23(10):1979– 1980, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i10/p1979_s1.

Srivastava:1982:QTF

[1192] S. K. Srivastava. Quantum tachyon in a Friedmann universe. Journal of Mathematical Physics, 23(10):1981–1983, October 1982. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1981_s1.

Matzner:1982:IDW

[1193] Richard A. Matzner, M. Rosenbaum, and M. P. Ryan, Jr. Initial data and wave propagation in one-dimensional inhomogeneous cosmologies. Journal of Mathematical Physics, 23(10):1984–1991, October 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/p1984_s1.

Bovier:1982:EFS

[1194] A. Bovier, M. Lüling, and D. Wyler. Erratum: “Finite subgroups of SU(3)”, [J. Math. Phys. 22 (1981), no. 8, 1543–1547; MR 83b:20009b]. Journal of Mathematical Physics, 23(10):1992, October 1982. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i10/

p1992_s1. See [684].

Gerry:1982:PIC

[1195] Christopher C. Gerry and Steven Silverman. Path integral for coherent states of the dynamical group SU(1, 1). Journal of Mathematical Physics, 23(11):1995–2003, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i11/p1995_s1.

Deenen:1982:DGMb

[1196] J. Deenen and C. Quesne. Dynamical group of microscopic collective states. II. Boson representations in d dimensions. Journal of Mathematical Physics, 23(11):2004–2015, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i11/p2004_s1.

Gaskell:1982:VWI

[1197] R. Gaskell and R. T. Sharp. G_{2} van der Waerden invariant. Journal of Mathematical Physics, 23(11):2016–2018, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/p2016_s1.

Wu:1982:CGD

[1198] Alfred C. T. Wu. Cartan–Gram determinants for the simple Lie groups. Journal of Mathematical Physics, 23(11):2019–2021, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/
p2019_s1.

OReilly:1982:CFP

[1199] Michael F. O’Reilly. A closed formula for the product of irreducible representations of SU(3). Journal of Mathematical Physics, 23(11):2022– 2028, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2022_s1.

Doria:1982:AFS

[1200] Mauro M. Doria. Addition formula for the Z_{N} × Z_{N} symmetric solutions of the factorization equations. Journal of Mathematical Physics, 23(11):2029–2032, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i11/p2029_s1.

Fokas:1982:UAT

[1201] A. S. Fokas and M. J. Ablowitz. On a unified approach to transformations and elementary solutions of Painlevé equations. Journal of Mathematical Physics, 23(11):2033–2042, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i11/p2033_s1.

Strampp:1982:EPH

[1202] W. Strampp. The equivalence problem for the heat equation. Journal of Mathematical Physics, 23(11):2043–2047, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/p2043_s1.

Nishigori:1982:STM

[1203] Takeo Nishigori. Space–time memory functions and solution of nonlinear evolution equations. Journal of Mathematical Physics, 23(11):2048– 2052, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2048_s1.

Gravel:1982:DSF

[1204] Pierre Gravel. Discontinuous solutions for first-order systems through a limiting process. Journal of Mathematical Physics, 23(11):2053–2058, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2053_s1.

Davies:1982:LAE

[1205] Ian Davies and Aubrey Truman. Laplace asymptotic expansions of conditional Wiener integrals and generalized Mehler kernel formulas. Journal of Mathematical Physics, 23(11):2059–2070, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/p2059_s1.

Estabrook:1982:MFP

[1206] Frank B. Estabrook. Moving frames and prolongation algebras. Journal of Mathematical Physics, 23(11):2071–2076, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/p2071_s1.

Hermann:1982:DGS

[1207] Robert Hermann. The differential geometric structure of general mechanical systems from the Lagrangian point of view. Journal of Mathematical Physics, 23(11):2077–2089, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i11/p2077_s1.

Johnson:1982:ETA

[1208] G. W. Johnson. The equivalence of two approaches to the Feynman integral. Journal of Mathematical Physics, 23(11):2090–2096, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i11/p2090_s1.

Kelley:1982:AMS

[1209] C. T. Kelley. Approximate methods for the solution of the Chandrasekhar H-equation. Journal of Mathematical Physics, 23(11):2097– 2100, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2097_s1.

Henneaux:1982:LSS

[1210] Marc Henneaux and L. C. Shepley. Lagrangians for spherically symmetric potentials. Journal of Mathematical Physics, 23(11):2101–2107, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2101_s1.

Lukacs:1982:STT

[1211] B. Lukács, Z. Perjés, Á. Sebestyén, E. T. Newman, and J. Porter. Structure of three-twistor particles. Journal of Mathematical Physics, 23 (11):2108–2115, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i11/p2108_s1.

MartinezAlonso:1982:TII

[1212] L. Martínez Alonso and E. Olmedilla. Trace identities in the inverse scattering transform method associated with matrix Schrödinger operators. Journal of Mathematical Physics, 23(11):2116–2121, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i11/p2116_s1.

Johnston:1982:LMW

[1213] George L. Johnston. Langer’s method for weakly bound states of the Helmholtz equation with symmetric profiles. Journal of Mathematical Physics, 23(11):2122–2126, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i11/p2122_s1.

Prosser:1982:FSI

[1214] Reese T. Prosser. Formal solutions of inverse scattering problems. IV. Error estimates. Journal of Mathematical Physics, 23(11):2127–2130, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2127_s1.

Bosanac:1982:PTI

[1215] S. Bosanac. Perturbation theory of inelastic resonances. Journal of Mathematical Physics, 23(11):2131–2139, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i11/p2131_s1.

Eftimiu:1982:DIS

[1216] C. Eftimiu. Direct and inverse scattering by a sphere of variable index of refraction. Journal of Mathematical Physics, 23(11):2140–2146, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2140_s1.

Kolassis:1982:NSS

[1217] Charalampos A. Kolassis. On the nonexistence of static solutions of the Einstein–Weyl equations in general relativity. Journal of Mathematical Physics, 23(11):2147–2150, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i11/p2147_s1.

Gowdy:1982:GPT

[1218] Robert H. Gowdy. Geometrical perturbation theory: action-principle surface terms in homogeneous cosmology. Journal of Mathematical Physics, 23(11):2151–2154, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i11/p2151_s1. See erratum [1429].

Ray:1982:CSG

[1219] Dipankar Ray. A class of solutions of the generalized Lund–Regge model. Journal of Mathematical Physics, 23(11):2155–2158, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/
p2155_s1.

Czapor:1982:OTI

[1220] S. R. Czapor and R. G. McLenaghan. Orthogonal transitivity, invertibility and null geodesic separability in type D vacuum solutions of Einstein’s field equations with cosmological constant. Journal of Mathematical Physics, 23(11):2159–2167, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i11/p2159_s1.

Ashtekar:1982:NMF

[1221] Abhay Ashtekar and Amitabha Sen. NUT 4-momenta are forever. Journal of Mathematical Physics, 23(11):2168–2178, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/
p2168_s1.

Wodkiewicz:1982:TNL

[1222] K. Wódkiewicz. Theory of nonlinear Langevin equation with quadratic noise. Journal of Mathematical Physics, 23(11):2179–2184, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i11/p2179_s1.

Hock:1982:OSI

[1223] J. L. Hock, P. E. Licato, and R. B. McQuistan. The occupational statistics for indistinguishable trimers on a 3 × N lattice space. Journal of Mathematical Physics, 23(11):2185–2189, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i11/p2185_s1.

Stefanescu:1982:CRA

[1224] I. Sabba Stefanescu. On the convergence of reflectionless approximants to confining potentials. Journal of Mathematical Physics, 23(11):2190– 2204, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2190_s1.

Cacuci:1982:FND

[1225] Dan G. Cacuci. On the finiteness of the number of discrete eigenvalues in neutron transport theory. Journal of Mathematical Physics, 23(11):2205– 2214, November 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i11/p2205_s1.

Bodine:1982:GFI

[1226] C. Bodine and R. W. Gaskell. Generating functions for IR multiplicities. Journal of Mathematical Physics, 23(12):2217–2220, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2217_s1.

Warchall:1982:IAG

[1227] Henry A. Warchall. Implementation of automorphism groups in certain representations of the canonical commutation relations. Journal of Mathematical Physics, 23(12):2221–2228, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2221_s1.

Bose:1982:DFD

[1228] A. K. Bose and A. Navon. Decomposition of the finite-dimensional fermion algebra into irreducible spaces. Journal of Mathematical Physics, 23(12):2229–2233, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2229_s1.

Fernandez:1982:LIW

[1229] Ariel Fernández and Oktay Sinanoğlu. The lifting of an Inönü–Wigner contraction at the level of universal coverings. Journal of Mathematical Physics, 23(12):2234–2235, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i12/p2234_s1.

Hurni:1982:IRS

[1230] J.-P. Hurni and B. Morel. Irreducible representations of the superalgebras type II. Journal of Mathematical Physics, 23(12):2236–2243, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2236_s1.

Schellekens:1982:RTP

[1231] A. N. Schellekens, In-Gyu Koh, and Kyungsik Kang. Reduction of tensor products with definite permutation symmetry: Embeddings of irreducible representations of Lie groups into fundamental representations of SU(M) and branchings. Journal of Mathematical Physics, 23(12):2244– 2256, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2244_s1.

Newton:1982:GHP

[1232] Roger G. Newton. On a generalized Hilbert problem. Journal of Mathematical Physics, 23(12):2257–2265, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2257_s1.

Schwartz:1982:CDI

[1233] Charles Schwartz. A class of discontinuous integrals involving Bessel functions. Journal of Mathematical Physics, 23(12):2266–2267, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2266_s1.

Nasr:1982:CMO

[1234] A. H. Nasr. The commutant of a multiplication operator. Journal of Mathematical Physics, 23(12):2268–2270, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/p2268_s1.

Andrade:1982:CEL

[1235] R. F. S. Andrade. Carleman embedding and Lyapunov exponents. Journal of Mathematical Physics, 23(12):2271–2275, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/p2271_s1.

Martin:1982:FAL

[1236] Pablo Martín and Jorge Zamudio-Cristi. Fractional approximations for linear first-order differential equations with polynomial coefficients — application to E_{1}(x). Journal of Mathematical Physics, 23(12):2276– 2280, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2276_s1.

Gascon:1982:PFI

[1237] F. G. Gascón, F. B. Ramos, and E. Aguirre-Daban. On the polynomial first integrals of certain second-order differential equations. Journal of Mathematical Physics, 23(12):2281–2285, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/p2281_s1.

Jaulent:1982:ISP

[1238] M. Jaulent. The inverse scattering problem for LCRG transmission lines. Journal of Mathematical Physics, 23(12):2286–2290, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2286_s1.

Manolitsakis:1982:TDS

[1239] I. Manolitsakis. Two-dimensional scattered fields: A description in terms of the zeros of entire functions. Journal of Mathematical Physics, 23 (12):2291–2298, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2291_s1.

Boffi:1982:GSN

[1240] V. C. Boffi and G. Spiga. Global solution to a nonlinear integral evolution problem in particle transport theory. Journal of Mathematical Physics, 23(12):2299–2303, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2299_s1.

Betounes:1982:NFS

[1241] David E. Betounes. Nonclassical fields with singularities on a moving surface. Journal of Mathematical Physics, 23(12):2304–2311, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2304_s1.

Bampi:1982:IPC

[1242] F. Bampi and A. Morro. The inverse problem of the calculus of variations applied to continuum physics. Journal of Mathematical Physics, 23(12):2312–2321, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2312_s1.

Sakmar:1982:MSFb

[1243] I. A. Sakmar. Maximum of the spin-flip cross section from unitarity and four constraints. Journal of Mathematical Physics, 23(12):2322– 2327, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2322_s1.

Rosengaus:1982:RLP

[1244] E. Rosengaus and R. L. Dewar. Renormalized Lie perturbation theory. Journal of Mathematical Physics, 23(12):2328–2338, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/
p2328_s1.

Herrera:1982:CNS

[1245] L. Herrera and J. Jiménez. The complexification of a nonrotating sphere: An extension of the Newman–Janis algorithm. Journal of Mathematical Physics, 23(12):2339–2345, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i12/p2339_s1.

Caviglia:1982:GDF

[1246] Giacomo Caviglia, Franco Salmistraro, and Clara Zordan. Geodesic deviation and first integrals of motion. Journal of Mathematical Physics, 23(12):2346–2352, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2346_s1.

Matsuo:1982:CPP

[1247] Kuniaki Matsuo, Bahaa E. A. Saleh, and Malvin Carl Teich. Cascaded Poisson processes. Journal of Mathematical Physics, 23(12):2353–2364, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2353_s1.

Hill:1982:GIM

[1248] Robert Nyden Hill and Barton D. Huxtable. A generating integral for matrix elements of the Coulomb Green’s function. Journal of Mathematical Physics, 23(12):2365–2370, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2365_s1.

Lewis:1982:DAF

[1249] H. Ralph Lewis and P. G. L. Leach. A direct approach to finding exact invariants for one-dimensional time-dependent classical Hamiltonians. Journal of Mathematical Physics, 23(12):2371–2374, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2371_s1.

Kocak:1982:LHS

[1250] Huseyin Kocak. Linear Hamiltonian systems are integrable with quadratics. Journal of Mathematical Physics, 23(12):2375–2380, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/
p2375_s1.

Gudder:1982:MHS

[1251] S. P. Gudder, G. T. Rüttimann, and R. J. Greechie. Measurements, Hilbert space and quantum logics. Journal of Mathematical Physics, 23 (12):2381–2386, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2381_s1.

Nasr:1982:OMS

[1252] A. H. Nasr. Observables measured simultaneously with the potential. Journal of Mathematical Physics, 23(12):2387–2388, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/
p2387_s1.

Sutherland:1982:JDI

[1253] Roderick I. Sutherland. Joint distribution indicated by the wave equations of quantum mechanics. Journal of Mathematical Physics, 23 (12):2389–2392, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2389_s1.

Bhakar:1982:IPR

[1254] B. S. Bhakar. Impact parameter representation for scattering involving arbitrary spins. Journal of Mathematical Physics, 23(12):2393–2401, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2393_s1.

Gerry:1982:SCR

[1255] Christopher C. Gerry and Akira Inomata. Semiclassical calculation of a radial path integral. Journal of Mathematical Physics, 23(12):2402– 2407, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2402_s1.

Ramm:1982:CMAb

[1256] A. G. Ramm. Convergence of the T -matrix approach in the potential scattering. Journal of Mathematical Physics, 23(12):2408–2409, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2408_s1.

Ashtekar:1982:LHN

[1257] Abhay Ashtekar and Jeffrey Winicour. Linkages and Hamiltonians at null infinity. Journal of Mathematical Physics, 23(12):2410–2417, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2410_s1.

Lugrin:1982:FIT

[1258] Ch. Lugrin and Ph. A. Martin. Functional integration treatment of one-dimensional ionic mixtures. Journal of Mathematical Physics, 23 (12):2418–2429, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2418_s1.

Hioe:1982:MMS

[1259] F. T. Hioe. On the mean motion and some statistical properties of a quasiperiodic observable in a fermion-boson model. Journal of Mathematical Physics, 23(12):2430–2438, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2430_s1.

Aguilera-Navarro:1982:FCS

[1260] V. C. Aguilera-Navarro, E. Ley Koo, M. de Llano, S. M. Peltier, and A. Plastino. Fermion clustering in a soluble one-dimensional many-body system. Journal of Mathematical Physics, 23(12):2439–2444, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2439_s1.

Lavenda:1982:OCD

[1261] B. H. Lavenda. Optimal control of the decay of nonequilibrium statistical correlations. Journal of Mathematical Physics, 23(12):2445–2451, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2445_s1.

Santamato:1982:ST

[1262] E. Santamato and B. H. Lavenda. The stochastic H-theorem. Journal of Mathematical Physics, 23(12):2452–2460, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v23/i12/p2452_s1.

Muller:1982:MFA

[1263] V. F. Müller and W. Rühl. Mean field approximations for U(N) and SU(N) lattice gauge theories. Journal of Mathematical Physics, 23 (12):2461–2466, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2461_s1.

Kyriakopoulos:1982:RSY

[1264] E. Kyriakopoulos. Real solutions of the Yang–Mills field equations in Minkowski space. Journal of Mathematical Physics, 23(12):2467–2471, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2467_s1.

Hogan:1982:ESS

[1265] P. A. Hogan. An embedding of some Stiefel bundles. Journal of Mathematical Physics, 23(12):2472–2474, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2472_s1.

Coleman:1982:SMY

[1266] Robert Alan Coleman and Robert B. Mann. Spacetime metric and Yang– Mills fields unified in a Galilean subspace structure. Journal of Mathematical Physics, 23(12):2475–2487, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2475_s1.

Hou:1982:AMO

[1267] Bo Yu Hou, Bo Yuan Hou, and Pei Wang. Angular momentum operator in a nonabelian gauge field of an N-dimensional monopole and an SO(N) top operator. Journal of Mathematical Physics, 23(12):2488– 2493, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2488_s1.

Krebs:1982:SSNa

[1268] Jerome Krebs. Scattering by slightly nonspherical particles in the high frequency limit. I. Impenetrable particles. Journal of Mathematical Physics, 23(12):2494–2501, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i12/p2494_s1.

Krebs:1982:SSNb

[1269] Jerome Krebs. Scattering by slightly nonspherical particles in the high frequency limit. II. Transparent particles. Journal of Mathematical Physics, 23(12):2502–2510, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v23/i12/p2502_s1.

Pohlmeyer:1982:LMB

[1270] K. Pohlmeyer. Large momentum behavior of the Feynman amplitudes in the φ_{4}^{4}-theory. Journal of Mathematical Physics, 23(12):2511–2519, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2511_s1.

Fradkin:1982:RAL

[1271] David M. Fradkin. Representation of the active Lorentz transformation for particle dynamics. Journal of Mathematical Physics, 23(12):2520– 2523, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2520_s1.

Jordan:1982:WHP

[1272] Thomas F. Jordan. What happens to the position, spin, and parity when the mass goes to zero. Journal of Mathematical Physics, 23(12):2524– 2528, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2524_s1.

Eichten:1982:AFC

[1273] Estia Eichten, Kyungsik Kang, and In-Gyu Koh. Anomaly-free complex representations in SU(N). Journal of Mathematical Physics, 23 (12):2529–2536, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2529_s1.

Castanos:1982:CSS

[1274] O. Castaños, A. Frank, E. Chacón, P. Hess, and M. Moshinsky. Complete set of states for microscopic nuclear collective models. Journal of Mathematical Physics, 23(12):2537–2553, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v23/i12/p2537_s1.

Raychaudhuri:1982:VFI

[1275] A. K. Raychaudhuri and S. K. Saha. Viscous fluid interpretation of electromagnetic fields. II. Journal of Mathematical Physics, 23(12):2554– 2557, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2554_s1.

Bagrov:1982:EML

[1276] V. G. Bagrov, D. M. Githman, and V. N. Shapovalov. Electron motion in longitudinal electromagnetic fields. Journal of Mathematical Physics, 23(12):2558–2561, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2558_s1.

Balanis:1982:ISD

[1277] George N. Balanis. Inverse scattering: determination of inhomogeneities in sound speed. Journal of Mathematical Physics, 23(12):2562–2568, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2562_s1.

Sinha:1982:SSM

[1278] S. K. Sinha, Y. S. Sainger, and Y. Singh. Semiclassical statistical mechanics of a two-dimensional fluid. Journal of Mathematical Physics, 23 (12):2569–2581, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i12/p2569_s1.

Rosen:1982:ETS

[1279] Gerald Rosen. Exact theory for the self-similarity and decay of homogeneous turbulence. Journal of Mathematical Physics, 23(12):2582–2584, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2582_s1.

Cade:1982:EDL

[1280] R. Cade. The electric double layer and electrohydrostatic surface deformation. Journal of Mathematical Physics, 23(12):2585–2594, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v23/i12/p2585_s1.

Baake:1982:ESR

[1281] M. Baake, B. Gemünden, and R. Oedingen. Erratum: “Structure and representations of the symmetry group of the four-dimensional cube”, [J. Math. Phys. 23, 944 (1982)]. Journal of Mathematical Physics, 23(12): 2595, December 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i12/p2595_s1. See [1038].

Kasperkovitz:1983:TAR

[1282] P. Kasperkovitz. Type-adapted representations of semidirect product groups. Journal of Mathematical Physics, 24(1):1–7, January 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p1_s1.

Kasperkovitz:1983:TAS

[1283] P. Kasperkovitz. Type-adapted subduction matrices. Journal of Mathematical Physics, 24(1):8–13, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i1/p8_s1.

Liu:1983:NMF

[1284] Zhe Ming Liu. A new mathematical function connected with boundary value problems in kinetic transport theory. Journal of Mathematical Physics, 24(1):14–21, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p14_s1.

Collins:1983:NFS

[1285] C. B. Collins. Null field solutions of the wave equation and certain generalizations. Journal of Mathematical Physics, 24(1):22–28, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p22_s1.

Hermann:1983:DLG

[1286] Robert Hermann. Deformation of Lie group representations and linear filters. I. Journal of Mathematical Physics, 24(1):29–32, January 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p29_s1.

Bahar:1983:PIN

[1287] E. Bahar. The paths of integration for the new generalized Bessel transform. Journal of Mathematical Physics, 24(1):33–35, January 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p33_s1.

Matsushita:1983:THI

[1288] Yasuo Matsushita. Thorpe–Hitchin inequality for compact Einstein 4manifolds of metric signature (+ + −−) and the generalized Hirzebruch index formula. Journal of Mathematical Physics, 24(1):36–40, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p36_s1.

Isaacson:1983:CBT

[1289] D. Isaacson, E. L. Isaacson, D. Marchesin, and P. J. Paes-Leme. Critical behavior of the two-state doubling algorithm. Journal of Mathematical Physics, 24(1):41–45, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p41_s1.

Zagrodzinski:1983:DAP

[1290] J. Zagrodziński. Direct approach to the periodic solutions of the multidimensional sine-Gordon equation. Journal of Mathematical Physics, 24(1):46–52, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p46_s1.

Liverani:1983:EAB

[1291] C. Liverani and G. Turchetti. Existence and asymptotic behavior of Padé approximants to the Korteweg–de Vries multisoliton solutions. Journal of Mathematical Physics, 24(1):53–64, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i1/p53_s1.

Burke:1983:MPI

[1292] William L. Burke. Manifestly parity invariant electromagnetic theory and twisted tensors. Journal of Mathematical Physics, 24(1):65–69, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p65_s1.

Lowdin:1983:SPP

[1293] Per-Olov Löwdin. On the stability problem of a pair of adjoint operators. Journal of Mathematical Physics, 24(1):70–87, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p70_s1.

Froelich:1983:HFS

[1294] Piotr Froelich and Per-Olov Löwdin. On the Hartree–Fock scheme for a pair of adjoint operators. Journal of Mathematical Physics, 24(1):88–96, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soto:1983:WWF

[1295] Francisco Soto and Pierre Claverie. When is the Wigner function of multidimensional systems nonnegative? Journal of Mathematical Physics, 24(1):97–100, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i1/p97_s1.

Demiralp:1983:CFA

[1296] Metin Demiralp. A characteristic function approach to the discrete spectrum of electrically charged particles. Journal of Mathematical Physics, 24(1):101–106, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p101_s1.

Bosanac:1983:PEM

[1297] S. Bosanac. Perturbation expansion of S matrix for background scattering. Journal of Mathematical Physics, 24(1):107–114, January 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p107_
s1.

Pechenick:1983:ISC

[1298] Kay R. Pechenick and Jeffrey M. Cohen. Inverse scattering with coinciding-pole reflection coefficients. Journal of Mathematical Physics, 24(1):115–119, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p115_s1.

Ferraris:1983:CAB

[1299] M. Ferraris, M. Francaviglia, and C. Reina. A constructive approach to bundles of geometric objects on a differentiable manifold. Journal of Mathematical Physics, 24(1):120–124, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i1/p120_s1.

Hojman:1983:ESS

[1300] R. Hojman and A. Smailagic. Exact solutions of strong gravity in generalized metrics. Journal of Mathematical Physics, 24(1):125–128, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p125_s1.

Chakraborty:1983:GTU

[1301] S. K. Chakraborty and N. Bandyopadhyay. Gödel-type universe with a perfect fluid and a scalar field. Journal of Mathematical Physics, 24(1):129–132, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i1/p129_s1.

Lai:1983:ESI

[1302] C. K. Lai. Existence of solutions of integral equations in the thermodynamics of one-dimensional fermions with repulsive delta function potential. Journal of Mathematical Physics, 24(1):133–137, January 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p133_
s1.

Pule:1983:FBG

[1303] J. V. Pulè. The free boson gas in a weak external potential. Journal of Mathematical Physics, 24(1):138–142, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i1/p138_s1.

Baker:1983:RGH

[1304] George A. Baker, Jr. Renormalization group hypothesis for critical phenomena theory. Journal of Mathematical Physics, 24(1):143–147, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p143_s1.

Lohe:1983:EUG

[1305] M. A. Lohe and John van der Hoek. Existence and uniqueness of generalized vortices. Journal of Mathematical Physics, 24(1):148–153, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p148_s1.

Stacey:1983:CPP

[1306] Richard Stacey. Compactification partly proven. Journal of Mathematical Physics, 24(1):154–156, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i1/p154_s1.

Hurni:1983:IR

[1307] J.-P. Hurni and B. Morel. Irreducible representations of SU(m/n). Journal of Mathematical Physics, 24(1):157–163, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p157_s1. See

erratum [1607].

Sitaram:1983:RGL

[1308] B. R. Sitaram and K. C. Tripathy. Representations of graded Lie groups. Journal of Mathematical Physics, 24(1):164–165, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/p164_s1.

Kleeman:1983:AMQ

[1309] R. Kleeman. Aspects of modular quantization. Journal of Mathematical Physics, 24(1):166–172, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i1/p166_s1.

Chang:1983:MPA

[1310] K. C. Chang, C. K. Chew, T. Y. Liang, K. K. Phua, and H. B. Low. Mathematical properties and asymptotic expansion of the generalized quark statistical function. Journal of Mathematical Physics, 24(1):173– 176, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i1/p173_s1.

Brannigan:1983:SPC

[1311] M. Brannigan and D. Eyre. Splines and the projection collocation method for solving integral equations in scattering theory. Journal of Mathematical Physics, 24(1):177–183, January 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i1/p177_s1. See erra-

tum [1722].

Conlon:1983:RTF

[1312] Joseph G. Conlon. A refinement of the Thomas–Fermi approximation for the N-body problem. Journal of Mathematical Physics, 24(1):184– 192, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i1/p184_s1.

Rosen:1983:FIN

[1313] Gerald Rosen. Functional integrals in Navier–Stokes incompressible fluid turbulence. Journal of Mathematical Physics, 24(1):193–195, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p193_s1.

Soundalgekar:1983:OMF

[1314] V. M. Soundalgekar and J. P. Bhat. Oscillatory magnetohydrodynamic flow and heat transfer between conducting plates. Journal of Mathematical Physics, 24(1):196–201, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i1/p196_s1.

Gumbs:1983:AES

[1315] G. Gumbs. Analysis of the effect of surfaces on the tricritical behavior of systems. Journal of Mathematical Physics, 24(1):202–208, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i1/p202_s1.

Hall:1983:EST

[1316] George L. Hall. Electrostatic structural transitions in a Yukawa–Wigner solid. Journal of Mathematical Physics, 24(1):209–214, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i1/
p209_s1.

Collins:1983:CSS

[1317] C. B. Collins. Comments on the static spherically symmetric cosmologies of Ellis, Maartens, and Nel. Journal of Mathematical Physics, 24(1):215– 219, January 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i1/p215_s1.

Wene:1983:GCI

[1318] G. P. Wene. A generalization of the construction of Ilamed and Salingaros. Journal of Mathematical Physics, 24(2):221–223, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/
p221_s1.

Klimyk:1983:RGB

[1319] Anatoli U. Klimyk. Representations of the groups Sp(n, R) and Sp(n) in a U(n) basis. Journal of Mathematical Physics, 24(2):224–232, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i2/p224_s1.

Nikam:1983:RIP

[1320] R. S. Nikam, K. V. Dinesha, and C. R. Sarma. Reduction of innerproduct representations of unitary groups. Journal of Mathematical Physics, 24(2):233–238, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i2/p233_s1. See erratum [1430].

Daubechies:1983:ITR

[1321] I. Daubechies, A. Grossmann, and J. Reignier. An integral transform related to quantization. II. Some mathematical properties. Journal of Mathematical Physics, 24(2):239–254, February 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i2/p239_s1.

Davies:1983:LAE

[1322] Ian Davies and Aubrey Truman. On the Laplace asymptotic expansion of conditional Wiener integrals and the Bender–Wu formula for x^{2N}-anharmonic oscillators. Journal of Mathematical Physics, 24(2):255– 266, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i2/p255_s1.

Volkov:1983:GFP

[1323] V. S. Volkov and V. N. Pokrovsky. Generalized Fokker–Planck equation for non-Markovian processes. Journal of Mathematical Physics, 24(2):267–270, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i2/p267_s1.

Chodos:1983:PUH

[1324] Alan Chodos and Charles M. Sommerfield. Practical use of the Hamilton– Jacobi equation. Journal of Mathematical Physics, 24(2):271–275, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i2/p271_s1.

Arnal:1983:CGI

[1325] D. Arnal, J. C. Cortet, P. Molin, and G. Pinczon. Covariance and geometrical invariance in ∗ quantization. Journal of Mathematical Physics, 24(2):276–283, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i2/p276_s1.

Kowalski:1983:ISS

[1326] K. L. Kowalski and A. Picklesimer. Interaction-set scattering equations. Journal of Mathematical Physics, 24(2):284–290, February 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/p284_
s1.

Chaichian:1983:IFM

[1327] Masud Chaichian, Jan Fischer, N. F. Nelipa, and Ivo Vrkoč. Improvement of the Froissart–Martin bound for complex scattering angles. Journal of Mathematical Physics, 24(2):291–299, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/p291_s1.

Caderni:1983:CEC

[1328] N. Caderni, M. Martellini, and F. Salmistraro. Conditions of existence for a class of self-dual solutions of the Einstein equations. Journal of Mathematical Physics, 24(2):300–302, February 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i2/p300_s1.

Moreschi:1983:RSC

[1329] Osvaldo M. Moreschi and George A. J. Sparling. On Riemannian spaces with conformal symmetries or a tool for the study of generalized Kaluza– Klein theories. Journal of Mathematical Physics, 24(2):303–310, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i2/p303_s1.

Dumcke:1983:CMC

[1330] R. Dümcke. Convergence of multitime correlation functions in the weak and singular coupling limits. Journal of Mathematical Physics, 24(2):311–315, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i2/p311_s1.

Boon:1983:RNL

[1331] M. Boon, J. Zak, and I. J. Zucker. Rational von Neumann lattices. Journal of Mathematical Physics, 24(2):316–323, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/p316_s1.

Hall:1983:GTE

[1332] Richard L. Hall. A geometrical theory of energy trajectories in quantum mechanics. Journal of Mathematical Physics, 24(2):324–335, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i2/p324_s1.

Hongler:1983:ERD

[1333] M.-O. Hongler and W. M. Zheng. Exact results for the diffusion in a class of asymmetric bistable potentials. Journal of Mathematical Physics, 24(2):336–340, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i2/p336_s1.

Garbaczewski:1983:QSF

[1334] Piotr Garbaczewski. Quantization of spinor fields. III. Fermions on coherent (Bose) domains. Journal of Mathematical Physics, 24(2):341– 346, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i2/p341_s1.

Gielerak:1983:VGM

[1335] Roman Gielerak. Verification of the global Markov property in some class of strongly coupled exponential interactions. Journal of Mathematical Physics, 24(2):347–355, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i2/p347_s1.

Kerner:1983:MFB

[1336] Richard Kerner. Multiple fiber bundles and gauge theories of higher order. Journal of Mathematical Physics, 24(2):356–360, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/
p356_s1.

Kerner:1983:GGT

[1337] Richard Kerner and Ema Manuela da Silva Maia. Graded gauge theories over supersymmetric space. Journal of Mathematical Physics, 24(2):361– 368, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i2/p361_s1.

Tomozawa:1983:NTB

[1338] Yukio Tomozawa. Normalization of the three-body Bethe–Salpeter wavefunction for protons. Journal of Mathematical Physics, 24(2):369–374, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i2/p369_s1.

Kalinowski:1983:DRP

[1339] M. W. Kalinowski, M. Seweryński, and L. Szymanowski. On the dimensional regularization procedure for massless Feynman integrals. Journal of Mathematical Physics, 24(2):375–376, February 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i2/p375_s1.

Choquet-Bruhat:1983:HFA

[1340] Yvonne Choquet-Bruhat and Antonio Greco. High frequency asymptotic solutions of Yang–Mills and associated fields. Journal of Mathematical Physics, 24(2):377–379, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i2/p377_s1.

Royer:1983:EAG

[1341] Antoine Royer. Expansions around the Gaussian and Lorentzian limits for randomly perturbed spectral lines and their shift, width, and asymmetry. Journal of Mathematical Physics, 24(2):380–396, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i2/p380_s1.

Tran:1983:RFM

[1342] Loc Binh Tran and J. N. Huffaker. Recursive formulas for Morseoscillator matrix elements of arbitrary powers of 1 − exp[−a(r − r_{e})]. Journal of Mathematical Physics, 24(2):397–398, February 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i2/p397_
s1.

Gross:1983:PIA

[1343] Eugene P. Gross. Path integral approach to multiple scattering. Journal of Mathematical Physics, 24(2):399–405, February 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i2/p399_s1.

Pechenick:1983:ESV

[1344] Kay R. Pechenick and Jeffrey M. Cohen. Exact solutions to the valley problem in inverse scattering. Journal of Mathematical Physics, 24(2):406–409, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i2/p406_s1.

Kim:1983:UTPa

[1345] Shoon K. Kim. A unified theory of the point groups. II. The general projective representations and their application to space groups. Journal of Mathematical Physics, 24(3):411–413, March 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i3/p411_s1.

Kim:1983:UTPb

[1346] Shoon K. Kim. A unified theory of the point groups. III. Classification and basis functions of improper point groups. Journal of Mathematical Physics, 24(3):414–418, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i3/p414_s1.

Kim:1983:UTPc

[1347] Shoon K. Kim. A unified theory of the point groups. IV. The general corepresentations of the crystallographic and noncrystallographic Shubnikov point groups. Journal of Mathematical Physics, 24(3):419– 423, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p419_s1.

Sullivan:1983:PCC

[1348] John J. Sullivan. Phase conventions consistent with duality between SUn and S_{L}. Journal of Mathematical Physics, 24(3):424–440, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i3/p424_s1.

Newmarch:1983:RAG

[1349] J. D. Newmarch and R. M. Golding. The Racah algebra for groups with time reversal symmetry. III. Journal of Mathematical Physics, 24(3): 441–450, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p441_s1.

Newmarch:1983:SS

[1350] J. D. Newmarch. On the symmetries of the 6j symbol. Journal of Mathematical Physics, 24(3):451–456, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p451_s1.

Pinczon:1983:MNR

[1351] Georges Pinczon and Jacques C. H. Simon. Models of nonlinear representations and examples of linearization techniques. Journal of Mathematical Physics, 24(3):457–462, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p457_s1.

Syed:1983:EAPa

[1352] Ansaruddin Syed. Euler angle parametrization of the complex rotation group and its subgroups. I. Journal of Mathematical Physics, 24(3):463– 470, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p463_s1.

Syed:1983:EAPb

[1353] Ansaruddin Syed. Euler angle parametrization of the complex rotation group and its subgroups. II. Journal of Mathematical Physics, 24(3): 471–477, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p471_s1.

Basu:1983:CGC

[1354] Debabrata Basu and Kurt Bernardo Wolf. The Clebsch–Gordan coefficients of the three-dimensional Lorentz algebra in the parabolic basis. Journal of Mathematical Physics, 24(3):478–500, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p478_s1.

Benguria:1983:ULB

[1355] Rafael Benguria and M. Cristina Depassier. Upper and lower bounds for eigenvalues of nonlinear elliptic equations. I. The lowest eigenvalue. Journal of Mathematical Physics, 24(3):501–503, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p501_s1.

Abellanas:1983:EEH

[1356] L. Abellanas and A. Galindo. Evolution equations with high order conservation laws. Journal of Mathematical Physics, 24(3):504–509, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i3/p504_s1.

Hermann:1983:GCP

[1357] Robert Hermann. Geometric construction and properties of some families of solutions of nonlinear partial differential equations. I. Journal of Mathematical Physics, 24(3):510–521, March 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i3/p510_s1.

Weiss:1983:PPPa

[1358] John Weiss, M. Tabor, and George Carnevale. The Painlevé property for partial differential equations. Journal of Mathematical Physics, 24(3): 522–526, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p522_s1.

Katehi:1983:RAI

[1359] P. B. Katehi and N. G. Alexopoulos. Real axis integration of Sommerfeld integrals with applications to printed circuit antennas. Journal of Mathematical Physics, 24(3):527–533, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p527_s1.

Gadella:1983:GWC

[1360] M. Gadella. A generalized Weyl correspondence. II. Some general results. Journal of Mathematical Physics, 24(3):534–538, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p534_s1.

Meiman:1983:UBA

[1361] N. N. Meĭman. An upper bound on the allowed bands of the Bloch spectrum of one-dimensional Schrödinger operators with periodic potentials. Journal of Mathematical Physics, 24(3):539–540, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p539_s1.

Hermann:1983:EHS

[1362] M. R. Hermann and P. W. Langhoff. Explicit Hilbert space representations of Schrödinger states: Definitions and properties of Stieltjes– Tchebycheff orbitals. Journal of Mathematical Physics, 24(3):541– 547, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p541_s1.

Prosser:1983:CBC

[1363] Reese T. Prosser. On the correspondence between classical and quantum mechanics. I. Journal of Mathematical Physics, 24(3):548–552, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i3/p548_s1.

Calvo:1983:CSP

[1364] Miguel Calvo and Simón Codriansky. A class of solvable Pauli– Schrödinger Hamiltonians. Journal of Mathematical Physics, 24(3):553– 559, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p553_s1.

Romo:1983:ISW

[1365] W. J. Romo. Irregular solutions of the s-wave radial Schrödinger equation for van der Waals potentials. Journal of Mathematical Physics, 24(3): 560–572, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p560_s1.

Datta:1983:AMA

[1366] D. P. Datta and S. Mukherjee. Angular momentum analyticity of Schrödinger energy levels for quark confining potentials. Journal of Mathematical Physics, 24(3):573–575, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p573_s1.

Evans:1983:STA

[1367] J. W. Evans, D. K. Hoffman, and D. J. Kouri. Scattering theory in arrangement channel quantum mechanics. Journal of Mathematical Physics, 24(3):576–587, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i3/p576_s1.

Abdel-Raouf:1983:POB

[1368] M. A. Abdel-Raouf. Partitioning “opposite” bounds for potential scattering phase shifts. Journal of Mathematical Physics, 24(3):588–589, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i3/p588_s1.

TorresdelCastillo:1983:NSB

[1369] G. F. Torres del Castillo. Null strings and Bianchi identities. Journal of Mathematical Physics, 24(3):590–596, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p590_s1.

Chakrabarti:1983:TCL

[1370] Sandip K. Chakrabarti, Robert Geroch, and Can Bin Liang. Timelike curves of limited acceleration in general relativity. Journal of Mathematical Physics, 24(3):597–598, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p597_s1.

Smallwood:1983:ODS

[1371] J. Smallwood. Orthogonal 2+2 decomposition of space–time. Journal of Mathematical Physics, 24(3):599–605, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p599_s1.

Nakamura:1983:SSA

[1372] Yoshimasa Nakamura. Symmetries of stationary axially symmetric vacuum Einstein equations and the new family of exact solutions. Journal of Mathematical Physics, 24(3):606–609, March 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i3/p606_s1.

De:1983:PSC

[1373] Utpal Kumar De and Dipankar Ray. Plane symmetric charged dust distribution. Journal of Mathematical Physics, 24(3):610–612, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i3/p610_s1.

Toller:1983:ETP

[1374] M. Toller. Extended test particles in geometric fields. Journal of Mathematical Physics, 24(3):613–619, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p613_s1.

Grensing:1983:BED

[1375] D. Grensing and G. Grensing. Boundary effects in the dimer problem on a non-Bravais lattice. Journal of Mathematical Physics, 24(3):620– 630, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p620_s1.

Bovier:1983:FMG

[1376] Anton Bovier. Factorized S matrices and generalized Baxter models. Journal of Mathematical Physics, 24(3):631–641, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p631_s1.

Raszillier:1983:SFD

[1377] H. Raszillier. Symmetry and fermion degeneracy on a lattice. Journal of Mathematical Physics, 24(3):642–650, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p642_s1.

Garbaczewski:1983:MFB

[1378] Piotr Garbaczewski. Mechanisms of the “fermion-boson reciprocity”. Journal of Mathematical Physics, 24(3):651–658, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/p651_s1.

Chakrabarti:1983:CSD

[1379] A. Chakrabarti, A. Comtet, and F. Koukiou. Complex, self-dual Euclidean SU(2) gauge fields with finite real actions. Journal of Mathematical Physics, 24(3):659–673, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i3/p659_s1.

Hogan:1983:SSY

[1380] P. A. Hogan. Some solutions of the Yang–Mills equations generalizing the Wu–Yang monopole. Journal of Mathematical Physics, 24(3):674– 676, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p674_s1.

Destri:1983:NBA

[1381] C. Destri and J. H. Lowenstein. Normalization of Bethe-ansatz states in the chiral-invariant Gross–Neveu model. Journal of Mathematical Physics, 24(3):677–682, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i3/p677_s1.

Talukdar:1983:CMN

[1382] B. Talukdar, S. Saha, and T. Sasakawa. Coulomb-modified nuclear scattering. II. Journal of Mathematical Physics, 24(3):683–686, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i3/
p683_s1.

Rahnema:1983:EBP

[1383] F. Rahnema and G. C. Pomraning. External boundary perturbations in neutron transport theory. Journal of Mathematical Physics, 24(3):687– 701, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p687_s1.

Salamo:1983:MEN

[1384] S. Salamó. On massive electromagnetism in the null-plane: their limit and “dual” invariance. Journal of Mathematical Physics, 24(3):702– 704, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p702_s1.

Druhl:1983:CPI

[1385] Kai Drühl and J. K. McIver. Charged particles in an intense, plane electromagnetic wave: Limitations of the nonrelativistic theory. Journal of Mathematical Physics, 24(3):705–711, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i3/p705_s1.

Hillion:1983:SFSa

[1386] P. Hillion and S. Quinnez. Strong fluctuations in the spinor wave optics. Mutual coherence function. Journal of Mathematical Physics, 24(3):712– 720, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p712_s1.

Hillion:1983:SFSb

[1387] P. Hillion and S. Quinnez. Strong fluctuations in the spinor wave optics. Intensity covariance function. Journal of Mathematical Physics, 24(3): 721–727, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i3/p721_s1.

Soni:1983:WVR

[1388] S. K. Soni. Weight vector representatives for SO(2N). Journal of Mathematical Physics, 24(4):729–735, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i4/p729_s1.

Applequist:1983:CP

[1389] Jon Applequist. Cartesian polytensors. Journal of Mathematical Physics, 24(4):736–741, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p736_s1.

Newmarch:1983:SCT

[1390] J. D. Newmarch. Some character theory for groups of linear and antilinear operators. Journal of Mathematical Physics, 24(4):742–756, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i4/p742_s1.

Newmarch:1983:S

[1391] J. D. Newmarch. On the 3j symmetries. Journal of Mathematical Physics, 24(4):757–764, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p757_s1.

Sartori:1983:TOS

[1392] G. Sartori. A theorem on orbit structures (strata) of compact linear Lie groups. Journal of Mathematical Physics, 24(4):765–768, April 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p765_
s1.

Seligman:1983:ILD

[1393] T. H. Seligman and R. T. Sharp. Internal labels of degenerate representations. Journal of Mathematical Physics, 24(4):769–771, April 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p769_
s1.

Somers:1983:AOP

[1394] L. J. Somers. Analysis of the outer product for the symmetric group. Journal of Mathematical Physics, 24(4):772–778, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p772_s1.

Couvreur:1983:CGR

[1395] G. Couvreur, J. Deenen, and C. Quesne. Complementary group with respect to SO(n). Journal of Mathematical Physics, 24(4):779–784, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i4/p779_s1.

Hecht:1983:SSW

[1396] K. T. Hecht and Y. Suzuki. Some special SU(3) ⊇ R(3) Wigner coefficients and their application. Journal of Mathematical Physics, 24(4): 785–792, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p785_s1.

Rosen:1983:RTO

[1397] Gerald Rosen. Reciprocal transformation for one-dimensional conservation equations. Journal of Mathematical Physics, 24(4):793–794, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i4/p793_s1.

Lakshmanan:1983:LTN

[1398] M. Lakshmanan and P. Kaliappan. Lie transformations, nonlinear evolution equations, and Painlevé forms. Journal of Mathematical Physics, 24(4):795–806, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p795_s1.

Percacci:1983:KKT

[1399] R. Percacci and S. Randjbar-Daemi. Kaluza–Klein theories on bundles with homogeneous fibers. I. Journal of Mathematical Physics, 24(4): 807–814, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p807_s1.

Ladde:1983:EES

[1400] G. S. Ladde and M. Sambandham. Error estimates of solutions and mean of solutions of stochastic differential systems. Journal of Mathematical Physics, 24(4):815–822, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p815_s1.

Channell:1983:ESD

[1401] Paul J. Channell. Explicit suspensions of diffeomorphisms — an inverse problem in classical dynamics. Journal of Mathematical Physics, 24(4): 823–827, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p823_s1.

Weston:1983:IPR

[1402] V. H. Weston. Inverse problem for the reduced wave equation with fixed incident field. III. Journal of Mathematical Physics, 24(4):828– 833, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p828_s1.

Janis:1983:OFE

[1403] V. Janiš, J. Souček, and V. Souček. Operator formalism equivalent to the Feynman quantization technique. Journal of Mathematical Physics, 24(4):834–838, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p834_s1.

Ptak:1983:WDF

[1404] Pavel Pták. Weak dispersion-free states and the hidden variables hypothesis. Journal of Mathematical Physics, 24(4):839–840, April 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p839_
s1.

Groblicki:1983:GCT

[1405] Roman Groblicki. A generalization of Cohen’s theorem. Journal of Mathematical Physics, 24(4):841–843, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i4/p841_s1.

Dhara:1983:PIA

[1406] A. K. Dhara, D. C. Khandekar, and S. V. Lawande. Path integration of an action related to an electron gas in a random potential. Journal of Mathematical Physics, 24(4):844–847, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i4/p844_s1.

Horwitz:1983:PIP

[1407] L. P. Horwitz and E. Katznelson. A partial inner product space of analytic functions for resonances. Journal of Mathematical Physics, 24(4): 848–859, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p848_s1.

Rampal:1983:GAO

[1408] Anita Rampal and K. Datta. The generalized anharmonic oscillator in three dimensions: Exact eigenvalues and eigenfunctions. Journal of Mathematical Physics, 24(4):860–866, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i4/p860_s1.

Khare:1983:ADP

[1409] Avinash Khare. The accidental degeneracy problem in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 24(4):867–873, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i4/p867_s1.

Gerry:1983:RCT

[1410] Christopher C. Gerry. Remarks on canonical transformations in phasespace path integrals. Journal of Mathematical Physics, 24(4):874–877, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v24/i4/p874_s1. See erratum [2218].

Sidharth:1983:RRP

[1411] B. G. Sidharth. A recurrence relation for the phase shifts of exponential and Yukawa potentials. Journal of Mathematical Physics, 24(4):878– 882, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p878_s1.

Swaminarayan:1983:CFD

[1412] N. S. Swaminarayan and J. L. Safko. A coordinate-free derivation of a generalized geodesic deviation equation. Journal of Mathematical Physics, 24(4):883–885, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p883_s1.

Kunstatter:1983:GIG

[1413] G. Kunstatter, J. W. Moffat, and J. Malzan. Geometrical interpretation of a generalized theory of gravitation. Journal of Mathematical Physics, 24(4):886–889, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p886_s1.

Mansouri:1983:LSE

[1414] Freydoon Mansouri. Local symmetries in extended conformal gravity and conformal supergravity in superspace. Journal of Mathematical Physics, 24(4):890–896, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p890_s1.

Royer:1983:QVC

[1415] Antoine Royer. Quantum virial coefficients as cumulants of imaginary time-ordered Mayer diagrams. Journal of Mathematical Physics, 24(4): 897–912, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p897_s1.

Marra:1983:RBP

[1416] Rossana Marra. On a relation between percolation and phase transition in the gauge invariant Ising model. Journal of Mathematical Physics, 24(4):913–916, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p913_s1.

Jaric:1983:OSM

[1417] Marko V. Jarić. Orbit space and minima of Landau–Higgs potentials. Journal of Mathematical Physics, 24(4):917–921, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p917_s1.

Ruijsenaars:1983:TPF

[1418] S. N. M. Ruijsenaars. On the two-point functions of some integrable relativistic quantum field theories. Journal of Mathematical Physics, 24(4): 922–931, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p922_s1.

Bishop:1983:PSS

[1419] Garner Bishop and Kurt Haller. Particle states and scattering theory in abelian gauge model with spontaneously broken symmetry. Journal of Mathematical Physics, 24(4):932–948, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i4/p932_s1.

Kent:1983:YMT

[1420] S. L. Kent and E. T. Newman. Yang–Mills theory in null-path space. Journal of Mathematical Physics, 24(4):949–959, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p949_s1.

Houston:1983:GHF

[1421] P. Houston. On the generalization of the ’t Hooft field strength in the Yang–Mills–Higgs model. Journal of Mathematical Physics, 24(4):960– 968, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p960_s1.

Bowtell:1983:PRK

[1422] Graham Bowtell and Allen E. G. Stuart. A particle representation for Korteweg–de Vries solitons. Journal of Mathematical Physics, 24(4): 969–981, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p969_s1.

MartinezAlonso:1983:GTA

[1423] L. Martínez Alonso. Group-theoretical analysis of the sine-Gordon equation as a relativistic dynamical system. Journal of Mathematical Physics, 24(4):982–989, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p982_s1.

Bender:1983:LTP

[1424] Carl M. Bender and David A. Williams. Lorentz transformation properties of nonconserved charges. Journal of Mathematical Physics, 24(4): 990–995, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i4/p990_s1.

Srivastava:1983:FUS

[1425] Sushil K. Srivastava. Friedmann Universe and superluminal Dirac particles. Journal of Mathematical Physics, 24(4):996–1000, April 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p996_
s1.

Gordon:1983:QAB

[1426] Paul Gordon. Quasisteady arterial blood flow. Journal of Mathematical Physics, 24(4):1001–1010, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i4/p1001_s1.

Ray:1983:NPW

[1427] Dipankar Ray. Nonlinear periodic waves in a self-gravitating fluid. Journal of Mathematical Physics, 24(4):1011–1012, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p1011_s1.

Datta:1983:ERT

[1428] D. P. Datta and S. Mukherjee. Erratum: “Regge trajectories in confining potentials”, [J. Math. Phys. 23, 665 (1982)]. Journal of Mathematical Physics, 24(4):1013, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i4/p1013_s3. See [1000].

Gowdy:1983:EGP

[1429] Robert H. Gowdy. Erratum: “Geometrical perturbation theory: actionprinciple surface terms in homogeneous cosmology”, [J. Math. Phys. 23, 2151 (1982)]. Journal of Mathematical Physics, 24(4):1013, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i4/p1013_s2. See [1218].

Nikam:1983:ERI

[1430] R. S. Nikam, K. V. Dinesha, and C. R. Sarma. Erratum: “Reduction of inner-product representations of unitary groups”, [J. Math. Phys. 24, 233 (1983)]. Journal of Mathematical Physics, 24(4):1013, April 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i4/p1013_

s1. See [1320].

Hughes:1983:SOR

[1431] J. W. B. Hughes. SU(2) × SU(2) shift operators and representations of SO(5). Journal of Mathematical Physics, 24(5):1015–1020, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/
p1015_s1.

Baake:1983:RBI

[1432] M. Baake, B. Gemünden, and R. Oedingen. On the relations between irreducible representations of the hyperoctahedral group and O(4) and SO(4). Journal of Mathematical Physics, 24(5):1021–1024, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/
p1021_s1.

VanderJeugt:1983:NES

[1433] J. Van der Jeugt, H. E. De Meyer, and G. Vanden Berghe. Nonscalar extension of shift operator techniques for SU(3) in an O(3) basis. III. Shift operators of second degree in the tensor components. Journal of Mathematical Physics, 24(5):1025–1031, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1025_s1.

Gruber:1983:IRP

[1434] B. Gruber and T. S. Santhanam. Indecomposable representations for para-Bose algebra. Journal of Mathematical Physics, 24(5):1032–1036, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1032_s1.

Ronveaux:1983:HPI

[1435] A. Ronveaux and Y. Saint-Aubin. Harmonic polynomials invariant under a finite subgroup of O(n). Journal of Mathematical Physics, 24(5):1037– 1040, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1037_s1.

Chen:1983:GMN

[1436] Feng Su Chen and D. Y. Hsieh. A Galerkin method and nonlinear oscillations and waves. Journal of Mathematical Physics, 24(5):1041–1046, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1041_s1.

Kalnins:1983:GTS

[1437] E. G. Kalnins and W. Miller, Jr. The general theory of R-separation for Helmholtz equations. Journal of Mathematical Physics, 24(5):1047– 1053, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1047_s1.

vanHaeringen:1983:SIS

[1438] H. van Haeringen. Some infinite series of products of Legendre and gamma functions. Journal of Mathematical Physics, 24(5):1054–1057, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1054_s1.

Bassetti:1983:SPT

[1439] B. Bassetti, E. Montaldi, and M. Raciti. A simple proof of a transformation formula for elliptic integrals. Journal of Mathematical Physics, 24(5): 1058–1059, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1058_s1.

Antone:1983:E

[1440] T. A. Antone. Expression for y_{LM}[(r_{1} ⋀ r_{2}) ⋀ r_{3}]. Journal of Mathematical Physics, 24(5):1060–1061, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i5/p1060_s1.

Harnad:1983:SPM

[1441] J. Harnad, P. Winternitz, and R. L. Anderson. Superposition principles for matrix Riccati equations. Journal of Mathematical Physics, 24(5): 1062–1072, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1062_s1.

Menyuk:1983:RMT

[1442] C. R. Menyuk, H. H. Chen, and Y. C. Lee. Restricted multiple three-wave interactions: Integrable cases of this system and other related systems. Journal of Mathematical Physics, 24(5):1073–1079, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1073_s1.

Fahnline:1983:MCC

[1443] Donald E. Fahnline. Manifestly covariant, coordinate-free dyadic expression for planar homogeneous Lorentz transformations. Journal of Mathematical Physics, 24(5):1080–1086, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1080_s1.

Tafel:1983:CPC

[1444] Jacek Tafel and Andrzej Trautman. Can poles change color? Journal of Mathematical Physics, 24(5):1087–1092, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1087_s1.

Osborn:1983:TEK

[1445] T. A. Osborn and Y. Fujiwara. Time evolution kernels: uniform asymptotic expansions. Journal of Mathematical Physics, 24(5):1093–1103, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1093_s1.

Soto-Eguibar:1983:TEW

[1446] Francisco Soto-Eguibar and Pierre Claverie. Time evolution of the Wigner function. Journal of Mathematical Physics, 24(5):1104–1109, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1104_s1.

Peres:1983:TMO

[1447] Asher Peres. Transfer matrices for one-dimensional potentials. Journal of Mathematical Physics, 24(5):1110–1119, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1110_s1.

Wenning:1983:PFQ

[1448] Thomas Wenning and Alexander Bach. A probabilistic formulation of quantum theory. III. Journal of Mathematical Physics, 24(5):1120–1122, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1120_s1.

Besnard:1983:NTD

[1449] D. Besnard, J. R. Burgan, A. Munier, M. R. Feix, and E. Fijalkow. Nonlinear time-dependent anharmonic oscillator: Asymptotic behavior and connected invariants. Journal of Mathematical Physics, 24(5):1123– 1128, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1123_s1.

Exner:1983:CPD

[1450] Pavel Exner. Complex-potential description of the damped harmonic oscillator. Journal of Mathematical Physics, 24(5):1129–1135, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/
p1129_s1.

Znojil:1983:ECF

[1451] M. Znojil. Extended continued fractions and energies of the anharmonic oscillators. Journal of Mathematical Physics, 24(5):1136–1141, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1136_s1.

Grosse:1983:NUP

[1452] H. Grosse and L. Pittner. On the number of unnatural parity bound states of the H^{−} ion. Journal of Mathematical Physics, 24(5):1142– 1147, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1142_s1.

Gray:1983:ISR

[1453] Samuel H. Gray. Inverse scattering for the reflectivity function. Journal of Mathematical Physics, 24(5):1148–1151, May 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1148_s1.

vanHaeringen:1983:CJS

[1454] H. van Haeringen. The Coulomb Jost states in the momentum representation for all partial waves in closed form. Journal of Mathematical Physics, 24(5):1152–1156, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1152_s1.

vanHaeringen:1983:URS

[1455] H. van Haeringen. The unitarity relations for the off-shell Coulomb T matrix for all partial waves. Journal of Mathematical Physics, 24(5):

1157–1159, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1157_s1.

Evans:1983:EUB

[1456] J. W. Evans. Existence and uniqueness of bound-state eigenvectors for some channel coupling Hamiltonians. Journal of Mathematical Physics, 24(5):1160–1162, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i5/p1160_s1.

Simon:1983:MSS

[1457] Walter Simon and Robert Beig. The multipole structure of stationary space–times. Journal of Mathematical Physics, 24(5):1163–1171, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1163_s1.

DelOlmo:1983:CRM

[1458] V. Del Olmo and J. Olivert. Compatibility of rigid motions with the relativistic incompressibility condition. Journal of Mathematical Physics, 24(5):1172–1181, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1172_s1.

Embacher:1983:RHC

[1459] Franz Embacher. Rotating hollow cylinders: General solution and Machian effects. Journal of Mathematical Physics, 24(5):1182–1186, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1182_s1.

Wu:1983:SDG

[1460] Yong Shi Wu and Mo Lin Ge. A simplified derivation of the Geroch group in two-dimensional reduced gravity. Journal of Mathematical Physics, 24(5):1187–1192, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1187_s1.

Winicour:1983:NGN

[1461] J. Winicour. Newtonian gravity on the null cone. Journal of Mathematical Physics, 24(5):1193–1198, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i5/p1193_s1.

Ivanovic:1983:FSD

[1462] I. D. Ivanović. Formal state determination. Journal of Mathematical Physics, 24(5):1199–1205, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1199_s1.

Yip:1983:STD

[1463] Ping Yip. Spinors in two dimensions. Journal of Mathematical Physics, 24(5):1206–1212, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i5/p1206_s1.

Gozzi:1983:CSM

[1464] Ennio Gozzi and Arunabha Guha. Canonical structure of the CP_{2}^{n−1} model in some noncovariant gauges. Journal of Mathematical Physics, 24(5):1213–1218, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i5/p1213_s1.

Legare:1983:CIC

[1465] Martin Légaré. Conformal invariance conditions for spinor fields with gauge freedom. Journal of Mathematical Physics, 24(5):1219–1223, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i5/p1219_s1.

Porter:1983:SDYa

[1466] John R. Porter. Self-dual Yang–Mills fields. Journal of Mathematical Physics, 24(5):1224–1232, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1224_s1.

Porter:1983:SDYb

[1467] John R. Porter. Self-dual Yang–Mills fields on Minkowski space–time. Journal of Mathematical Physics, 24(5):1233–1239, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1233_s1.

Howland:1983:CSA

[1468] James S. Howland. Complex scaling of ac Stark Hamiltonians. Journal of Mathematical Physics, 24(5):1240–1244, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1240_s1.

Coll:1983:CSO

[1469] Bartolomé Coll and Carlos Moreno. On a class of synchronized observers attached to the Lorentzian structures. Journal of Mathematical Physics, 24(5):1245–1255, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1245_s1.

Garber:1983:LMS

[1470] Wolf-Dieter Garber. Local S-matrix symmetries in Galilean field theories. Journal of Mathematical Physics, 24(5):1256–1263, May 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1256_
s1.

Garber:1983:FMS

[1471] Wolf-Dieter Garber. Finitely many symmetries of the S-matrix. Journal of Mathematical Physics, 24(5):1264–1266, May 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1264_s1.

vanHaeringen:1983:PWP

[1472] H. van Haeringen. The partial-wave projected Coulomb T matrix for all l in closed hypergeometric form. Journal of Mathematical Physics, 24(5): 1267–1273, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1267_s1.

vanHaeringen:1983:SMC

[1473] H. van Haeringen. Off-shell T matrix for Coulomb plus simple separable potentials for all l in closed form. Journal of Mathematical Physics, 24(5): 1274–1281, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1274_s1.

Dita:1983:GFR

[1474] P. Diţă. On a generalized F. Riesz problem. Journal of Mathematical Physics, 24(5):1282–1285, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1282_s1.

Beckers:1983:MECa

[1475] J. Beckers and V. Hussin. Minimal electromagnetic coupling schemes. I. Symmetries of potentials and gauge transformations. Journal of Mathematical Physics, 24(5):1286–1294, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i5/p1286_s1.

Beckers:1983:MECb

[1476] J. Beckers and V. Hussin. Minimal electromagnetic coupling schemes. II. Relativistic and nonrelativistic Maxwell groups. Journal of Mathematical Physics, 24(5):1295–1298, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1295_s1.

Beau:1983:GHG

[1477] Daniel Beau. Generalized Hooke groups and the mass-spectrum problem. Journal of Mathematical Physics, 24(5):1299–1310, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1299_s1.

Srivastava:1983:IST

[1478] Sushil K. Srivastava. Interaction of a spin-1/2 tachyon with a spin-1/2 bradyon and a tachyon–bradyon model. Journal of Mathematical Physics, 24(5):1311–1316, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i5/p1311_s1.

Srivastava:1983:PTS

[1479] Sushil K. Srivastava. Production of tachyons from Schwarzschild black holes. Journal of Mathematical Physics, 24(5):1317–1320, May 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1317_
s1.

Pao:1983:SAN

[1480] C. V. Pao. Stability analysis of the neutron transport equation with temperature feedback. Journal of Mathematical Physics, 24(5):1321– 1325, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i5/p1321_s1.

Chakraborty:1983:RTI

[1481] B. B. Chakraborty, H. K. S. Iyengar, and A. R. Nayak. Rayleigh–Taylor instability of a fluid supported against gravity by an oscillating magnetic field. Journal of Mathematical Physics, 24(5):1326–1331, May 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1326_
s1.

Brandt:1983:SCS

[1482] Howard E. Brandt. Symmetry of the complete second-order nonlinear conductivity tensor for an unmagnetized relativistic turbulent plasma. Journal of Mathematical Physics, 24(5):1332–1338, May 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1332_

s1. See erratum [1606].

VanderJeugt:1983:SOTa

[1483] J. Van der Jeugt, H. E. De Meyer, and G. Vanden Berghe. Shift operator techniques for the classification of multipole-phonon states. XI. properties of mixed type quadratic product operators in R(7). Journal of Mathematical Physics, 24(5):1339–1344, May 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1339_s1.

VanderJeugt:1983:SOTb

[1484] J. Van der Jeugt, G. Vanden Berghe, and H. E. De Meyer. Shift operator techniques for the classification of multipole-phonon states. XII. O_{l}^{0} eigenstate and eigenvalue determination in R(7). Journal of Mathematical Physics, 24(5):1345–1349, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i5/p1345_s1.

Suffern:1983:ECP

[1485] Kevin G. Suffern, Edward D. Fackerell, and Christopher M. Cosgrove. Eigenvalues of the Chandrasekhar–Page angular functions. Journal of Mathematical Physics, 24(5):1350–1358, May 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i5/p1350_s1.

Chakraborty:1983:NMS

[1486] S. K. Chakraborty. A non-Machian solution in Brans–Dicke theory. Journal of Mathematical Physics, 24(5):1359–1360, May 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i5/p1359_s1.

Djokovic:1983:NFE

[1487] D. Ž. Djoković, J. Patera, P. Winternitz, and H. Zassenhaus. Normal forms of elements of classical real and complex Lie and Jordan algebras. Journal of Mathematical Physics, 24(6):1363–1374, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1363_s1.

Patera:1983:VDE

[1488] J. Patera and C. Rousseau. Versal deformations of elements of classical Jordan algebras. Journal of Mathematical Physics, 24(6):1375–1380, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1375_s1.

Perroud:1983:FII

[1489] M. Perroud. The fundamental invariants of inhomogeneous classical groups. Journal of Mathematical Physics, 24(6):1381–1391, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/
p1381_s1.

Balster:1983:RM

[1490] G. J. Balster, O. S. van Roosmalen, and A. E. L. Dieperink. Representation matrices for U(4). Journal of Mathematical Physics, 24(6): 1392–1394, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1392_s1.

Ray:1983:SCW

[1491] Alak Ray. Sufficient conditions for weak minimum of a functional depending on n functions and their derivatives. Journal of Mathematical Physics, 24(6):1395–1398, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i6/p1395_s1.

Beylkin:1983:FII

[1492] G. Beylkin. The fundamental identity for iterated spherical means and the inversion formula for diffraction tomography and inverse scattering. Journal of Mathematical Physics, 24(6):1399–1400, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1399_s1.

Wodkiewicz:1983:ESN

[1493] K. Wódkiewicz and M. S. Zubairy. Exact solution of a nonlinear Langevin equation with applications to photoelectron counting and noise-induced instability. Journal of Mathematical Physics, 24(6):1401–1404, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1401_s1.

Weiss:1983:PPPb

[1494] John Weiss. The Painlevé property for partial differential equations. II: Bäcklund transformation, Lax pairs, and the Schwarzian derivative. Journal of Mathematical Physics, 24(6):1405–1413, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1405_s1.

Bruschi:1983:HSN

[1495] M. Bruschi and O. Ragnisco. The Hamiltonian structure of the nonabelian Toda hierarchy. Journal of Mathematical Physics, 24(6):1414– 1421, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1414_s1.

Aaberg:1983:EIC

[1496] Sven Åberg. On the expansion of integrals containing Fermi distributions. Journal of Mathematical Physics, 24(6):1422–1424, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1422_s1.

Robinson:1983:CGF

[1497] Ivor Robinson and Andrzej Trautman. Conformal geometry of flows in n dimensions. Journal of Mathematical Physics, 24(6):1425–1429, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1425_s1.

Wollenberg:1983:HQI

[1498] L. S. Wollenberg. Homogeneous quadratic invariants for one-dimensional classical homogeneous quadratic Hamiltonians: Some properties of the auxiliary equation. Journal of Mathematical Physics, 24(6):1430–1438, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1430_s1.

Segur:1983:WKS

[1499] Harvey Segur. Wobbling kinks in φ^{4} and sine-Gordon theory. Journal of Mathematical Physics, 24(6):1439–1443, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i6/p1439_s1.

Keys:1983:GLI

[1500] R. G. Keys and A. B. Weglein. Generalized linear inversion and the first Born theory for acoustic media. Journal of Mathematical Physics, 24(6): 1444–1449, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1444_s1.

Navara:1983:RNP

[1501] Mirko Navara and Pavel Pták. On the Radon–Nikodym property for σ-classes. Journal of Mathematical Physics, 24(6):1450, June 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1450_
s1.

Sutherland:1983:CUJ

[1502] Roderick I. Sutherland. Compatibility of the uncorrelated joint distribution with the classical limit. Journal of Mathematical Physics, 24(6): 1451–1452, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1451_s1.

Daubechies:1983:CSC

[1503] Ingrid Daubechies. Continuity statements and counterintuitive examples in connection with Weyl quantization. Journal of Mathematical Physics, 24(6):1453–1461, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i6/p1453_s1.

Gadella:1983:RHS

[1504] M. Gadella. A rigged Hilbert space of Hardy-class functions: Applications to resonances. Journal of Mathematical Physics, 24(6):1462–1469, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1462_s1.

Dudley:1983:OTW

[1505] David Dudley and Alan P. Wang. Operator theory on the WKB method and Bremmer series. Journal of Mathematical Physics, 24(6):1470–1476, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1470_s1.

Isaacson:1983:LCB

[1506] David Isaacson. Large coupling behavior of the energy eigenvalues of two anharmonic oscillators. Journal of Mathematical Physics, 24(6):1477– 1481, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1477_s1.

Sun:1983:DSM

[1507] Chia Chung Sun and Zong Hao Zeng. On the dynamical symmetry of the Morse potential problem. Journal of Mathematical Physics, 24(6): 1482–1486, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1482_s1.

Osborn:1983:SSK

[1508] T. A. Osborn and R. Wong. Schrödinger spectral kernels: High order asymptotic expansions. Journal of Mathematical Physics, 24(6):1487– 1501, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1487_s1.

deCastro:1983:IST

[1509] Helena M. A. de Castro and Walter F. Wreszinski. Inverse scattering theory for systems of the Zakharov–Shabat type. Journal of Mathematical Physics, 24(6):1502–1508, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i6/p1502_s1.

Kroger:1983:SAW

[1510] Helmut Kröger. Strong approximation of wave operators for short range and Coulomb interactions. Journal of Mathematical Physics, 24(6):1509– 1515, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1509_s1.

Bruning:1983:CWS

[1511] E. Brüning and F. Gesztesy. Continuity of wave and scattering operators with respect to interactions. Journal of Mathematical Physics, 24(6): 1516–1528, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1516_s1.

Bolle:1983:TDL

[1512] D. Bollé, F. Gesztesy, and H. Grosse. Time delay for long-range interactions. Journal of Mathematical Physics, 24(6):1529–1541, June 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1529_
s1.

Exner:1983:OAT

[1513] P. Exner and I. Úlehla. On the optical approximation in two-channel systems. Journal of Mathematical Physics, 24(6):1542–1547, June 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1542_
s1.

Brannigan:1983:SGM

[1514] M. Brannigan and D. Eyre. Splines and the Galerkin method for solving the integral equations of scattering theory. Journal of Mathematical Physics, 24(6):1548–1554, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i6/p1548_s1. See erratum [1721].

Bolle:1983:LES

[1515] D. Bollé and S. F. J. Wilk. Low-energy sum rules for two-particle scattering. Journal of Mathematical Physics, 24(6):1555–1563, June 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1555_
s1.

Chen:1983:CSM

[1516] Y. Chen, D. S. Guo, and F. J. Ernst. Charged spinning mass field involving rational functions. Journal of Mathematical Physics, 24(6):1564– 1567, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1564_s1.

Iyer:1983:MAS

[1517] B. R. Iyer and C. V. Vishveshwara. Magnetization of all stationary cylindrically symmetric vacuum metrics. Journal of Mathematical Physics, 24(6):1568–1573, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i6/p1568_s1.

Krori:1983:SES

[1518] K. D. Krori, R. Bhattacharya, and D. Nandy. Some exact solutions of Einstein–Dirac–Maxwell zero mass scalar equations. Journal of Mathematical Physics, 24(6):1574–1576, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i6/p1574_s1.

Tsamparlis:1983:SCG

[1519] M. Tsamparlis and D. P. Mason. On spacelike congruences in general relativity. Journal of Mathematical Physics, 24(6):1577–1593, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1577_s1.

Amigo:1983:SIM

[1520] J. M. Amigó and H. Reeh. Summational invariants in the mechanics of mass points. Journal of Mathematical Physics, 24(6):1594–1602, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1594_s1.

Levine:1983:GPI

[1521] R. Y. Levine and Y. Tomozawa. Generalized parastatistics and internal symmetry. Journal of Mathematical Physics, 24(6):1603–1607, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1603_s1.

Buffet:1983:FPI

[1522] E. Buffet and J. V. Pulè. Fluctuation properties of the imperfect Bose gas. Journal of Mathematical Physics, 24(6):1608–1616, June 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1608_
s1.

Nakagomi:1983:CSS

[1523] Teruaki Nakagomi. Cycle statistics of steady states for master-equation systems. Journal of Mathematical Physics, 24(6):1617–1624, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/
p1617_s1.

Boffi:1983:EHT

[1524] V. C. Boffi and G. Spiga. An equation of Hammerstein type arising in particle transport theory. Journal of Mathematical Physics, 24(6):1625– 1629, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1625_s1.

Turner:1983:TDS

[1525] Leaf Turner and Carl M. Bender. Two-dimensional, steady-state diffusion of a passive scalar in a flow field. Journal of Mathematical Physics, 24(6): 1630–1632, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i6/p1630_s1.

Wichmann:1983:SLO

[1526] Eyvind H. Wichmann. On systems of local operators and the duality condition. Journal of Mathematical Physics, 24(6):1633–1644, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1633_s1.

Farrukh:1983:PDQ

[1527] M. O. Farrukh. Potential densities for quantum field theories in the scattering picture. Journal of Mathematical Physics, 24(6):1645–1651, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i6/p1645_s1.

Schuch:1983:NST

[1528] D. Schuch, K.-M. Chung, and H. Hartmann. Nonlinear Schrödinger-type field equation for the description of dissipative systems. I. Derivation of the nonlinear field equation and one-dimensional example. Journal of Mathematical Physics, 24(6):1652–1660, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i6/p1652_s1.

Hogan:1983:SYM

[1529] P. A. Hogan. Some Yang–Mills fields constructed from Maxwell fields. Journal of Mathematical Physics, 24(6):1661–1664, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/p1661_s1.

Iranzo:1983:PPC

[1530] V. Iranzo, J. Llosa, F. Marqués, and A. Molina. The problem of physical coordinates in predictive Hamiltonian systems. Journal of Mathematical Physics, 24(6):1665–1671, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i6/p1665_s1.

Hebert:1983:HLS

[1531] D. J. Hebert. A Hamiltonian–Lagrangian system of vortex boundary curves. Journal of Mathematical Physics, 24(6):1672–1678, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i6/
p1672_s1.

Green:1983:CIC

[1532] H. S. Green and P. D. Jarvis. Casimir invariants, characteristic identities, and Young diagrams for color algebras and superalgebras. Journal of Mathematical Physics, 24(7):1681–1687, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1681_s1.

Baclawski:1983:CGU

[1533] Kenneth Baclawski. Character generators for unitary and symplectic groups. Journal of Mathematical Physics, 24(7):1688–1694, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i7/
p1688_s1.

Bincer:1983:MLO

[1534] Adam M. Bincer. Missing label operators in the reduction O(p) ↓ O(p − 2) × O(2). Journal of Mathematical Physics, 24(7):1695–1699, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1695_s1.

Shintani:1983:CDI

[1535] Meiun Shintani. Construction of N-dimensional indecomposable representations for scale and special conformal transformations. Journal of Mathematical Physics, 24(7):1700–1704, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1700_s1.

Jarvis:1983:CIC

[1536] P. D. Jarvis and M. K. Murray. Casimir invariants, characteristic identities, and tensor operators for “strange” superalgebras. Journal of Mathematical Physics, 24(7):1705–1710, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i7/p1705_s1.

Price:1983:ILU

[1537] John F. Price. Inequalities and local uncertainty principles. Journal of Mathematical Physics, 24(7):1711–1714, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1711_s1.

Strobel:1983:SRN

[1538] George L. Strobel. Superposition rules for nonlinear coupled first-order differential equations. Journal of Mathematical Physics, 24(7):1715– 1719, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1715_s1. See comments [2004].

Leo:1983:NEE

[1539] M. Leo, R. A. Leo, G. Soliani, L. Solombrino, and L. Martina. Nonlinear evolution equations and nonabelian prolongations. Journal of Mathematical Physics, 24(7):1720–1730, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i7/p1720_s1.

Nakajima:1983:CCB

[1540] Hideo Nakajima. Characterization of canonical Bose–Fermi systems by “anti-Hermitian” symplectic forms. Journal of Mathematical Physics, 24(7):1731–1751, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i7/p1731_s1.

Burdet:1983:CSG

[1541] G. Burdet, C. Duval, and M. Perrin. Cartan structures on Galilean manifolds: The chronoprojective geometry. Journal of Mathematical Physics, 24(7):1752–1760, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i7/p1752_s1.

Katzin:1983:TDV

[1542] Gerald H. Katzin and Jack Levine. Time-dependent vector constants of motion, symmetries, and orbit equations for the dynamical system r̈ = l̂_{r}{[Ü(t)/U(t)]r−[μ_{0}/U(t)]r^{−2}}. Journal of Mathematical Physics, 24(7): 1761–1771, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1761_s1.

TabordaDuarte:1983:DHD

[1543] J. Taborda Duarte and R. Vilela Mendes. Deformation of Hamiltonian dynamics and constants of motion in dissipative systems. Journal of Mathematical Physics, 24(7):1772–1778, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1772_s1.

Gisin:1983:IQD

[1544] N. Gisin. Irreversible quantum dynamics and the Hilbert space structure of quantum kinematics. Journal of Mathematical Physics, 24(7):1779– 1782, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1779_s1.

Takahashi:1983:FIP

[1545] Yasushi Takahashi. The Fierz identities — a passage between spinors and tensors. Journal of Mathematical Physics, 24(7):1783–1790, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1783_s1.

Weiner:1983:SPT

[1546] Brian Weiner. Superoperator perturbation theory for propagators. Journal of Mathematical Physics, 24(7):1791–1796, July 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i7/p1791_s1.

Corbett:1983:STD

[1547] J. V. Corbett. Scattering theory for the dilation group. I. Simple quantum mechanical scattering. Journal of Mathematical Physics, 24(7):1797– 1805, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1797_s1. See erratum [2102].

Garbaczewski:1983:NGQ

[1548] Piotr Garbaczewski. Non-Grassmann quantization of the massive Thirring model. Journal of Mathematical Physics, 24(7):1806–1809, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1806_s1.

Sudarshan:1983:ESQ

[1549] E. C. G. Sudarshan and Xerxes Tata. An exactly solvable quantum field theory in three dimensions. Journal of Mathematical Physics, 24(7): 1810–1813, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1810_s1.

Ross:1983:DGM

[1550] D. K. Ross. Dirac gravitational magnetic monopoles do not exist. Journal of Mathematical Physics, 24(7):1814–1816, July 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i7/p1814_s1.

Hoenselaers:1983:RTS

[1551] Cornelius Hoenselaers and Frederick J. Ernst. Remarks on the Tomimatsu–Sato metrics. Journal of Mathematical Physics, 24(7):1817– 1820, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1817_s1.

Brown:1983:RGP

[1552] L. S. Brown and A. Zee. Response to gravitational probes and induced Newton’s constant. Journal of Mathematical Physics, 24(7):1821–1823, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1821_s1.

Isaacson:1983:NCC

[1553] R. A. Isaacson, J. S. Welling, and J. Winicour. Null cone computation of gravitational radiation. Journal of Mathematical Physics, 24(7):1824– 1834, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1824_s1.

Kalinowski:1983:NKK

[1554] M. W. Kalinowski. The nonsymmetric Kaluza–Klein theory. Journal of Mathematical Physics, 24(7):1835–1845, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1835_s1.

Noakes:1983:IVF

[1555] David R. Noakes. The initial value formulation of higher derivative gravity. Journal of Mathematical Physics, 24(7):1846–1850, July 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i7/p1846_
s1.

Geroch:1983:LNT

[1556] Robert Geroch. The local nonsingularity theorem. Journal of Mathematical Physics, 24(7):1851–1858, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i7/p1851_s1.

Hock:1983:OSI

[1557] J. L. Hock and R. B. McQuistan. The occupation statistics for indistinguishable dumbbells on a 2 × 2 × N lattice space. Journal of Mathematical Physics, 24(7):1859–1865, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i7/p1859_s1.

Farrukh:1983:FTM

[1558] M. O. Farrukh. Field theory models in the scattering picture. Journal of Mathematical Physics, 24(7):1866–1870, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1866_s1.

Fulp:1983:SCG

[1559] R. O. Fulp and L. K. Norris. Splitting of the connection in gauge theories with broken symmetry. Journal of Mathematical Physics, 24(7):1871– 1887, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1871_s1.

doAmaral:1983:HFB

[1560] A. F. Furtado do Amaral, F. A. Doria, and M. Gleiser. Higgs fields as Bargmann–Wigner fields and classical symmetry breaking. Journal of Mathematical Physics, 24(7):1888–1890, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i7/p1888_s1.

King:1983:ABI

[1561] Frederick W. King. Application of Benson’s inequalities to the atomic electronic density. Journal of Mathematical Physics, 24(7):1891–1895, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1891_s1.

Arai:1983:RTS

[1562] Asao Arai. Rigorous theory of spectra and radiation for a model in quantum electrodynamics. Journal of Mathematical Physics, 24(7):1896– 1910, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1896_s1.

Davidson:1983:EDM

[1563] Russell Davidson and John J. Kozak. Exact dynamics of a model for a three-level quantum system interacting with a continuous spectrum. Journal of Mathematical Physics, 24(7):1911–1922, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i7/p1911_s1.

Chanda:1983:SWE

[1564] Pranab Krishna Chanda and Dipankar Ray. Stationary wave envelopes in nonlinear optics. Journal of Mathematical Physics, 24(7):1923–1926, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i7/p1923_s1.

Pathria:1983:CBS

[1565] R. K. Pathria. Critical behavior of the spin van der Waals model. Journal of Mathematical Physics, 24(7):1927–1931, July 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i7/p1927_s1.

Beals:1983:EPR

[1566] Richard Beals. Erratum: “Partial-range completeness and existence of solutions to two-way diffusion equations”, [J. Math. Phys. 22 (1981), no. 5, 954–960; MR 83j:35030]. Journal of Mathematical Physics, 24 (7):1932, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i7/p1932_s1. See [586].

Dirl:1983:SGR

[1567] R. Dirl. Symmetry and generating relations for Clebsch–Gordan coefficients arising from automorphisms. Journal of Mathematical Physics, 24(8):1935–1944, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p1935_s1.

Rao:1983:QRP

[1568] K. N. Srinivasa Rao, A. V. Gopala Rao, and B. S. Narahari. On the quaternion representation of the proper Lorentz group SO(3, 1). Journal of Mathematical Physics, 24(8):1945–1954, August 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i8/p1945_s1.

Koh:1983:CGC

[1569] In Gyu Koh, J. Patera, and C. Rousseau. Clebsch–Gordan coefficients for SU(5) ⊇ SU(3) × SU(2) × U(1) theories. Journal of Mathematical Physics, 24(8):1955–1967, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i8/p1955_s1.

Brehamet:1983:TAH

[1570] Lionel Bréhamet. Tensorial analysis of the hyperfine interaction operator by extensive use of Racah algebra and of translational invariance. Journal of Mathematical Physics, 24(8):1968–1972, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p1968_s1.

Patera:1983:MAS

[1571] J. Patera, P. Winternitz, and H. Zassenhaus. Maximal abelian subalgebras of real and complex symplectic Lie algebras. Journal of Mathematical Physics, 24(8):1973–1985, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i8/p1973_s1.

Sudbery:1983:ODE

[1572] A. Sudbery. Octonionic description of exceptional Lie superalgebras. Journal of Mathematical Physics, 24(8):1986–1988, August 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p1986_
s1.

Niukkanen:1983:NCG

[1573] A. W. Niukkanen. On a new class of gradient formulas in the angular momentum theory. Journal of Mathematical Physics, 24(8):1989– 1991, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p1989_s1.

Arribas:1983:HOH

[1574] E. Arribas, J. Navarro, and M. Fabre de la Ripelle. On the higher orders of hyperspherical harmonics. Journal of Mathematical Physics, 24(8):1992–1996, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p1992_s1.

Houard:1983:NAP

[1575] J. C. Houard and M. Irac-Astaud. A new approach to perturbation theory: star diagrams. Journal of Mathematical Physics, 24(8):1997– 2005, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p1997_s1.

Gonzalez-Gascon:1983:SDE

[1576] F. González-Gascón and A. González-López. Symmetries of differential equations. IV. Journal of Mathematical Physics, 24(8):2006–2021, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2006_s1.

Boyer:1983:CIR

[1577] C. P. Boyer, E. G. Kalnins, and P. Winternitz. Completely integrable relativistic Hamiltonian systems and separation of variables in Hermitian hyperbolic spaces. Journal of Mathematical Physics, 24(8):2022– 2034, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2022_s1.

Boiti:1983:NHH

[1578] M. Boiti, C. Laddomada, F. Pempinelli, and G. Z. Tu. On a new hierarchy of Hamiltonian soliton equations. Journal of Mathematical Physics, 24(8):2035–2041, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2035_s1.

Cornille:1983:CBE

[1579] H. Cornille and A. Gervois. Connection between the existence of bisolitons for quadratic nonlinearities and a factorization of the associate linear operator. Journal of Mathematical Physics, 24(8):2042–2055, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2042_s1.

Yoshimura:1983:ANO

[1580] Tetz Yoshimura. Application of nonlinear operator theory to the Edwards–Freed equations in the theory of polymer solutions. Journal of Mathematical Physics, 24(8):2056–2060, August 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2056_s1.

Urani:1983:TAO

[1581] John R. Urani and Marilyn H. Kemp. Tetrads and arbitrary observers. Journal of Mathematical Physics, 24(8):2061–2064, August 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2061_
s1.

Caviglia:1983:DSA

[1582] Giacomo Caviglia. Dynamical symmetries: An approach to Jacobi fields and to constants of geodesic motion. Journal of Mathematical Physics, 24(8):2065–2069, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2065_s1.

Luttinger:1983:AEC

[1583] J. M. Luttinger. The asymptotic evaluation of a class of path integrals. II. Journal of Mathematical Physics, 24(8):2070–2073, August 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2070_
s1.

Masson:1983:RHOa

[1584] D. Masson. The rotating harmonic oscillator eigenvalue problem. I. Continued fractions and analytic continuation. Journal of Mathematical Physics, 24(8):2074–2088, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i8/p2074_s1.

Masson:1983:RHOb

[1585] D. Masson. The rotating harmonic oscillator eigenvalue problem. II. Analytic perturbation theory. Journal of Mathematical Physics, 24(8):2089– 2094, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2089_s1.

vanderSchaft:1983:SCL

[1586] A. J. van der Schaft. Symmetries, conservation laws, and time reversibility for Hamiltonian systems with external forces. Journal of Mathematical Physics, 24(8):2095–2101, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i8/p2095_s1.

Cushman:1983:NFR

[1587] Richard Cushman, André Deprit, and Richard Mosak. Normal form and representation theory. Journal of Mathematical Physics, 24(8):2102– 2117, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2102_s1.

delaLlave:1983:SPP

[1588] R. de la Llave. A simple proof of a particular case of C. Siegel’s center theorem. Journal of Mathematical Physics, 24(8):2118–2121, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2118_s1.

Kuchar:1983:MMC

[1589] Karel Kuchař. Measure for measure: Covariant skeletonizations of phase space path integrals for systems moving on Riemannian manifolds. Journal of Mathematical Physics, 24(8):2122–2141, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2122_s1.

Gadella:1983:DVS

[1590] M. Gadella. A description of virtual scattering states in the rigged Hilbert space formulation of quantum mechanics. Journal of Mathematical Physics, 24(8):2142–2145, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i8/p2142_s1.

Moses:1983:GLE

[1591] Harry E. Moses and Reese T. Prosser. The Gel’fand–Levitan equation on a finite interval. Journal of Mathematical Physics, 24(8):2146–2151, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2146_s1.

Newton:1983:ISL

[1592] Roger G. Newton. Inverse scattering by a local impurity in a periodic potential in one dimension. Journal of Mathematical Physics, 24(8):2152– 2162, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2152_s1.

Dammert:1983:TTS

[1593] Örjan Dammert. Transmission through a system of potential barriers. I. Transmission coefficient. Journal of Mathematical Physics, 24(8):2163– 2175, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2163_s1.

Accioly:1983:SCI

[1594] A. J. Accioly, A. N. Vaidya, and M. M. Som. Static conformally invariant scalar field. Journal of Mathematical Physics, 24(8):2176–2177, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2176_s1.

Harrison:1983:UEE

[1595] B. Kent Harrison. Unification of Ernst-equation Bäcklund transformations using a modified Wahlquist–Estabrook technique. Journal of Mathematical Physics, 24(8):2178–2187, August 1983. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i8/p2178_s1.

Knutsen:1983:AHS

[1596] Henning Knutsen. On apparent horizons and the Schwarzschild surface for a uniform fluid sphere in general relativity. Journal of Mathematical Physics, 24(8):2188–2190, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i8/p2188_s1.

SalazarI:1983:SNT

[1597] Humberto Salazar I., Alberto García D., and Jerzy F. Plebański. Symmetries of the nontwisting type-N solutions with cosmological constant. Journal of Mathematical Physics, 24(8):2191–2196, August 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2191_
s1.

Baumgartner:1983:GSG

[1598] Bernhard Baumgartner. On the group structure, GKS and FKG inequalities for Ising models. Journal of Mathematical Physics, 24(8):2197– 2199, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2197_s1.

Hattori:1983:GPT

[1599] Tetsuya Hattori. A generalization of the proof of the triviality of scalar field theories. Journal of Mathematical Physics, 24(8):2200–2203, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2200_s1.

Snyder:1983:TPM

[1600] M. A. Snyder. A two-parameter matrix Riccati equation pair for a class of nonlinear sigma models based on a symmetric space. Journal of Mathematical Physics, 24(8):2204–2207, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i8/p2204_s1.

Krause:1983:RGC

[1601] Jorge Krause. Radiation gauge covariance. II. Quantization of the free electromagnetic field. Journal of Mathematical Physics, 24(8):2208– 2211, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i8/p2208_s1.

Carey:1983:FCD

[1602] A. L. Carey, C. A. Hurst, and D. M. O’Brien. Fermion currents in 1 + 1 dimensions. Journal of Mathematical Physics, 24(8):2212–2221, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2212_s1.

Larsson:1983:CRFa

[1603] Jonas Larsson. Current responses of first and second order in a collisionless plasma. III. Three-wave interaction in periodic plasmas. Journal of Mathematical Physics, 24(8):2222–2230, August 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i8/p2222_s1.

Larsson:1983:CRFb

[1604] Jonas Larsson. Current responses of first and second order in a collisionless plasma. IV. Applications. Journal of Mathematical Physics, 24(8):2231–2239, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2231_s1.

Ott:1983:CEB

[1605] R. H. Ott and D. J. Fang. Comparison of the extended boundary condition method and perturbation method for scattering from a sinusoidal sea surface. Journal of Mathematical Physics, 24(8):2240–2249, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i8/p2240_s1.

Brandt:1983:ESC

[1606] Howard E. Brandt. Erratum: “Symmetry of the complete second-order nonlinear conductivity tensor for an unmagnetized relativistic turbulent plasma”, [J. Math. Phys. 24, 1332 (1983)]. Journal of Mathematical Physics, 24(8):2250, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2250_s3. See [1482].

Hurni:1983:EIR

[1607] J.-P. Hurni and B. Morel. Erratum: “Irreducible representations of SU(m/n)”, [J. Math. Phys. 24, 157 (1983)]. Journal of Mathematical Physics, 24(8):2250, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2250_s1. See [1307].

Newberger:1983:ENS

[1608] Barry S. Newberger. Erratum: “New sum rule for products of Bessel functions with application to plasma physics”, [J. Math. Phys. 23 (1982), no. 7, 1278–1281; MR 83m:33017]. Journal of Mathematical Physics, 24(8):2250, August 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i8/p2250_s2. See [1089, 1886].

Bars:1983:KDD

[1609] I. Bars, B. Morel, and H. Ruegg. Kac–Dynkin diagrams and supertableaux. Journal of Mathematical Physics, 24(9):2253–2262, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i9/p2253_s1.

DeWilde:1983:PBS

[1610] P. De Wilde, H. De Meyer, and G. Vanden Berghe. The SO(7) polynomial basis for symmetric representations. Journal of Mathematical Physics, 24(9):2263–2265, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i9/p2263_s1.

Zhu:1983:NAP

[1611] Cheng Jiu Zhu and Jin Quan Chen. A new approach to permutation group representations. II. Journal of Mathematical Physics, 24(9):2266– 2267, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i9/p2266_s1.

Hermann:1983:GTD

[1612] Robert Hermann. The geometric theory of deformation and linearization of Pfaffian systems and its applications to system theory and mathematical physics. I. Journal of Mathematical Physics, 24(9):2268–2276, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i9/p2268_s1.

deCastro:1983:IPF

[1613] Helena M. A. de Castro and Walter F. Wreszinski. The isospectral property for a family of non-self-adjoint operators. Journal of Mathematical Physics, 24(9):2277–2281, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i9/p2277_s1.

Dorizzi:1983:NCI

[1614] B. Dorizzi, B. Grammaticos, and A. Ramani. A new class of integrable systems. Journal of Mathematical Physics, 24(9):2282–2288, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i9/p2282_s1.

Grammaticos:1983:IHT

[1615] B. Grammaticos, B. Dorizzi, and A. Ramani. Integrability of Hamiltonians with third- and fourth-degree polynomial potentials. Journal of Mathematical Physics, 24(9):2289–2295, September 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i9/p2289_s1.

Kupershmidt:1983:DGZ

[1616] Boris Kupershmidt and Dale Peterson. A deformation of the general zero-curvature equations associated to simple Lie algebras. Journal of Mathematical Physics, 24(9):2296–2300, September 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i9/p2296_s1.

Corones:1983:HOP

[1617] J. P. Corones and R. J. Krueger. Higher-order parabolic approximations to time-independent wave equations. Journal of Mathematical Physics, 24(9):2301–2304, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i9/p2301_s1.

Grundland:1983:SWQa

[1618] Alfred M. Grundland and Roman Żelazny. Simple waves in quasilinear hyperbolic systems. I. Theory of simple waves and simple states. Examples of applications. Journal of Mathematical Physics, 24(9):2305–2314, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i9/p2305_s1.

Grundland:1983:SWQb

[1619] Alfred M. Grundland and Roman Żelazny. Simple waves in quasilinear hyperbolic systems. II. Riemann invariants for the problem of simple wave interactions. Journal of Mathematical Physics, 24(9):2315–2328, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i9/p2315_s1.

Lupieri:1983:GSP

[1620] Giancarlo Lupieri. Generalized stochastic processes and continual observations in quantum mechanics. Journal of Mathematical Physics, 24 (9):2329–2339, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i9/p2329_s1.

Odundun:1983:OTD

[1621] Olúwolé Òdúndún. Operators for the two-dimensional harmonic oscillator in an angular momentum basis. Journal of Mathematical Physics, 24(9):2340–2344, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i9/p2340_s1. See erratum [2224].

Kranold:1983:TDS

[1622] H. U. Kranold. A time-dependent Schrödinger equation. Journal of Mathematical Physics, 24(9):2345–2347, September 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i9/p2345_s1.

Umezawa:1983:LCE

[1623] M. Umezawa. Lorentz covariance of an extended object in the tree approximation. II. Nonspherical object in 3 + 1 dimensions. Journal of Mathematical Physics, 24(9):2348–2365, September 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i9/p2348_s1.

Hostler:1983:IRD

[1624] Levere C. Hostler. An SL(2, C)-invariant representation of the Dirac equation. II. Coulomb Green’s function. Journal of Mathematical Physics, 24(9):2366–2373, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i9/p2366_s1.

Kersten:1983:ISC

[1625] Paul H. M. Kersten. Infinitesimal symmetries and conserved currents for nonlinear Dirac equations. Journal of Mathematical Physics, 24(9):2374– 2376, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i9/p2374_s1.

Gaskell:1983:CGC

[1626] R. W. Gaskell. Character generators for compact semisimple Lie groups. Journal of Mathematical Physics, 24(10):2379–2386, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2379_s1.

Moody:1983:CGE

[1627] R. V. Moody, J. Patera, and R. T. Sharp. Character generators for elements of finite order in simple Lie groups A_{1}, A_{2}, A_{3}, B_{2}, and G_{2}. Journal of Mathematical Physics, 24(10):2387–2396, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i10/p2387_s1.

Rao:1983:SPI

[1628] K. N. Srinivasa Rao, A. V. Gopala Rao, and B. S. Narahari. Special properties of the irreducible representations of the proper Lorentz group. Journal of Mathematical Physics, 24(10):2397–2403, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2397_s1.

Berdjis:1983:CFD

[1629] F. Berdjis. A class of finite-dimensional Lie algebras, the Casimir operators of which are not of finite type. Journal of Mathematical Physics, 24(10):2404–2406, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2404_s1.

Urani:1983:DGC

[1630] John R. Urani and George P. Barker. Dirac general covariance and tetrads. I. Clifford and Lie bundles and torsion. Journal of Mathematical Physics, 24(10):2407–2410, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i10/p2407_s1.

Bordley:1983:MQB

[1631] Robert F. Bordley. Modeling quantum behavior with standard (nonquantum) probability theory. Journal of Mathematical Physics, 24(10):2411– 2421, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i10/p2411_s1.

Hermann:1983:GFI

[1632] Robert Hermann. The geometric foundations of the integrability property of differential equations and physical systems. I. Lie’s “function groups”. Journal of Mathematical Physics, 24(10):2422–2432, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2422_s1.

Reid:1983:HOS

[1633] James L. Reid and John R. Ray. Higher-order special self-adjoint equations and particle dynamics. Journal of Mathematical Physics, 24 (10):2433–2437, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2433_s1.

Shieh:1983:LES

[1634] S. Y. Shieh. Lagrange equations for a spinning gas cloud. Journal of Mathematical Physics, 24(10):2438–2440, October 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i10/p2438_s1.

Aerts:1983:CTN

[1635] Dirk Aerts. Classical theories and nonclassical theories as special cases of a more general theory. Journal of Mathematical Physics, 24(10):2441– 2453, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i10/p2441_s1.

Ghirardi:1983:VPQ

[1636] G. C. Ghirardi, A. Rimini, and T. Weber. Value preserving quantum measurements: Impossibility theorems and lower bounds for the distortion. Journal of Mathematical Physics, 24(10):2454–2460, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2454_s1.

Rosensteel:1983:AFB

[1637] G. Rosensteel and D. J. Rowe. An analytic formula for u(3)-boson matrix elements. Journal of Mathematical Physics, 24(10):2461–2463, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2461_s1.

Royer:1983:BCE

[1638] Antoine Royer. On the binary collision expansion and resummations of it. Journal of Mathematical Physics, 24(10):2464–2466, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i10/
p2464_s1.

vanHaeringen:1983:CPS

[1639] H. van Haeringen. Charged-particle off-shell scattering: common structure of all two-body off-shell scattering quantities for Coulomb plus rational separable potentials. Journal of Mathematical Physics, 24(10):2467– 2470, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i10/p2467_s1.

Dusek:1983:NML

[1640] J. Dušek. Nonsingular modified Lippmann–Schwinger equation for two charged particles. Journal of Mathematical Physics, 24(10):2471–2480, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i10/p2471_s1.

Kozameh:1983:TLC

[1641] Carlos N. Kozameh and Ezra T. Newman. Theory of light cone cuts of null infinity. Journal of Mathematical Physics, 24(10):2481–2489, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2481_s1.

Joshi:1983:LCC

[1642] P. S. Joshi, C. N. Kozameh, and E. T. Newman. Light cone cuts of null infinity in Schwarzschild geometry. Journal of Mathematical Physics, 24(10):2490–2497, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2490_s1.

GarciaD:1983:ANO

[1643] Alberto García D. and Humberto Salazar I. All null orbit type D electrovac solutions with cosmological constant. Journal of Mathematical Physics, 24(10):2498–2503, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i10/p2498_s1.

Sinha:1983:CED

[1644] S. K. Sinha and Y. Singh. Cluster expansion of the distribution functions for a ground state Fermi system. Journal of Mathematical Physics, 24(10):2504–2511, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2504_s1.

Lee:1983:DGL

[1645] M. Howard Lee. Derivation of the generalized Langevin equation by a method of recurrence relations. Journal of Mathematical Physics, 24 (10):2512–2514, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2512_s1.

Salamon:1983:GCT

[1646] Peter Salamon, Ed Ihrig, and R. Stephen Berry. A group of coordinate transformations which preserve the metric of Weinhold. Journal of Mathematical Physics, 24(10):2515–2520, October 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i10/p2515_s1.

Sharma:1983:SDE

[1647] L. K. Sharma, Sheilly Mehta, and V. P. Iyer. Solution of the Dirac equation for the general even power potential with application to particle spectroscopy. Journal of Mathematical Physics, 24(10):2521–2527, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i10/p2521_s1.

Cronstrom:1983:TBC

[1648] C. Cronström and J. Mickelsson. On topological boundary characteristics in nonabelian gauge theory. Journal of Mathematical Physics, 24 (10):2528–2531, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i10/p2528_s1. See erratum [2222].

Chi:1983:HSS

[1649] L. K. Chi. On the hydrodynamic self-similar cosmological models. Journal of Mathematical Physics, 24(10):2532–2533, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i10/

p2532_s1. See erratum [2214].

Parker:1983:GBC

[1650] C. N. Parker and B. D. Winter. On z_{n} graded Baker–Campbell–Hausdorff formulas and coset space parametrizations. Journal of Mathematical Physics, 24(11):2537–2541, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i11/p2537_s1.

Sullivan:1983:DCP

[1651] John J. Sullivan. A duality consistent phase convention for complex conjugation in SUn. Journal of Mathematical Physics, 24(11):2542–2545, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i11/p2542_s1.

Bincer:1983:ECO

[1652] Adam M. Bincer. Eigenvalues of Casimir operators for the general linear and orthosymplectic Lie superalgebras. Journal of Mathematical Physics, 24(11):2546–2549, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i11/p2546_s1.

Moses:1983:UCF

[1653] Harry E. Moses. The use of comparison filters in linear filter theory. Journal of Mathematical Physics, 24(11):2550–2552, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/
p2550_s1.

Weniger:1983:NRS

[1654] E. Joachim Weniger and E. Otto Steinborn. New representations for the spherical tensor gradient and the spherical delta function. Journal of Mathematical Physics, 24(11):2553–2563, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/p2553_s1.

Gallatin:1983:UFE

[1655] G. M. Gallatin. A useful formula for evaluating commutators. Journal of Mathematical Physics, 24(11):2564–2565, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/p2564_s1.

Betounes:1983:DTD

[1656] David E. Betounes. Dirac tensor distributions for moving submanifolds of R^{n}. Journal of Mathematical Physics, 24(11):2566–2572, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i11/p2566_s1.

Kree:1983:RGR

[1657] Paul Krée. Realization of Gaussian random fields. Journal of Mathematical Physics, 24(11):2573–2580, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i11/p2573_s1.

Skinner:1983:FOE

[1658] Ray Skinner. First-order equations of motion for classical mechanics. Journal of Mathematical Physics, 24(11):2581–2588, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/
p2581_s1.

Skinner:1983:GHDa

[1659] Ray Skinner and Raymond Rusk. Generalized Hamiltonian dynamics. I. formulation on T^{∗}Q⊕TQ. Journal of Mathematical Physics, 24(11):2589– 2594, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i11/p2589_s1.

Skinner:1983:GHDb

[1660] Ray Skinner and Raymond Rusk. Generalized Hamiltonian dynamics. II. Gauge transformations. Journal of Mathematical Physics, 24(11):2595– 2601, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i11/p2595_s1.

Marx:1983:TFD

[1661] Egon Marx. Transient fields in dispersive media. Journal of Mathematical Physics, 24(11):2602–2607, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i11/p2602_s1.

Huang:1983:CQM

[1662] Garng M. Huang, T. J. Tarn, and John W. Clark. On the controllability of quantum-mechanical systems. Journal of Mathematical Physics, 24 (11):2608–2618, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i11/p2608_s1.

Kristensson:1983:CMA

[1663] G. Kristensson, A. G. Ramm, and S. Ström. Convergence of the T -matrix approach in scattering theory. II. Journal of Mathematical Physics, 24 (11):2619–2631, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i11/p2619_s1.

Lorenz-Petzold:1983:PDS

[1664] Dieter Lorenz-Petzold. Positive-definite self-dual solutions of Einstein’s field equations. Journal of Mathematical Physics, 24(11):2632–2634, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i11/p2632_s1.

Mondaini:1983:SEY

[1665] R. P. Mondaini and N. O. Santos. Solutions of Einstein–Yang–Mills equations with plane symmetry. Journal of Mathematical Physics, 24 (11):2635–2636, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i11/p2635_s1.

Nishimori:1983:SML

[1666] Hidetoshi Nishimori and Robert B. Griffiths. Structure and motion of the Lee–Yang zeros. Journal of Mathematical Physics, 24(11):2637–2647, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i11/p2637_s1. See erratum [2147].

Arik:1983:GTD

[1667] M. Arik and F. Neyzi. Generalized two-dimensional O(3) sigma model. Journal of Mathematical Physics, 24(11):2648–2651, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/
p2648_s1.

MartinezAlonso:1983:ISV

[1668] L. Martínez Alonso. Inverse scattering variables of the KdV equation from the point of view of Galilean mechanics. Journal of Mathematical Physics, 24(11):2652–2657, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i11/p2652_s1.

Scheunert:1983:GTC

[1669] M. Scheunert. Graded tensor calculus. Journal of Mathematical Physics, 24(11):2658–2670, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i11/p2658_s1.

Scheunert:1983:CEL

[1670] M. Scheunert. Casimir elements of ε Lie algebras. Journal of Mathematical Physics, 24(11):2671–2680, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i11/p2671_s1.

Scheunert:1983:ECO

[1671] M. Scheunert. Eigenvalues of Casimir operators for the general linear, the special linear, and the orthosymplectic Lie superalgebras. Journal of Mathematical Physics, 24(11):2681–2688, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i11/p2681_s1.

Banerjee:1983:ACM

[1672] A. Banerjee and N. O. Santos. Anisotropic cosmological model with viscous fluid. Journal of Mathematical Physics, 24(11):2689–2691, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i11/p2689_s1.

Chen:1983:TCP

[1673] Jin Quan Chen, David F. Collinson, and Mei Juan Gao. Transformation coefficients of permutation groups. Journal of Mathematical Physics, 24 (12):2695–2705, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2695_s1.

Moylan:1983:URS

[1674] P. Moylan. Unitary representations of the (4+1)-de Sitter group on irreducible representation spaces of the Poincaré group. Journal of Mathematical Physics, 24(12):2706–2721, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i12/p2706_s1.

Okubo:1983:GIS

[1675] S. Okubo and J. Patera. General indices of simple Lie algebras and symmetrized product representations. Journal of Mathematical Physics, 24(12):2722–2733, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2722_s1.

Dragt:1983:CNB

[1676] Alex J. Dragt and Etienne Forest. Computation of nonlinear behavior of Hamiltonian systems using Lie algebraic methods. Journal of Mathematical Physics, 24(12):2734–2744, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i12/p2734_s1.

Asorey:1983:GCT

[1677] Manuel Asorey, José F. Cariñena, and Luis A. Ibort. Generalized canonical transformations for time-dependent systems. Journal of Mathematical Physics, 24(12):2745–2750, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i12/p2745_s1.

delaPena:1983:SEF

[1678] L. de la Peña and A. Jáuregui. Stochastic electrodynamics for the free particle. Journal of Mathematical Physics, 24(12):2751–2761, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i12/p2751_s1.

Froelich:1983:SPE

[1679] Piotr Froelich. Some properties of the eigenfunctions of the dilated model Hamiltonians with complex potentials. Journal of Mathematical Physics, 24(12):2762–2763, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2762_s1.

Laedke:1983:ETC

[1680] E. W. Laedke, K. H. Spatschek, and L. Stenflo. Evolution theorem for a class of perturbed envelope soliton solutions. Journal of Mathematical Physics, 24(12):2764–2769, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i12/p2764_s1.

Dutt:1983:EFS

[1681] Ranabir Dutt and Y. P. Varshni. Extension of Fuda’s off-shell analysis to screened Coulomb potentials for arbitrary l and limiting relations. Journal of Mathematical Physics, 24(12):2770–2775, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i12/p2770_s1.

Barwick:1983:NSI

[1682] John T. F. Barwick. A new semiclassical interpretation of the Lamb shift. Journal of Mathematical Physics, 24(12):2776–2779, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i12/p2776_s1.

Soffer:1983:ARQ

[1683] A. Soffer and L. P. Horwitz. B^{∗}-algebra representations in a quaternionic Hilbert module. Journal of Mathematical Physics, 24(12):2780– 2782, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i12/p2780_s1.

Krisch:1983:OFP

[1684] J. P. Krisch. A one-fixed-point Killing parameter transform. Journal of Mathematical Physics, 24(12):2783–2785, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i12/p2783_s1.

Szekeres:1983:SEE

[1685] Peter Szekeres. Solutions of Einstein’s equations involving arbitrary functions. Journal of Mathematical Physics, 24(12):2786–2792, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i12/p2786_s1.

Teyssandier:1983:CPT

[1686] P. Teyssandier and Ph. Tourrenc. The Cauchy problem for the R + R^{2} theories of gravity without torsion. Journal of Mathematical Physics, 24 (12):2793–2799, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2793_s1.

Ronen:1983:PRT

[1687] Y. Ronen and A. Dubi. A probabilistic rejection test for multivariable sensitivity analysis. Journal of Mathematical Physics, 24(12):2800–2808, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v24/i12/p2800_s1.

Driessler:1983:DSC

[1688] Wulf Driessler and Stephen J. Summers. A dense set of cyclic vectors for quantum field polynomial algebras. Journal of Mathematical Physics, 24(12):2809–2819, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2809_s1.

Garber:1983:GFT

[1689] Wolf-Dieter Garber. Galilean field theories and nonlocal S-matrix symmetries. Journal of Mathematical Physics, 24(12):2820–2822, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v24/i12/p2820_s1.

Lorente:1983:LBW

[1690] M. Lorente and M. A. Rodríguez. A Lagrangian of Bargmann–Wigner equations for massive particles of spin 2. Journal of Mathematical Physics, 24(12):2823–2827, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v24/i12/p2823_s1.

Binegar:1983:CQ

[1691] B. Binegar, C. Fronsdal, and W. Heidenreich. Conformal QED. Journal of Mathematical Physics, 24(12):2828–2846, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i12/p2828_s1.

Anders:1983:CSP

[1692] Till B. Anders and Wolfgang Jachmann. Cross sections with polarized spin-1/2 particles in terms of helicity amplitudes. Journal of Mathematical Physics, 24(12):2847–2854, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v24/i12/p2847_s1.

Bagrov:1983:QTC

[1693] V. G. Bagrov, V. V. Belov, and I. M. Ternov. Quasiclassical trajectorycoherent states of a particle in an arbitrary electromagnetic field. Journal of Mathematical Physics, 24(12):2855–2859, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v24/i12/p2855_s1.

Yoshida:1983:ONS

[1694] Zensho Yoshida and Yoshikazu Giga. On the Ohm–Navier–Stokes system in magnetohydrodynamics. Journal of Mathematical Physics, 24 (12):2860–2864, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2860_s1.

Jaric:1983:SPT

[1695] Marko V. Jarić. Structural phase transitions in crystals: Brokensymmetry (isotropy) groups. Journal of Mathematical Physics, 24 (12):2865–2882, December 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v24/i12/p2865_s1.

DHoker:1984:DRB

[1696] E. D’Hoker. Decomposition of representations into basis representations for the classical groups. Journal of Mathematical Physics, 25(1): 1–12, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i1/p1_s1.

Weiss:1984:CIS

[1697] John Weiss. On classes of integrable systems and the Painlevé property. Journal of Mathematical Physics, 25(1):13–24, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/p13_s1.

Gerdjikov:1984:ESS

[1698] V. S. Gerdjikov, M. I. Ivanov, and P. P. Kulish. Expansions over the “squared” solutions and difference evolution equations. Journal of Mathematical Physics, 25(1):25–34, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p25_s1.

Gilain:1984:NIE

[1699] C. Gilain and D. Lévy. A new integral equation for summing Feynman graph series (general scalar Lagrangian case). Journal of Mathematical Physics, 25(1):35–47, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i1/p35_s1.

Harrell:1984:HOM

[1700] Evans M. Harrell II. Hamiltonian operators with maximal eigenvalues. Journal of Mathematical Physics, 25(1):48–51, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/p48_s1. See

erratum [2223].

Potthoff:1984:SPO

[1701] Jürgen Potthoff. Stochastic path-ordered exponentials. Journal of Mathematical Physics, 25(1):52–56, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p52_s1.

Hartle:1984:PIP

[1702] James B. Hartle and Karel V. Kuchař. Path integrals in parametrized theories: Newtonian systems. Journal of Mathematical Physics, 25(1): 57–75, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i1/p57_s1.

Fannes:1984:QEE

[1703] M. Fannes, P. Vanheuverzwijn, and A. Verbeure. Quantum energyentropy inequalities: A new method for proving the absence of symmetry breaking. Journal of Mathematical Physics, 25(1):76–78, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i1/p76_s1.

Ozawa:1984:QMP

[1704] Masanao Ozawa. Quantum measuring processes of continuous observables. Journal of Mathematical Physics, 25(1):79–87, January 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/p79_s1.

Zheng:1984:DTS

[1705] W. M. Zheng. The Darboux transformation and solvable double-well potential models for Schrödinger equations. Journal of Mathematical Physics, 25(1):88–90, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i1/p88_s1.

Carroll:1984:TMP

[1706] Robert Carroll and Stanley Dolzycki. Transmutation as a minimizing procedure. Journal of Mathematical Physics, 25(1):91–93, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/
p91_s1.

Cheney:1984:ISD

[1707] Margaret Cheney. Inverse scattering in dimension two. Journal of Mathematical Physics, 25(1):94–107, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p94_s1.

Moses:1984:EEA

[1708] Harry E. Moses and Reese T. Prosser. Eigenvalues and eigenfunctions associated with the Gel’fand–Levitan equation. Journal of Mathematical Physics, 25(1):108–112, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i1/p108_s1.

Lester:1984:CAS

[1709] J. A. Lester. The causal automorphism of de Sitter and Einstein cylinder spacetimes. Journal of Mathematical Physics, 25(1):113–116, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i1/p113_s1.

Kalinowski:1984:SSS

[1710] M. W. Kalinowski and G. Kunstatter. Spherically symmetric solution in the nonsymmetric Kaluza–Klein theory. Journal of Mathematical Physics, 25(1):117–130, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i1/p117_s1.

Larsen:1984:SMI

[1711] Edward W. Larsen. Solution of multidimensional inverse transport problems. Journal of Mathematical Physics, 25(1):131–135, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/
p131_s1.

Brown:1984:SHS

[1712] M. R. Brown. Symmetric Hadamard series. Journal of Mathematical Physics, 25(1):136–140, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i1/p136_s1.

Capri:1984:PSM

[1713] Anton Z. Capri and Ruggero Ferrari. Properties of the Schwinger model. Journal of Mathematical Physics, 25(1):141–149, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/p141_s1.

Kerner:1984:NME

[1714] Edward H. Kerner. A novel mass-eigenvalue problem for spinors in de Sitter space. Journal of Mathematical Physics, 25(1):150–153, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i1/p150_s1.

vanderHoek:1984:VPF

[1715] John van der Hoek and M. A. Lohe. Vortex properties in first- and second-order formulations of abelian gauge theories. Journal of Mathematical Physics, 25(1):154–160, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p154_s1.

McKellar:1984:GPG

[1716] R. J. McKellar. A gravitational Poincaré gauge theory and Higgs mechanism. Journal of Mathematical Physics, 25(1):161–166, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/
p161_s1.

Balachandran:1984:UGA

[1717] A. P. Balachandran, G. Marmo, N. Mukunda, J. S. Nilsson, A. Simoni, E. C. G. Sudarshan, and F. Zaccaria. Unified geometrical approach to relativistic particle dynamics. Journal of Mathematical Physics, 25(1):167– 176, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i1/p167_s1.

Smith:1984:STF

[1718] P. D. Smith. A smooth transonic flow in the plane. Journal of Mathematical Physics, 25(1):177–184, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p177_s1.

Larsson:1984:SCS

[1719] Jonas Larsson. Symmetry of the complete second-order conductivity tensor in a Vlasov plasma. Journal of Mathematical Physics, 25(1):185– 186, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i1/p185_s1.

Asorey:1984:ESR

[1720] M. Asorey. Erratum: “Some remarks on the classical vacuum structure of gauge field theories”, [J. Math. Phys. 22 (1981), 179–184; MR 82d:81094]. Journal of Mathematical Physics, 25(1):187, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i1/

p187_s1. See [461].

Brannigan:1984:ESG

[1721] M. Brannigan and D. Eyre. Erratum: “Splines and the Galerkin method for solving the integral equations of scattering theory”. Journal of Mathematical Physics, 25(1):187, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1514].

Brannigan:1984:ESP

[1722] M. Brannigan and D. Eyre. Erratum: “Splines and the projection collocation method for solving integral equations in scattering theory”, [J. Math. Phys. 24 (1983), 177–183; MR 84b:65125]. Journal of Mathematical Physics, 25(1):187, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i1/p187_s2. See [1311].

Kim:1984:UTPa

[1723] Shoon K. Kim. A unified theory of the point groups. V. The general projective corepresentations of the magnetic point groups and their applications to magnetic space groups. Journal of Mathematical Physics, 25(2):189–196, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p189_s1.

Kim:1984:GIC

[1724] Shoon K. Kim. General irreducibility condition for vector and projective corepresentations of antiunitary groups. Journal of Mathematical Physics, 25(2):197–199, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i2/p197_s1.

Klein:1984:SIG

[1725] D. J. Klein. Semiregular induction of group representations. Journal of Mathematical Physics, 25(2):200–203, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p200_s1.

Nash:1984:ISG

[1726] Patrick L. Nash. Identities satisfied by the generators of the Dirac algebra. Journal of Mathematical Physics, 25(2):204–209, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i2/
p204_s1.

Aguirre:1984:IST

[1727] M. Aguirre and J. Krause. Infinitesimal symmetry transformations of some one-dimensional linear systems. Journal of Mathematical Physics, 25(2):210–218, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p210_s1.

Okubo:1984:GIR

[1728] S. Okubo and J. Patera. General indices of representations and Casimir invariants. Journal of Mathematical Physics, 25(2):219–227, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p219_s1.

Kalikstein:1984:VBT

[1729] Kalman Kalikstein, Larry Spruch, and Alberto Baider. Variational bounds on the temperature distribution. Journal of Mathematical Physics, 25(2):228–236, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i2/p228_s1.

Steeb:1984:LBV

[1730] W.-H. Steeb. On Lie–Bäcklund vector fields of the evolution equations ∂^{2}u/∂x∂t = f(u)and∂u/∂t = ∂^{2}u/∂x^{2} + f(u). Journal of Mathematical Physics, 25(2):237–239, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i2/p237_s1.

Morris:1984:SSA

[1731] Hedley C. Morris and Roger K. Dodd. Singular solutions of the axially symmetric Bogomolny equations. Journal of Mathematical Physics, 25(2):240–244, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p240_s1.

Gaffet:1984:SEU

[1732] B. Gaffet. SU(3) symmetry of the equations of unidimensional gas flow, with arbitrary entropy distribution. Journal of Mathematical Physics, 25(2):245–255, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p245_s1.

Lund:1984:CHC

[1733] Fernando Lund. The continuous Heisenberg chain and constrained harmonic motion. Journal of Mathematical Physics, 25(2):256–260, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p256_s1.

McQuistan:1984:ERC

[1734] R. B. McQuistan and J. L. Hock. An exact recursion for the composite nearest-neighbor degeneracy for a 2 × N lattice space. Journal of Mathematical Physics, 25(2):261–265, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p261_s1.

Braunss:1984:CSS

[1735] Günter Braunss. On the construction of state spaces for classical dynamical systems with a time-dependent Hamilton function. Journal of Mathematical Physics, 25(2):266–270, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p266_s1.

Cantrijn:1984:SAN

[1736] F. Cantrijn. Symplectic approach to nonconservative mechanics. Journal of Mathematical Physics, 25(2):271–276, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p271_s1.

Kaup:1984:FTL

[1737] D. J. Kaup. The forced Toda lattice: An example of an almost integrable system. Journal of Mathematical Physics, 25(2):277–281, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p277_s1.

Kaup:1984:SBR

[1738] D. J. Kaup and D. H. Neuberger. The soliton birth rate in the forced Toda lattice. Journal of Mathematical Physics, 25(2):282–284, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p282_s1.

Fishman:1984:DAEa

[1739] Louis Fishman and John J. McCoy. Derivation and application of extended parabolic wave theories. I. The factorized Helmholtz equation. Journal of Mathematical Physics, 25(2):285–296, February 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i2/p285_
s1.

Fishman:1984:DAEb

[1740] Louis Fishman and John J. McCoy. Derivation and application of extended parabolic wave theories. II. Path integral representations. Journal of Mathematical Physics, 25(2):297–308, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p297_s1.

Censor:1984:DEN

[1741] Dan Censor and D. M. Le Vine. The Doppler effect: Now you see it, now you don’t. Journal of Mathematical Physics, 25(2):309–316, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p309_s1.

Yamani:1984:MRK

[1742] Hashim A. Yamani. The J-matrix reproducing kernel: Numerical weights at the Harris energy eigenvalues. Journal of Mathematical Physics, 25(2):317–322, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p317_s1.

Talukdar:1984:CMN

[1743] B. Talukdar, D. K. Ghosh, and T. Sasakawa. Coulomb-modified nuclear scattering. III. Journal of Mathematical Physics, 25(2):323–326, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p323_s1.

Gering:1984:FCS

[1744] M. Z. I. Gering and W. D. Heiss. A fully crossing symmetric vertex function. Journal of Mathematical Physics, 25(2):327–329, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p327_s1.

Mallett:1984:NFW

[1745] Ronald L. Mallett. A note on fluctuations of weak gravitational radiation and the classical Hanbury Brown–Twiss effect. Journal of Mathematical Physics, 25(2):330–331, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i2/p330_s1.

White:1984:CSF

[1746] A. J. White and C. B. Collins. A class of shear-free perfect fluids in general relativity. I. Journal of Mathematical Physics, 25(2):332–337, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i2/p332_s1.

Stachel:1984:GFS

[1747] John Stachel. The gravitational fields of some rotating and nonrotating cylindrical shells of matter. Journal of Mathematical Physics, 25(2):338– 341, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i2/p338_s1.

TorresdelCastillo:1984:NSH

[1748] G. F. Torres del Castillo. Null strings and Hertz potentials. Journal of Mathematical Physics, 25(2):342–346, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p342_s1.

Moffat:1984:HDR

[1749] J. W. Moffat. Higher-dimensional Riemannian geometry and quaternion and octonion spaces. Journal of Mathematical Physics, 25(2):347– 350, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i2/p347_s1. See erratum [2215].

Rajput:1984:UGE

[1750] B. S. Rajput. Unification of generalized electromagnetic and gravitational fields. Journal of Mathematical Physics, 25(2):351–353, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p351_s1.

Sancho:1984:SPD

[1751] J. M. Sancho. Stochastic processes driven by dichotomous Markov noise: Some exact dynamical results. Journal of Mathematical Physics, 25(2):354–359, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p354_s1.

Castagnino:1984:SPF

[1752] Mario A. Castagnino, Luis Chimento, Diego D. Harari, and Carmen A. Núñez. A spin-1/2 particle formalism in curved space–time. Journal of Mathematical Physics, 25(2):360–367, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p360_s1.

Harnad:1984:SSB

[1753] J. Harnad, Y. Saint-Aubin, and S. Shnider. Superposition of solutions to Bäcklund transformations for the SU(n) principal σ-model. Journal of Mathematical Physics, 25(2):368–375, February 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i2/p368_s1.

Heidenreich:1984:GTD

[1754] W. Heidenreich. Group theory of the dipole ghost. Journal of Mathematical Physics, 25(2):376–379, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i2/p376_s1.

Luehr:1984:UGT

[1755] C. P. Luehr and M. Rosenbaum. Unified gauge theory for electromagnetism and gravitation based on twistor bundles. Journal of Mathematical Physics, 25(2):380–387, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i2/p380_s1.

Castanos:1984:HSC

[1756] O. Castaños and A. Frank. The U(6)∧SU(3) hidden symmetry in collective excitations of many-body systems. Journal of Mathematical Physics, 25(2):388–395, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p388_s1.

Hameiri:1984:BSR

[1757] Eliezer Hameiri and Peter Laurence. The ballooning spectrum of rotating plasmas. Journal of Mathematical Physics, 25(2):396–405, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i2/p396_s1.

Fernandez:1984:GAG

[1758] Ariel Fernández and Oktay Sinanoğlu. Global attractors and global stability for closed chemical systems. Journal of Mathematical Physics, 25(2):406–409, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p406_s1.

Karadayi:1984:ANC

[1759] Hasan R. Karadayı. On the anomaly number of the classical groups. Journal of Mathematical Physics, 25(3):411–413, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/p411_s1.

Wene:1984:AAB

[1760] G. P. Wene. Algebras with anticommuting basal elements, space–time symmetries, and quantum theory. Journal of Mathematical Physics, 25(3):414–416, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p414_s1.

Nikam:1984:DGB

[1761] R. S. Nikam and C. R. Sarma. Double-Gel’fand boson polynomials and the permutation group. Journal of Mathematical Physics, 25(3):417– 421, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p417_s1.

Klimyk:1984:IOSa

[1762] A. U. Klimyk and A. M. Gavrilik. Infinitesimal operators and structure of the representations of the groups SO^{∗}(2n) and SO(2n) in a U(n) basis and of the groups SU^{∗}(2n) and SU(2n) in an Sp(n) basis. Journal of Mathematical Physics, 25(3):422–430, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i3/p422_s1.

Dundarer:1984:DRT

[1763] Resit Dündarer and Feza Gürsey. Dyson representation of SU(3) in terms of five boson operators. Journal of Mathematical Physics, 25(3):431– 432, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p431_s1.

Taormina:1984:COS

[1764] A. Taormina. Charge operators in simple Lie groups. Journal of Mathematical Physics, 25(3):433–442, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p433_s1.

Mitra:1984:CLA

[1765] Bani Mitra, B. R. Sitaram, and K. C. Tripathy. Cohomology of Lie algebras with a nontrivial center. Journal of Mathematical Physics, 25(3): 443–444, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p443_s1.

GuilGuerrero:1984:SIS

[1766] F. Guil Guerrero. Specializations of integrable systems and affine Lie algebras. Journal of Mathematical Physics, 25(3):445–450, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/
p445_s1.

Cianci:1984:SAS

[1767] Roberto Cianci. Supermanifolds and automorphisms of super Lie groups. Journal of Mathematical Physics, 25(3):451–455, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/p451_s1.

Gudder:1984:FQP

[1768] Stanley P. Gudder. Finite quantum processes. Journal of Mathematical Physics, 25(3):456–465, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p456_s1.

Peratt:1984:CFE

[1769] Anthony L. Peratt. Continued fraction expansions for the complete, incomplete, and relativistic plasma dispersion functions. Journal of Mathematical Physics, 25(3):466–468, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p466_s1.

Ramm:1984:SPB

[1770] A. G. Ramm. Scattering by a penetrable body. Journal of Mathematical Physics, 25(3):469–471, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p469_s1.

dePrunele:1984:TTR

[1771] E. de Prunelé. Three-term recursion relations for hydrogen wave functions: Exact calculations and semiclassical approximations. Journal of Mathematical Physics, 25(3):472–480, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i3/p472_s1.

Dorizzi:1984:EIH

[1772] B. Dorizzi, B. Grammaticos, and A. Ramani. Explicit integrability for Hamiltonian systems and the Painlevé conjecture. Journal of Mathematical Physics, 25(3):481–485, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p481_s1.

Leach:1984:FIS

[1773] P. G. L. Leach, H. R. Lewis, and W. Sarlet. First integrals for some nonlinear time-dependent Hamiltonian systems. Journal of Mathematical Physics, 25(3):486–490, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p486_s1.

May:1984:GVP

[1774] Helge-Otmar May. Generalized variational principles and nondifferentiable potentials in analytical mechanics. Journal of Mathematical Physics, 25(3):491–493, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i3/p491_s1.

Mazuruk:1984:DCF

[1775] Konstanty Mazuruk. The direct correlation function for the hard-core potentials. Journal of Mathematical Physics, 25(3):494–498, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/
p494_s1.

Robinson:1984:LCM

[1776] Ivor Robinson and Krzysztof Rózga. Lightlike contractions on Minkowski space–time. Journal of Mathematical Physics, 25(3):499–505, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p499_s1.

Aldaya:1984:QSN

[1777] V. Aldaya, J. A. de Azcárraga, and K. B. Wolf. Quantization, symmetry, and natural polarization. Journal of Mathematical Physics, 25(3):506– 512, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p506_s1.

Cattaneo:1984:OOA

[1778] Gianpiero Cattaneo and Giuseppe Nisticò. Orthogonality and orthocomplementations in the axiomatic approach to quantum mechanics: Remarks about some critiques. Journal of Mathematical Physics, 25(3): 513–531, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p513_s1.

Pechukas:1984:SIQ

[1779] Philip Pechukas. Sharpening an inequality in quantum ergodic theory. Journal of Mathematical Physics, 25(3):532–534, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/p532_s1.

Leaf:1984:QTD

[1780] Boris Leaf. Quantization of three-dimensional space. Journal of Mathematical Physics, 25(3):535–540, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p535_s1.

Curtright:1984:EPQ

[1781] T. L. Curtright and C. B. Thorn. The effective potential in quantum mechanics. Journal of Mathematical Physics, 25(3):541–548, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p541_s1.

Zinn-Justin:1984:IQM

[1782] J. Zinn-Justin. Instantons in quantum mechanics: Numerical evidence for a conjecture. Journal of Mathematical Physics, 25(3):549–555, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p549_s1.

Iyer:1984:PST

[1783] V. P. Iyer and L. K. Sharma. Perturbative solutions in two-channel Schrödinger and Klein–Gordon equations. Journal of Mathematical Physics, 25(3):556–563, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i3/p556_s1.

Papiez:1984:SFF

[1784] Lech Papież. Stochastic formulation of Feynman path integrals from the least action point of view. Journal of Mathematical Physics, 25(3): 564–569, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p564_s1.

Asano:1984:NSA

[1785] Naruyoshi Asano and Yusuke Kato. Non-self-adjoint Zakharov–Shabat operator with a potential of the finite asymptotic values. II. Inverse problem. Journal of Mathematical Physics, 25(3):570–588, March 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/p570_
s1.

Robinson:1984:SFC

[1786] Ivor Robinson and Krzysztof Rózga. On some family of congruences of null strings. Journal of Mathematical Physics, 25(3):589–596, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p589_s1.

Parker:1984:STG

[1787] Phillip E. Parker. On some theorems of Geroch and Stiefel. Journal of Mathematical Physics, 25(3):597–599, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i3/p597_s1.

Bicak:1984:ICG

[1788] J. Bičák and B. G. Schmidt. Isometries compatible with gravitational radiation. Journal of Mathematical Physics, 25(3):600–606, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/
p600_s1.

Krori:1984:ABH

[1789] K. D. Krori, Sumita Chaudhury, and S. Dowerah. Accelerating black hole in a magnetic field. Journal of Mathematical Physics, 25(3):607– 611, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p607_s1.

Post:1984:MDF

[1790] E. J. Post. The metric dependence of four-dimensional formulations of electromagnetism. Journal of Mathematical Physics, 25(3):612– 613, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p612_s1.

Majewski:1984:DBC

[1791] W. A. Majewski. The detailed balance condition in quantum statistical mechanics. Journal of Mathematical Physics, 25(3):614–616, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p614_s1.

Jakobczyk:1984:BAF

[1792] Lech Jakóbczyk. Borchers algebra formulation of an indefinite inner product quantum field theory. Journal of Mathematical Physics, 25(3): 617–622, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p617_s1.

Najmi:1984:SQF

[1793] Amir-Homayoon Najmi. Spin-1/2 quantum field theory in nonstatic space–times. I. construction of states. Journal of Mathematical Physics, 25(3):623–628, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p623_s1.

Gallatin:1984:ZFR

[1794] G. M. Gallatin. On zeta function regularization for operators with continuous spectra. Journal of Mathematical Physics, 25(3):629–632, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p629_s1.

Osipov:1984:PAC

[1795] Edward P. Osipov. A possible approach to the construction of the : exp ξ:_{d} quantum field theory in a finite volume. Journal of Mathematical Physics, 25(3):633–648, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p633_s1.

Rudiger:1984:SSD

[1796] Roland Rüdiger. Separable systems for the Dirac equation in curved space–times. Journal of Mathematical Physics, 25(3):649–654, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p649_s1.

Hussin:1984:WGL

[1797] V. Hussin and J. Beckers. On the Weyl group, its little groups, and associated massless particle descriptions. Journal of Mathematical Physics, 25(3):655–659, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p655_s1.

Oh:1984:BMM

[1798] C. H. Oh. Bifurcation and the magnetic multipole solutions of the Yang– Mills equations with sources. Journal of Mathematical Physics, 25(3): 660–662, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p660_s1.

Mielke:1984:RPG

[1799] Eckehard W. Mielke. Reduction of the Poincaré gauge field equations by means of duality rotations. Journal of Mathematical Physics, 25(3): 663–668, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p663_s1.

Dodd:1984:BTS

[1800] Roger K. Dodd and Hedley C. Morris. Bäcklund transformations for spherically symmetric solutions to the self-dual Yang–Mills equations. Journal of Mathematical Physics, 25(3):669–672, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i3/p669_s1.

Saliu:1984:CGG

[1801] L. Saliu and L. Tătaru. Compact gauge group representations without triangular anomalies. Journal of Mathematical Physics, 25(3):673– 677, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p673_s1.

Ekhaguere:1984:TSR

[1802] G. O. S. Ekhaguere. The theory of superselection rules. II. Sectors of the free electromagnetic field. Journal of Mathematical Physics, 25(3): 678–683, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i3/p678_s1.

deMaag:1984:CMS

[1803] J. W. de Maag, L. P. Kok, and H. van Haeringen. Coulomb-modified scattering parameters for Coulomb-plus-separable potentials for all l. Journal of Mathematical Physics, 25(3):684–692, March 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i3/p684_s1.

Srivastava:1984:SPT

[1804] Sushil K. Srivastava. On the survival of primordial tachyons up to the present epoch. Journal of Mathematical Physics, 25(3):693–697, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i3/p693_s1.

Niukkanen:1984:GDO

[1805] A. W. Niukkanen. Generating differential operators for the basis functions in the variational LCAO-type methods. Journal of Mathematical Physics, 25(3):698–705, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p698_s1.

Salingaros:1984:RMC

[1806] Nikos Salingaros. Relativistic motion of a charged particle, the Lorentz group, and the Thomas precession. Journal of Mathematical Physics, 25(3):706–716, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i3/p706_s1.

Eringen:1984:TNP

[1807] A. Cemal Eringen. Theory of nonlocal piezoelectricity. Journal of Mathematical Physics, 25(3):717–727, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p717_s1.

Walter:1984:EPA

[1808] O. Walter, L. S. Cederbaum, and J. Schirmer. The eigenvalue problem for “arrow” matrices. Journal of Mathematical Physics, 25(4):729– 737, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p729_s1.

Salingaros:1984:RBF

[1809] Nikos Salingaros. The relationship between finite groups and Clifford algebras. Journal of Mathematical Physics, 25(4):738–742, April 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p738_
s1.

Klimyk:1984:IOSb

[1810] A. U. Klimyk and Bruno Gruber. Infinitesimal operators and structure of the most degenerate representations of the groups Sp(p+q) and Sp(p, q) in an Sp(p) × Sp(q) basis. Journal of Mathematical Physics, 25(4):743– 750, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p743_s1.

DeMeyer:1984:RBS

[1811] H. De Meyer, G. Vanden Berghe, and J. Van der Jeugt. Realizations of F_{4} in SO(3)⊗SO(3) bases and structural zeros of the 6j-symbol. Journal of Mathematical Physics, 25(4):751–754, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p751_s1.

Gruber:1984:MEI

[1812] Bruno Gruber and Anatoli U. Klimyk. Matrix elements for indecomposable representations of complex su(2). Journal of Mathematical Physics, 25(4):755–764, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p755_s1.

Taflin:1984:FLN

[1813] E. Taflin. Formal linearization of nonlinear massive representations of the connected Poincaré group. Journal of Mathematical Physics, 25(4): 765–771, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p765_s1.

VandenBerghe:1984:LAB

[1814] G. Vanden Berghe and H. De Meyer. On Lie algebras built from SO(3) tensor operators. Journal of Mathematical Physics, 25(4):772–773, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p772_s1.

Figotin:1984:PLE

[1815] A. L. Figotin and L. A. Pastur. The positivity of the Lyapunov exponent and the absence of the absolutely continuous spectrum for the almost-Mathieu equation. Journal of Mathematical Physics, 25(4):774– 777, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p774_s1.

Hermann:1984:GFI

[1816] Robert Hermann. The geometric foundations of the integrability property of differential equations and physical systems. II. Mechanics on affinely-connected manifolds and the work of Kowalewski and Painlevé. Journal of Mathematical Physics, 25(4):778–785, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p778_s1.

Brull:1984:SLE

[1817] L. Brüll and H. Lange. The Schrödinger–Langevin equation: Special solutions and nonexistence of solitary waves. Journal of Mathematical Physics, 25(4):786–790, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p786_s1.

Grundland:1984:SRN

[1818] A. M. Grundland, J. Harnad, and P. Winternitz. Symmetry reduction for nonlinear relativistically invariant equations. Journal of Mathematical Physics, 25(4):791–806, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p791_s1.

Ramm:1984:RSH

[1819] A. G. Ramm. Representation of solutions to Helmholtz’s equation. Journal of Mathematical Physics, 25(4):807–809, April 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i4/p807_s1.

Singh:1984:STD

[1820] G. S. Singh and L. S. Kothari. On the solution of the two-dimensional Helmholtz equation. Journal of Mathematical Physics, 25(4):810–811, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p810_s1.

Davis:1984:CCS

[1821] E. D. Davis and W. D. Heiss. Convergence of a crossing-symmetric perturbation series for the four-point vertex. Journal of Mathematical Physics, 25(4):812–813, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p812_s1.

Verdiyev:1984:ITL

[1822] Yi. A. Verdiyev. An integral transformation in Lobachevsky space. Journal of Mathematical Physics, 25(4):814, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p814_s1.

Fuller:1984:PCS

[1823] William Roger Fuller. Planar coherent states. Journal of Mathematical Physics, 25(4):815–819, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p815_s1.

Nam:1984:GTM

[1824] Kiwon Nam and Michael J. Moravcsik. General transformation matrix for Dirac spinors and the calculation of spinorial amplitudes. Journal of Mathematical Physics, 25(4):820–827, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p820_s1.

Hasselbarth:1984:ODC

[1825] Werner Hässelbarth. Orthogonal decompositions of certain finitedimensional vector spaces. Journal of Mathematical Physics, 25(4):828– 832, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p828_s1.

Hoyos:1984:GSS

[1826] J. Hoyos, M. Quirós, J. Ramírez Mittelbrunn, and F. J. de Urríes. Generalized supermanifolds. I. Superspaces and linear operators. Journal of Mathematical Physics, 25(4):833–840, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p833_s1.

Hoyos:1984:GSIa

[1827] J. Hoyos, M. Quirós, J. Ramírez Mittelbrunn, and F. J. de Urríes. Generalized supermanifolds. II. Analysis on superspaces. Journal of Mathematical Physics, 25(4):841–846, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p841_s1.

Hoyos:1984:GSIb

[1828] J. Hoyos, M. Quirós, J. Ramírez Mittelbrunn, and F. J. de Urríes. Generalized supermanifolds. III. ρ-supermanifolds. Journal of Mathematical Physics, 25(4):847–854, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p847_s1.

VilelaMendes:1984:SSP

[1829] R. Vilela Mendes. Symmetries and stable periodic orbits for onedimensional maps. Journal of Mathematical Physics, 25(4):855–857, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p855_s1.

Applebaum:1984:FD

[1830] D. Applebaum and R. L. Hudson. Fermion diffusions. Journal of Mathematical Physics, 25(4):858–861, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p858_s1.

Garbaczewski:1984:QSF

[1831] Piotr Garbaczewski. Quantization of spinor fields. IV. Joint Bose–Fermi spectral problems. Journal of Mathematical Physics, 25(4):862–871, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p862_s1.

Ullah:1984:EII

[1832] Nazakat Ullah. Evaluation of an integral involving associated Legendre polynomials and inverse powers of (1 − X^{2}). Journal of Mathematical Physics, 25(4):872–873, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p872_s1.

Amorim:1984:PCV

[1833] R. Amorim. A projective constrained variational principle for a classical particle with spin. Journal of Mathematical Physics, 25(4):874– 877, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p874_s1.

Ramani:1984:IPP

[1834] A. Ramani, B. Dorizzi, B. Grammaticos, and T. Bountis. Integrability and the Painlevé property for low-dimensional systems. Journal of Mathematical Physics, 25(4):878–883, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p878_s1.

Verosky:1984:HOS

[1835] John Verosky. Higher-order symmetries of the compressible onedimensional isentropic fluid equations. Journal of Mathematical Physics, 25(4):884–888, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p884_s1.

Chen:1984:GMC

[1836] Feng Su Chen. A Galerkin method to cylindrical KdV equations. Journal of Mathematical Physics, 25(4):889–891, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p889_s1.

Santini:1984:LIL

[1837] P. M. Santini, M. J. Ablowitz, and A. S. Fokas. On the limit from the intermediate long wave equation to the Benjamin–Ono equation. Journal of Mathematical Physics, 25(4):892–899, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p892_s1.

Satsuma:1984:BTS

[1838] Junkichi Satsuma, Thiab R. Taha, and Mark J. Ablowitz. On a Bäcklund transformation and scattering problem for the modified intermediate long wave equation. Journal of Mathematical Physics, 25(4):900–904, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p900_s1.

Blinder:1984:PIR

[1839] S. M. Blinder. Propagators from integral representations of Green’s functions for the N-dimensional free-particle, harmonic oscillator and Coulomb problems. Journal of Mathematical Physics, 25(4):905–909, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p905_s1.

Flessas:1984:ATO

[1840] George P. Flessas, R. R. Whitehead, and A. Rigas. Analytic treatment of the αx^{2} + βx^{4} oscillator. Journal of Mathematical Physics, 25(4): 910–917, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p910_s1.

Oevel:1984:IMC

[1841] W. Oevel and A. S. Fokas. Infinitely many commuting symmetries and constants of motion in involution for explicitly time-dependent evolution equations. Journal of Mathematical Physics, 25(4):918–922, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p918_s1.

Hamoui:1984:GDS

[1842] Adnan Hamoui and André Lichnerowicz. Geometry of the dynamical systems with time-dependent constraints and time-dependent Hamiltonian: An approach towards quantization. Journal of Mathematical Physics, 25(4):923–931, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p923_s1.

Arteca:1984:EAO

[1843] Gustavo A. Arteca, Francisco M. Fernández, and Eduardo A. Castro. Eigenvalues of anharmonic oscillators from a variational functional method. Journal of Mathematical Physics, 25(4):932–938, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/
p932_s1.

Cutkosky:1984:HFM

[1844] R. E. Cutkosky. Harmonic functions and matrix elements for hyperspherical quantum field models. Journal of Mathematical Physics, 25(4): 939–942, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p939_s1.

Mlodinow:1984:SSE

[1845] Leonard D. Mlodinow and Michael P. Shatz. Solving the Schrödinger equation with use of 1/N perturbation theory. Journal of Mathematical Physics, 25(4):943–950, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p943_s1.

Ben-Artzi:1984:UES

[1846] Matania Ben-Artzi. Unitary equivalence and scattering theory for Starklike Hamiltonians. Journal of Mathematical Physics, 25(4):951–964, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p951_s1.

Sidharth:1984:MPS

[1847] B. G. Sidharth. The minimum potential strength for the existence of bound states. Journal of Mathematical Physics, 25(4):965–968, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p965_s1.

Garbaczewski:1984:QST

[1848] Piotr Garbaczewski. On quantum solitons and their classical relatives: Reducible quantum fields and infinite constituent “elementary” systems. Journal of Mathematical Physics, 25(4):969–981, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p969_s1.

Dusek:1984:SSS

[1849] Jan Dušek. On-shell singularities of the off-shell Jost functions and Jost solutions for long-range potentials. Journal of Mathematical Physics, 25(4):982–986, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p982_s1.

Narnhofer:1984:TDD

[1850] Heide Narnhofer. Time delay and dilation properties in scattering theory. Journal of Mathematical Physics, 25(4):987–991, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p987_s1.

Cohen:1984:TEM

[1851] Jeffrey M. Cohen and F. de Felice. The total effective mass of the Kerr– Newman metric. Journal of Mathematical Physics, 25(4):992–994, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i4/p992_s1.

Collins:1984:SFP

[1852] C. B. Collins. Shear-free perfect fluids with zero magnetic Weyl tensor. Journal of Mathematical Physics, 25(4):995–1000, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p995_s1.

McNamara:1984:BTE

[1853] Paul McNamara and John Shepanski. On bilinear and trilinear extensions to the form of the Einstein tensor. Journal of Mathematical Physics, 25(4):1001–1003, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i4/p1001_s1.

Reula:1984:PBE

[1854] O. Reula and K. P. Tod. Positivity of the Bondi energy. Journal of Mathematical Physics, 25(4):1004–1008, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i4/p1004_s1.

Kunzle:1984:HFG

[1855] H. P. Künzle and J. M. Nester. Hamiltonian formulation of gravitating perfect fluids and the Newtonian limit. Journal of Mathematical Physics, 25(4):1009–1018, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i4/p1009_s1.

Kamran:1984:SVS

[1856] N. Kamran and R. G. McLenaghan. Separation of variables and symmetry operators for the neutrino and Dirac equations in the space–times admitting a two-parameter abelian orthogonally transitive isometry group and a pair of shearfree geodesic null congruences. Journal of Mathematical Physics, 25(4):1019–1027, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p1019_s1.

Kulkarni:1984:TAI

[1857] Anil D. Kulkarni. Time-asymmetric initial data for the N black hole problem in general relativity. Journal of Mathematical Physics, 25(4):1028–1034, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i4/p1028_s1.

Simon:1984:MES

[1858] Walter Simon. The multipole expansion of stationary Einstein–Maxwell fields. Journal of Mathematical Physics, 25(4):1035–1038, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/
p1035_s1.

Toller:1984:CME

[1859] M. Toller and R. Vaia. A complete multipole expansion for a test particle in geometric fields. Journal of Mathematical Physics, 25(4):1039– 1044, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i4/p1039_s1.

Kalinowski:1984:MSN

[1860] M. W. Kalinowski. Material sources in the nonsymmetric Kaluza–Klein theory. Journal of Mathematical Physics, 25(4):1045–1049, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/
p1045_s1.

Frigerio:1984:DPQ

[1861] Alberto Frigerio and Vittorio Gorini. Diffusion processes, quantum dynamical semigroups, and the classical KMS condition. Journal of Mathematical Physics, 25(4):1050–1065, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p1050_s1.

Hernandez-Machado:1984:DPN

[1862] A. Hernández-Machado and M. San Miguel. Dynamical properties of non-Markovian stochastic differential equations. Journal of Mathematical Physics, 25(4):1066–1075, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i4/p1066_s1.

Ganapol:1984:GFS

[1863] B. D. Ganapol and E. B. Parlette. The generalized function solution of the Fokker–Planck equation for monoenergetic charged particle transport. Journal of Mathematical Physics, 25(4):1076–1079, April 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1076_
s1.

Mathews:1984:ICU

[1864] P. M. Mathews and B. Vijayalakshmi. On inequivalent classes of uniquemass-spin relativistic wave equations involving repeated irreducible representations with arbitrary multiplicities. Journal of Mathematical Physics, 25(4):1080–1087, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i4/p1080_s1.

Reifler:1984:VWE

[1865] Frank Reifler. A vector wave equation for neutrinos. Journal of Mathematical Physics, 25(4):1088–1092, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p1088_s1.

Teitelboim:1984:EEG

[1866] Claudio Teitelboim. Explicit evaluation of group-invariant measure as by-product of path integration over Yang–Mills fields. Journal of Mathematical Physics, 25(4):1093–1101, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p1093_s1.

vonRieth:1984:HJT

[1867] J. von Rieth. The Hamilton–Jacobi theory of De Donder and Weyl applied to some relativistic field theories. Journal of Mathematical Physics, 25(4):1102–1115, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i4/p1102_s1.

Castrigiano:1984:CTW

[1868] D. P. L. Castrigiano. Causal transformations of Wightman localizations. Journal of Mathematical Physics, 25(4):1116–1121, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1116_s1.

Agudin:1984:BTE

[1869] J. L. Agudin and A. M. Platzeck. Bradyons and tachyons in an expansion of the spherical wave exp[i(ω/c)(r − ct)]/r into cylindrical waves. Journal of Mathematical Physics, 25(4):1122–1124, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1122_s1.

Hylton:1984:RDG

[1870] D. J. Hylton. The reduced Dirac Green function for the Coulomb potential. Journal of Mathematical Physics, 25(4):1125–1132, April 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1125_
s1.

Suzuki:1984:SCL

[1871] Noboru Suzuki. Simple calculation of Löwdin’s alpha-function. Journal of Mathematical Physics, 25(4):1133–1138, April 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1133_s1. See er-

ratum [2148].

Lewis:1984:EIQ

[1872] H. Ralph Lewis. Exact invariants quadratic in the momentum for a particle in a three-dimensional electromagnetic field. Journal of Mathematical Physics, 25(4):1139–1143, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i4/p1139_s1.

Fried:1984:GFZ

[1873] H. M. Fried and J. Tessendorf. Green’s functions at zero viscosity. Journal of Mathematical Physics, 25(4):1144–1154, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i4/p1144_s1.

Royer:1984:SPH

[1874] Antoine Royer. On some partitions of hypergraphs and cumulants having applications in statistical mechanics. Journal of Mathematical Physics, 25(5):1157–1175, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i5/p1157_s1.

Conte:1984:DPG

[1875] R. Conte, Jacques Raynal, and E. Soulié. Determination of point group harmonics for arbitrary j by a projection method. I. Cubic group, quantization along an axis of order 4. Journal of Mathematical Physics, 25(5): 1176–1186, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1176_s1.

Raynal:1984:DPGa

[1876] Jacques Raynal. Determination of point group harmonics for arbitrary j by a projection method. II. Icosahedral group, quantization along an axis of order 5. Journal of Mathematical Physics, 25(5):1187–1194, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1187_s1.

DeWilde:1984:CSR

[1877] P. De Wilde, H. De Meyer, J. Van der Jeugt, and G. Vanden Berghe. On a class of spinor representations of SO(7). Journal of Mathematical Physics, 25(5):1195–1198, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i5/p1195_s1.

Nikam:1984:MBR

[1878] R. S. Nikam and C. R. Sarma. Multispinor basis for representations of SO(N). Journal of Mathematical Physics, 25(5):1199–1204, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/
p1199_s1.

vandenBroek:1984:GRI

[1879] P. M. van den Broek. Group representations in indefinite metric spaces. Journal of Mathematical Physics, 25(5):1205–1210, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1205_s1.

Castanos:1984:AEM

[1880] O. Castaños, E. Chacón, and M. Moshinsky. Analytic expressions for the matrix elements of generators of Sp(6) in an Sp(6) ⊇ U(3) basis. Journal of Mathematical Physics, 25(5):1211–1218, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1211_s1.

VanderJeugt:1984:PCM

[1881] Joris Van der Jeugt. A pair of commuting missing label operators for G ⊇ [SU(2)]^{n}. Journal of Mathematical Physics, 25(5):1219–1221, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1219_s1.

Feldman:1984:SCL

[1882] G. Feldman, T. Fulton, and P. T. Matthews. Structure constants for Lie algebras. Journal of Mathematical Physics, 25(5):1222–1229, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1222_s1.

Feldman:1984:SSL

[1883] G. Feldman, T. Fulton, and P. T. Matthews. The semisimple subalgebras of Lie algebras. Journal of Mathematical Physics, 25(5):1230– 1252, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1230_s1.

Gruber:1984:UDR

[1884] Bruno Gruber, T. S. Santhanam, and Raj Wilson. A unified description of the representations of the graded Lie algebra Gsl(2). Journal of Mathematical Physics, 25(5):1253–1261, May 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1253_s1.

Rabsztyn:1984:CPC

[1885] Szymon Rabsztyn. On the Cauchy problem for the coupled Schrödinger– Klein–Gordon equations in one space dimension. Journal of Mathematical Physics, 25(5):1262–1265, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i5/p1262_s1.

Bakker:1984:SRP

[1886] M. Bakker and N. M. Temme. Sum rule for products of Bessel functions: comments on a paper: “New sum rule for products of Bessel functions with application to plasma physics”, [J. Math. Phys. 23 (1982), no. 7, 1278–1281; MR 83m:33017 ] by B. S. Newberger. Journal of Mathematical Physics, 25(5):1266–1267, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i5/p1266_s1. See [1089, 1608].

Chalbaud:1984:FAS

[1887] Eduardo Chalbaud and Pablo Martín. Fractional approximations for the spherically symmetric Coulomb scattering wave functions. Journal of Mathematical Physics, 25(5):1268–1273, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1268_s1.

Brittin:1984:POT

[1888] Wesley E. Brittin and Arthur Y. Sakakura. Projection operator techniques in physics. Journal of Mathematical Physics, 25(5):1274–1295, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1274_s1.

Nottale:1984:FNA

[1889] L. Nottale and J. Schneider. Fractals and nonstandard analysis. Journal of Mathematical Physics, 25(5):1296–1300, May 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1296_s1.

Fauth:1984:MCR

[1890] D. Fauth and M. Sorg. Mean curvature and radiation field in SO(2) electrodynamics. Journal of Mathematical Physics, 25(5):1301–1307, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1301_s1.

Kishida:1984:ERF

[1891] Kuniharu Kishida. Equivalent random force and time-series model in systems far from equilibrium. Journal of Mathematical Physics, 25(5): 1308–1313, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1308_s1.

Zambrini:1984:MPL

[1892] J.-C. Zambrini. Maupertuis’ principle of least action in stochastic calculus of variations. Journal of Mathematical Physics, 25(5):1314–1322, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1314_s1.

Johnson:1984:BCT

[1893] G. W. Johnson. A bounded convergence theorem for the Feynman integral. Journal of Mathematical Physics, 25(5):1323–1326, May 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1323_
s1.

Albeverio:1984:SIS

[1894] S. Albeverio, R. Høegh-Krohn, and M. Mebkhout. Scattering by impurities in a solvable model of a three-dimensional crystal. Journal of Mathematical Physics, 25(5):1327–1334, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1327_s1.

Cornille:1984:CNE

[1895] Henri Cornille. On a class of nonintegrable equations in 1 + 1 dimensions with factorized associated linear operators. Journal of Mathematical Physics, 25(5):1335–1347, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i5/p1335_s1.

Gomis:1984:LHC

[1896] J. Gomis, J. Llosa, and N. Román. Lee Hwa Chung theorem for presymplectic manifolds. Canonical transformations for constrained systems. Journal of Mathematical Physics, 25(5):1348–1355, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1348_s1.

Murmann:1984:DHM

[1897] Michael G. Mürmann. On the derivation of hydrodynamics from molecular dynamics. Journal of Mathematical Physics, 25(5):1356–1363, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1356_s1.

Au:1984:KSP

[1898] C. Au and P. C. W. Fung. A KdV soliton propagating with varying velocity. Journal of Mathematical Physics, 25(5):1364–1369, May 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1364_
s1.

Fung:1984:SNA

[1899] P. C. W. Fung and C. Au. A series of new analytical solutions to the nonlinear equation y_{t} + y_{xxx} − 6y^{2}y_{x} + 6λy_{x} = 0. I. Journal of Mathematical Physics, 25(5):1370–1371, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i5/p1370_s1.

Tagami:1984:ESS

[1900] Yukiko Tagami. Exact solution of some nonlinear evolution equations. Journal of Mathematical Physics, 25(5):1372–1377, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1372_s1.

Frisch:1984:BLH

[1901] Uriel Frisch and Jean-Luc Gautero. Backscattering and localization of high-frequency waves in a one-dimensional random medium. Journal of Mathematical Physics, 25(5):1378–1381, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1378_s1.

Weston:1984:MIPa

[1902] V. H. Weston. Multifrequency inverse problem for the reduced wave equation with sparse data. Journal of Mathematical Physics, 25(5):1382– 1390, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1382_s1.

Vasilopoulos:1984:LRT

[1903] Pagiotis Vasilopoulos and Carolyn M. Van Vliet. Linear response theory revisited. IV. applications. Journal of Mathematical Physics, 25(5):1391– 1403, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1391_s1. See erratum [2181].

Werner:1984:QHA

[1904] R. Werner. Quantum harmonic analysis on phase space. Journal of Mathematical Physics, 25(5):1404–1411, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1404_s1.

Meiman:1984:FII

[1905] N. N. Meĭman. Functional integrals as integrals on locally noncompact groups with generalized measures. Journal of Mathematical Physics, 25(5):1412–1433, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i5/p1412_s1. See erratum [2179].

Aerts:1984:CTP

[1906] Dirk Aerts. Construction of the tensor product for the lattices of properties of physical entities. Journal of Mathematical Physics, 25(5):1434– 1441, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1434_s1.

Chacon:1984:SDO

[1907] E. Chacón, O. Castaños, and A. Frank. A six-dimensional oscillator basis classified by O(6) ⊇ SO(2) × SU(3) ⊇ SO(3). Journal of Mathematical Physics, 25(5):1442–1448, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i5/p1442_s1.

Cheney:1984:TDS

[1908] Margaret Cheney. Two-dimensional scattering: The number of bound states from scattering data. Journal of Mathematical Physics, 25(5): 1449–1455, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1449_s1.

Roy:1984:SMM

[1909] S. R. Roy and Raj Bali. Some magnetohydrostatic models of cylindrical symmetry in general relativity. Journal of Mathematical Physics, 25(5): 1456–1459, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1456_s1.

Collins:1984:CSF

[1910] C. B. Collins and A. J. White. A class of shear-free perfect fluids in general relativity. II. Journal of Mathematical Physics, 25(5):1460–1472, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1460_s1.

Ashtekar:1984:PSG

[1911] Abhay Ashtekar and Gary T. Horowitz. Phase space of general relativity revisited: A canonical choice of time and simplification of the Hamiltonian. Journal of Mathematical Physics, 25(5):1473–1480, May 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1473_
s1.

Tikekar:1984:SCF

[1912] Ramesh Tikekar. Spherical charged fluid distributions in general relativity. Journal of Mathematical Physics, 25(5):1481–1483, May 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1481_
s1.

Embacher:1984:NCE

[1913] Franz Embacher. A new class of exact pp-wave solutions in simple supergravity. Journal of Mathematical Physics, 25(5):1484–1488, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1484_s1.

Eris:1984:SSP

[1914] Ahmet Eriş, Metin Gürses, and Atalay Karasu. Symmetric space property and an inverse scattering formulation of the SAS Einstein–Maxwell field equations. Journal of Mathematical Physics, 25(5):1489–1495, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1489_s1.

Dundarer:1984:GVP

[1915] Resit Dündarer, Feza Gürsey, and Chia Hsiung Tze. Generalized vector products, duality, and octonionic identities in D = 8 geometry. Journal of Mathematical Physics, 25(5):1496–1506, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1496_s1.

Beretta:1984:RBC

[1916] Gian Paolo Beretta. On the relation between classical and quantumthermodynamic entropy. Journal of Mathematical Physics, 25(5):1507– 1510, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1507_s1.

Daniel:1984:LTN

[1917] R. Daniel, G. C. Pomraning, and Edward W. Larsen. Linear transport in nonhomogeneous media. III. Journal of Mathematical Physics, 25(5):

1511–1518, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1511_s1.

Manoukian:1984:PDT

[1918] Edward B. Manoukian. Proof of the decoupling theorem of field theory in Minkowski space. Journal of Mathematical Physics, 25(5):1519– 1523, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1519_s1.

Burzlaff:1984:RBC

[1919] Jürgen Burzlaff and J. E. M. Hornos. On the relation between complex manifolds and soliton theoretic constructions for self-dual fields. Journal of Mathematical Physics, 25(5):1524–1527, May 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1524_s1.

Gutierrez:1984:CGS

[1920] W. R. Gutierrez and L. F. Saez. Composite gluons and scalar fields. Journal of Mathematical Physics, 25(5):1528–1530, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1528_s1.

Grigore:1984:LGI

[1921] D. R. Grigore. Local gauge invariant Lagrangians in classical field theories. Journal of Mathematical Physics, 25(5):1531–1538, May 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1531_
s1.

Magg:1984:SSC

[1922] M. Magg. Static solutions of the coupled Yang–Mills–Weyl equations. Journal of Mathematical Physics, 25(5):1539–1541, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1539_s1.

Legare:1984:RYM

[1923] M. Légaré and J. Harnad. SO(4) reduction of the Yang–Mills equations for the classical gauge groups. Journal of Mathematical Physics, 25(5): 1542–1547, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1542_s1.

Pennington:1984:LML

[1924] M. R. Pennington. Leading and misleading logs in perturbative QCD. Journal of Mathematical Physics, 25(5):1548–1554, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1548_s1.

Moshinsky:1984:CGG

[1925] Marcos Moshinsky. Collectivity and geometry. I. General approach. Journal of Mathematical Physics, 25(5):1555–1564, May 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1555_s1.

Chacon:1984:CGI

[1926] E. Chacón, P. Hess, and M. Moshinsky. Collectivity and geometry. II. The two-dimensional case. Journal of Mathematical Physics, 25(5):1565– 1576, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1565_s1.

Hill:1984:ASW

[1927] Robert Nyden Hill. On the analytic structure of the wave function for a hydrogen atom in an analytic potential. Journal of Mathematical Physics, 25(5):1577–1583, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i5/p1577_s1.

Jackson:1984:RRD

[1928] E. Atlee Jackson. Radiation reaction dynamics in an electromagnetic wave and constant electric field. Journal of Mathematical Physics, 25(5): 1584–1591, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1584_s1.

Nieto-Vesperinas:1984:RTD

[1929] M. Nieto-Vesperinas. On the representation of two-dimensional scalar wave fields in the complex plane. Journal of Mathematical Physics, 25(5): 1592–1598, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i5/p1592_s1.

Koiller:1984:MSW

[1930] Jair Koiller. A mechanical system with a “wild” horseshoe. Journal of Mathematical Physics, 25(5):1599–1604, May 1984. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i5/p1599_s1.

Maggs:1984:SFO

[1931] J. E. Maggs, Alfredo Baños, Jr., and G. J. Morales. Solutions of a fourth-order differential equation describing mode conversion in a magnetized plasma. Journal of Mathematical Physics, 25(5):1605–1618, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i5/p1605_s1.

Bernard:1984:ESD

[1932] Douglas A. Bernard and Jaan Noolandi. Exact solution for the discrete end-to-end intrachain memory function of a reptating polymer chain. Journal of Mathematical Physics, 25(5):1619–1630, May 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i5/p1619_s1.

Amundsen:1984:ESU

[1933] P. A. Amundsen and P. H. Damgaard. Exact solutions of unitarily invariant matrix models in zero dimensions. Journal of Mathematical Physics, 25(6):1633–1637, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1633_s1.

Deenen:1984:DGM

[1934] J. Deenen and C. Quesne. Dynamical group of microscopic collective states. III. Coherent state representations in d dimensions. Journal of Mathematical Physics, 25(6):1638–1650, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i6/p1638_s1.

Delduc:1984:MSS

[1935] F. Delduc and M. Gourdin. Mixed supertableaux of the superunitary groups. I. SU(n|1). Journal of Mathematical Physics, 25(6):1651–1661, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1651_s1.

Ek:1984:DVS

[1936] Bengt Ek and Bengt Nagel. Differentiable vectors and sharp momentum states of helicity representations of the Poincaré group. Journal of Mathematical Physics, 25(6):1662–1670, June 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1662_s1.

Gaeta:1984:GPB

[1937] Giuseppe Gaeta and Pietro Rossi. Gradient property of bifurcation equation for systems with rotational symmetry. Journal of Mathematical Physics, 25(6):1671–1673, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1671_s1.

Lorente:1984:CEV

[1938] Miguel Lorente and Bruno Gruber. Construction of extremal vectors for Verma submodules of Verma modules. Journal of Mathematical Physics, 25(6):1674–1681, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1674_s1.

Raszillier:1984:LDF

[1939] H. Raszillier. Lattice degeneracies of fermions. Journal of Mathematical Physics, 25(6):1682–1693, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1682_s1.

Kim:1984:OSL

[1940] Jai Sam Kim. Orbit spaces of low-dimensional representations of simple compact connected Lie groups and extrema of a group-invariant scalar potential. Journal of Mathematical Physics, 25(6):1694–1717, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/
p1694_s1.

Navelet:1984:HTG

[1941] H. Navelet. On the Hankel transform of a generalized Laguerre polynomial and on the convolution involving special Bessel functions. Journal of Mathematical Physics, 25(6):1718–1720, June 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1718_s1.

Blaive:1984:EEG

[1942] B. Blaive and J. Metzger. Explicit expressions of the nth gradient of 1/r. Journal of Mathematical Physics, 25(6):1721–1724, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1721_s1.

Boiti:1984:RGL

[1943] Marco Boiti, Jérôme J.-P. Leon, and Flora Pempinelli. A recursive generation of local higher-order sine-Gordon equations and their Bäcklund transformation. Journal of Mathematical Physics, 25(6):1725–1734, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1725_s1.

MartinezAlonso:1984:QSN

[1944] L. Martínez Alonso. Quantum solitons of the nonlinear Schrödinger field as Galilean particles. Journal of Mathematical Physics, 25(6):1735–1741, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1735_s1.

Grossmann:1984:FPH

[1945] Alexander Grossmann and Tai Tsun Wu. Fermi pseudopotential in higher dimensions. Journal of Mathematical Physics, 25(6):1742–1745, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1742_s1.

Ebrahim:1984:GNR

[1946] Abdul Ebrahim and Mustafa Savci. General nonlinear realization of chiral SU_{2} × SU_{2} symmetry in curved isospin space. Journal of Mathematical Physics, 25(6):1746–1751, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p1746_s1.

Brosens:1984:EEP

[1947] F. Brosens and J. T. Devreese. Explicit evaluation of a path integral with memory kernel. Journal of Mathematical Physics, 25(6):1752–1755, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1752_s1.

Phares:1984:OSI

[1948] Alain J. Phares. The occupation statistics for indistinguishable dumbbells on a rectangular lattice space. I. Journal of Mathematical Physics, 25(6):1756–1770, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i6/p1756_s1.

Yoshida:1984:WSQ

[1949] Zensho Yoshida, Nobuyuki Inoue, and Taijiro Uchida. Weak solutions of a quasistatic model of plasmas. Journal of Mathematical Physics, 25(6): 1771–1775, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1771_s1.

Hojman:1984:FOE

[1950] Sergio Hojman and Javier Gómez. First-order equivalent Lagrangians and conservation laws. Journal of Mathematical Physics, 25(6):1776– 1779, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1776_s1.

Mandal:1984:SOI

[1951] Birendranath Mandal and Sudip Kumar Goswami. The scattering of an obliquely incident surface wave by a submerged fixed vertical plate. Journal of Mathematical Physics, 25(6):1780–1783, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1780_s1.

Brownstein:1984:VPE

[1952] K. R. Brownstein. Variational principle for electromagnetic problems in a linear, static, inhomogeneous anisotropic medium. Journal of Mathematical Physics, 25(6):1784–1786, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p1784_s1.

Monroe:1984:LTD

[1953] David K. Monroe. Lorentz transformation of Debye potentials. Journal of Mathematical Physics, 25(6):1787–1790, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i6/p1787_s1.

Ramm:1984:DDN

[1954] A. G. Ramm. Description of the degree of nonuniqueness in inverse source problems. Journal of Mathematical Physics, 25(6):1791–1793, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1791_s1.

Busch:1984:JLB

[1955] Paul Busch. On joint lower bounds of position and momentum observables in quantum mechanics. Journal of Mathematical Physics, 25(6): 1794–1797, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1794_s1.

McKenna:1984:RCG

[1956] I. H. McKenna and K. K. Wan. The role of the connection in geometric quantization. Journal of Mathematical Physics, 25(6):1798–1803, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1798_s1.

Cheng:1984:PTD

[1957] Bin Kang Cheng. The propagator of the time-dependent forced harmonic oscillator with constant damping. Journal of Mathematical Physics, 25(6):1804–1809, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i6/p1804_s1.

Ichinose:1984:PDP

[1958] Takashi Ichinose and Hiroshi Tamura. Propagation of a Dirac particle. A path integral approach. Journal of Mathematical Physics, 25(6):1810– 1819, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1810_s1.

Gerry:1984:PIQ

[1959] Christopher C. Gerry. On the path integral quantization of the damped harmonic oscillator. Journal of Mathematical Physics, 25(6):1820–1822, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1820_s1.

Hamilton:1984:DEH

[1960] J. Dwayne Hamilton. The Dirac equation and Hestenes’ geometric algebra. Journal of Mathematical Physics, 25(6):1823–1832, June 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1823_
s1.

Hietarinta:1984:CVQ

[1961] Jarmo Hietarinta. Classical versus quantum integrability. Journal of Mathematical Physics, 25(6):1833–1840, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i6/p1833_s1.

Chandler:1984:BQS

[1962] Colston Chandler and A. G. Gibson. N-body quantum scattering theory in two Hilbert spaces. IV. Approximate equations. Journal of Mathematical Physics, 25(6):1841–1856, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p1841_s1.

Coen:1984:VDT

[1963] Shimon Coen, Margaret Cheney, and Arthur Weglein. Velocity and density of a two-dimensional acoustic medium from point source surface data. Journal of Mathematical Physics, 25(6):1857–1861, June 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1857_
s1.

Combes:1984:QSA

[1964] J. M. Combes and J. M. Ghez. The quantum scattering amplitude in R^{n} for potentials decreasing faster than |x|^{−(n+1)/2}. Journal of Mathematical Physics, 25(6):1862–1874, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1862_s1.

Kroger:1984:FRA

[1965] Helmut Kröger. Finite rank approximation of wave operators for shortrange and Coulomb interactions. Journal of Mathematical Physics, 25(6): 1875–1880, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1875_s1.

Muthuramalingam:1984:SPV

[1966] Pl. Muthuramalingam. Spectral properties of vaguely elliptic pseudodifferential operators with momentum-dependent long-range potentials using time-dependent scattering theory. Journal of Mathematical Physics, 25(6):1881–1899, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p1881_s1.

Bava:1984:ISO

[1967] G. P. Bava and G. Ghione. Inverse scattering for optical couplers. Exact solution of Marchenko equations. Journal of Mathematical Physics, 25(6):1900–1904, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i6/p1900_s1. See erratum [2180].

Moses:1984:TDI

[1968] Harry E. Moses. The time-dependent inverse source problem for the acoustic and electromagnetic equations in the one- and three-dimensional cases. Journal of Mathematical Physics, 25(6):1905–1923, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/
p1905_s1.

Prosser:1984:SGL

[1969] Reese T. Prosser. On the solutions of the Gel’fand–Levitan equation. Journal of Mathematical Physics, 25(6):1924–1929, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1924_s1.

Plebanski:1984:ONS

[1970] Jerzy F. Plebañski and Krzysztof Rózga. The optics of null strings. Journal of Mathematical Physics, 25(6):1930–1940, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1930_s1.

Robinson:1984:CNS

[1971] Ivor Robinson and Krzysztof Rózga. Congruences of null strings in complex space–times and some Cauchy–Kovalevski-like problems. Journal of Mathematical Physics, 25(6):1941–1946, June 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1941_s1.

Anderson:1984:ACS

[1972] James L. Anderson. Approximate causal solutions for a class of wave equations with backscatter. Journal of Mathematical Physics, 25(6): 1947–1950, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1947_s1.

GarciaDiaz:1984:ETS

[1973] Alberto García Díaz. Electrovac type D solutions with cosmological constant. Journal of Mathematical Physics, 25(6):1951–1954, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/

p1951_s1. See erratum [2213].

Debever:1984:EIS

[1974] R. Debever, N. Kamran, and R. G. McLenaghan. Exhaustive integration and a single expression for the general solution of the type D vacuum and electrovac field equations with cosmological constant for a nonsingular aligned Maxwell field. Journal of Mathematical Physics, 25(6):1955– 1972, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1955_s1.

Hojman:1984:ESP

[1975] Roberto Hojman and Jorge Santamarina. Exact solutions of plane symmetric cosmological models. Journal of Mathematical Physics, 25(6): 1973–1974, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1973_s1.

Sanyal:1984:CSE

[1976] Abhik Kumar Sanyal and Dipankar Ray. Cosmological solutions of the Einstein equation with heat flow. Journal of Mathematical Physics, 25(6): 1975–1976, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p1975_s1.

Sanyal:1984:SMF

[1977] Abhik Kumar Sanyal and Dipankar Ray. Scalar meson field in a conformally flat space. Journal of Mathematical Physics, 25(6):1977–1979, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p1977_s1.

TorresdelCastillo:1984:KVA

[1978] G. F. Torres del Castillo. Killing vectors in algebraically special space– times. Journal of Mathematical Physics, 25(6):1980–1984, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/
p1980_s1.

Bellezza:1984:CEG

[1979] V. Bellezza and V. Ferrari. Coupled electromagnetic and gravitational linearized perturbations of the Kerr–Newman metric. Journal of Mathematical Physics, 25(6):1985–1990, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p1985_s1.

Mansouri:1984:IDR

[1980] Freydoon Mansouri and Louis Witten. Isometries and dimensional reduction. Journal of Mathematical Physics, 25(6):1991–1994, June 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1991_
s1.

Mellein:1984:SCS

[1981] B. Mellein, A. Maltz, and E. E. Mola. Stochastic convergence of the saturation coverage of one-dimensional arrays of β-bell particles. Journal of Mathematical Physics, 25(6):1995–1999, June 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i6/p1995_s1.

Sen:1984:EFC

[1982] Siddhartha Sen. An explicit formula for the Casimir energy of an arbitrary compact smooth surface without boundary. Journal of Mathematical Physics, 25(6):2000–2006, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p2000_s1.

Nutku:1984:HFK

[1983] Y. Nutku. Hamiltonian formulation of the KdV equation. Journal of Mathematical Physics, 25(6):2007–2008, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i6/p2007_s1.

Arik:1984:GPI

[1984] M. Arik and F. Neyzi. Geometrical properties and invariances of the generalized sigma model. Journal of Mathematical Physics, 25(6):2009– 2012, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p2009_s1.

Gunther:1984:IMO

[1985] Marian Günther. Indefinite metric and other-than-Feynman propagators as tools in a relativistic two-body problem. Journal of Mathematical Physics, 25(6):2013–2024, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p2013_s1.

Fairlie:1984:ICF

[1986] D. B. Fairlie and J. Nuyts. Integration conditions for first-order differential linear equations in higher-dimensional gauge theories. Journal of Mathematical Physics, 25(6):2025–2027, June 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i6/p2025_s1.

Balantekin:1984:CEG

[1987] A. B. Balantekin. Character expansion for U(N) groups and U(N/M) supergroups. Journal of Mathematical Physics, 25(6):2028–2030, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p2028_s1.

Hasiewicz:1984:SCA

[1988] Z. Hasiewicz, A. K. Kwaśniewski, and P. Morawiec. Supersymmetry and Clifford algebras. Journal of Mathematical Physics, 25(6):2031– 2036, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p2031_s1.

Kim:1984:IES

[1989] Jaegu Kim. Irreducible SO(2) extended superfields. Journal of Mathematical Physics, 25(6):2037–2051, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i6/p2037_s1.

Stefanescu:1984:CAM

[1990] I. Sabba Stefanescu. On the construction of amplitudes with Mandelstam analyticity from observable quantities. Journal of Mathematical Physics, 25(6):2052–2086, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p2052_s1.

Roux:1984:EDH

[1991] Jean-François Roux. Exact diagonalization of the Hamiltonian of a Dirac particle with anomalous moment, interacting with an external magnetic field. Journal of Mathematical Physics, 25(6):2087–2089, June 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p2087_
s1.

Maia:1984:CSC

[1992] M. D. Maia. Combined symmetries in curved space–times. Journal of Mathematical Physics, 25(6):2090–2094, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i6/p2090_s1.

Atkinson:1984:NCQ

[1993] D. Atkinson, H. Boelens, S. J. Hiemstra, P. W. Johnson, W. J. Schoenmaker, and K. Stam. Nonperturbative confinement in quantum chromodynamics. III. Improved gluon propagator. Journal of Mathematical Physics, 25(6):2095–2098, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i6/p2095_s1.

Ruskai:1984:BLH

[1994] Mary Beth Ruskai and Frank H. Stillinger. Binding limit in the Hartree approximation. Journal of Mathematical Physics, 25(6):2099–2103, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i6/p2099_s1.

Nash:1984:LTC

[1995] Patrick L. Nash. A Lagrangian theory of the classical spinning electron. Journal of Mathematical Physics, 25(6):2104–2108, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p2104_s1.

Nieto-Vesperinas:1984:ISP

[1996] M. Nieto-Vesperinas. Inverse scattering problems: A study in terms of the zeros of entire functions. Journal of Mathematical Physics, 25(6): 2109–2115, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i6/p2109_s1.

Willis:1984:SRA

[1997] J. R. Willis. Some remarks on the application of the QCA to the determination of the overall elastic response of a matrix/inclusion composite. Journal of Mathematical Physics, 25(6):2116–2120, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i6/p2116_s1.

Kim:1984:UTPb

[1998] Shoon K. Kim. A unified theory of point groups. VI. The projective corepresentations of the magnetic point groups of infinite order. Journal of Mathematical Physics, 25(7):2125–2127, July 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2125_s1.

Patera:1984:SFR

[1999] J. Patera and R. T. Sharp. Signatures of finite SU(p, q) representations. Journal of Mathematical Physics, 25(7):2128–2131, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2128_s1.

Syed:1984:RME

[2000] Ansaruddin Syed. The representation matrix elements of the group O^{+}(2, 2). Journal of Mathematical Physics, 25(7):2132–2135, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/
p2132_s1.

Rzewuski:1984:SGC

[2001] Jan Rzewuski. Some geometrical consequences of physical symmetries. Journal of Mathematical Physics, 25(7):2136–2139, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2136_s1.

Amar:1984:PAE

[2002] V. Amar, U. Dozzio, and C. Oleari. Proof of an algorithm for the evaluation of the branching multiplicity SO(2n) → SO(2n − 2) ⊗ U(1). Journal of Mathematical Physics, 25(7):2140–2143, July 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2140_s1.

Gaskell:1984:FSF

[2003] R. W. Gaskell and R. T. Sharp. Fixed symmetry and fixed class generating functions. Journal of Mathematical Physics, 25(7):2144–2148, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i7/p2144_s1.

Winternitz:1984:CSR

[2004] P. Winternitz. Comments on superposition rules for nonlinear coupled first-order differential equations. Journal of Mathematical Physics, 25(7): 2149–2150, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2149_s1. See [1538].

Evangelidis:1984:SBP

[2005] E. A. Evangelidis. The summation of Bessel products. Journal of Mathematical Physics, 25(7):2151–2153, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2151_s1.

Gotay:1984:PRQ

[2006] Mark J. Gotay. Poisson reduction and quantization for the n+1 photon. Journal of Mathematical Physics, 25(7):2154–2159, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2154_s1.

Tikochinsky:1984:EIT

[2007] Y. Tikochinsky and R. D. Levine. Estimation of inverse temperature and other Lagrange multipliers: The dual distribution. Journal of Mathematical Physics, 25(7):2160–2168, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2160_s1.

Phares:1984:DSP

[2008] Alain J. Phares. Decoupling of a system of partial difference equations with constant coefficients and application. Journal of Mathematical Physics, 25(7):2169–2173, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2169_s1.

Matsuo:1984:PBP

[2009] Kuniaki Matsuo, Malvin Carl Teich, and Bahaa E. A. Saleh. Poisson branching point processes. Journal of Mathematical Physics, 25(7):2174– 2185, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2174_s1.

Mach:1984:OPE

[2010] R. Mach. Orthogonal polynomials with exponential weight in a finite interval and application to the optical model. Journal of Mathematical Physics, 25(7):2186–2193, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i7/p2186_s1.

Grammaticos:1984:TDT

[2011] B. Grammaticos and B. Dorizzi. Two-dimensional time-dependent Hamiltonian systems with an exact invariant. Journal of Mathematical Physics, 25(7):2194–2199, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2194_s1.

Dorizzi:1984:IMT

[2012] B. Dorizzi, B. Grammaticos, R. Padjen, and V. Papageorgiou. Integrals of motion for Toda systems with unequal masses. Journal of Mathematical Physics, 25(7):2200–2211, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2200_s1.

Popowicz:1984:LBT

[2013] Ziemowit Popowicz. The Liouville–Bäcklund transformation for the twodimensional SU(N) Toda lattice. Journal of Mathematical Physics, 25(7): 2212–2216, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2212_s1.

Mishra:1984:DIT

[2014] S. C. Mishra, R. S. Kaushal, and K. C. Tripathy. Dynamical invariants for two-dimensional time-dependent classical systems. Journal of Mathematical Physics, 25(7):2217–2221, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2217_s1.

Molino:1984:GPD

[2015] Pierre Molino. General prolongations and (x, t)-depending pseudopotentials for the KdV equation. Journal of Mathematical Physics, 25(7): 2222–2225, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2222_s1.

Weiss:1984:SGE

[2016] John Weiss. The sine-Gordon equations: Complete and partial integrability. Journal of Mathematical Physics, 25(7):2226–2235, July 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2226_
s1.

Matzner:1984:ENS

[2017] Richard A. Matzner and Michael P. Ryan, Jr. An exact nonghost solution for a plane-symmetric cosmology containing a classical spinor field. Journal of Mathematical Physics, 25(7):2236–2239, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2236_s1.

Janssen:1984:PPP

[2018] A. J. E. M. Janssen. Positivity properties of phase-plane distribution functions. Journal of Mathematical Physics, 25(7):2240–2252, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/
p2240_s1.

Vujicic:1984:QMT

[2019] Milan Vujičić and Fedor Herbut. A quantum-mechanical theory of distant correlations. Journal of Mathematical Physics, 25(7):2253–2259, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i7/p2253_s1.

Gisin:1984:PSP

[2020] N. Gisin. Propensities and the state-property structure of classical and quantum systems. Journal of Mathematical Physics, 25(7):2260–2265, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i7/p2260_s1.

Fry:1984:BTO

[2021] M. P. Fry. Bound on the Nth order term of the partition function of the massive Schwinger model. Journal of Mathematical Physics, 25(7): 2266–2275, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2266_s1.

Henneaux:1984:EFI

[2022] Marc Henneaux. Electromagnetic fields invariant up to a duality rotation under a group of isometries. Journal of Mathematical Physics, 25(7): 2276–2283, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2276_s1.

Guo:1984:NMC

[2023] Dong Sheng Guo. Noniterative method for constructing many-parameter solutions of the Einstein and Einstein–Maxwell field equations. Journal of Mathematical Physics, 25(7):2284–2292, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i7/p2284_s1.

Radhakrishna:1984:PRF

[2024] L. Radhakrishna and N. Ibohal Singh. Pure radiation fields admitting nontrivial null symmetries. Journal of Mathematical Physics, 25(7):2293– 2300, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2293_s1.

Hogan:1984:KKT

[2025] P. A. Hogan. Kaluza–Klein theory derived from a Riemannian submersion. Journal of Mathematical Physics, 25(7):2301–2305, July 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2301_
s1.

Case:1984:PCQ

[2026] K. M. Case. “polynomial constants” for the quantized NLS equation. Journal of Mathematical Physics, 25(7):2306–2314, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i7/p2306_s1.

McLennan:1984:CIW

[2027] James A. McLennan. Conformally invariant wave equations for massless particles. Journal of Mathematical Physics, 25(7):2315–2320, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i7/p2315_s1.

Urbantke:1984:IPY

[2028] H. Urbantke. On integrability properties of SU(2) Yang–Mills fields. I. Infinitesimal part. Journal of Mathematical Physics, 25(7):2321–2324, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i7/p2321_s1.

Hou:1984:SSI

[2029] Bo Yu Hou. Some series of infinitely many symmetry generators in symmetric space chiral models. Journal of Mathematical Physics, 25(7): 2325–2330, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2325_s1.

Steeb:1984:IND

[2030] W.-H. Steeb, W. Oevel, and W. Strampp. On the integrability of nonlinear Dirac equations. Journal of Mathematical Physics, 25(7):2331– 2335, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i7/p2331_s1.

Gilmore:1984:CPP

[2031] R. Gilmore. Critical properties of pseudospin Hamiltonians. Journal of Mathematical Physics, 25(7):2336–2338, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i7/p2336_s1.

Chatterjee:1984:CFE

[2032] Ibha Chatterjee. Crystal field effect in an Ising model. Journal of Mathematical Physics, 25(7):2339–2344, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i7/p2339_s1.

Froelich:1984:EMC

[2033] John Froelich and Nikos Salingaros. The exponential mapping in Clifford algebras. Journal of Mathematical Physics, 25(8):2347–2350, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2347_s1.

Wene:1984:CRC

[2034] G. P. Wene. A construction relating Clifford algebras and Cayley– Dickson algebras. Journal of Mathematical Physics, 25(8):2351–2353, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2351_s1.

Deenen:1984:PCS

[2035] J. Deenen and C. Quesne. Partially coherent states of the real symplectic group. Journal of Mathematical Physics, 25(8):2354–2366, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2354_s1.

Schmidt:1984:KPF

[2036] Jeffrey R. Schmidt and Adam M. Bincer. The Kostant partition function for simple Lie algebras. Journal of Mathematical Physics, 25(8):2367– 2373, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2367_s1.

Luning:1984:SFS

[2037] C. D. Luning and W. L. Perry. Shape functions for separable solutions to cross-field diffusion problems. Journal of Mathematical Physics, 25(8):2374–2376, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2374_s1.

Arteca:1984:NMS

[2038] Gustavo A. Arteca, Francisco M. Fernández, and Eduardo A. Castro. A new method for summation of divergent power series. Journal of Mathematical Physics, 25(8):2377–2389, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i8/p2377_s1.

Spencer:1984:PSE

[2039] Richard G. Spencer. Poisson structure of the equations of ideal multispecies fluid electrodynamics. Journal of Mathematical Physics, 25 (8):2390–2396, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2390_s1.

Gudder:1984:PM

[2040] Stanley P. Gudder. Probability manifolds. Journal of Mathematical Physics, 25(8):2397–2401, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i8/p2397_s1.

Cohen:1984:PDG

[2041] Leon Cohen. Probability distributions with given multivariate marginals. Journal of Mathematical Physics, 25(8):2402–2403, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2402_
s1.

Mead:1984:MEP

[2042] Lawrence R. Mead and N. Papanicolaou. Maximum entropy in the problem of moments. Journal of Mathematical Physics, 25(8):2404–2417, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2404_s1.

Bampi:1984:CBV

[2043] F. Bampi and A. Morro. The connection between variational principles in Eulerian and Lagrangian descriptions. Journal of Mathematical Physics, 25(8):2418–2421, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2418_s1.

Anastasselou:1984:GSB

[2044] E. G. Anastasselou and N. I. Ioakimidis. A generalization of the Siewert– Burniston method for the determination of zeros of analytic functions. Journal of Mathematical Physics, 25(8):2422–2425, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2422_
s1.

Batalin:1984:CAO

[2045] I. A. Batalin and E. S. Fradkin. Closing and abelizing operatorial gauge algebra generated by first class constraints. Journal of Mathematical Physics, 25(8):2426–2429, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i8/p2426_s1.

Carinena:1984:BRS

[2046] José F. Cariñena and Manuel F. Rañada. Blow-up regularization of singular Lagrangians. Journal of Mathematical Physics, 25(8):2430–2435, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2430_s1.

Codaccioni:1984:AOG

[2047] J. P. Codaccioni and R. Caboz. Anharmonic oscillators and generalized hypergeometric functions. Journal of Mathematical Physics, 25(8):2436– 2438, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2436_s1.

Dominici:1984:HJT

[2048] D. Dominici, G. Longhi, J. Gomis, and J. M. Pons. Hamilton–Jacobi theory for constrained systems. Journal of Mathematical Physics, 25 (8):2439–2452, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2439_s1.

Werth:1984:EPN

[2049] J.-E. Werth. The equivalence problem for nonconservative mechanics. Journal of Mathematical Physics, 25(8):2453–2455, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2453_
s1.

Bette:1984:PRM

[2050] Andreas Bette. On a pointlike relativistic massive and spinning particle. Journal of Mathematical Physics, 25(8):2456–2460, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2456_
s1.

Green:1984:EEE

[2051] A. E. Green and P. M. Naghdi. Electromagnetic effects in an elastic circular cylindrical waveguide. Journal of Mathematical Physics, 25(8):2461– 2466, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2461_s1.

Kaup:1984:SES

[2052] D. J. Kaup. The squared eigenstates of the sine-Gordon eigenvalue problem. Journal of Mathematical Physics, 25(8):2467–2471, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2467_s1.

Lochak:1984:ASS

[2053] Pierre Lochak. On the adiabatic stability of solitons and the matching of conservation laws. Journal of Mathematical Physics, 25(8):2472– 2476, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2472_s1.

Santamato:1984:SIK

[2054] E. Santamato. Statistical interpretation of the Klein–Gordon equation in terms of the space–time Weyl curvature. Journal of Mathematical Physics, 25(8):2477–2480, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v25/i8/p2477_s1.

Gadella:1984:RDR

[2055] M. Gadella. On the RHS description of resonances and virtual states. Journal of Mathematical Physics, 25(8):2481–2485, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2481_
s1.

Moretti:1984:MCF

[2056] P. Moretti and M. Mancini. A method of calculating the function of two noncommuting operators. Journal of Mathematical Physics, 25(8):2486– 2489, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2486_s1.

Hoffmann-Ostenhof:1984:EEO

[2057] Maria Hoffmann-Ostenhof and Thomas Hoffmann-Ostenhof. On embedded eigenvalues for the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 25(8):2490–2493, August 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2490_s1.

Fokas:1984:IST

[2058] A. S. Fokas and M. J. Ablowitz. On the inverse scattering transform of multidimensional nonlinear equations related to first-order systems in the plane. Journal of Mathematical Physics, 25(8):2494–2505, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2494_s1.

Winicour:1984:NIQ

[2059] Jeffrey Winicour. Null infinity from a quasi-Newtonian view. Journal of Mathematical Physics, 25(8):2506–2514, August 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2506_s1.

Krori:1984:SGI

[2060] K. D. Krori and D. Nandy. Self-gravitating isotropic and anisotropic fluid distributions. Journal of Mathematical Physics, 25(8):2515–2518, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2515_s1.

Wolf:1984:ESI

[2061] N. O. Wolf, J. W. Evans, and D. K. Hoffman. Exactly solvable irreversible processes on one-dimensional lattices. Journal of Mathematical Physics, 25(8):2519–2526, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2519_s1.

Evans:1984:ESI

[2062] J. W. Evans. Exactly solvable irreversible processes on Bethe lattices. Journal of Mathematical Physics, 25(8):2527–2532, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2527_
s1.

Glasser:1984:DEL

[2063] M. L. Glasser and N. Shawagfeh. Dirac equation for a linear potential. Journal of Mathematical Physics, 25(8):2533–2537, August 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i8/p2533_
s1.

Kolar:1984:VPA

[2064] P. Kolář and J. Fischer. On the validity and practical applicability of derivative analyticity relations. Journal of Mathematical Physics, 25(8):2538–2544, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i8/p2538_s1.

Lukierski:1984:SFE

[2065] J. Lukierski and A. Nowicki. On superfield formulation of Euclidean supersymmetry. Journal of Mathematical Physics, 25(8):2545–2549, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2545_s1.

Mitra:1984:CGL

[2066] Bani Mitra and K. C. Tripathy. The cohomology of the generalized Lie algebras. Journal of Mathematical Physics, 25(8):2550–2556, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2550_s1.

Chanda:1984:SES

[2067] Pranab Krishna Chanda, Dipankar Ray, and Utpal Kumar De. Some exact solutions of nonlinear chiral field equations. Journal of Mathematical Physics, 25(8):2557–2562, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i8/p2557_s1.

Dutt:1984:TBC

[2068] Ranabir Dutt and Y. P. Varshni. Tight bounds for the critical screening parameters of the generalized exponential cosine screened Coulomb potential. Journal of Mathematical Physics, 25(8):2563–2566, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2563_s1.

Hahne:1984:KCM

[2069] G. E. Hahne. Kinematics of the CS method for the treatment of molecular collisions. Journal of Mathematical Physics, 25(8):2567–2575, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i8/p2567_s1.

Fernandez:1984:LAN

[2070] Ariel Fernández and Oktay Sinanoğlu. Locally attractive normal modes for chemical process. Journal of Mathematical Physics, 25(8):2576– 2581, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2576_s1.

VandenBerghe:1984:TOR

[2071] G. Vanden Berghe, H. De Meyer, and J. Van der Jeugt. Tensor operator realizations of E_{6} and structural zeros of the 6j-symbol. Journal of Mathematical Physics, 25(9):2585–2588, September 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2585_s1.

Perez:1984:SGP

[2072] J. Fernando Perez, C. P. Malta, and F. A. B. Coutinho. On some general properties of the point spectrum of three particles moving in one dimension. Journal of Mathematical Physics, 25(9):2589–2592, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v25/i9/p2589_s1. See erratum [2217].

Rieckers:1984:CPM

[2073] Alfred Rieckers. On the classical part of the mean field dynamics for quantum lattice systems in grand canonical representations. Journal of Mathematical Physics, 25(9):2593–2601, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2593_s1.

Wong:1984:GTT

[2074] M. K. F. Wong and Hsin Yang Yeh. Group theoretic treatment of the Dirac–Coulomb equation and matrix elements of its tensor operators. Journal of Mathematical Physics, 25(9):2602–2613, September 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2602_
s1.

Santini:1984:DLC

[2075] P. M. Santini, M. J. Ablowitz, and A. S. Fokas. The direct linearization of a class of nonlinear evolution equations. Journal of Mathematical Physics, 25(9):2614–2619, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i9/p2614_s1.

Kalinowski:1984:ONM

[2076] M. W. Kalinowski. On the old-new method of solving nonlinear equations. Journal of Mathematical Physics, 25(9):2620–2632, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i9/p2620_s1.

Epifanio:1984:APC

[2077] Giuseppina Epifanio and Camillo Trapani. V^{∗}-algebras: A particular class of unbounded operator algebras. Journal of Mathematical Physics, 25(9):2633–2637, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i9/p2633_s1.

Dullemond:1984:CMM

[2078] C. Dullemond. On the center-of-mass motion of geometrically confined classical particles. Journal of Mathematical Physics, 25(9):2638–2647, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2638_s1.

Grinberg:1984:HCF

[2079] Horacio Grinberg, Julio Marañon, and Hector Vucetich. Homogeneous canonical formulation of the nonrelativistic hydrogen atom. Journal of Mathematical Physics, 25(9):2648–2650, September 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2648_s1.

Skorupski:1984:DPI

[2080] Andrzej A. Skorupski. Double phase-integral approximations: A systematic simplification technique for wave equations with cutoffs and resonances. Journal of Mathematical Physics, 25(9):2651–2654, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i9/p2651_s1.

Skorupski:1984:TTC

[2081] Andrzej A. Skorupski. Transmission through cutoffs and resonances in the double phase-integral approximation. Journal of Mathematical Physics, 25(9):2655–2661, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i9/p2655_s1.

Rowe:1984:CST

[2082] D. J. Rowe. Coherent state theory of the noncompact symplectic group. Journal of Mathematical Physics, 25(9):2662–2671, September 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2662_
s1.

Ramm:1984:RAI

[2083] A. G. Ramm. Remarks about inverse diffraction problem. Journal of Mathematical Physics, 25(9):2672–2674, September 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2672_s1.

Letelier:1984:CSS

[2084] P. S. Letelier. Cylindrically symmetric solitary wave solutions to the Einstein equations. Journal of Mathematical Physics, 25(9):2675–2681, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2675_s1.

Ashtekar:1984:DSI

[2085] Abhay Ashtekar and Anne Magnon. From i^{∘} to the 3 + 1 description of spatial infinity. Journal of Mathematical Physics, 25(9):2682–2690, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2682_s1.

Kunstatter:1984:FEE

[2086] G. Kunstatter. Finite energy electric monopoles in an extended theory of gravitation. Journal of Mathematical Physics, 25(9):2691–2695, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2691_s1.

Wu:1984:MFK

[2087] Yong Shi Wu and A. Zee. Massless fermions and Kaluza–Klein theory with torsion. Journal of Mathematical Physics, 25(9):2696–2703, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2696_s1.

Naudts:1984:ICI

[2088] J. Naudts. An infinite chain of inequalities for correlation functions of classical lattice systems. Journal of Mathematical Physics, 25(9):2704– 2707, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2704_s1.

Hall:1984:KPQ

[2089] Richard L. Hall. Kinetic potentials in quantum mechanics. Journal of Mathematical Physics, 25(9):2708–2715, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2708_s1.

Dacol:1984:SAS

[2090] D. K. Dacol and H. Rabitz. Sensitivity analysis of stochastic kinetic models. Journal of Mathematical Physics, 25(9):2716–2727, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i9/p2716_s1.

Takahashi:1984:GWE

[2091] Y. Takahashi. The Gauss and the Weingarten equations for extended objects and their spinorization. Journal of Mathematical Physics, 25 (9):2728–2735, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i9/p2728_s1.

Actor:1984:TBC

[2092] Alfred Actor. Twisted boundary conditions for gauge theories on a torus. Journal of Mathematical Physics, 25(9):2736–2740, September 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2736_
s1.

OCarroll:1984:APC

[2093] Michael O’Carroll and Gastão de Almeida Braga. Analyticity properties and a convergent expansion for the glueball mass and dispersion curve of strongly coupled Euclidean 2+1 lattice gauge theories. Journal of Mathematical Physics, 25(9):2741–2743, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i9/p2741_s1.

Kostelecky:1984:SAL

[2094] V. Alan Kostelecký and Jeffrey M. Rabin. A superspace approach to lattice supersymmetry. Journal of Mathematical Physics, 25(9):2744– 2748, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i9/p2744_s1.

Zivkovic:1984:SPC

[2095] Tomislav P. Živković. Symmetry properties of the configuration interaction space in relation to one- and two-particle operators: The splitting theorem. Journal of Mathematical Physics, 25(9):2749–2766, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i9/p2749_s1.

Szu:1984:IPI

[2096] Harold H. Szu. Inverse problem of image processing. Journal of Mathematical Physics, 25(9):2767–2772, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i9/p2767_s1.

Szu:1984:DPI

[2097] Harold H. Szu. Direct problem of image processing. Journal of Mathematical Physics, 25(9):2773–2779, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i9/p2773_s1.

Mavroyannis:1984:RFS

[2098] Constantine Mavroyannis. Resonance fluorescence spectra of a strong bichromatic field interacting with a three-level atom in the “V ” configuration. Journal of Mathematical Physics, 25(9):2780–2787, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i9/p2780_s1.

Abi-Ghanem:1984:ERS

[2099] Georges Abi-Ghanem and V. V. Nguyen. Extension of Ruelle’s stability criteria to continuous systems. Journal of Mathematical Physics, 25 (9):2788–2790, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i9/p2788_s1.

McIlroy:1984:EBC

[2100] Douglas K. McIlroy and David P. Mason. Effect of the boundary condition at distance of closest approach on Wien dissociation of a weak electrolyte: A singular perturbation solution. Journal of Mathematical Physics, 25(9):2791–2799, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i9/p2791_s1.

Singh:1984:PFM

[2101] T. Singh and Tarkeshwar Singh. Perfect fluid models of Bianchi type-I in modified Brans–Dicke cosmology. Journal of Mathematical Physics, 25(9):2800–2802, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i9/p2800_s1.

Corbett:1984:EST

[2102] J. V. Corbett. Erratum: “Scattering theory for the dilation group. I. Simple quantum mechanical scattering”, [J. Math. Phys. 24, 1797 (1983)]. Journal of Mathematical Physics, 25(9):2803, September 1984. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i9/p2803_

s1. See [1547].

Madivanane:1984:NGI

[2103] S. Madivanane. A note on general involutional transformations. Journal of Mathematical Physics, 25(10):2807, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i10/p2807_s1.

Bickerstaff:1984:SRS

[2104] R. P. Bickerstaff. Simultaneity and reality of U(n) and SU(n) 3jm and 6j symbols. Journal of Mathematical Physics, 25(10):2808–2814, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2808_s1.

Castanos:1984:CGI

[2105] O. Castaños, E. Chacón, and M. Moshinsky. Collectivity and geometry. III. the three-dimensional case in the Sp(6) ⊇ Sp(2) × O(3) chain for closed shells. Journal of Mathematical Physics, 25(10):2815–2825, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2815_s1.

Fujiwara:1984:ADG

[2106] Y. Fujiwara, Y. C. Tang, and H. Horiuchi. Application of double Gel’fand polynomials to the symmetric group and spin-isospin wave functions of cluster systems. Journal of Mathematical Physics, 25(10):2826–2837, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2826_s1.

Moody:1984:GCC

[2107] R. V. Moody and J. Patera. General charge conjugation operators in simple Lie groups. Journal of Mathematical Physics, 25(10):2838–2847, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2838_s1.

Gould:1984:EAA

[2108] M. D. Gould and S. A. Edwards. Enveloping algebra annihilators and projection techniques for finite-dimensional cyclic modules of a semisimple Lie algebra. Journal of Mathematical Physics, 25(10):2848–2855, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2848_s1.

Iosifescu:1984:PBR

[2109] Mircea Iosifescu and Horia Scutaru. Poisson bracket realizations of Lie algebras and subrepresentations of (ad^{⊗k})_{s}. Journal of Mathematical Physics, 25(10):2856–2862, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p2856_s1.

Koh:1984:CGC

[2110] In Guy Koh, J. Patera, and C. Rousseau. Clebsch–Gordan coefficients for E_{6} and SO(10) unification models. Journal of Mathematical Physics, 25(10):2863–2872, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p2863_s1.

Royer:1984:FSR

[2111] Antoine Royer. On the Fourier series representations of path integrals. Journal of Mathematical Physics, 25(10):2873–2884, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i10/
p2873_s1.

Antone:1984:ETV

[2112] T. A. Antone. The expression for the triple vector product solid harmonic. Journal of Mathematical Physics, 25(10):2885–2887, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2885_s1.

Rubin:1984:EDD

[2113] Mark A. Rubin and Carlos R. Ordóñez. Eigenvalues and degeneracies for n-dimensional tensor spherical harmonics. Journal of Mathematical Physics, 25(10):2888–2894, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p2888_s1.

Connor:1984:DEC

[2114] J. N. L. Connor and P. R. Curtis. Differential equations for the cuspoid canonical integrals. Journal of Mathematical Physics, 25(10):2895– 2902, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p2895_s1.

Keller:1984:DTC

[2115] Joseph B. Keller. Discriminant, transmission, coefficient, and stability bands of Hill’s equation. Journal of Mathematical Physics, 25(10):2903– 2904, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p2903_s1.

Strampp:1984:LPS

[2116] Walter Strampp. Lax-pairs, spectral problems, and recursion operators. Journal of Mathematical Physics, 25(10):2905–2909, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i10/
p2905_s1.

Fassari:1984:SOP

[2117] S. Fassari. On the Schrödinger operator with periodic point interactions in the three-dimensional case. Journal of Mathematical Physics, 25(10):2910–2917, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p2910_s1.

CattaneoGasparini:1984:GRD

[2118] Ida Cattaneo Gasparini. Global reduction of a dynamical system on a foliated manifold. Journal of Mathematical Physics, 25(10):2918–2921, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2918_s1.

Williams:1984:HIL

[2119] D. Williams and J. F. Cornwell. The Haar integral for Lie supergroups. Journal of Mathematical Physics, 25(10):2922–2932, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2922_s1.

Glasser:1984:NCB

[2120] M. L. Glasser. A novel class of Bessel function integrals. Journal of Mathematical Physics, 25(10):2933–2934, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v25/i10/p2933_s1. See erratum

[2219].

Argueso:1984:PNE

[2121] F. Argüeso and J. L. Sanz. Post-Newtonian extensions of the Runge– Lenz vector. Journal of Mathematical Physics, 25(10):2935–2938, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2935_s1.

Deliyannis:1984:GMQ

[2122] Platon C. Deliyannis. Geometrical models for quantum logics with conditioning. Journal of Mathematical Physics, 25(10):2939–2946, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2939_s1.

Sharma:1984:AOG

[2123] G. S. Sharma and L. K. Sharma. Anharmonic oscillator with general polynomial potential. Journal of Mathematical Physics, 25(10):2947– 2952, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p2947_s1.

VanderJeugt:1984:SOT

[2124] J. Van der Jeugt and P. De Wilde. The shift operator technique for SO(7) in an [SU(2)]^{3} basis. I. Theory. Journal of Mathematical Physics, 25(10):2953–2957, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p2953_s1.

DeWilde:1984:SOT

[2125] P. De Wilde and J. Van der Jeugt. The shift operator technique for SO(7) in an [SU(2)]^{3} basis. II. Applications. Journal of Mathematical Physics, 25(10):2958–2965, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p2958_s1.

Tang:1984:FRB

[2126] Dun-Sang Tang. Formal relations between classical superalgebras and fermion-boson creation and annihilation operators. Journal of Mathematical Physics, 25(10):2966–2973, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i10/p2966_s1.

Leach:1984:ESS

[2127] P. G. L. Leach. Exact solutions of the Schrödinger equation for a class of three-dimensional isotropic anharmonic oscillators. Journal of Mathematical Physics, 25(10):2974–2978, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i10/p2974_s1.

Znojil:1984:CMT

[2128] M. Znojil. Chain models and their effective Hamiltonians. Journal of Mathematical Physics, 25(10):2979–2987, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i10/p2979_s1.

Cheney:1984:RDM

[2129] Margaret Cheney. A rigorous derivation of the “miracle” identity of three-dimensional inverse scattering. Journal of Mathematical Physics, 25(10):2988–2990, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p2988_s1.

Newton:1984:RIS

[2130] Roger G. Newton. Remarks on inverse scattering in one dimension. Journal of Mathematical Physics, 25(10):2991–2994, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i10/p2991_s1.

Rose:1984:CBT

[2131] James H. Rose, Margaret Cheney, and Brian DeFacio. The connection between time- and frequency-domain three-dimensional inverse scattering methods. Journal of Mathematical Physics, 25(10):2995–3000, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p2995_s1.

vanHaeringen:1984:TFO

[2132] H. van Haeringen. Twenty-four optimal inequalities and several new representations for the Coulomb T matrices in momentum space. Journal of Mathematical Physics, 25(10):3001–3032, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i10/p3001_s1.

vanHaeringen:1984:NIC

[2133] H. van Haeringen and L. P. Kok. New inequalities for the Coulomb T matrix in momentum space. Journal of Mathematical Physics, 25(10):3033– 3038, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p3033_s1. See erratum [2220].

Beck:1984:KLR

[2134] Johanna Friedl Beck and Akira Inomata. Kruskal-like representation of the de Sitter metric. Journal of Mathematical Physics, 25(10):3039– 3040, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p3039_s1.

Banerjee:1984:SSS

[2135] A. Banerjee, S. B. Dutta Choudhury, and N. Banerjee. Static spherically symmetric solutions in a general class of scalar tensor theories. Journal of Mathematical Physics, 25(10):3041–3046, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i10/p3041_s1.

Maia:1984:AHD

[2136] Marcos D. Maia and W. Mecklenburg. Aspects of high-dimensional theories in embedding spaces. Journal of Mathematical Physics, 25(10):3047– 3050, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p3047_s1.

Evans:1984:IRC

[2137] J. W. Evans, D. R. Burgess, and D. K. Hoffman. Irreversible random and cooperative processes on lattices: Spatial correlations. Journal of Mathematical Physics, 25(10):3051–3063, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i10/p3051_s1.

Bruning:1984:NCW

[2138] Erwin Brüning. A new class of Wick powers of generalized free fields. Journal of Mathematical Physics, 25(10):3064–3075, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i10/
p3064_s1.

Matsumoto:1984:ECQ

[2139] H. Matsumoto, Y. Nakano, and H. Umezawa. An equivalence class of quantum field theories at finite temperature. Journal of Mathematical Physics, 25(10):3076–3085, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p3076_s1.

Schuch:1984:NST

[2140] D. Schuch, K.-M. Chung, and H. Hartmann. Nonlinear Schrödinger-type field equation for the description of dissipative systems. III. Frictionally damped free motion as an example for an aperiodic motion. Journal of Mathematical Physics, 25(10):3086–3092, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i10/p3086_s1.

Hudson:1984:HFS

[2141] L. B. Hudson and R. Kantowski. Higgs fields from symmetric connections — the bundle picture. Journal of Mathematical Physics, 25(10):3093– 3104, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p3093_s1.

Saito:1984:ABM

[2142] Yoshimi Saitō. An asymptotic behavior of the S-matrix and the inverse scattering problem. Journal of Mathematical Physics, 25(10):3105– 3111, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i10/p3105_s1.

Cade:1984:PSD

[2143] R. Cade. Plane-surface deformation in electrohydrostatics with electric double layer. Journal of Mathematical Physics, 25(10):3112–3119, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i10/p3112_s1.

Newman:1984:LFE

[2144] William I. Newman. A Lyapunov functional for the evolution of solutions to the porous medium equation to self-similarity. I. Journal of Mathematical Physics, 25(10):3120–3123, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v25/i10/p3120_s1.

Ralston:1984:LFE

[2145] James Ralston. A Lyapunov functional for the evolution of solutions to the porous medium equation to self-similarity. II. Journal of Mathematical Physics, 25(10):3124–3127, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p3124_s1.

Thomas:1984:LTE

[2146] Lawrence E. Thomas and Zhong Yin. Low temperature expansions for the Gibbs states of quantum Ising lattice systems. Journal of Mathematical Physics, 25(10):3128–3134, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p3128_s1.

Nishimori:1984:ESM

[2147] H. Nishimori and R. B. Griffiths. Erratum: “Structure and motion of the Lee–Yang zeros”, [J. Math. Phys. 24, 2637 (1983)]. Journal of Mathematical Physics, 25(10):3135, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i10/p3135_s1. See [1666].

Suzuki:1984:ESC

[2148] Noboru Suzuki. Erratum: “Simple calculation of Löwdin’s alphafunction”, [J. Math. Phys. 25, 1133 (1984)]. Journal of Mathematical Physics, 25(10):3135, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i10/p3135_s2. See [1871].

Fujimoto:1984:RFH

[2149] Yasushi Fujimoto and Robert Grigjanis. Reduction formulas for higherorder indices and anomaly. Journal of Mathematical Physics, 25(11): 3141–3147, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3141_s1.

Jaric:1984:SGT

[2150] Marko V. Jarić and Marjorie Senechal. Space groups and their isotropy subgroups. Journal of Mathematical Physics, 25(11):3148–3154, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3148_s1.

Shnider:1984:CSN

[2151] S. Shnider and P. Winternitz. Classification of systems of nonlinear ordinary differential equations with superposition principles. Journal of Mathematical Physics, 25(11):3155–3165, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/p3155_s1.

Syed:1984:NSE

[2152] Ansaruddin Syed. A new set of Euler angles for the generalized Lorentz group. Journal of Mathematical Physics, 25(11):3166–3170, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3166_s1.

Baake:1984:SRH

[2153] M. Baake. Structure and representations of the hyperoctahedral group. Journal of Mathematical Physics, 25(11):3171–3182, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/
p3171_s1.

Trostel:1984:CAV

[2154] Robert Trostel. Color analysis, variational self-adjointness, and color Poisson (super)algebras. Journal of Mathematical Physics, 25(11):3183– 3189, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3183_s1.

Barakat:1984:NCU

[2155] Richard Barakat and Garry Newsam. Necessary conditions for a unique solution to two-dimensional phase recovery. Journal of Mathematical Physics, 25(11):3190–3193, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i11/p3190_s1.

Ciulli:1984:BCP

[2156] M. Ciulli, S. Ciulli, and T. D. Spearman. Bounds for the continuation of perturbative results to the spectral region. Journal of Mathematical Physics, 25(11):3194–3203, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i11/p3194_s1.

Shabani:1984:CFO

[2157] J. Shabani. Commutants of a family of operators on a partial inner product space. Journal of Mathematical Physics, 25(11):3204–3208, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3204_s1.

Castellani:1984:HGS

[2158] L. Castellani, R. D’Auria, P. Fré, and P. van Nieuwenhuizen. Holonomy groups, sesquidual torsion fields, and SU(8) in d = 11 supergravity. Journal of Mathematical Physics, 25(11):3209–3213, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/
p3209_s1.

Saclioglu:1984:LPE

[2159] Cihan K. Saçlioğlu. Liouville and Painlevé equations and Yang–Mills strings. Journal of Mathematical Physics, 25(11):3214–3220, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3214_s1.

Leach:1984:FIP

[2160] P. G. L. Leach. First integrals via polynomial canonical transformations. Journal of Mathematical Physics, 25(11):3221–3226, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3221_s1.

Aiyer:1984:FMR

[2161] Raju N. Aiyer. Factorization of the 2×2 matrix recursion operator of the coupled KdV equation. Journal of Mathematical Physics, 25(11):3227– 3230, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3227_s1.

Ramm:1984:ISG

[2162] A. G. Ramm and Arthur B. Weglein. Inverse scattering for geophysical problems. II. Inversion of acoustical data. Journal of Mathematical Physics, 25(11):3231–3234, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i11/p3231_s1.

Eringen:1984:EMD

[2163] A. Cemal Eringen. Electrodynamics of memory-dependent nonlocal elastic continua. Journal of Mathematical Physics, 25(11):3235–3249, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3235_s1.

Datta:1984:PRA

[2164] D. P. Datta and S. Mukherjee. Proof of Regge analyticity for power law potentials. Journal of Mathematical Physics, 25(11):3250–3251, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3250_s1.

Paul:1984:FAH

[2165] Thierry Paul. Functions analytic on the half-plane as quantum mechanical states. Journal of Mathematical Physics, 25(11):3252–3263, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3252_s1.

Rezende:1984:QST

[2166] Jorge Rezende. Quantum systems with time-dependent harmonic part and the Morse index. Journal of Mathematical Physics, 25(11):3264– 3269, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3264_s1.

Barab:1984:NAF

[2167] Jacqueline E. Barab. Nonexistence of asymptotically free solutions for a nonlinear Schrödinger equation. Journal of Mathematical Physics, 25 (11):3270–3273, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3270_s1.

Herrera:1984:AFC

[2168] L. Herrera, J. Jiménez, L. Leal, J. Ponce de León, M. Esculpi, and V. Galina. Anisotropic fluids and conformal motions in general relativity. Journal of Mathematical Physics, 25(11):3274–3278, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3274_s1.

Weber:1984:KSC

[2169] Edmond Weber. Kantowski–Sachs cosmological models approaching isotropy. Journal of Mathematical Physics, 25(11):3279–3285, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3279_s1.

Kandrup:1984:SMG

[2170] Henry E. Kandrup. Statistical mechanics of the gravitational field in a conformally static setting. Journal of Mathematical Physics, 25 (11):3286–3296, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3286_s1.

Wesson:1984:CSE

[2171] Paul S. Wesson. A cosmological solution of Einstein’s equations. Journal of Mathematical Physics, 25(11):3297–3298, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/p3297_s1.

Banerjee:1984:SST

[2172] N. Banerjee, S. B. Dutta Choudhury, and A. Banerjee. Static space– time in general scalar tensor theory. Journal of Mathematical Physics, 25(11):3299–3302, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3299_s1.

Peruggi:1984:PMHa

[2173] Fulvio Peruggi. Probability measures and Hamiltonian models on Bethe lattices. I. Properties and construction of MRT probability measures. Journal of Mathematical Physics, 25(11):3303–3315, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/
p3303_s1.

Peruggi:1984:PMHb

[2174] Fulvio Peruggi. Probability measures and Hamiltonian models on Bethe lattices. II. The solution of thermal and configurational problems. Journal of Mathematical Physics, 25(11):3316–3323, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3316_s1.

Sharpe:1984:GCF

[2175] Bruce Sharpe. Gribov copies and the Faddeev–Popov formula in lattice gauge theories. Journal of Mathematical Physics, 25(11):3324–3330, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3324_s1.

Leveille:1984:CIE

[2176] M. Léveillé. Casimir invariants for the eight-dimensional subgroups of the Poincaré group P(1, 4). Journal of Mathematical Physics, 25(11):3331– 3333, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i11/p3331_s1.

VanderJeugt:1984:FID

[2177] Joris Van der Jeugt. Finite- and infinite-dimensional representations of the orthosymplectic superalgebra OSP(3, 2). Journal of Mathematical Physics, 25(11):3334–3349, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i11/p3334_s1.

Gervois:1984:SII

[2178] A. Gervois and H. Navelet. Some integrals involving three Bessel functions when their arguments satisfy the triangle inequalities. Journal of Mathematical Physics, 25(11):3350–3356, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/p3350_s1.

Meiman:1984:EFI

[2179] N. N. Meĭman. Erratum: “Functional integrals as integrals on locally noncompact groups with generalized measures”, [J. Math. Phys. 25, 1412 (1984)]. Journal of Mathematical Physics, 25(11):3357, November 1984.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/

p3357_s1. See [1905].

Bava:1984:EIS

[2180] G. P. Bava and G. Ghione. Erratum: “Inverse scattering for optical couplers. Exact solution of Marchenko equations”, [J. Math. Phys. 25, 1900 (1984)]. Journal of Mathematical Physics, 25(11):3358, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i11/p3358_s1. See [1967].

Vasilopoulos:1984:ELR

[2181] Pagiotis Vasilopoulos and Carolyn M. Van Vliet. Erratum: “Linear response theory revisited. IV. Applications”, [J. Math Phys. 25, 1391 (1984)]. Journal of Mathematical Physics, 25(11):3358, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i11/

p3358_s2. See [1903].

Jaric:1984:PPG

[2182] Marko V. Jarić. Properties of Poincaré generating functions for polynomial covariants (tensors). Journal of Mathematical Physics, 25(12):3363– 3366, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3363_s1.

Kobayashi:1984:LGL

[2183] Yuji Kobayashi and Shigeaki Nagamachi. Lie groups and Lie algebras with generalized supersymmetric parameters. Journal of Mathematical Physics, 25(12):3367–3374, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3367_s1.

Raynal:1984:DPGb

[2184] Jacques Raynal and Robert Conte. Determination of point group harmonics for arbitrary j by a projection method. III. Cubic group, quantization along a ternary axis. Journal of Mathematical Physics, 25(12):3375– 3381, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3375_s1.

Glass:1984:NPR

[2185] E. N. Glass and D. C. Robinson. A nilpotent prolongation of the Robinson–Trautman equation. Journal of Mathematical Physics, 25 (12):3382–3386, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3382_s1.

Mielnik:1984:FMN

[2186] Bogdan Mielnik. Factorization method and new potentials with the oscillator spectrum. Journal of Mathematical Physics, 25(12):3387–3389, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3387_s1.

Boyd:1984:DCW

[2187] John P. Boyd. The double cnoidal wave of the Korteweg–de Vries equation: An overview. Journal of Mathematical Physics, 25(12):3390–3401, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3390_s1.

Boyd:1984:PSD

[2188] John P. Boyd. Perturbation series for the double cnoidal wave of the Korteweg–de Vries equation. Journal of Mathematical Physics, 25 (12):3402–3414, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3402_s1.

Boyd:1984:SMT

[2189] John P. Boyd. The special modular transformation for polycnoidal waves of the Korteweg–de Vries equation. Journal of Mathematical Physics, 25 (12):3415–3423, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3415_s1.

David:1984:ECT

[2190] Daniel David. On an extension of the classical Thirring model. Journal of Mathematical Physics, 25(12):3424–3432, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i12/p3424_s1.

Kundu:1984:LLH

[2191] A. Kundu. Landau–Lifshitz and higher-order nonlinear systems gauge generated from nonlinear Schrödinger-type equations. Journal of Mathematical Physics, 25(12):3433–3438, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i12/p3433_s1.

Grecchi:1984:GLB

[2192] V. Grecchi and M. Maioli. Generalized logarithmic Borel summability. Journal of Mathematical Physics, 25(12):3439–3443, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i12/
p3439_s1.

Spiga:1984:SCN

[2193] G. Spiga, R. L. Bowden, and V. C. Boffi. On the solutions to a class of nonlinear integral equations arising in transport theory. Journal of Mathematical Physics, 25(12):3444–3450, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i12/p3444_s1.

Houard:1984:TTS

[2194] J. C. Houard and M. Irac-Astaud. Two theorems on star diagrams. Journal of Mathematical Physics, 25(12):3451–3454, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v25/i12/p3451_s1.

Pilch:1984:FEL

[2195] K. Pilch and A. N. Schellekens. Formulas for the eigenvalues of the Laplacian on tensor harmonics on symmetric coset spaces. Journal of Mathematical Physics, 25(12):3455–3459, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i12/p3455_s1.

Cieciura:1984:CCS

[2196] Grzegorz Cieciura and Alfred Grundland. A certain class of solutions of the nonlinear wave equation. Journal of Mathematical Physics, 25 (12):3460–3469, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3460_s1.

Grammaticos:1984:HHO

[2197] B. Grammaticos, B. Dorizzi, and A. Ramani. Hamiltonians with highorder integrals and the “weak-Painlevé” concept. Journal of Mathematical Physics, 25(12):3470–3473, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3470_s1.

Thompson:1984:PCM

[2198] Gerard Thompson. Polynomial constants of motion in flat space. Journal of Mathematical Physics, 25(12):3474–3478, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v25/i12/

p3474_s1. See [2221, 2225].

Obata:1984:VKK

[2199] Tsunehiro Obata and Jiro Chiba. A viewpoint of Kaluza–Klein type in elasticity theory. Journal of Mathematical Physics, 25(12):3479–3482, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3479_s1.

Weston:1984:MIPb

[2200] V. H. Weston. Multifrequency inverse problem for the reduced wave equation: Resolution cell and stability. Journal of Mathematical Physics, 25(12):3483–3488, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3483_s1.

Battaglia:1984:RTN

[2201] Franco Battaglia and Thomas F. George. A rule for the total number of topologically distinct Feynman diagrams. Journal of Mathematical Physics, 25(12):3489–3491, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3489_s1. See erratum [2216].

Arteca:1984:SSD

[2202] Gustavo A. Arteca, Francisco M. Fernández, and Eduardo A. Castro. Summation of strongly divergent perturbation series. Journal of Mathematical Physics, 25(12):3492–3496, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i12/p3492_s1.

Ulehla:1984:BSE

[2203] I. Úlehla and D. Adamová. Bound state energy eigenvalues for a general class of one-dimensional problems on the whole axis (−∞, ∞) via the Prüfer transformation. Journal of Mathematical Physics, 25(12):3497– 3502, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3497_s1.

Lopuszanski:1984:CMP

[2204] Jan T. Łopuszański. On charges of massless particles. Journal of Mathematical Physics, 25(12):3503–3509, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v25/i12/p3503_s1.

Gupta:1984:SGF

[2205] Y. K. Gupta, S. P. Sharma, and R. S. Gupta. Self-gravitating fluids of class one with nonvanishing Weyl tensor. Journal of Mathematical Physics, 25(12):3510–3512, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3510_s1.

Coleman:1984:CNI

[2206] Robert Alan Coleman and Herbert Korte. Constraints on the nature of inertial motion arising from the universality of free fall and the conformal causal structure of space–time. Journal of Mathematical Physics, 25 (12):3513–3526, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i12/p3513_s1.

Hobill:1984:ASG

[2207] David W. Hobill. Axially symmetric gravitational radiation from isolated sources. Journal of Mathematical Physics, 25(12):3527–3537, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v25/i12/p3527_s1.

Zhong:1984:LGS

[2208] Zai Zhe Zhong. On the local GL(4, R) gauge symmetry of hyperbolic complex metrics. Journal of Mathematical Physics, 25(12):3538–3539, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3538_s1.

Reid:1984:GFC

[2209] J. H. Reid and N. N. Trofimenkoff. A generating function for Chew– Mandelstam functions. Journal of Mathematical Physics, 25(12):3540– 3544, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3540_s1.

Ogura:1984:BNS

[2210] H. Ogura and D. J. Rowe. An Sp(3, R) × O(N − 1) basis for the nuclear shell model. Journal of Mathematical Physics, 25(12):3545–3550, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3545_s1.

Nogueira:1984:EMC

[2211] Arnaldo C. R. Nogueira. Electrons moving in a crystal weakly coupled to a random reservoir. Journal of Mathematical Physics, 25(12):3551– 3557, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i12/p3551_s1.

Choy:1984:TIM

[2212] T. C. Choy. Triangular Ising model: Complete expressions for even spin correlations in an exact evaluation of P (h, T). Journal of Mathematical Physics, 25(12):3558–3562, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3558_s1.

GarciaDiaz:1984:EET

[2213] Alberto García Díaz. Erratum: “Electrovac type D solutions with cosmological constant”, [J. Math. Phys. 25, 1951 (1984)]. Journal of Mathematical Physics, 25(12):3563, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i12/p3563_s1. See [1973].

Chi:1985:EHS

[2214] L. K. Chi. Erratum: “On the hydrodynamic self-similar cosmological models”, [J. Math. Phys. 24 (1983), no. 10, 2532–2533; MR 84j:83047]. Journal of Mathematical Physics, 26(1):214, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p214_s1. See

Moffat:1985:EHD

[2215] J. W. Moffat. Erratum: “Higher-dimensional Riemannian geometry and quaternion and octonion spaces”, [J. Math. Phys. 25 (1984), no. 2, 347–350; MR 85d:83072]. Journal of Mathematical Physics, 26(1): 214, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p214_s2. See [1749].

Battaglia:1985:ERT

[2216] Franco Battaglia and Thomas F. George. Erratum: “A rule for the total number of topologically distinct Feynman diagrams”, [J. Math. Phys. 25, 3489 (1984)]. Journal of Mathematical Physics, 26(4):882, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p882_s2. See [2201].

Perez:1985:ESG

[2217] J. Fernando Perez, C. P. Malta, and F. A. B. Coutinho. Erratum: “On some general properties of the point spectrum of three particles moving in one dimension”, [J. Math. Phys. 25, 2589 (1984)]. Journal of Mathematical Physics, 26(4):882, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p882_s1. See [2072].

Gerry:1985:ERC

[2218] Christopher C. Gerry. Erratum: “Remarks on canonical transformations in phase-space path integrals”, [J. Math. Phys. 24 (1983), no. 4, 874–877; MR 84g:81016]. Journal of Mathematical Physics, 26(6):1436, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1410].

Glasser:1985:ENC

[2219] M. L. Glasser. Erratum: “A novel class of Bessel function integrals”, [J. Math. Phys. 25 (1984), no. 10, 2933–2934; MR 85i:33007]. Journal of Mathematical Physics, 26(8):2082, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i8/p2082_s1. See [2120].

vanHaeringen:1985:ENI

[2220] H. van Haeringen and L. P. Kok. Erratum: “New inequalities for the Coulomb T matrix in momentum space”, [J. Math. Phys. 25 (1984), no. 10, 3033–3038; MR 85m:81129]. Journal of Mathematical Physics, 26(11):3016, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p3016_s1. See [2133].

Leach:1986:CAP

[2221] P. G. L. Leach. Comment on an aspect of a paper: “Polynomial constants of motion in flat space”, [J. Math. Phys. 25 (1984), no. 12, 3474–3478; MR 86e:70007 ] by G. Thompson. Journal of Mathematical Physics, 27 (1):153–156, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2198] and erratum [2225].

Cronstrom:1986:ETB

[2222] C. Cronström and J. Mickelsson. Erratum: “On topological boundary characteristics in nonabelian gauge theory”, [J. Math. Phys. 24 (1983), no. 10, 2528–2531; MR 85e:81083]. Journal of Mathematical Physics, 27 (1):419, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p419_s1. See [1648].

Harrell:1986:EHO

[2223] E. M. Harrell II. Erratum: “Hamiltonian operators with maximal eigenvalues”, [J. Math. Phys. 25 (1984), no. 1, 48–51; MR 85e:35089]. Journal of Mathematical Physics, 27(1):419, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p419_s2. See [1700].

Odundun:1986:EOT

[2224] Olúwolé Òdúndún. Erratum: “Operators for the two-dimensional harmonic oscillator in an angular momentum basis”, [J. Math. Phys. 24 (1983), no. 9, 2340–2344; MR 84m:81072]. Journal of Mathematical Physics, 27(5):1492, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i5/p1492_s1. See [1621].

Leach:1986:ECA

[2225] P. G. L. Leach. Erratum: “Comment on an aspect of a paper”. Journal of Mathematical Physics, 27(9):2445–2446, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2198, 2221].